### v1

# Another look at inference after prediction

Abstract
Prediction-based (PB) inference is increasingly used in applications where the outcome of interest is difficult to obtain, but its predictors are readily available. Unlike traditional inference, PB inference performs statistical inference using a partially observed outcome and a set of covariates by leveraging a prediction of the outcome generated from a machine learning (ML) model. Motwani and Witten (2023) recently revisited two innovative PB inference approaches for ordinary least squares. They found that the method proposed by Wang et al. (2020) yields a consistent estimator for the association of interest when the ML model perfectly captures the underlying regression function. Conversely, the prediction-powered inference (PPI) method proposed by Angelopoulos et al. (2023a) yields valid inference regardless of the model’s accuracy. In this paper, we study the statistical efficiency of the PPI estimator. Our analysis reveals that a more efficient estimator, proposed 25 years ago by Chen and Chen (2000), can be obtained by simply adding a weight to the PPI estimator. We also contextualize PB inference with methods from the economics and statistics literature dating back to the 1960s. Our extensive theoretical and numerical analyses indicate that the Chen and Chen (CC) estimator offers a balance between robustness to ML model specification and statistical efficiency, making it the preferred choice for use in practice.
Keywords: Missing data, prediction-based inference, semi-parametric inference
1 Introduction
In the rapidly evolving landscape of data science, accurately predicting an outcome is often just an initial step in scientific inquiry. Researchers are increasingly turning to machine learning (ML) models to obtain predictions for outcomes that are prohibitively time-consuming or expensive to directly measure, and then drawing statistical inferences from the ML model’s predictions (Hoffman et al., 2024; Fan et al., 2024; Gao et al., 2024; Miao and Lu, 2024). This prediction-based (PB) inference scenario was introduced by Wang et al in the context of global health and genomics applications, and is comprised of two steps outlined in Figure 1 (Wang et al., 2020). First, the outcome of interest is imputed using a prediction from a pre-trained model. Then, the predictions are used to draw statistical inference on the parameter of interest, such as the association between an outcome and a set of covariates.
Figure 1: Prediction-based (PB) inference. $Y$ is the outcome of interest, which is only partially observed, ${\bf{Z}}$ is a set of predictors of $Y$, $\widehat{Y}$ is the prediction of $Y$ based on ${\bf{Z}}$, and ${\bf{X}}$ is a set of covariates. The focus of scientific inquiry is on the association between ${\bf{X}}$ and $Y$. Traditional statistical inference relies solely on the data on ${\bf{X}}$ and $Y$ while PB inference incorporates $\widehat{Y}$ to maximize the use of all available data. Dark red represents labeled observations while light red represents unlabeled observations.
In the last several years, applications of PB inference have arisen in diverse fields, spanning genetics, biology, and medicine. For example, population biobanks, such as the UK Biobank ($N\approx$ 500K individuals), provide an unprecedented data source for high-throughput genetic research. However, many traits and disease outcomes are only partially observed due to the time and cost of measurement. ML models are commonly used to predict missing outcomes based on fully observed baseline data and the predictions are utilized for genetic discovery (Tong et al., 2020; Alipanahi et al., 2021; Cosentino et al., 2023; Dahl et al., 2023; An et al., 2023; McCaw et al., 2024; Liu et al., 2024; Miao et al., 2024). Similarly, in the field of structural biology, AlphaFold provides highly accurate structural predictions without the need for expensive and labor-intensive experiments for proteomic studies (Jumper et al., 2021). In the analysis of electronic health records data, large language models can rapidly ascertain disease status at an accuracy competitive with manual chart review by experts, enabling downstream epidemiological and clinical studies (Yang et al., 2022; Zhou et al., 2022; Alsentzer et al., 2023).
Given the relevance of PB inference, Motwani and Witten (2023) recently revisited two innovative methods in this area. In the context of least squares estimation, they showed that the approach proposed by Wang et al. (2020) provides a consistent estimator for the association between an outcome and a set of covariates only in restrictive scenarios. For example, when the pre-trained model perfectly captures the regression function. This is a significant practical limitation as it is essential that PB inference methods yield valid inference in the likely scenario that the prediction model is imperfect. In contrast, the prediction powered inference (PPI) approach proposed by Angelopoulos et al. (2023a) possesses this necessary property and yields valid inference regardless of the quality of the model.
Here we build on the work on Motwani and Witten to study the statistical efficiency of the PPI method. We highlight that a more efficient estimator can be obtained by simply adding a weight to the PPI estimator, an approach that was proposed by Chen & Chen (CC) more than 25 years ago in the context of double sampling designs (Chen and Chen, 2000). We begin by formally introducing the PB inference problem and reviewing the contributions of Motwani and Witten in Section 2. We then study the asymptotic properties of the PPI estimator and introduce the CC estimator in Section 3. Our key finding is that the CC estimator is optimal with respect to statistical efficiency in a class of estimators to which the PPI estimator belongs. We also point to earlier literature from the statistics and economics literature dating back to the 1960s that provides important historical context for the PB inference problem in Section 4. In Section 5, we evaluate our findings with extensive simulation studies. Our results show that the CC estimator offers a balance between robustness to ML model specification and statistical efficiency, making it the preferred choice for use in practice. We close with additional discussion in Section 6.
2 Problem setting & background
2.1 Notation and Assumptions
We let $Y$ denote the outcome of interest and ${\bf{Z}}$ a set of predictors. $Y$ is only partially observed and a pre-trained ML model is available to predict it on the basis of ${\bf{Z}}$. The ML model, $\hat{f}:\mathcal{Z}\rightarrow\mathcal{Y}$, maps from the predictor space, $\mathcal{Z}$, to the outcome space, $\mathcal{Y}$. The predictions of $Y$ are denoted as $\widehat{Y}=\hat{f}({\bf{Z}})$. We let ${\bf{X}}=(1,X_{1},\dots,X_{p})^{{\sf\scriptscriptstyle{T}}}$ be a set of covariates that are potentially associated with $Y$ and that are the focus of scientific inquiry. ${\bf{X}}$ may also be predictive of $Y$ and these covariates may or may not be included in ${\bf{Z}}$. We assume that ${\bf{X}}$ and ${\bf{Z}}$ are both fully observed. The available data therefore consists of a labeled dataset and an independent unlabeled dataset, respectively denoted as
=$\mathscr{L}=\{(Y_{i},{\bf{Z}}_{i}^{{\sf\scriptscriptstyle{T}}},{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}})^{{\sf\scriptscriptstyle{T}}}\mid i=1,\dots,n_{\text{\tiny{lab}}}\}\quad\mbox{and}\quad\mathscr{U}=\{({\bf{Z}}_{i}^{{\sf\scriptscriptstyle{T}}},{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}})^{{\sf\scriptscriptstyle{T}}}\mid i=n_{\text{\tiny{lab}}}+1,\dots,n\},$
where $n=n_{\text{\tiny{lab}}}+n_{\text{\tiny{unlab}}}$, $n_{\text{\tiny{unlab}}}$ is the number of unlabeled observations, and $n_{\text{\tiny{lab}}}/n\to\pi$ as $n\to\infty$ for $\pi\in(0,1)$.
We focus on the case in which $\mathscr{L}$ and $\mathscr{U}$ are independent and identically distributed samples from the population of interest. More complex scenarios have been the focus of recent work (Angelopoulos et al., 2023a; Zrnic and Candès, 2024a). Additionally, $\hat{f}$ is assumed to be pre-trained with a dataset that is independent of $\mathscr{L}$ and $\mathscr{U}$ and is therefore treated as “fixed” (i.e., the variation in estimating $\hat{f}$ is not of interest in downstream inference). In fact, $\hat{f}$ may be a black box function in that its training data and specification are not be available to the users. Approaches that address uncertainty in the estimation of $\hat{f}$ have recently been proposed in Zrnic and Candès (2024b).
2.2 Parameter of interest
Following Motwani and Witten (2023), we consider least squares estimation and aim to make inference on the population parameter
$\bm{\beta}^{*}\coloneq\operatorname*{arg\,min}_{\bm{\beta}}\mathbb{E}[(Y-{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\beta})^{2}]=[\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})]^{-1}\mathbb{E}({\bf{X}}Y),$
which quantifies the association between $Y$ and ${\bf{X}}$. Throughout, we use superscripts to denote different estimators of $\bm{\beta}^{*}$. For example, the simplest unbiased estimator of $\bm{\beta}^{*}$, denoted as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, uses only $\mathscr{L}$ for estimation and is the solution to the estimating equation
$\sum_{i=1}^{n_{\text{\tiny{lab}}}}\bm{\phi}(Y_{i},{\bf{X}}_{i};\bm{\beta})\coloneq\sum_{i=1}^{n_{\text{\tiny{lab}}}}{\bf{X}}_{i}(Y_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\bm{\beta})=\bm{0}.$
2.3 Motivation for PB inference
The goal of PB inference is to utilize the abundant realizations of $\widehat{Y}$, which can be computed in the unlabeled data, to provide valid and efficient inference on the association between $Y$ and ${\bf{X}}$. Motwani and Witten consider the question of validity and we study efficiency in Section 3. With respect to validity, PB inference should enable users to draw correct conclusions regarding the magnitude and direction of $\bm{\beta}^{*}$ and to perform reliable uncertainty quantification. Incorporating information from $\widehat{Y}$ should not change the parameter being estimated and the credibility of inferences should not depend on the quality of predictions from $\hat{f}$.
The need for PB inference is highlighted by limitations of “naive” inference, in which inference on $\widehat{Y}$ is performed in lieu of inference on $Y$ (Wang et al., 2020). The parameter targeted by naive inference is
$\bm{\gamma}^{*}\coloneq\operatorname*{arg\,min}_{\bm{\gamma}}\mathbb{E}[(\widehat{Y}-{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma})^{2}]=[\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})]^{-1}\mathbb{E}({\bf{X}}\widehat{Y}).$
A least squares estimator for $\bm{\gamma}^{*}$, denoted as $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$, is the solution to
$\sum_{i=n_{\text{\tiny{lab}}}+1}^{n}\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma})\coloneq\sum_{i=n_{\text{\tiny{lab}}}+1}^{n}{\bf{X}}_{i}(\widehat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma})=\bm{0}.$
Critically, $\bm{\gamma}^{*}$ bears no necessary correspondence with $\bm{\beta}^{*}$ except under highly restrictive scenarios, such as when ${\bf{Z}}$ contains ${\bf{X}}$ and the ML model captures the true regression function. The fundamental problem is that $(\widehat{Y},{\bf{X}}^{{\sf\scriptscriptstyle{T}}})^{{\sf\scriptscriptstyle{T}}}$ has a different distribution than $(Y,{\bf{X}}^{{\sf\scriptscriptstyle{T}}})^{{\sf\scriptscriptstyle{T}}}$. To overcome these limitations, PB inference methods utilize $\mathscr{L}$ together with $\mathscr{U}$ to essentially balance the relative contributions of $Y$ and $\widehat{Y}$ in making inference on $\bm{\beta}^{*}$.
2.4 PB inference methods
Motwani and Witten revisited two innovative strategies for PB inference. The first is the proposal of Wang et al. (2020), which is a two-step procedure to incorporate $\widehat{Y}$ into estimation. First, $\mathscr{L}$ is used to model the relationship between $Y$ and $\widehat{Y}$ to address potential inaccuracies in the predictions. Then, this model is used together with $\mathscr{U}$ to make inference on $\bm{\beta}^{*}$, with analytical or bootstrap approaches described. In their illuminating work, Motwani and Witten show that the estimator of Wang et al is consistent for $\bm{\beta}^{*}$ under the strong and unrealistic assumption that the ML model captures the true regression function. In other scenarios, the Wang et al estimator may not be consistent for $\bm{\beta}^{*}$ even while its standard error converges to 0 as $(n_{\text{\tiny{lab}}},n_{\text{\tiny{unlab}}})\to\infty$. Consequently, the familiar Wald statistic will not converge in distribution under the null, resulting in type I error that is not controlled at the nominal level and confidence intervals with inadequate coverage.
Motwani and Witten also study the PPI method, which has the crucial property that inference remains valid regardless of the quality of the predictions. Stated otherwise, the PPI method does not require that $\hat{f}$ capture the true regression function for the estimator to maintain consistency or to yield reliable hypothesis testing and interval estimation. The proposed estimator utilizes a common statistical technique involving the augmentation of an initial consistent estimator for the parameter of interest with an additional term that converges in probability to 0. This approach dates back to at least Robins et al. (1994), and has been applied in multiple settings to improve estimation efficiency outside of the PB inference problem (e.g., Chen and Chen (2000); Tsiatis (2006); Tian et al. (2012); Wang and Wang (2015)). We return to this point in Section 4.
The PPI estimator is
$\hat{\bm{\beta}}^{\text{\tiny{PPI}}}=\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}\right),$
where $\hat{\bm{\gamma}}^{\text{\tiny{lab}}}$ is the solution to $\sum_{i=1}^{n_{\text{\tiny{lab}}}}\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma})=\boldsymbol{0}$ [^1] Angelopoulos et al. (2023a) express $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ as $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}-\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\beta}}^{\text{\tiny{lab}}}\right)$ where $\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\beta}}^{\text{\tiny{lab}}}\right)$ is referred to as a “rectifier.” The rectifier serves the purpose of correcting potential bias in $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ due to inaccuracies in $\widehat{Y}$ for $Y$. We reframe the PPI method to study its statistical efficiency.. In contrast to the approach of Wang et al, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is consistent for $\bm{\beta}^{*}$ and $n^{\frac{1}{2}}(\hat{\bm{\beta}}^{\text{\tiny{PPI}}}-\bm{\beta}^{*})$ converges weakly to a zero-mean multivariate normal distribution regardless of the quality of the prediction. The proposed method therefore asymptotically controls the type I error for Wald-based hypothesis testing and provides confidence intervals with the nominal coverage level.
3 Efficiency analysis
As PB inference methods suppose the availability of at least a small labeled data set, a natural baseline estimator for comparison is $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, which simply regresses $Y$ on ${\bf{X}}$ within $\mathscr{L}$. For a PB inference method to be preferred over this baseline, it should not only yield valid inference, but also improve statistical efficiency relative to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when the prediction is informative of $Y$, and perform no worse than estimation based on $\mathscr{L}$ when $\widehat{Y}$ is uninformative of $Y$. For the remainder of this paper, we study the latter property and analyze an estimator very similar to $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ given by
$\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}=\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right),$
where $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$ is the solution to $\sum_{i=1}^{n}\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma})=\boldsymbol{0}$. The only difference between $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is that all of the data, rather than the unlabeled data, is used to estimate $\bm{\gamma}^{*}$ in the augmentation term (i.e., $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ is replaced by $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$). When $\pi\approx 0$, $\hat{\bm{\gamma}}^{\text{\tiny{all}}}\approx\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ so that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}\approx\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$. In general, it can be shown that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ gains efficiency over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ whenever $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ gains over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ can be more efficient than $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ when $\widehat{Y}$ is informative of $Y$ (see Appendix B, specifically Section B.3, for a detailed analysis). Additionally, estimators of this form are commonly used in the semi-parametric inference literature, which we draw upon in the subsequent section in our efficiency analysis.
3.1 Asymptotic analysis of the PPI estimator
To study the asymptotic variance of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, Theorem 1 provides the influence function expansion of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$. All of the asymptotic results for $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ are detailed in Appendix A.
Theorem 1 (Influence function expansion of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$)
Let $R_{i}$ be the indicator of whether the $i$th observation is labeled. Then,
$n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}-\bm{\beta}^{*}\right)=n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})+o_{p}(1),$
where ${\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})={\bf{A}}\left\{\left[\ \bm{\phi}(Y_{i},{\bf{X}}_{i};\bm{\beta}^{*})\right]\frac{R_{i}}{\pi}+\left[\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma}^{*})\right]\frac{\pi-R_{i}}{\pi}\right\}$ and ${\bf{A}}=[\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})]^{-1}$. Therefore, $n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}-\bm{\beta}^{*}\right)$ converges weakly to a zero-mean multivariate normal distribution with variance-covariance
$\mathbb{E}\left[{\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})\hskip 2.84526pt{\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})^{{\sf\scriptscriptstyle{T}}}\right].$
We use Theorem 1 to compare the asymptotic variance-covariance of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ with that of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. Corollary 1.1 shows that the benefit of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ scales directly with the predictiveness of $\widehat{Y}$ for $Y$ and inversely with the proportion of labeled data, $\pi$.
Corollary 1.1 (Comparison of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ with $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$)
Let $\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}=\mathbb{E}\left(\epsilon_{\widehat{Y}}^{2}{\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\right)$, $\boldsymbol{\Sigma}_{Y,\widehat{Y}}=\mathbb{E}\left(\epsilon_{Y}\epsilon_{\widehat{Y}}{\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\right)$, $\epsilon_{Y}=Y-{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\beta}^{*}$ and $\epsilon_{\widehat{Y}}=\widehat{Y}-{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma}^{*}$. The difference in the asymptotic variance-covariance (aVar) of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is
$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})$ $=n\left[\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{lab}}})-\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})\right]=\left(\frac{1}{\pi}-1\right){\bf{A}}\left(2\boldsymbol{\Sigma}_{Y,\widehat{Y}}-\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}\right){\bf{A}}.$
$\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is therefore more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if $2\boldsymbol{\Sigma}_{Y,\widehat{Y}}\succ\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}$. A sufficient condition for $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ to have lower variance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ is that
$\mathbb{E}\left[\left(\epsilon_{\widehat{Y}}-\epsilon_{Y}\right)^{2}\mid{\bf{X}}\right]<\mathbb{E}\left(\epsilon_{Y}^{2}\mid{\bf{X}}\right).$
More specifically, the sufficient condition in Corollary 1.1 coincides with the intuition that incorporating unlabeled data into estimation can improve efficiency when $\widehat{Y}$ has additional predictive power for $Y$ beyond what is already explained by the regression model of $Y$ on ${\bf{X}}$. This condition is similar to the result for mean estimation provided in the Supplementary Materials of Angelopoulos et al. (2023a) and can be used to describe several scenarios when $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is guaranteed to have lower asymptotic variance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when the ML model captures the true regression function (see Appendix A.4). Here we consider a simple example to demonstrate when $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ gains efficiency over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ under homoskedasticity.
Example 1 (Efficiency gain of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ under homoskedasticity)
In the homoskedastic setting where $\mathbb{E}(\epsilon_{Y}^{2}\mid{\bf{X}})=\sigma_{Y}^{2}$, $\mathbb{E}(\epsilon_{\widehat{Y}}^{2}\mid{\bf{X}})=\sigma_{\widehat{Y}}^{2}$, and $\mathbb{E}(\epsilon_{Y}\epsilon_{\widehat{Y}}\mid{\bf{X}})=\sigma_{Y,\widehat{Y}}$,
$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})=\left(\frac{1}{\pi}-1\right)\left[2\sigma_{Y,\widehat{Y}}-\sigma_{\widehat{Y}}^{2}\right]{\bf{A}},$
and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if
$\rho_{Y,\widehat{Y}}>\frac{1}{2}\frac{\sigma_{\widehat{Y}}}{\sigma_{Y}},$
where $\rho_{Y,\widehat{Y}}=\sigma_{Y,\widehat{Y}}/\sigma_{Y}\sigma_{\widehat{Y}}$ is the correlation of residuals for $Y$ and $\widehat{Y}$ given ${\bf{X}}$.
The improvement of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ is therefore stronger with increasing correlation of residuals between $Y$ and $\widehat{Y}$ and decreasing variability in the residuals of $\widehat{Y}$. Therefore, $\widehat{Y}$ must be a strong predictor for $Y$ after adjusting for ${\bf{X}}$, though it does not necessarily need to be derived from a correctly specified ML model. In particular, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ gains efficiency if and only if the correlation of residuals is at least half of the ratio of residual standard deviations of $\widehat{Y}$ to that of $Y$.
3.2 Asymptotic analysis of the CC estimator
To further analyze the efficiency of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ we draw on a landmark paper that introduced a class of augmented inverse probability weighted (AIPW) estimators which includes all regular asymptotic linear estimators (Robins et al., 1994). In the current setting, where the goal is to estimate linear regression coefficients $\bm{\beta}^{*}$ with $Y$ partially observed, but $\widehat{Y}$ fully observed, these estimators have influence functions of the form
$\text{IF}(Y,\widehat{Y},{\bf{X}};\bm{\omega})={\bf{A}}\left[\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\frac{R}{\pi}+\bm{\omega}(\widehat{Y},{\bf{X}})\frac{\pi-R}{\pi}\right],$
with $\bm{\omega}(\widehat{Y},{\bf{X}})$ being a function of $\widehat{Y}$ and ${\bf{X}}$. The optimal choice of $\bm{\omega}(\widehat{Y},{\bf{X}})$ with respect to asymptotic variance is
$\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})=\mathbb{E}[\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\mid\widehat{Y},{\bf{X}}].$
By Theorem 1, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is an AIPW estimator and can be seen as an approximation to the optimal estimator with $\bm{\omega}(\widehat{Y},{\bf{X}})=\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})$. The PPI estimator is therefore not the most efficient estimator in its class and, from Corollary 1.1, can even be less efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when $\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}\succ 2\boldsymbol{\Sigma}_{Y,\widehat{Y}}$. A PB inference strategy should enhance statistical efficiency relative to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ whenever possible and, at the very least, should not be negatively impacted by incorporating unlabeled data into estimation. This raises concerns about the practical utility of the PPI estimator.
To improve upon $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, one approach is to assume a parametric model for $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$ that can be estimated from $\mathscr{L}$ for analysis. However, it is often difficult to specify an appropriate parametric model for $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$, particularly when the relationship between $Y$ and $\widehat{Y}$ is unknown. Moreover, when the model for $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$ is misspecified, an efficiency gain is not guaranteed. To overcome these difficulties, Chen and Chen (2000) proposed a particular choice of $\bm{\omega}(\widehat{Y},{\bf{X}})$ that is guaranteed to be at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and that does not require modeling of $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$. Their approach is motivated by the double-sampling design which, apart from differences in terminology and notation (i.e., they regard $\widehat{Y}$ as a mismeasured version of $Y$), is identical to the partially labeled data setting considered here. The CC estimator approximates $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$ with
$\bm{\omega}^{\text{\tiny{CC}}}(\widehat{Y},{\bf{X}})=\text{Cov}[\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*}),\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})]\{\text{Var}[\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})]\}^{-1}\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*}).$ (1)
The estimator with this influence function can be written as
$\hat{\bm{\beta}}^{\text{\tiny{CC}}}=\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\widehat{{\bf{W}}}^{\text{\tiny{CC}}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right),$
where $\widehat{{\bf{W}}}^{\text{\tiny{CC}}}$ is a consistent estimator of
${{\bf{W}}}^{\text{\tiny{CC}}}={\bf{A}}{\text{Cov}}[\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*}),\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})]\{{\text{Var}}[\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})]\}^{-1}{\bf{A}}^{-1}$
obtained with $\mathscr{L}$. Therefore, a simple way to ensure that an augmented estimator like the PPI estimator has asymptotic variance at least as low as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ is to add a simple weight function in front of the augmentation term. In fact, the framework of Robins et al. (1994) implies that all regular asymptotically linear estimators for $\bm{\beta}^{*}$ augment $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ with an augmentation term that incorporates information from $\widehat{Y}$. We present the influence function expansion for $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ in Theorem 2 to further study its asymptotic variance and compare it with $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$. All of the asymptotic results provided for $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ are detailed in Appendix C.
Theorem 2 (Influence function expansion of $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$)
Following Chen and Chen (2000),
$n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{CC}}}-\bm{\beta}^{*}\right)=n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{IF}}^{\text{\tiny{CC}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})+o_{p}(1),$
where ${\bf{IF}}^{\text{\tiny{CC}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})={\bf{A}}\left[\bm{\phi}(Y_{i},{\bf{X}}_{i};\bm{\beta}^{*})\frac{R_{i}}{\pi}+\bm{\omega}^{\text{\tiny{CC}}}(\widehat{Y}_{i},{\bf{X}}_{i})\frac{\pi-R_{i}}{\pi}\right]$ and $\bm{\omega}^{\text{\tiny{CC}}}(\widehat{Y},{\bf{X}})$ is defined in (1). Therefore, $n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{CC}}}-\bm{\beta}^{*}\right)$ converges weakly to a zero-mean multivariate normal distribution with variance-covariance
$\hskip 2.84526pt\mathbb{E}\left[{\bf{IF}}^{\text{\tiny{CC}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})\hskip 2.84526pt{\bf{IF}}^{\text{\tiny{CC}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})^{{\sf\scriptscriptstyle{T}}}\right]\hskip 2.84526pt.$
The subsequent corollaries apply Theorem 2 to show our two key results: (1) $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and (2) $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ has the highest asymptotic efficiency within the class of estimators with a weighted linear augmentation term. Together these results illustrate that the CC method is both robust to ML model specification and optimizes statistical efficiency, making it preferable to the PPI method in practice.
Corollary 2.1 ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$)
The difference in the asymptotic variance-covariance of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is
$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{CC}}})=\left(\frac{1}{\pi}-1\right){\bf{A}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}\boldsymbol{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}{\bf{A}}\succeq{\bf{0}}.$
Therefore, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is guaranteed to be at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
Corollary 2.2 ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is the most efficient estimator in its class)
Let the mean squared error (MSE) of an estimator $\hat{\bm{\beta}}$ for $\bm{\beta}^{*}$ be defined as $\mathbb{E}[(\hat{\bm{\beta}}-\bm{\beta}^{*})^{{\sf\scriptscriptstyle{T}}}(\hat{\bm{\beta}}-\bm{\beta}^{*})]$. Analogously to the well-known result for the best linear predictor (e.g., Bickel and Doksum (2015), page 40), $\widehat{{\bf{W}}}^{\text{\tiny{CC}}}$ is the weight, ${\bf{W}}$, that minimizes the asymptotic MSE of estimators of the form $\hat{\bm{\beta}}^{\text{\tiny{lab}}}-{{\bf{W}}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right)$ for estimating $\bm{\beta}^{*}$.
From Corollary 2.1, we note that in contrast to $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})$, the covariance term in $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{CC}}})$ is normalized by the residual variance of the $\widehat{Y}$ to scale the covariance according to the precision of $\widehat{Y}$. This scaling essentially enables $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ to always be as least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. Coming back to the example from the previous section, we can gain more intuition as to why $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is guaranteed to be at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ while $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ does not possess this property.
Example 1 continued (Efficiency gain of $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ under homoskedasticity) We revisit the case of homoskedasticity and recall that
$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})$ $=\left(\frac{1}{\pi}-1\right)\left(2\sigma_{Y,\widehat{Y}}-\sigma_{\widehat{Y}}^{2}\right){\bf{A}}$
$=\left(\frac{1}{\pi}-1\right)\left[-\left(\sigma_{Y}\rho_{Y,\widehat{Y}}-\sigma_{\widehat{Y}}\right)^{2}+\sigma_{Y}^{2}\rho_{Y,\widehat{Y}}^{2}\right]{\bf{A}}.$
Theorem 2 can be used to show that
$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{CC}}})$ $=\left(\frac{1}{\pi}-1\right)\left(\frac{\sigma_{Y,\widehat{Y}}^{2}}{\sigma_{\widehat{Y},\widehat{Y}}^{2}}\right){\bf{A}}=\left(\frac{1}{\pi}-1\right)\sigma_{Y}^{2}\rho_{Y,\widehat{Y}}^{2}{\bf{A}}.$
$\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ provided that $\rho_{Y,\widehat{Y}}>0$ and the gain in efficiency scales directly with $\rho_{Y,\widehat{Y}}$ and $\sigma_{Y}^{2}$. In contrast, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ will only be more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ provided that $\rho_{Y,\widehat{Y}}>\frac{1}{2}\frac{\sigma_{\widehat{Y}}}{\sigma_{Y}}$ and the gain in efficiency will only scale in this manner with the additional condition that $\rho_{Y,\widehat{Y}}=\frac{\sigma_{\widehat{Y}}}{\sigma_{Y}}$.
4 Related literature
Existing PB inference literature has established connections with semi-supervised inference, semi-parametric inference, missing data, and measurement error (Angelopoulos et al., 2023a, b; Gan and Liang, 2023; Miao et al., 2023; Zrnic and Candès, 2024a, b). Here we draw connections between PB inference methods and methods from the economics and statistics literature dating back to the 1960s. These connections reveal that aspects of the PB inference problem have been explored in various areas, albeit with differences in motivation, terminology, and notation. Given the close relationship between the PPI and CC methods, future research in PB inference could benefit from a more comprehensive review of these and other related approaches than what is provided here.
4.1 Seemingly unrelated regression & its use in statistical genetics
Seemingly unrelated regression (SUR) was introduced by economist Arnold Zellner in 1962 for simultaneous estimation of multiple related regression models (Zellner, 1962; Swamy and Mehta, 1975; Schmidt, 1977). The regression models have different outcome variables, but are assumed to have correlated errors. SUR leverages the correlation of the error terms to produce more efficient estimators compared to estimating each regression independently. In the current setting, the SUR problem can be stated formally as
$Y={\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\beta}^{*}+\epsilon\quad\mbox{and}\quad\widehat{Y}={\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma}^{*}+\epsilon^{\prime}\quad$
where $\mathbb{E}(\epsilon\mid{\bf{X}})=0$ and $\mathbb{E}(\epsilon^{\prime}\mid{\bf{X}})=0.$ An estimator for $\bm{\beta}^{*}$ that is based on augmenting $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ was proposed by Conniffe (1985) and is given by:
$\hat{\bm{\beta}}^{\text{\tiny{SUR}}}=\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\widehat{W}^{\text{\tiny{SUR}}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right)\quad\mbox{where}\quad\widehat{W}^{\text{\tiny{SUR}}}=\frac{\sum_{i=1}^{n_{\text{\tiny{lab}}}}(Y_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\hat{\bm{\beta}}^{\text{\tiny{lab}}})(\widehat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\hat{\bm{\gamma}}^{\text{\tiny{lab}}})}{\sum_{i=1}^{n_{\text{\tiny{lab}}}}(\widehat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\hat{\bm{\gamma}}^{\text{\tiny{lab}}})^{2}}.$
Note that $\widehat{W}^{\text{\tiny{SUR}}}$ is an estimator of the scalar $W^{\text{\tiny{SUR}}}=\frac{\mathbb{E}[\text{Cov}(\epsilon,\epsilon^{\prime}\mid{\bf{X}})]}{\mathbb{E}[\text{Var}(\epsilon^{\prime}\mid{\bf{X}})]}$. Similar to the PPI and CC methods, the contribution of the augmentation term to estimation increases with the predictiveness of $\widehat{Y}$ for $Y$. In the extreme case where $\widehat{Y}$ is a perfect predictor for $Y$ such that $\widehat{Y}=Y$, $\widehat{W}^{\text{\tiny{SUR}}}=1$ and $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ is simply $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$. In the homoskedastic case when $\text{Var}(\epsilon\mid{\bf{X}})$ and $\text{Cov}(\epsilon,\epsilon^{\prime}\mid{\bf{X}})$ are constant and do not depend on ${\bf{X}}$, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is asymptotically equivalent to $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ (Chen and Chen, 2000). Moreover, when the conditional distribution of $(\epsilon,\epsilon^{\prime})$ given ${\bf{X}}$ is bivariate normal with mean $\boldsymbol{0}$ and has constant variance-covariance, $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ (and thus $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ as well) is the maximum likelihood estimator of $\bm{\beta}^{*}$ and fully efficient under the assumed parametric model (Conniffe, 1985). In other scenarios, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ achieves a lower asymptotic variance than $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$.
The SUR estimator was recently applied for a PB inference problem inspired by genome-wide association studies conducted with population biobanks. These studies often rely on pre-trained models to impute outcomes of interest (i.e., phenotypes or traits) as they are partially missing due to the time and expense of ascertainment (McCaw et al., 2024). The proposed method, “SynSurr”, assumes the true and predicted outcomes follow a bivariate normal distribution. The SynSurr estimator is equivalent to $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ and asymptotically equivalent to $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ when $\text{Var}(\epsilon\mid{\bf{X}})$ and $\text{Cov}(\epsilon,\epsilon^{\prime}\mid{\bf{X}})$ are constant. Building on the SynSurr estimator, Miao et al. (2024) proposed the “POP-GWAS” estimator for a variety of regression models utilizing the augmentation technique. The POP-GWAS estimator is
$\hat{\bm{\beta}}^{\text{\tiny{POP}}}=\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\widehat{W}^{\text{\tiny{POP}}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}\right)\quad\mbox{where}\quad\widehat{W}^{\text{\tiny{POP}}}=\frac{n_{\text{\tiny{unlab}}}}{n}\widehat{\text{Corr}}(Y,\widehat{Y}\mid{\bf{X}}).$
In contrast to the SUR estimator, $\hat{\bm{\beta}}^{\text{\tiny{POP}}}$ uses $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ instead of $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$ for the augmentation, and the weight, $\widehat{W}^{\text{\tiny{POP}}}$ is not generally a consistent estimator of $W^{\text{SUR}}$. $\hat{\bm{\beta}}^{\text{\tiny{POP}}}$ is therefore less efficient than $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ when the residuals are bivariate normal with constant variance-covariance and less efficient than $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ in general.
4.2 Augmented estimation for quantifying a treatment effect
Outside of statistical genetics, the augmented estimation approach used by the PPI and CC methods has been widely used for treatment effect estimation in the statistics and biostatistics literature (e.g., Leon et al. (2003); Lu and Tsiatis (2008); Tsiatis et al. (2008); Gilbert et al. (2009); Zhang and Gilbert (2010); Tian et al. (2012); Yang and Ding (2020)). Many of these approaches build on methods from the survey sampling literature (e.g., Cassel et al. (1976); Särndal et al. (2003)).
For example, Tsiatis et al. (2008) consider covariate adjusted analyses for comparing two treatments in randomized clinical trials. Letting $X$ be an indicator for randomization to the treated or control arms, the focus of inference is on $\beta$ in the model
$\mathbb{E}(Y\mid X)=\alpha+\beta X$
in order to estimate the treatment effect $\Delta=\mathbb{E}(Y\mid X=1)-\mathbb{E}(Y\mid X=0)$. A standard estimator of $\Delta$ is $\widehat{\Delta}=\overline{Y}^{1}-\overline{Y}^{0}$, where $\overline{Y}^{1}=n_{1}^{-1}\sum_{i=1}^{n}Y_{i}X_{i}$ and $\overline{Y}^{0}=n_{0}^{-1}\sum_{i=1}^{n}Y_{i}(1-X_{i})$ are the empirical means of $Y$ in the treated and control groups of sizes $n_{1}$ and $n_{0}$, respectively. To utilize information in baseline covariates, ${\bf{Z}}$, Tsiatis et al. (2008) propose an augmented estimator of the treatment effect given by
$\widehat{\Delta}-\sum_{i=1}^{n}(X_{i}-\bar{X})[n_{0}^{-1}h^{(0)}({\bf{Z}}_{i})+n_{1}^{-1}h^{(1)}({\bf{Z}}_{i})],$ (2)
where $\bar{X}=n^{-1}\sum_{i=1}^{n}X_{i}$ and $h^{(k)}({\bf{Z}})$ for $k=0,1$ are arbitrary scalar functions of ${\bf{Z}}$. As $X\perp{\bf{Z}}$ are independent due to randomization, the augmentation term in 2 converges in probability to 0, analogously to the augmentation terms in the PPI and CC estimators. In the context of PB inference, pre-trained models could be used for predicting $Y$ on the basis of ${\bf{Z}}$ for $h^{(k)}({\bf{Z}})$ for $k=0,1$.
4.3 Improving precision of Monte Carlo simulations with control variates
The augmentation technique has also been utilized to improve the efficiency of Monte Carlo estimation. Specifically, Rothery (1982) considers using a control variate, a statistic whose properties are known, to improve precision when estimating power via simulation. For exposition, suppose interest lies in estimating the power of a test $T$ of $H_{0}:\theta=0$ for some scalar parameter $\theta$. Let $S$ denote an existing test of the same hypothesis whose power $\mu_{Z}=\mathbb{E}_{\theta}(Z)$ is known. Define $Y$ as an indicator that test $T$ rejects and $Z$ as an indicator that test $S$ rejects. For a set of $n$ simulation replicates, the baseline estimator of power for $T$ is simply $\overline{Y}=n^{-1}\sum_{i=1}^{n}Y_{i}$ while a control variate augmented estimator is given by:
$\frac{1}{n}\sum_{i=1}^{n}\left[Y_{i}-\lambda\{Z_{i}-\mu_{Z}\}\right].$ (3)
The variance of the augmented estimator is minimized at $\lambda^{*}=\frac{\text{Cov}(Y,Z)}{\text{Var}(Z)}$ Rothery (1982); Davidson and MacKinnon (1992). The augmentation term reduces the variance of the unadjusted power estimate $\overline{Y}$ in proportion to the correlation between $Y$ and $Z$. Building on work from the survey sampling literature, a similar idea was later used for mean estimation in the semi-supervised setting by Zhang et al. (2019).
5 Numerical studies
5.1 Data generation
We perform several simulation studies to compare the performance of the PPI and CC methods. In all settings, we generate two datasets: (i) a training dataset with realizations of $(Y,{\bf{Z}}^{{\sf\scriptscriptstyle{T}}},X_{1})^{{\sf\scriptscriptstyle{T}}}$ to train the ML model and (ii) an independent analysis dataset for evaluating the PB inference methods containing a labeled dataset with realizations of $(Y,{\bf{Z}}^{{\sf\scriptscriptstyle{T}}},X_{1})^{{\sf\scriptscriptstyle{T}}}$ and an unlabeled dataset with realizations of $({\bf{Z}}^{{\sf\scriptscriptstyle{T}}},X_{1})^{{\sf\scriptscriptstyle{T}}}$. We use 10,000 replications of the analysis dataset for each scenario and a single training set as $\hat{f}$ is assumed to be fixed. The size of the training set is fixed at 10,000 and the size of the analysis dataset is also set to $n=10,000$. We let $n_{\text{\tiny{lab}}}=1000,2000,3000,4000,5000$ and $n_{\text{\tiny{unlab}}}=n-n_{\text{\tiny{lab}}}$. For both the training and analysis datasets, $({\bf{Z}}^{{\sf\scriptscriptstyle{T}}},X_{1})^{{\sf\scriptscriptstyle{T}}}$ are generated from a multivariate normal distribution. Each element of $({\bf{Z}}^{{\sf\scriptscriptstyle{T}}},X_{1})^{{\sf\scriptscriptstyle{T}}}$ has a mean of 0, variance of 1, and covariance of 0.4 with the remaining variables. The goal is to conduct inference on the marginal association of $Y$ and $X_{1}$ in a linear model. The parameter of interest is therefore $\beta^{*}_{1}$ defined as
$(\beta^{*}_{0},\beta^{*}_{1})=\operatorname*{arg\,min}_{\bm{\beta}}\mathbb{E}[(Y-(\beta_{0}+\beta_{1}X_{1}))^{2}].$
We consider 2 primary scenarios for evaluating the PB inference methods. In both scenarios, the outcome is generated as
$Y=X_{1}+\theta_{\text{\tiny{INT}}}X_{1}W_{1}+{\bf{W}}^{{\sf\scriptscriptstyle{T}}}\bm{\theta}+\epsilon$
where ${\bf{W}}^{{\sf\scriptscriptstyle{T}}}=(W_{1},\dots,W_{q})$ and the error term $\epsilon$ is simulated independently from a standard normal distribution. We varied the coefficients of $\bm{\theta}$ and $\theta_{\text{\tiny{INT}}}$ as well as $q$ and the composition of ${\bf{Z}}$.
Scenario 1
We consider two sub-scenarios to illustrate Corollary 1.1, that is, that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when the sufficient condition holds. In both sub-scenarios, we let $p=10$, $\bm{\theta}=(1,1,0.5,0.5,0.25,{\bf{0}}_{1\times 5})^{{\sf\scriptscriptstyle{T}}}$, and $\theta_{\text{\tiny{INT}}}=0$. In Scenario 1a, we obtain a strong prediction model by regressing $Y$ on ${\bf{Z}}=(X_{1},W_{1},\dots,W_{10})^{{\sf\scriptscriptstyle{T}}}$ while in Scenario 1b we obtain a weaker prediction model by regressing $Y$ on ${\bf{Z}}=(X_{1},W_{5},X_{1}W_{5})^{{\sf\scriptscriptstyle{T}}}$.
Scenario 2
We consider two sub-scenarios to illustrate the results in Corollary 1.1 and Corollary 2.1, which show that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ can be less efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ with a weak prediction model while $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is at least as efficient $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. In both scenarios, we mimic a common scenario where the prediction model does not contain $X_{1}$ (e.g., in a genome-wide association study). In Scenario 2a, we let $p=9$, $\bm{\theta}=(0.2,0.1\times{\bf{1}}_{1\times 8})^{{\sf\scriptscriptstyle{T}}}$, and $\theta_{\text{\tiny{INT}}}=0$. We obtain the prediction model by regressing $Y$ on ${\bf{Z}}=(W_{1},\dots,W_{9})^{{\sf\scriptscriptstyle{T}}}$. In Scenario 2b, we let $p=3$, $\bm{\theta}=(0.25,0.25,0.25)^{{\sf\scriptscriptstyle{T}}}$, and $\theta_{\text{\tiny{INT}}}=0.25$. We obtain the prediction model by regressing $Y$ on ${\bf{Z}}=(W_{1},W_{2},W_{3},W_{2}W_{1},W_{3}W_{1},W_{3}W_{2})^{{\sf\scriptscriptstyle{T}}}$. Scenario 2b uses a weaker prediction model than Scenario 2a.
We evaluate the performance of the PPI and CC methods: $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, and $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$. We also include two comparison methods: $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, serving as the baseline, and $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$, highlighting the limitations of naive inference. We present the percent bias for each estimator. For the PPI and CC methods, we present the confidence interval (CI) length ratio, computed as the mean of the ratio of the length of the PB inference estimator’s normal-based CI relative to the length of the normal-based CI for $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ (which is simply the ratio of the asymptotic standard errors). The CI length ratio has a baseline value of 1, which indicates that the CI of the PB inference method is the same length as the CI for $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. A CI length ratio less than 1 indicates that the PB inference estimator has lower standard error than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
5.2 Results
Table 1 presents the percent bias of the various estimators across the different scenarios and labeled dataset sizes. With the exception of $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$, the methods exhibit minimal bias as they are all consistent estimators for $\bm{\beta}^{*}_{1}$. The limitations of naive inference are particularly apparent in Scenarios 2a and 2b, which do not include $X_{1}$ in ML model training. $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$ exhibits an average percent bias of -47% in Scenario 2a and -64% in Scenario 2b.
Figure 2 displays the CI length ratio across the different settings. In Scenario 1, we observe that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ and $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ perform similarly, with CI lengths much less 1, indicating that the PB inference methods have reduced variance relative to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. Comparing Scenario 1a and 1b, we note a significant decrease in the gains in statistical efficiency of the PB inference methods due to the weaker prediction model used in Scenario 1b. However, all methods are at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, as the sufficient condition in Corollary 1.1 is satisfied. Additionally, as the labeled set size increases and hence $\pi=n_{\text{\tiny{lab}}}/n$ increases, the efficiency gain decreases as expected. $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is always more efficient than $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, particularly when the labeled dataset size is larger, due to the difference in the augmentation terms of these two estimators. In this scenario, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is asymptotically equivalent to $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$.
Scenario 2 highlights the differences in performance between $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ and $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$, which underscores our findings in Corollary 1.1 and Corollary 2.1. In both sub-scenarios, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ consistently outperforms both $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ and is at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. In Scenario 2a, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ achieves an average CI length ratio of 0.9 across the various sample sizes, indicating a gain in efficiency relative to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. In Scenario 2b, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is comparable to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ for the larger labeled dataset sizes. In contrast, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ are outperformed by $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and have confidence intervals up to 30% longer than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ for the larger labeled dataset sizes. This behavior is exacerbated in Scenario 2b, where the CI of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is up to 50% longer than the CI for $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and up to 72% longer for $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$.
Labeled set size
Method 1000 2000 3000 4000 5000
Scenario 1a
     $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$0.181 0.124 0.650 0.430 0.165
     $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ 0.035 -0.042 0.015 0.005 -0.005
     $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ 0.004 0.010 -0.004 -0.018 -0.001
     $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ 0.007 0.000 0.002 -0.009 -0.003
     $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ 0.007 0.000 0.002 -0.009 -0.003
Scenario 1b
     $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$1.070 0.181 0.426 -0.423 -1.205
     $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ 0.008 0.002 -0.016 -0.001 0.004
     $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ 0.007 -0.001 -0.003 -0.008 -0.010
     $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ 0.007 $0$0.007 -0.006 -0.003
     $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ 0.006 $0$0.007 -0.006 -0.003
Scenario 2a
     $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$46.980 -47.523 -45.946 -46.725 -47.695
     $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$0.033 -0.023 0.027 0.004 0.005
     $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$0.023 -0.009 -0.003 -0.032 -0.003
     $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$0.024 -0.012 0.006 -0.018 0.001
     $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$0.030 -0.016 0.016 -0.007 0.003
Scenario 2b
     $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$64.272 -64.115 -63.307 -63.759 -64.627
     $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$0.012 -0.003 0.002 0.012 0.013
     $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ 0.000 0.001 -0.006 0.012 0.009
     $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$0.001 $0$0.004 0.012 0.010
     $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$0.014 -0.006 -0.003 0.011 0.010
Table 1: Percent biases of the various estimators. Negative values indicate underestimation, while positive values indicate overestimation of the true parameter.
Figure 2: Confidence interval length ratio of the PB inference methods compared to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. A ratio less than 1 indicates that the PB inference method has a confidence interval that is shorter than the confidence interval for $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
6 Discussion
In this paper, we showed that adding a simple weight to the PPI estimator provides valid inference and is guaranteed to be at least as efficient as estimation using only labeled data. The weight is drawn from the work of Chen and Chen (2000) in the context of double sampling. In contrast, the original PPI estimator can be less efficient than estimation using only the labeled data, a property that can limit its practical utility. By incorporating the weight derived from Chen and Chen (2000), the augmented estimation approach principally balances the contribution of labeled and unlabeled data and extends the applicability of the PPI method to scenarios where leveraging all available data is essential for making valid and efficient inference. In addition to theoretical results that support these findings, our simulation studies demonstrate that the PPI method can result in substantial efficiency loss when the ML model is inaccurate. Additionally, we provide connections to several existing approaches from the economics and statistics literature that provide some interesting context for the PB inference problem. While we focused on least squares estimation, these results extend to other settings. The work of Chen and Chen (2000) and Angelopoulos et al. (2023a) apply to a wide variety of models beyond linear regression, such as parameters defining generalized linear models or parameters derived from estimating equations.
The work of Angelopoulos et al. (2023a) was pivotal in advancing the growing body of literature in PB inference. Recent work has considered settings where the ML model is not black box and where active learning is utilized to obtain the labeled data Zrnic and Candès (2024b, b). Additionally, two concurrent pre-prints have noted potential losses in efficiency of PB inference approaches. The authors of the PPI paper introduced a new method, PPI++, which aims to address this limitation of the PPI estimator (Angelopoulos et al., 2023b). The PPI++ method is designed to adaptively minimize the trace of the asymptotic variance-covariance of the proposed estimator to ensure that it yields an improvement over an approach based solely on labeled data. In contrast, the approach of Chen and Chen (2000) minimizes the mean square error among a class of estimators with a linear augmentation term. Simultaneously, Miao et al. (2023) introduced the PoSt-Prediction Adaptive inference (PSPA) estimator that is both robust to the accuracy of the ML model and guaranteed to be at least as efficient as estimation using only labeled data. Unlike the PPI++ estimator, the PSPA estimator is designed to achieve element-wise variance reduction and was shown to outperform the PPI++ approach. The PSPA approach is based on augmenting estimating equations with information from $\widehat{Y}$, while the CC method provides a simple and natural extension to the PPI method by adding a weight to the augmentation term.
Code availability
Our analyses can be replicated with the code at: https://github.com/SelinaS37/PBInference.
Acknowledgments
The authors gratefully acknowledge funding from an NSERC Discovery Grant (RGPIN-2021-03734), the Connaught Fund, the University of Toronto Data Science Institute, and the McLaughlin Center for Genetic Research.
Appendix A Asymptotic properties of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$
A.1 Preliminaries
Notation. We denote the outer product of a vector $\boldsymbol{v}$ as $\boldsymbol{v}^{\otimes 2}=\boldsymbol{v}\boldsymbol{v}^{{\sf\scriptscriptstyle{T}}}$. We denote that a matrix, ${\bf{M}}$, is positive semi-definite as ${\bf{M}}\succeq{\bf{0}},$ and positive definite as ${\bf{M}}\succ{\bf{0}}.$ For any estimator $\hat{\bm{\beta}}$, we let $\text{aVar}(\hat{\bm{\beta}})$ represent its asymptotic variance-covariance matrix. We define the difference in asymptotic variance-covariance matrices between two estimators $\hat{\bm{\beta}}_{1}$ and $\hat{\bm{\beta}}_{2}$ as $\Delta_{\text{aVar}}(\hat{\bm{\beta}}_{1},\hat{\bm{\beta}}_{2})=n[\text{aVar}(\hat{\bm{\beta}}_{1})-\text{aVar}(\hat{\bm{\beta}}_{2})].$
Basic assumptions. We make the standard assumption that $\text{Var}({\bf{X}})\succ{\bf{0}}$. We also assume that ${\bf{X}}$ has compact support and that $(Y_{i},\widehat{Y}_{i},{\bf{X}}^{{\sf\scriptscriptstyle{T}}})^{{\sf\scriptscriptstyle{T}}}$ have finite first and second moments.
A.2 Proof of Theorem 1
Theorem 1 (Influence function expansion of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$)
Let $R_{i}$ be the indicator of whether the $i$th observation is labeled. Then,
$n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}-\bm{\beta}^{*}\right)=n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})+o_{p}(1)$
where ${\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})={\bf{A}}\left\{\left[\ \bm{\phi}(Y_{i},{\bf{X}}_{i};\bm{\beta}^{*})\right]\frac{R_{i}}{\pi}+\left[\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma}^{*})\right]\frac{\pi-R_{i}}{\pi}\right\}$ and ${\bf{A}}=[\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})]^{-1}$. Therefore, $n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}-\bm{\beta}^{*}\right)$ converges weakly to a zero-mean multivariate normal distribution with variance-covariance
$\mathbb{E}[{\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})\hskip 2.84526pt{\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})^{{\sf\scriptscriptstyle{T}}}].$
Proof  First, note that $n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}-\bm{\beta}^{*}\right)$ may be decomposed as
$n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}-\bm{\beta}^{*}\right)=n^{\frac{1}{2}}\left[\left(\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\bm{\beta}^{*}\right)-\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\bm{\gamma}^{*}\right)+\left(\hat{\bm{\gamma}}^{\text{\tiny{all}}}-\bm{\gamma}^{*}\right)\right].$
We consider the asymptotic linear expansions of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, $\hat{\bm{\gamma}}^{\text{\tiny{lab}}}$, and $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$. Since $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ is the familiar least squares estimator,
$n_{\text{\tiny{lab}}}^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\bm{\beta}^{*}\right)=\left(n_{\text{\tiny{lab}}}^{-1}\sum_{i=1}^{n_{\text{\tiny{lab}}}}{\bf{X}}_{i}{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\right)^{-1}n_{\text{\tiny{lab}}}^{-\frac{1}{2}}\sum_{i=1}^{n_{\text{\tiny{lab}}}}{\bf{X}}_{i}(Y_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\bm{\beta}^{*}).$ (4)
By Slutsky’s theorem and the assumption that $\lim_{n\to\infty}n_{\text{\tiny{lab}}}/n=\pi$, equation (4) can be rewritten as
$\begin{split}&n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\bm{\beta}^{*}\right)=n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{A}}\left\{\left[{\bf{X}}_{i}\left(Y_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\bm{\beta}^{*}\right)\right]\frac{R_{i}}{\pi}\right\}+o_{p}(1).\end{split}$ (5)
Similar arguments can also be used to show that
$n^{\frac{1}{2}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\bm{\gamma}^{*}\right)=n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{A}}\left\{\left[{\bf{X}}_{i}\left(\widehat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma}^{*}\right)\right]\frac{R_{i}}{\pi}\right\}+o_{p}(1)$
and
$n^{\frac{1}{2}}\left(\hat{\bm{\gamma}}^{\text{\tiny{all}}}-\bm{\gamma}^{*}\right)=n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{A}}\left\{\left[{\bf{X}}_{i}\left(\widehat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma}^{*}\right)\right]\right\}+o_{p}(1).$
Consequently,
$n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}-\bm{\beta}^{*}\right)$ $=n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{A}}\left\{\left[\bm{\phi}(Y_{i},{\bf{X}}_{i};\bm{\beta}^{*})\frac{R_{i}}{\pi}+\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i},\bm{\gamma}^{*})\frac{\pi-R_{i}}{\pi}\right]\right\}+o_{p}(1)$
$=n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})+o_{p}(1).$
It follows from the standard central limit theorem (CLT) that $n^{\frac{1}{2}}(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}-\bm{\beta}^{*})$ converges weakly to a zero mean multivariate Gaussian distribution with variance-covariance given by $\mathbb{E}\left[{\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})\hskip 2.84526pt{\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})^{{\sf\scriptscriptstyle{T}}}\right].$
A.3 Proof of Corollary 1.1
Corollary 1.1 (Comparison of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ with $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$)
The difference in the asymptotic variance-covariance of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is
$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})=\left(\frac{1}{\pi}-1\right){\bf{A}}\left(2\boldsymbol{\Sigma}_{Y,\widehat{Y}}-\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}\right){\bf{A}}.$
$\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is therefore more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if $2\boldsymbol{\Sigma}_{Y,\widehat{Y}}\succ\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}$. A sufficient condition for $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ to have lower variance-covariance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ is that
$\mathbb{E}\left[\left(\epsilon_{\widehat{Y}}-\epsilon_{Y}\right)^{2}\mid{\bf{X}}\right]<\mathbb{E}\left(\epsilon_{Y}^{2}\mid{\bf{X}}\right).$
Proof  Equation (5) in the proof of Theorem 1 gives the asymptotic expansion of $n^{\frac{1}{2}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\bm{\beta})$. The standard CLT implies $n^{\frac{1}{2}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\bm{\beta})$ converges weakly to a zero-mean multivariate Gaussian distribution with variance-covariance given by
$\mathbb{E}\left\{\left[{\bf{A}}\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\right]^{\otimes 2}\frac{R^{2}}{\pi^{2}}\right\}=\frac{1}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{Y,Y}{\bf{A}}.$
Similarly, by Theorem 1
$\begin{split}n[\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})]=&\mathbb{E}\left\{\left[{\bf{A}}\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\right]^{\otimes 2}\frac{R^{2}}{\pi^{2}}\right\}+\mathbb{E}\left\{\left[{\bf{A}}\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})\right]^{\otimes 2}\frac{(\pi-R)^{2}}{\pi^{2}}\right\}\\ +&2\mathbb{E}\left\{\left[{\bf{A}}\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\right]\left[{\bf{A}}\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})\right]^{{\sf\scriptscriptstyle{T}}}\frac{R(\pi-R)}{\pi^{2}}\right\}\\ =&\frac{1}{\pi}{\bf{A}}\left[\boldsymbol{\Sigma}_{Y,Y}+(1-\pi)\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}-2(1-\pi)\boldsymbol{\Sigma}_{Y,\widehat{Y}}\right]{\bf{A}}.\end{split}$
Thus,
$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})=\left(\frac{1}{\pi}-1\right){\bf{A}}\left(2\boldsymbol{\Sigma}_{Y,\widehat{Y}}-\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}\right){\bf{A}}.$
$\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is therefore more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if $2\boldsymbol{\Sigma}_{Y,\widehat{Y}}\succ\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}$. Additionally, if $2\mathbb{E}\left(\epsilon_{Y}\epsilon_{\widehat{Y}}\mid{\bf{X}}\right)>\mathbb{E}\left(\epsilon_{\widehat{Y}}^{2}\mid{\bf{X}}\right)$ or equivalently if $\mathbb{E}\left[\left(\epsilon_{\widehat{Y}}-\epsilon_{Y}\right)^{2}\mid{\bf{X}}\right]<\mathbb{E}\left(\epsilon_{Y}^{2}\mid{\bf{X}}\right),$ the diagonal entries of $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})$ are positive and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ will have a lower variance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
A.4 Examples for Corollary 1.1
We provide several scenarios where the sufficient condition from Corollary 1.1 is satisfied and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. These scenarios characterize settings where the predictions contain more information about $Y$ beyond what is explained by the linear regression model based on ${\bf{X}}$ (i.e., the focus of statistical inference.)
Examples of Corollary 1.1. Let ${\bf{Z}}=({\bf{X}}^{{\sf\scriptscriptstyle{T}}},{\bf{W}}^{{\sf\scriptscriptstyle{T}}})^{{\sf\scriptscriptstyle{T}}}$ where ${\bf{W}}$ are predictors of $Y$ beyond ${\bf{X}}$. $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ in the following scenarios where $\hat{f}$ perfectly estimates the true regression function:
  1. 1.
    $\hat{f}({\bf{X}})=\mathbb{E}(Y\mid{\bf{X}})$ and $\mathbb{E}(Y\mid{\bf{X}})\neq{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\beta}^{*}$
  2. 2.
    $\hat{f}({\bf{Z}})=\mathbb{E}(Y\mid{\bf{Z}})$, ${\bf{W}}\perp Y\mid{\bf{X}}$, and $\mathbb{E}(Y\mid{\bf{X}})\neq{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\beta}^{*}$
  3. 3.
    $\hat{f}({\bf{Z}})=\mathbb{E}(Y\mid{\bf{Z}})$ and ${\bf{W}}\not\perp Y\mid{\bf{X}}$
  4. 4.
    $\hat{f}({\bf{W}})=\mathbb{E}(Y\mid{\bf{W}})$ and ${\bf{X}}\perp Y\mid{\bf{W}}$
Proof  For scenario 1, assume that $\hat{f}({\bf{X}})=\mathbb{E}(Y\mid{\bf{X}})$. It follows that
$\bm{\gamma}^{*}=\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})\mathbb{E}[{\bf{X}}\mathbb{E}(Y\mid{\bf{X}})]=\bm{\beta}^{*}$
so that
$\mathbb{E}\left[\left(\epsilon_{\widehat{Y}}-\epsilon_{Y}\right)^{2}\mid{\bf{X}}\right]$ $=\mathbb{E}\left\{[\mathbb{E}(Y\mid{\bf{X}})-Y]^{2}\mid{\bf{X}}\right\}$ $<\mathbb{E}\left(\epsilon_{Y}^{2}\mid{\bf{X}}\right)=\mathbb{E}\left[(Y-{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\beta}^{*})^{2}\mid{\bf{X}}\right].$
This inequality is strict provided that $\mathbb{E}(Y\mid{\bf{X}})\neq{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\beta}^{*}$. If the regression model is correctly specified (i.e., $\mathbb{E}(Y\mid{\bf{X}})={\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\beta}^{*}$), then $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ are asymptotically equivalent.
For scenario 2, assume that $\hat{f}({\bf{Z}})=\mathbb{E}(Y\mid{\bf{Z}})$ where ${\bf{Z}}=({\bf{X}}^{{\sf\scriptscriptstyle{T}}},{\bf{W}}^{{\sf\scriptscriptstyle{T}}})^{{\sf\scriptscriptstyle{T}}}$ and ${\bf{W}}\perp Y\mid{\bf{X}}$. Similar to scenario 1,
$\bm{\gamma}^{*}=\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})\mathbb{E}[{\bf{X}}\mathbb{E}(Y\mid{\bf{X}},{\bf{W}})]=\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})\mathbb{E}[{\bf{X}}\mathbb{E}(Y\mid{\bf{X}})]=\bm{\beta}^{*}.$
Therefore,
$\mathbb{E}\left[\left(\epsilon_{\widehat{Y}}-\epsilon_{Y}\right)^{2}\mid{\bf{X}}\right]$ $=\mathbb{E}\left\{[\mathbb{E}(Y\mid{\bf{Z}})-Y]^{2}\mid{\bf{X}}\right\}$
$=\mathbb{E}\left\{[\mathbb{E}(Y\mid{\bf{X}})-Y]^{2}\mid{\bf{X}}\right\}=\text{Var}(Y\mid{\bf{X}})\leq\mathbb{E}\left(\epsilon_{Y}^{2}\mid{\bf{X}}\right).$
The inequality is strict when $\mathbb{E}(Y\mid{\bf{X}})\neq{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\beta}^{*}$ and is an equality if the regression model is correctly specified.
For scenario 3, again assume that $\hat{f}({\bf{Z}})=\mathbb{E}(Y\mid{\bf{Z}})$ so that $\bm{\gamma}^{*}=\bm{\beta}^{*}$. It then follows that
$\mathbb{E}\left[\left(\epsilon_{\widehat{Y}}-\epsilon_{Y}\right)^{2}\mid{\bf{X}}\right]$ $=\mathbb{E}\left\{[\mathbb{E}(Y\mid{\bf{Z}})-Y]^{2}\mid{\bf{X}}\right\}$
$=\text{Var}(Y\mid{\bf{X}})-\text{Var}[\mathbb{E}(Y\mid{\bf{X}},{\bf{W}})\mid{\bf{X}}]$
$<\text{Var}(Y\mid{\bf{X}})\leq\mathbb{E}\left(\epsilon_{Y}^{2}\mid{\bf{X}}\right).$
The equality is strict as we assume ${\bf{W}}\not\perp Y\mid{\bf{X}}$.
Lastly, for scenario 4, assume that $\hat{f}({\bf{W}})=E(Y\mid{\bf{W}})$ and ${\bf{X}}\perp Y\mid{\bf{W}}$. Then,
$\bm{\gamma}^{*}=\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})\mathbb{E}[{\bf{X}}\mathbb{E}(Y\mid{\bf{W}})]=\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})\mathbb{E}[{\bf{X}}\mathbb{E}(Y\mid{\bf{W}},{\bf{X}})]=\bm{\beta}^{*}.$
Analogously to the argument for scenario 3,
$\mathbb{E}\left[\left(\epsilon_{\widehat{Y}}-\epsilon_{Y}\right)^{2}\mid{\bf{X}}\right]$ $=\mathbb{E}\left\{[\mathbb{E}(Y\mid{\bf{W}})-Y]^{2}\mid{\bf{X}}\right\}$
$=\text{Var}(Y\mid{\bf{X}})-\text{Var}[\mathbb{E}(Y\mid{\bf{W}})\mid{\bf{X}}]$
$<\text{Var}(Y\mid{\bf{X}})\leq\mathbb{E}\left(\epsilon_{Y}^{2}\mid{\bf{X}}\right).$
Appendix B Asymptotic properties of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$
B.1 Asymptotic expansion of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$
Theorem S1 (Influence function expansion of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$)
Let $R_{i}$ be the indicator of whether the $i$th observation is labeled. Then,
$n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{PPI}}}-\bm{\beta}^{*}\right)=n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{IF}}^{\text{\tiny{PPI}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})+o_{p}(1)$
where ${\bf{IF}}^{\text{\tiny{PPI}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})={\bf{A}}\left\{\left[\bm{\phi}(Y_{i},{\bf{X}}_{i};\bm{\beta}^{*})\frac{R_{i}}{\pi}+\bm{\omega}^{\text{\tiny{PPI}}}(\widehat{Y}_{i},{\bf{X}}_{i},\bm{\gamma}^{*})\frac{\pi-R_{i}}{\pi}\right]\right\}$, and
$\bm{\omega}^{\text{\tiny{PPI}}}(\widehat{Y}_{i},{\bf{X}}_{i},\bm{\gamma}^{*})=\frac{\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma}^{*})}{1-\pi}$. Therefore, $n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{PPI}}}-\bm{\beta}^{*}\right)$ converges weakly to a zero-mean multivariate normal distribution with variance-covariance
$\mathbb{E}\left[{\bf{IF}}^{\text{\tiny{PPI}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})\hskip 2.84526pt{\bf{IF}}^{\text{\tiny{PPI}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})^{{\sf\scriptscriptstyle{T}}}\right].$
Proof  The only difference in studying the asymptotic properties of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ compared to $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is that we rely on the asymptotic linear expansion of $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ rather than $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$. Using the same argument as in the proof of Theorem 1,
$\begin{split}n^{\frac{1}{2}}\left(\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}-\bm{\gamma}^{*}\right)&=n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{A}}\left\{\left[{\bf{X}}_{i}\left(\widehat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma}^{*}\right)\right]\frac{1-R_{i}}{1-\pi}\right\}+o_{p}(1).\end{split}$
Consequently,
$n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{PPI}}}-\bm{\beta}^{*}\right)$ $=n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{A}}\left\{\left[\bm{\phi}(Y_{i},{\bf{X}}_{i};\bm{\beta}^{*})\frac{R_{i}}{\pi}+\frac{\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma}^{*})}{1-\pi}\frac{\pi-R_{i}}{\pi}\right]\right\}+o_{p}(1)$
$=n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{IF}}^{\text{\tiny{PPI}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})+o_{p}(1).$
The standard CLT implies $n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{PPI}}}-\bm{\beta}^{*}\right)$ converges weakly to a zero-mean multivariate normal with variance-covariance $\mathbb{E}\left[{\bf{IF}}^{\text{\tiny{PPI}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})\hskip 2.84526pt{\bf{IF}}^{\text{\tiny{PPI}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})^{{\sf\scriptscriptstyle{T}}}\right].$
B.2 Comparison of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$
Corollary S1.1 (Comparison of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ with $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$)
The difference in the asymptotic variance-covariance of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is
$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}})$ $=\frac{1}{\pi}{\bf{A}}\left(2\boldsymbol{\Sigma}_{Y,\widehat{Y}}-\frac{1}{1-\pi}\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}\right){\bf{A}}.$
$\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is therefore more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if $2\boldsymbol{\Sigma}_{Y,\widehat{Y}}\succ\frac{1}{1-\pi}\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}$. A sufficient condition for $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ to have lower variance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ is that
$\mathbb{E}\left[\left(\epsilon_{\widehat{Y}}-\epsilon_{Y}\right)^{2}\mid{\bf{X}}\right]<\mathbb{E}\left(\epsilon_{Y}^{2}\mid{\bf{X}}\right)+\left(1-\frac{1}{1-\pi}\right)\mathbb{E}\left(\epsilon_{\widehat{Y}}^{2}\mid{\bf{X}}\right).$
Proof  From the proof for Corollary 1.1, $n[\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{lab}}})]=\frac{1}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{Y,Y}{\bf{A}}.$ By Theorem S1,
$n[\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{PPI}}})]=$ $\mathbb{E}\left\{\left[{\bf{A}}\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\right]^{\otimes 2}\frac{R^{2}}{\pi^{2}}\right\}+\mathbb{E}\left\{\left[{\bf{A}}\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})\right]^{\otimes 2}\frac{(\pi-R)^{2}}{\pi^{2}(1-\pi)^{2}}\right\}$
$+2\mathbb{E}\left\{\left[{\bf{A}}\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\right]\left[{\bf{A}}\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})\right]^{{\sf\scriptscriptstyle{T}}}\frac{R(\pi-R)}{\pi^{2}(1-\pi)}\right\}$
$=\frac{1}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{Y,Y}{\bf{A}}+\frac{1}{\pi(1-\pi)}{\bf{A}}\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}{\bf{A}}-\frac{2}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}{\bf{A}}$
$=\frac{1}{\pi}{\bf{A}}\left[\boldsymbol{\Sigma}_{Y,Y}+\frac{1}{1-\pi}\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}-2\boldsymbol{\Sigma}_{Y,\widehat{Y}}\right]{\bf{A}}.$
Thus,
$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}})=\frac{1}{\pi}{\bf{A}}\left[2\boldsymbol{\Sigma}_{Y,\widehat{Y}}-\frac{1}{1-\pi}\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}\right]{\bf{A}}.$
$\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is therefore more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if $2\boldsymbol{\Sigma}_{Y,\widehat{Y}}\succ\frac{1}{1-\pi}\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}$. Note that if $2\mathbb{E}\left(\epsilon_{Y}\epsilon_{\widehat{Y}}\mid{\bf{X}}\right)>\frac{1}{1-\pi}\mathbb{E}\left(\epsilon_{\widehat{Y}}^{2}\mid{\bf{X}}\right)$ or equivalently if $\mathbb{E}\left[\left(\epsilon_{\widehat{Y}}-\epsilon_{Y}\right)^{2}\mid{\bf{X}}\right]<\mathbb{E}\left(\epsilon_{Y}^{2}\mid{\bf{X}}\right)+\left(1-\frac{1}{1-\pi}\right)\mathbb{E}\left(\epsilon_{\widehat{Y}}^{2}\mid{\bf{X}}\right),$ the diagonal entries of $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}})$ are positive and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ will have a lower variance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
B.3 Comparison of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$
Recall from Theorem 1 and S1 that
$n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}-\bm{\beta}^{*}\right)=n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{A}}\left\{\left[\bm{\phi}(Y_{i},{\bf{X}}_{i};\bm{\beta}^{*})\frac{R_{i}}{\pi}+\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i},\bm{\gamma}^{*})\frac{\pi-R_{i}}{\pi}\right]\right\}+o_{p}(1)\quad\mbox{and}$
$n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{PPI}}}-\bm{\beta}^{*}\right)=n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{A}}\left\{\left[\bm{\phi}(Y_{i},{\bf{X}}_{i};\bm{\beta}^{*})\frac{R_{i}}{\pi}+\bm{\omega}^{\text{\tiny{PPI}}}(\widehat{Y}_{i},{\bf{X}}_{i},\bm{\gamma}^{*})\frac{\pi-R_{i}}{\pi}\right]\right\}+o_{p}(1).$
From Corollary 1.1 and Corollary S1.1, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if
$2\boldsymbol{\Sigma}_{Y,\widehat{Y}}\succ\frac{1}{1-\pi}\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}$
whereas $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if
$2\boldsymbol{\Sigma}_{Y,\widehat{Y}}\succ\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}.$
Therefore, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. Moreover, the difference between asymptotic variance-covariance can be derived from Corollary 1.1 and S1.1 as
$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{PPI}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})=\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})-\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}})$
$=\left(\frac{1}{\pi}-1\right){\bf{A}}\left(2\boldsymbol{\Sigma}_{Y,\widehat{Y}}-\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}\right){\bf{A}}-\frac{1}{\pi}{\bf{A}}\left(2\boldsymbol{\Sigma}_{Y,\widehat{Y}}-\frac{1}{1-\pi}\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}\right){\bf{A}}$
$={\bf{A}}\left(\frac{2-\pi}{1-\pi}\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}-2\boldsymbol{\Sigma}_{Y,\widehat{Y}}\right){\bf{A}}.$
Note that if $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is less efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ then $\frac{1}{1-\pi}\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}\succ 2\boldsymbol{\Sigma}_{Y,\widehat{Y}}$, implying
$\frac{2-\pi}{1-\pi}\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}\succ 2\boldsymbol{\Sigma}_{Y,\widehat{Y}}.$
This confirms that when $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is less efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, it will also be less efficient than $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$. A sufficient condition for $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ to be more efficient than $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is
$\mathbb{E}[\epsilon_{Y}^{2}\mid{\bf{X}}]<\mathbb{E}[(\epsilon_{\widehat{Y}}-\epsilon_{Y})^{2}\mid{\bf{X}}]+\frac{1}{1-\pi}\mathbb{E}[\epsilon_{\widehat{Y}}^{2}\mid{\bf{X}}].$
Appendix C Asymptotic properties of $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$
C.1 Proof of Theorem 2
Theorem 2 (Influence function expansion of $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$)
Following Chen and Chen (2000),
$n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{CC}}}-\bm{\beta}^{*}\right)=n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{IF}}^{\text{\tiny{CC}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})+o_{p}(1)$
where ${\bf{IF}}^{\text{\tiny{CC}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})={\bf{A}}\left[\bm{\phi}(Y_{i},{\bf{X}}_{i};\bm{\beta}^{*})\frac{R_{i}}{\pi}+\bm{\omega}^{\text{\tiny{CC}}}(\widehat{Y}_{i},{\bf{X}}_{i})\frac{\pi-R_{i}}{\pi}\right]$ and $\bm{\omega}^{\text{\tiny{CC}}}(\widehat{Y},{\bf{X}})$ is defined in (1). Therefore, $n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{CC}}}-\bm{\beta}^{*}\right)$ converges weakly to a zero-mean multivariate normal distribution with variance-covariance
$\hskip 2.84526pt\mathbb{E}\left[{\bf{IF}}^{\text{\tiny{CC}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})\hskip 2.84526pt{\bf{IF}}^{\text{\tiny{CC}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})^{{\sf\scriptscriptstyle{T}}}\right]\hskip 2.84526pt.$
Proof  The proof follows from Chen and Chen (2000), but is detailed for completeness. Recall that
$\begin{split}\widehat{{\bf{W}}}^{\text{\tiny{CC}}}&=\widehat{{\bf{A}}}\left\{\widehat{\text{Cov}}\left[\phi(Y,{\bf{X}};\bm{\beta}^{*}),\phi(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})\right]\right\}\{\widehat{\text{Var}}[\phi(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})]\}^{-1}\widehat{{\bf{A}}}^{-1}\\ &=\widehat{{\bf{A}}}\widehat{\boldsymbol{\Sigma}}_{Y,\widehat{Y}}\widehat{\boldsymbol{\Sigma}}^{-1}_{\widehat{Y},\widehat{Y}}\widehat{{\bf{A}}}^{-1}\end{split}$
where
$\begin{split}&\widehat{{\bf{A}}}=n_{\text{\tiny{lab}}}\left(\sum_{i=1}^{n_{\text{\tiny{lab}}}}{\bf{X}}_{i}{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\right)^{-1},\\ &\widehat{\boldsymbol{\Sigma}}_{Y,\widehat{Y}}=n_{\text{\tiny{lab}}}^{-1}\sum_{i=1}^{n_{\text{\tiny{lab}}}}\left(Y_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\hat{\bm{\beta}}^{\text{\tiny{lab}}}\right)\left(\hat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\hat{\bm{\gamma}}^{\text{\tiny{lab}}}\right){\bf{X}}_{i}{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}},\\ &\mbox{and }\widehat{\boldsymbol{\Sigma}}_{\widehat{Y},\widehat{Y}}=n_{\text{\tiny{lab}}}^{-1}\sum_{i=1}^{n_{\text{\tiny{lab}}}}\left(\hat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\hat{\bm{\gamma}}^{\text{\tiny{lab}}}\right)^{2}{\bf{X}}_{i}{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}.\end{split}$
It follows that
$n^{\frac{1}{2}}\left[\widehat{{\bf{W}}}^{\text{\tiny{CC}}}(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\bm{\gamma}^{*})\right]={\bf{A}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}\boldsymbol{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\left\{n^{-\frac{1}{2}}\sum_{i=1}^{n}\left[{\bf{X}}_{i}\left(\widehat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma}^{*}\right)\right]\frac{R_{i}}{\pi}\right\}+o_{p}(1)$
and
$n^{\frac{1}{2}}\left[\widehat{{\bf{W}}}^{\text{\tiny{CC}}}(\hat{\bm{\gamma}}^{\text{\tiny{all}}}-\bm{\gamma}^{*})\right]={\bf{A}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}\boldsymbol{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\left\{n^{-\frac{1}{2}}\sum_{i=1}^{n}\left[{\bf{X}}_{i}\left(\widehat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma}^{*}\right)\right]\right\}+o_{p}(1).$
Consequently,
$n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{CC}}}-\bm{\beta}^{*}\right)=$ $n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{A}}\left[\bm{\phi}(Y_{i},{\bf{X}}_{i};\bm{\beta}^{*})\frac{R_{i}}{\pi}+\bm{\omega}^{\text{\tiny{CC}}}(\widehat{Y}_{i},{\bf{X}}_{i})\frac{\pi-R_{i}}{\pi}\right]+o_{p}(1).$
$=$ $n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{IF}}^{\text{\tiny{CC}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})+o_{p}(1).$
The standard CLT implies $n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{CC}}}-\bm{\beta}^{*}\right)$ converges weakly to a zero-mean multivariate normal distribution with variance-covariance
$\mathbb{E}\left[{\bf{IF}}^{\text{\tiny{CC}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})\hskip 2.84526pt{\bf{IF}}^{\text{\tiny{CC}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})^{{\sf\scriptscriptstyle{T}}}\right].$
C.2 Proof of Corollary 2.1
Corollary 2.1 ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$)
The difference in the asymptotic covariance of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is
$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{CC}}})=\left(\frac{1}{\pi}-1\right){\bf{A}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}\boldsymbol{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}{\bf{A}}\succeq{\bf{0}}.$
Therefore, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is guaranteed to be at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
Proof  From the proof for Corollary 1.1, $n[\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{lab}}})]=\frac{1}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{Y,Y}{\bf{A}}.$ By Theorem 2.1, $n[\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{CC}}})]$ is given by
$\begin{split}&\mathbb{E}\left\{\left[{\bf{A}}\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\right]^{\otimes 2}\frac{R^{2}}{\pi^{2}}\right\}+\mathbb{E}\left\{\left[{\bf{A}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}\boldsymbol{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})\right]^{\otimes 2}\frac{(\pi-R)^{2}}{\pi^{2}}\right\}\\ +&2\mathbb{E}\left\{\left[{\bf{A}}\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\right]\left[{\bf{A}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}\boldsymbol{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})\right]^{{\sf\scriptscriptstyle{T}}}\frac{R(\pi-R)}{\pi^{2}}\right\}\\ =&\frac{1}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{Y,Y}{\bf{A}}+\frac{1-\pi}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}\boldsymbol{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}{\bf{A}}-2\frac{1-\pi}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}\boldsymbol{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}{\bf{A}}\\ =&\frac{1}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{Y,Y}{\bf{A}}-\frac{1-\pi}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}\boldsymbol{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}{\bf{A}}.\end{split}$
Hence, $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{CC}}})=\frac{1-\pi}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}\boldsymbol{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}{\bf{A}}\succeq 0$.
C.3 Proof of Corollary 2.2
Corollary 2.2 ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is the most efficient estimator in its class)
Let the mean squared error (MSE) of $\hat{\bm{\beta}}$ for estimating $\bm{\beta}^{*}$ be defined as $\mathbb{E}[(\hat{\bm{\beta}}-\bm{\beta}^{*})^{{\sf\scriptscriptstyle{T}}}(\hat{\bm{\beta}}-\bm{\beta}^{*})]$. Analogously to the well-known result for the best linear predictor (e.g., Bickel and Doksum (2015), page 40), $\widehat{{\bf{W}}}^{\text{\tiny{CC}}}$ is the weight, ${\bf{W}}$, that minimizes the asymptotic MSE of estimators of the form $\hat{\bm{\beta}}^{\text{\tiny{lab}}}-{{\bf{W}}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right)$ for estimating $\bm{\beta}^{*}$.
Proof  Letting $\hat{\bm{\beta}}_{{\bf{W}}}=\hat{\bm{\beta}}^{\text{\tiny{lab}}}-{\bf{W}}(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}})$, the MSE of estimating $\hat{\bm{\beta}}_{{\bf{W}}}$ is
$\mathbb{Q}({\bf{W}})=\mathbb{E}\left\{\left[\hat{\bm{\beta}}^{\text{\tiny{lab}}}-{\bf{W}}(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}})-\bm{\beta}^{*}\right]^{\otimes 2}\right\}.$
Thus, the optimal choice of ${\bf{W}}$ that minimizes the MSE is
${\bf{W}}^{*}=\mathbb{E}[(\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\bm{\beta}^{*})(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}})]\left\{\mathbb{E}\left[(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}})^{\otimes 2}\right]\right\}^{-1}.$
Note that the first term is the covariance between $\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\bm{\beta}^{*}$ and $\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}$, and the second term is the inverse variance-covariance of $\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}$. Since
$n^{\frac{1}{2}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right)={\bf{A}}\left[n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{X}}_{i}(\widehat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma}^{*})\frac{R_{i}-\pi}{\pi}\right]+o_{p}(1),$
we have $\text{Cov}(\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\bm{\beta}^{*},\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}})=\frac{1-\pi}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}{\bf{A}}$ and $\text{Var}(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}})=\frac{1-\pi}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}{\bf{A}}$. Thus,
${\bf{W}}^{*}=\frac{1-\pi}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}{\bf{A}}\left[\frac{1-\pi}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}{\bf{A}}\right]^{-1}={\bf{A}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}\boldsymbol{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}{\bf{A}}^{-1}={\bf{W}}^{CC}.$

### v2

@@
 Here we build on the work on Motwani and Witten to study the statistical efficiency of the PPI method.
-We highlight that a more efficient estimator can be obtained by simply adding a weight to the PPI estimator, an approach that was proposed by Chen & Chen (CC) more than 25 years ago in the context of double sampling designs (Chen and Chen, 2000).
+We show that a more efficient estimator can be obtained by simply adding a weight to the PPI estimator, an approach that was proposed by Chen & Chen (CC) more than 25 years ago in the context of double sampling designs (Chen and Chen, 2000).
 We begin by formally introducing the PB inference problem and reviewing the contributions of Motwani and Witten in Section 2.
@@
 Consequently, the familiar Wald statistic will not converge in distribution under the null, resulting in type I error that is not controlled at the nominal level and confidence intervals with inadequate coverage.
-Motwani and Witten also study the PPI method, which has the crucial property that inference remains valid regardless of the quality of the predictions.
+Motwani and Witten also studied the PPI method, which has the crucial property that inference remains valid regardless of the quality of the predictions.
 Stated otherwise, the PPI method does not require that $\hat{f}$ capture the true regression function for the estimator to maintain consistency or to yield reliable hypothesis testing and interval estimation.
@@
 $\mathbb{E}\left[\left(\epsilon_{\widehat{Y}}-\epsilon_{Y}\right)^{2}\mid{\bf{X}}\right]<\mathbb{E}\left(\epsilon_{Y}^{2}\mid{\bf{X}}\right).$
-More specifically, the sufficient condition in Corollary 1.1 coincides with the intuition that incorporating unlabeled data into estimation can improve efficiency when $\widehat{Y}$ has additional predictive power for $Y$ beyond what is already explained by the regression model of $Y$ on ${\bf{X}}$.
+The sufficient condition in Corollary 1.1 coincides with the intuition that incorporating unlabeled data into estimation can improve efficiency when $\widehat{Y}$ has additional predictive power for $Y$ beyond what is already explained by the regression model of $Y$ on ${\bf{X}}$.
 This condition is similar to the result for mean estimation provided in the Supplementary Materials of Angelopoulos et al.
@@
 $\rho_{Y,\widehat{Y}}>\frac{1}{2}\frac{\sigma_{\widehat{Y}}}{\sigma_{Y}},$
-where $\rho_{Y,\widehat{Y}}=\sigma_{Y,\widehat{Y}}/\sigma_{Y}\sigma_{\widehat{Y}}$ is the correlation of residuals for $Y$ and $\widehat{Y}$ given ${\bf{X}}$.
-The improvement of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ is therefore stronger with increasing correlation of residuals between $Y$ and $\widehat{Y}$ and decreasing variability in the residuals of $\widehat{Y}$.
-Therefore, $\widehat{Y}$ must be a strong predictor for $Y$ after adjusting for ${\bf{X}}$, though it does not necessarily need to be derived from a correctly specified ML model.
-In particular, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ gains efficiency if and only if the correlation of residuals is at least half of the ratio of residual standard deviations of $\widehat{Y}$ to that of $Y$.
+where $\rho_{Y,\widehat{Y}}=\sigma_{Y,\widehat{Y}}/\sigma_{Y}\sigma_{\widehat{Y}}$ is the correlation of the residuals for $Y$ and $\widehat{Y}$ given ${\bf{X}}$.
+$\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ therefore gains efficiency if and only if the correlation of the residuals is at least half of the ratio of residual standard deviations of $\widehat{Y}$ to that of $Y$.
+Put another way, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ improves over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ with respect to efficiency when $\widehat{Y}$ is a strong predictor for $Y$ after adjusting for ${\bf{X}}$, though $\widehat{Y}$ does not necessarily need to be derived from a ML model that captures the true regression function.
+The improvement is stronger with increasing correlation of the residuals between $Y$ and $\widehat{Y}$ and decreasing variability in the residuals of $\widehat{Y}$.
 3.2 Asymptotic analysis of the CC estimator
@@
 To overcome these difficulties, Chen and Chen (2000) proposed a particular choice of $\bm{\omega}(\widehat{Y},{\bf{X}})$ that is guaranteed to be at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and that does not require modeling of $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$.
-Their approach is motivated by the double-sampling design which, apart from differences in terminology and notation (i.e., they regard $\widehat{Y}$ as a mismeasured version of $Y$), is identical to the partially labeled data setting considered here.
+Their approach is motivated by the double-sampling design which, apart from differences in terminology and notation (i.e., they regard $\widehat{Y}$ as a mismeasured version of $Y$), is analogous to the partially labeled data setting considered here.
 The CC estimator approximates $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$ with
@@
 obtained with $\mathscr{L}$.
-Therefore, a simple way to ensure that an augmented estimator like the PPI estimator has asymptotic variance at least as low as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ is to add a simple weight function in front of the augmentation term.
+Therefore, a simple way to ensure that an augmented estimator like the PPI estimator has asymptotic variance at least as low as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ is to add a weight function in front of the augmentation term.
 In fact, the framework of Robins et al.
@@
 Let the mean squared error (MSE) of an estimator $\hat{\bm{\beta}}$ for $\bm{\beta}^{*}$ be defined as $\mathbb{E}[(\hat{\bm{\beta}}-\bm{\beta}^{*})^{{\sf\scriptscriptstyle{T}}}(\hat{\bm{\beta}}-\bm{\beta}^{*})]$.
-Analogously to the well-known result for the best linear predictor (e.g., Bickel and Doksum (2015), page 40), $\widehat{{\bf{W}}}^{\text{\tiny{CC}}}$ is the weight, ${\bf{W}}$, that minimizes the asymptotic MSE of estimators of the form $\hat{\bm{\beta}}^{\text{\tiny{lab}}}-{{\bf{W}}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right)$ for estimating $\bm{\beta}^{*}$.
-From Corollary 2.1, we note that in contrast to $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})$, the covariance term in $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{CC}}})$ is normalized by the residual variance of the $\widehat{Y}$ to scale the covariance according to the precision of $\widehat{Y}$.
-This scaling essentially enables $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ to always be as least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
-Coming back to the example from the previous section, we can gain more intuition as to why $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is guaranteed to be at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ while $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ does not possess this property.
+Analogously to the well-known result for the best linear predictor (e.g., Bickel and Doksum (2015), page 40), ${\bf{W}}^{\text{\tiny{CC}}}$ is the weight, ${\bf{W}}$, that minimizes the asymptotic MSE of estimators of the form $\hat{\bm{\beta}}^{\text{\tiny{lab}}}-{{\bf{W}}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right)$ for estimating $\bm{\beta}^{*}$.
+From Corollary 2.1, we note that in contrast to $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})$, the covariance term, $\boldsymbol{\Sigma}_{Y,\widehat{Y}}$, in $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{CC}}})$ is normalized by the residual variance of $\widehat{Y}$, $\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}$.
+This normalization effectively scales the covariance according to the precision of $\widehat{Y}$ and enables $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ to always be at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
+Coming back to the example from the previous section, we can gain more intuition as to why $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ possesses this property while $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ does not.
 Example 1 continued (Efficiency gain of $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ under homoskedasticity) We revisit the case of homoskedasticity and recall that
@@
 4 Related literature
-Existing PB inference literature has established connections with semi-supervised inference, semi-parametric inference, missing data, and measurement error (Angelopoulos et al., 2023a, b; Gan and Liang, 2023; Miao et al., 2023; Zrnic and Candès, 2024a, b).
+Existing PB inference literature has established connections with semi-supervised inference, semi-parametric inference, missing data, and measurement error (Angelopoulos et al., 2023a, b; Miao et al., 2023; Zrnic and Candès, 2024a, b).
 Here we draw connections between PB inference methods and methods from the economics and statistics literature dating back to the 1960s.
@@
 The SynSurr estimator is equivalent to $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ and asymptotically equivalent to $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ when $\text{Var}(\epsilon\mid{\bf{X}})$ and $\text{Cov}(\epsilon,\epsilon^{\prime}\mid{\bf{X}})$ are constant.
-Building on the SynSurr estimator, Miao et al.
+Building on SynSurr, Miao et al.
 (2024) proposed the “POP-GWAS” estimator for a variety of regression models utilizing the augmentation technique.
@@
 $\widehat{\Delta}-\sum_{i=1}^{n}(X_{i}-\bar{X})[n_{0}^{-1}h^{(0)}({\bf{Z}}_{i})+n_{1}^{-1}h^{(1)}({\bf{Z}}_{i})],$ (2)
-where $\bar{X}=n^{-1}\sum_{i=1}^{n}X_{i}$ and $h^{(k)}({\bf{Z}})$ for $k=0,1$ are arbitrary scalar functions of ${\bf{Z}}$.
-As $X\perp{\bf{Z}}$ are independent due to randomization, the augmentation term in 2 converges in probability to 0, analogously to the augmentation terms in the PPI and CC estimators.
-In the context of PB inference, pre-trained models could be used for predicting $Y$ on the basis of ${\bf{Z}}$ for $h^{(k)}({\bf{Z}})$ for $k=0,1$.
+where $\bar{X}=n^{-1}\sum_{i=1}^{n}X_{i}$ and $h^{(k)}({\bf{Z}})$ are arbitrary scalar functions of ${\bf{Z}}$ for $k=0,1$.
+As $X$ and ${\bf{Z}}$ are independent due to randomization, the augmentation term in 2 converges in probability to 0, analogously to the augmentation terms in the PPI and CC estimators.
+In the context of PB inference, pre-trained models could be used for predicting $Y$ on the basis of ${\bf{Z}}$ in the augmentation term.
 4.3 Improving precision of Monte Carlo simulations with control variates
@@
 $\frac{1}{n}\sum_{i=1}^{n}\left[Y_{i}-\lambda\{Z_{i}-\mu_{Z}\}\right].$ (3)
-The variance of the augmented estimator is minimized at $\lambda^{*}=\frac{\text{Cov}(Y,Z)}{\text{Var}(Z)}$ Rothery (1982); Davidson and MacKinnon (1992).
+The variance of the augmented estimator is minimized at $\lambda^{*}=\frac{\text{Cov}(Y,Z)}{\text{Var}(Z)}$ (Rothery, 1982; Davidson and MacKinnon, 1992).
 The augmentation term reduces the variance of the unadjusted power estimate $\overline{Y}$ in proportion to the correlation between $Y$ and $Z$.
@@
 We use 10,000 replications of the analysis dataset for each scenario and a single training set as $\hat{f}$ is assumed to be fixed.
-The size of the training set is fixed at 10,000 and the size of the analysis dataset is also set to $n=10,000$.
+The size of the training set is set to 10,000 and the size of the analysis dataset is also $n=10,000$.
 We let $n_{\text{\tiny{lab}}}=1000,2000,3000,4000,5000$ and $n_{\text{\tiny{unlab}}}=n-n_{\text{\tiny{lab}}}$.
@@
 $(\beta^{*}_{0},\beta^{*}_{1})=\operatorname*{arg\,min}_{\bm{\beta}}\mathbb{E}[(Y-(\beta_{0}+\beta_{1}X_{1}))^{2}].$
-We consider 2 primary scenarios for evaluating the PB inference methods. In both scenarios, the outcome is generated as
-$Y=X_{1}+\theta_{\text{\tiny{INT}}}X_{1}W_{1}+{\bf{W}}^{{\sf\scriptscriptstyle{T}}}\bm{\theta}+\epsilon$
+We consider two primary scenarios for evaluating the PB inference methods. In both scenarios, the outcome is generated as
+$Y=\theta_{X_{1}}X_{1}+\theta_{\text{\tiny{INT}}}X_{1}W_{1}+{\bf{W}}^{{\sf\scriptscriptstyle{T}}}\bm{\theta}+\epsilon$
 where ${\bf{W}}^{{\sf\scriptscriptstyle{T}}}=(W_{1},\dots,W_{q})$ and the error term $\epsilon$ is simulated independently from a standard normal distribution.
-We varied the coefficients of $\bm{\theta}$ and $\theta_{\text{\tiny{INT}}}$ as well as $q$ and the composition of ${\bf{Z}}$.
+We varied the coefficients of $\theta_{X_{1}}$, $\theta_{\text{\tiny{INT}}}$, and $\bm{\theta}$ as well as $q$ and the composition of ${\bf{Z}}$.
 Scenario 1
 We consider two sub-scenarios to illustrate Corollary 1.1, that is, that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when the sufficient condition holds.
-In both sub-scenarios, we let $p=10$, $\bm{\theta}=(1,1,0.5,0.5,0.25,{\bf{0}}_{1\times 5})^{{\sf\scriptscriptstyle{T}}}$, and $\theta_{\text{\tiny{INT}}}=0$.
+In both sub-scenarios, we let $q=10$, $\theta_{X_{1}}=1$, $\theta_{\text{\tiny{INT}}}=0$, and $\bm{\theta}=(1,1,0.5,0.5,0.25,{\bf{0}}_{1\times 5})^{{\sf\scriptscriptstyle{T}}}$.
 In Scenario 1a, we obtain a strong prediction model by regressing $Y$ on ${\bf{Z}}=(X_{1},W_{1},\dots,W_{10})^{{\sf\scriptscriptstyle{T}}}$ while in Scenario 1b we obtain a weaker prediction model by regressing $Y$ on ${\bf{Z}}=(X_{1},W_{5},X_{1}W_{5})^{{\sf\scriptscriptstyle{T}}}$.
+The $R^{2}$ of $\widehat{Y}$ for $Y$ is 0.9 and 0.6 in Scenarios 1a and 1b, respectively.
 Scenario 2
 We consider two sub-scenarios to illustrate the results in Corollary 1.1 and Corollary 2.1, which show that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ can be less efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ with a weak prediction model while $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is at least as efficient $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
-In both scenarios, we mimic a common scenario where the prediction model does not contain $X_{1}$ (e.g., in a genome-wide association study).
-In Scenario 2a, we let $p=9$, $\bm{\theta}=(0.2,0.1\times{\bf{1}}_{1\times 8})^{{\sf\scriptscriptstyle{T}}}$, and $\theta_{\text{\tiny{INT}}}=0$.
+In both scenarios, we mimic a common scenario where the prediction model does not contain $X_{1}$.
+In Scenario 2a, we let $q=9$, $\theta_{X_{1}}=1$, $\theta_{\text{\tiny{INT}}}=0$, and $\bm{\theta}=(0.2,0.1\times{\bf{1}}_{1\times 8})^{{\sf\scriptscriptstyle{T}}}$.
 We obtain the prediction model by regressing $Y$ on ${\bf{Z}}=(W_{1},\dots,W_{9})^{{\sf\scriptscriptstyle{T}}}$.
-In Scenario 2b, we let $p=3$, $\bm{\theta}=(0.25,0.25,0.25)^{{\sf\scriptscriptstyle{T}}}$, and $\theta_{\text{\tiny{INT}}}=0.25$.
+In Scenario 2b, we let $q=3$, $\theta_{X_{1}}=2$, $\theta_{\text{\tiny{INT}}}=0.25$, and $\bm{\theta}=(0.25,0.25,0.25)^{{\sf\scriptscriptstyle{T}}}$.
 We obtain the prediction model by regressing $Y$ on ${\bf{Z}}=(W_{1},W_{2},W_{3},W_{2}W_{1},W_{3}W_{1},W_{3}W_{2})^{{\sf\scriptscriptstyle{T}}}$.
-Scenario 2b uses a weaker prediction model than Scenario 2a.
+The $R^{2}$ of $\widehat{Y}$ for $Y$ is 0.5 and 0.4 in Scenarios 2a and 2b, respectively.
 We evaluate the performance of the PPI and CC methods: $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, and $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$.
-We also include two comparison methods: $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, serving as the baseline, and $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$, highlighting the limitations of naive inference.
+We also include two comparison methods: $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, serving as the baseline, and $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$, highlighting the limitations of naive inference.
 We present the percent bias for each estimator.
@@
 Table 1 presents the percent bias of the various estimators across the different scenarios and labeled dataset sizes.
-With the exception of $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$, the methods exhibit minimal bias as they are all consistent estimators for $\bm{\beta}^{*}_{1}$.
+With the exception of $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$, the methods exhibit minimal bias as they are all consistent estimators for $\beta^{*}_{1}$.
 The limitations of naive inference are particularly apparent in Scenarios 2a and 2b, which do not include $X_{1}$ in ML model training.
-$\hat{\bm{\gamma}}^{\text{\tiny{all}}}$ exhibits an average percent bias of -47% in Scenario 2a and -64% in Scenario 2b.
+$\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ exhibits an average percent bias of -47% in Scenario 2a and -64% in Scenario 2b.
 Figure 2 displays the CI length ratio across the different settings.
-In Scenario 1, we observe that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ and $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ perform similarly, with CI lengths much less 1, indicating that the PB inference methods have reduced variance relative to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
+In Scenario 1, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ and $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ perform similarly, with CI length ratios below 1, indicating that the PB inference methods have reduced standard error relative to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
 Comparing Scenario 1a and 1b, we note a significant decrease in the gains in statistical efficiency of the PB inference methods due to the weaker prediction model used in Scenario 1b.
@@
 In this scenario, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is asymptotically equivalent to $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$.
-Scenario 2 highlights the differences in performance between $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ and $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$, which underscores our findings in Corollary 1.1 and Corollary 2.1.
+Scenario 2 highlights the differences in performance between $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ and $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$, which underscore our findings in Corollary 1.1 and Corollary 2.1.
 In both sub-scenarios, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ consistently outperforms both $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ and is at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
 In Scenario 2a, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ achieves an average CI length ratio of 0.9 across the various sample sizes, indicating a gain in efficiency relative to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
-In Scenario 2b, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is comparable to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ for the larger labeled dataset sizes.
-In contrast, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ are outperformed by $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and have confidence intervals up to 30% longer than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ for the larger labeled dataset sizes.
+In Scenario 2b, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is comparable to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ for the larger labeled dataset sizes as there is not much to gain from the unlabeled data due to its relatively small size and the fact that $\widehat{Y}$ is a weak prediction of $Y$.
+In contrast, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ are outperformed by $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ in nearly all scenarios.
+In Scenario 2a, these estimators have CIs up to 30% longer than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ for the larger labeled dataset sizes.
 This behavior is exacerbated in Scenario 2b, where the CI of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is up to 50% longer than the CI for $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and up to 72% longer for $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$.
@@
 Scenario 1a
-     $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$0.181 0.124 0.650 0.430 0.165
+     $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$0.181 0.124 0.650 0.430 0.165
      $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ 0.035 -0.042 0.015 0.005 -0.005
@@
 Scenario 1b
-     $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$1.070 0.181 0.426 -0.423 -1.205
+     $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$1.070 0.181 0.426 -0.423 -1.205
      $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ 0.008 0.002 -0.016 -0.001 0.004
@@
 Scenario 2a
-     $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$46.980 -47.523 -45.946 -46.725 -47.695
+     $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$46.980 -47.523 -45.946 -46.725 -47.695
      $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$0.033 -0.023 0.027 0.004 0.005
@@
 Scenario 2b
-     $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$64.272 -64.115 -63.307 -63.759 -64.627
+     $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$64.272 -64.115 -63.307 -63.759 -64.627
      $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$0.012 -0.003 0.002 0.012 0.013
@@
 6 Discussion
-In this paper, we showed that adding a simple weight to the PPI estimator provides valid inference and is guaranteed to be at least as efficient as estimation using only labeled data.
+In this paper, we showed that adding a simple weight to the PPI estimator provides valid PB inference that is guaranteed to be at least as efficient as estimation using only labeled data.
 The weight is drawn from the work of Chen and Chen (2000) in the context of double sampling.
 In contrast, the original PPI estimator can be less efficient than estimation using only the labeled data, a property that can limit its practical utility.
-By incorporating the weight derived from Chen and Chen (2000), the augmented estimation approach principally balances the contribution of labeled and unlabeled data and extends the applicability of the PPI method to scenarios where leveraging all available data is essential for making valid and efficient inference.
+By incorporating the proposed weight, the augmented estimation approach principally balances the contribution of labeled and unlabeled data and extends the applicability of the PPI method to scenarios where leveraging all available data is essential for making valid and efficient inference.
 In addition to theoretical results that support these findings, our simulation studies demonstrate that the PPI method can result in substantial efficiency loss when the ML model is inaccurate.
-Additionally, we provide connections to several existing approaches from the economics and statistics literature that provide some interesting context for the PB inference problem.
-While we focused on least squares estimation, these results extend to other settings.
+Additionally, we provide connections to several existing approaches from the economics and statistics literature that provide interesting context for the PB inference problem.
+While we focused on least squares estimation, our results extend to other settings.
 The work of Chen and Chen (2000) and Angelopoulos et al.
-(2023a) apply to a wide variety of models beyond linear regression, such as parameters defining generalized linear models or parameters derived from estimating equations.
+(2023a) apply to a wide variety of settings beyond linear regression, such as generalized linear models.
 The work of Angelopoulos et al.
 (2023a) was pivotal in advancing the growing body of literature in PB inference.
-Recent work has considered settings where the ML model is not black box and where active learning is utilized to obtain the labeled data Zrnic and Candès (2024b, b).
+Recent work has considered settings where the ML model is not black box and where active learning is utilized to obtain the labeled data (Zrnic and Candès, 2024b, b).
 Additionally, two concurrent pre-prints have noted potential losses in efficiency of PB inference approaches.
 The authors of the PPI paper introduced a new method, PPI++, which aims to address this limitation of the PPI estimator (Angelopoulos et al., 2023b).
-The PPI++ method is designed to adaptively minimize the trace of the asymptotic variance-covariance of the proposed estimator to ensure that it yields an improvement over an approach based solely on labeled data.
+The PPI++ method is designed to adaptively minimize the trace of the asymptotic variance-covariance of the proposed estimator to ensure that it yields an improvement over an approach based on labeled data.
 In contrast, the approach of Chen and Chen (2000) minimizes the mean square error among a class of estimators with a linear augmentation term.
@@
 Unlike the PPI++ estimator, the PSPA estimator is designed to achieve element-wise variance reduction and was shown to outperform the PPI++ approach.
-The PSPA approach is based on augmenting estimating equations with information from $\widehat{Y}$, while the CC method provides a simple and natural extension to the PPI method by adding a weight to the augmentation term.
+The PSPA approach is based on augmenting estimating equations with information from $\widehat{Y}$, while the CC method provides a simple and natural extension to the PPI method by adding a weight to the linear augmentation term.
 Code availability
@@
 We denote that a matrix, ${\bf{M}}$, is positive semi-definite as ${\bf{M}}\succeq{\bf{0}},$ and positive definite as ${\bf{M}}\succ{\bf{0}}.$ For any estimator $\hat{\bm{\beta}}$, we let $\text{aVar}(\hat{\bm{\beta}})$ represent its asymptotic variance-covariance matrix.
-We define the difference in asymptotic variance-covariance matrices between two estimators $\hat{\bm{\beta}}_{1}$ and $\hat{\bm{\beta}}_{2}$ as $\Delta_{\text{aVar}}(\hat{\bm{\beta}}_{1},\hat{\bm{\beta}}_{2})=n[\text{aVar}(\hat{\bm{\beta}}_{1})-\text{aVar}(\hat{\bm{\beta}}_{2})].$
+We define the difference in asymptotic variance-covariance matrices between two estimators, $\hat{\bm{\beta}}_{1}$ and $\hat{\bm{\beta}}_{2}$, as $\Delta_{\text{aVar}}(\hat{\bm{\beta}}_{1},\hat{\bm{\beta}}_{2})=n[\text{aVar}(\hat{\bm{\beta}}_{1})-\text{aVar}(\hat{\bm{\beta}}_{2})].$
 Basic assumptions.
@@
 Proof  For scenario 1, assume that $\hat{f}({\bf{X}})=\mathbb{E}(Y\mid{\bf{X}})$. It follows that
-$\bm{\gamma}^{*}=\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})\mathbb{E}[{\bf{X}}\mathbb{E}(Y\mid{\bf{X}})]=\bm{\beta}^{*}$
+$\bm{\gamma}^{*}=[\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})]^{-1}\mathbb{E}[{\bf{X}}\mathbb{E}(Y\mid{\bf{X}})]=\bm{\beta}^{*}$
 so that
@@
 This inequality is strict provided that $\mathbb{E}(Y\mid{\bf{X}})\neq{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\beta}^{*}$.
-If the regression model is correctly specified (i.e., $\mathbb{E}(Y\mid{\bf{X}})={\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\beta}^{*}$), then $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ are asymptotically equivalent.
+If the regression model is also correctly specified (i.e., $\mathbb{E}(Y\mid{\bf{X}})={\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\beta}^{*}$), then $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ are asymptotically equivalent.
 For scenario 2, assume that $\hat{f}({\bf{Z}})=\mathbb{E}(Y\mid{\bf{Z}})$ where ${\bf{Z}}=({\bf{X}}^{{\sf\scriptscriptstyle{T}}},{\bf{W}}^{{\sf\scriptscriptstyle{T}}})^{{\sf\scriptscriptstyle{T}}}$ and ${\bf{W}}\perp Y\mid{\bf{X}}$.
 Similar to scenario 1,
-$\bm{\gamma}^{*}=\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})\mathbb{E}[{\bf{X}}\mathbb{E}(Y\mid{\bf{X}},{\bf{W}})]=\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})\mathbb{E}[{\bf{X}}\mathbb{E}(Y\mid{\bf{X}})]=\bm{\beta}^{*}.$
+$\bm{\gamma}^{*}=[\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})]^{-1}\mathbb{E}[{\bf{X}}\mathbb{E}(Y\mid{\bf{X}},{\bf{W}})]=[\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})]^{-1}\mathbb{E}[{\bf{X}}\mathbb{E}(Y\mid{\bf{X}})]=\bm{\beta}^{*}.$
 Therefore,
@@
 Lastly, for scenario 4, assume that $\hat{f}({\bf{W}})=E(Y\mid{\bf{W}})$ and ${\bf{X}}\perp Y\mid{\bf{W}}$. Then,
-$\bm{\gamma}^{*}=\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})\mathbb{E}[{\bf{X}}\mathbb{E}(Y\mid{\bf{W}})]=\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})\mathbb{E}[{\bf{X}}\mathbb{E}(Y\mid{\bf{W}},{\bf{X}})]=\bm{\beta}^{*}.$
+$\bm{\gamma}^{*}=[\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})]^{-1}\mathbb{E}[{\bf{X}}\mathbb{E}(Y\mid{\bf{W}})]=[\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})]^{-1}\mathbb{E}[{\bf{X}}\mathbb{E}(Y\mid{\bf{W}},{\bf{X}})]=\bm{\beta}^{*}.$
 Analogously to the argument for scenario 3,
@@
 Let the mean squared error (MSE) of $\hat{\bm{\beta}}$ for estimating $\bm{\beta}^{*}$ be defined as $\mathbb{E}[(\hat{\bm{\beta}}-\bm{\beta}^{*})^{{\sf\scriptscriptstyle{T}}}(\hat{\bm{\beta}}-\bm{\beta}^{*})]$.
-Analogously to the well-known result for the best linear predictor (e.g., Bickel and Doksum (2015), page 40), $\widehat{{\bf{W}}}^{\text{\tiny{CC}}}$ is the weight, ${\bf{W}}$, that minimizes the asymptotic MSE of estimators of the form $\hat{\bm{\beta}}^{\text{\tiny{lab}}}-{{\bf{W}}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right)$ for estimating $\bm{\beta}^{*}$.
+Analogously to the well-known result for the best linear predictor (e.g., Bickel and Doksum (2015), page 40), ${\bf{W}}^{\text{\tiny{CC}}}$ is the weight, ${\bf{W}}$, that minimizes the asymptotic MSE of estimators of the form $\hat{\bm{\beta}}^{\text{\tiny{lab}}}-{{\bf{W}}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right)$ for estimating $\bm{\beta}^{*}$.
 Proof  Letting $\hat{\bm{\beta}}_{{\bf{W}}}=\hat{\bm{\beta}}^{\text{\tiny{lab}}}-{\bf{W}}(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}})$, the MSE of estimating $\hat{\bm{\beta}}_{{\bf{W}}}$ is

### v3

@@
 Abstract
-Prediction-based (PB) inference is increasingly used in applications where the outcome of interest is difficult to obtain, but its predictors are readily available.
-Unlike traditional inference, PB inference performs statistical inference using a partially observed outcome and a set of covariates by leveraging a prediction of the outcome generated from a machine learning (ML) model.
-Motwani and Witten (2023) recently revisited two innovative PB inference approaches for ordinary least squares.
-They found that the method proposed by Wang et al.
-(2020) yields a consistent estimator for the association of interest when the ML model perfectly captures the underlying regression function.
-Conversely, the prediction-powered inference (PPI) method proposed by Angelopoulos et al.
-(2023a) yields valid inference regardless of the model’s accuracy.
-In this paper, we study the statistical efficiency of the PPI estimator.
-Our analysis reveals that a more efficient estimator, proposed 25 years ago by Chen and Chen (2000), can be obtained by simply adding a weight to the PPI estimator.
-We also contextualize PB inference with methods from the economics and statistics literature dating back to the 1960s.
-Our extensive theoretical and numerical analyses indicate that the Chen and Chen (CC) estimator offers a balance between robustness to ML model specification and statistical efficiency, making it the preferred choice for use in practice.
-Keywords: Missing data, prediction-based inference, semi-parametric inference
+From structural biology to epidemiology, predictions from machine learning (ML) models increasingly complement costly gold-standard data to enable faster, more affordable, and scalable scientific inquiry.
+In response, prediction-based (PB) inference has emerged to accommodate statistical analysis using a large volume of predictions together with a small amount of gold-standard data.
+The goals of PB inference are two-fold: (i) to mitigate bias from errors in predictions and (ii) to improve efficiency relative to traditional inference using only the gold-standard data.
+Motwani and Witten [2023] recently revisited two key PB inference approaches and found that only one method, Prediction-powered Inference (PPI) proposed by Angelopoulos et al.
+[2023a], achieves (i).
+In this paper, we find that PPI does not achieve (ii).
+We revisit the double sampling literature and show that, with a simple modification, PPI can be adjusted to provide theoretically justified improvements in efficiency.
+We also contextualize PB inference with economics and statistics literature dating back to the 1960s to highlight the utility of classical methods in this contemporary problem.
+Our extensive theoretical analyses, along with an analysis of UK Biobank data, indicate that our proposal effectively mitigates bias and improves efficiency, making it preferable for use in practice.
+Keywords: Machine learning, measurement error, missing data, prediction-based inference, semi-parametric inference
 1 Introduction
 In the rapidly evolving landscape of data science, accurately predicting an outcome is often just an initial step in scientific inquiry.
-Researchers are increasingly turning to machine learning (ML) models to obtain predictions for outcomes that are prohibitively time-consuming or expensive to directly measure, and then drawing statistical inferences from the ML model’s predictions (Hoffman et al., 2024; Fan et al., 2024; Gao et al., 2024; Miao and Lu, 2024).
-This prediction-based (PB) inference scenario was introduced by Wang et al in the context of global health and genomics applications, and is comprised of two steps outlined in Figure 1 (Wang et al., 2020).
+Researchers are increasingly turning to machine learning (ML) models to obtain predictions for outcomes that are prohibitively time-consuming or expensive to directly measure, and then drawing statistical inferences from the ML model’s predictions [Hoffman et al., 2024, Fan et al., 2024, Gao et al., 2024, Miao and Lu, 2024].
+This prediction-based (PB) inference scenario was introduced by Wang et al in the context of global health and genomics applications, and is comprised of two steps outlined in Figure 1 [Wang et al., 2020].
 First, the outcome of interest is imputed using a prediction from a pre-trained model.
-Then, the predictions are used to draw statistical inference on the parameter of interest, such as the association between an outcome and a set of covariates.
+Then, the predictions are used to draw statistical inference on the parameter of interest, such as the association between an outcome and a set of covariates, together with a small amount of gold-standard data  [^1] PB inference has focused on outcomes that are difficult to measure.
+Recent work has considered predicted covariates (e.g., Kluger et al.
+[2025])..
+In contrast to traditional inference that relies solely on gold-standard data, PB inference aims to leverage the predictions to improve statistical efficiency without introducing bias.
 Figure 1: Prediction-based (PB) inference.
 $Y$ is the outcome of interest, which is only partially observed, ${\bf{Z}}$ is a set of predictors of $Y$, $\widehat{Y}$ is the prediction of $Y$ based on ${\bf{Z}}$, and ${\bf{X}}$ is a set of covariates.
-The focus of scientific inquiry is on the association between ${\bf{X}}$ and $Y$.
-Traditional statistical inference relies solely on the data on ${\bf{X}}$ and $Y$ while PB inference incorporates $\widehat{Y}$ to maximize the use of all available data.
+The focus of scientific inquiry is on the association between $Y$ and ${\bf{X}}$.
+Traditional statistical inference relies solely on the data on $Y$ and ${\bf{X}}$ while PB inference incorporates $\widehat{Y}$ to maximize the use of all available data.
 Dark red represents labeled observations while light red represents unlabeled observations.
@@
 However, many traits and disease outcomes are only partially observed due to the time and cost of measurement.
-ML models are commonly used to predict missing outcomes based on fully observed baseline data and the predictions are utilized for genetic discovery (Tong et al., 2020; Alipanahi et al., 2021; Cosentino et al., 2023; Dahl et al., 2023; An et al., 2023; McCaw et al., 2024; Liu et al., 2024; Miao et al., 2024).
-Similarly, in the field of structural biology, AlphaFold provides highly accurate structural predictions without the need for expensive and labor-intensive experiments for proteomic studies (Jumper et al., 2021).
-In the analysis of electronic health records data, large language models can rapidly ascertain disease status at an accuracy competitive with manual chart review by experts, enabling downstream epidemiological and clinical studies (Yang et al., 2022; Zhou et al., 2022; Alsentzer et al., 2023).
-Given the relevance of PB inference, Motwani and Witten (2023) recently revisited two innovative methods in this area.
+ML models are commonly used to predict missing outcomes based on fully observed baseline data and the predictions are utilized for genetic discovery [Tong et al., 2020, Alipanahi et al., 2021, Cosentino et al., 2023, Dahl et al., 2023, An et al., 2023, McCaw et al., 2024, Liu et al., 2024, Miao et al., 2024].
+Similarly, in the field of structural biology, AlphaFold provides structural predictions without the need for expensive and labor-intensive experiments for proteomic studies [Jumper et al., 2021, Gan et al., 2022].
+In the analysis of electronic health records data, large language models can rapidly ascertain disease status at an accuracy competitive with manual chart review by experts, enabling downstream epidemiological and clinical studies [Yang et al., 2022, Zhou et al., 2022, Alsentzer et al., 2023].
+Given the relevance of PB inference, Motwani and Witten [2023] recently revisited two innovative methods in this field.
 In the context of least squares estimation, they showed that the approach proposed by Wang et al.
-(2020) provides a consistent estimator for the association between an outcome and a set of covariates only in restrictive scenarios.
+[2020] provides a consistent estimator for the association between an outcome and a set of covariates only in restrictive scenarios.
 For example, when the pre-trained model perfectly captures the regression function.
@@
 In contrast, the prediction powered inference (PPI) approach proposed by Angelopoulos et al.
-(2023a) possesses this necessary property and yields valid inference regardless of the quality of the model.
-Here we build on the work on Motwani and Witten to study the statistical efficiency of the PPI method.
-We show that a more efficient estimator can be obtained by simply adding a weight to the PPI estimator, an approach that was proposed by Chen & Chen (CC) more than 25 years ago in the context of double sampling designs (Chen and Chen, 2000).
+[2023a] possesses this necessary property and yields valid inference regardless of the quality of the model.
+This finding is critical given that the PPI paper has been cited over 100 times in its first year of publication in Science.
+In this paper, we study the statistical efficiency of the PPI estimator and show that it can be outperformed by traditional inference.
+We propose a more efficient estimator that can be obtained by simply adding a weight to the PPI estimator, an approach introduced by Chen & Chen (CC) 25 years ago in the context of double sampling designs [Chen and Chen, 2000].
 We begin by formally introducing the PB inference problem and reviewing the contributions of Motwani and Witten in Section 2.
@@
 Our key finding is that the CC estimator is optimal with respect to statistical efficiency in a class of estimators to which the PPI estimator belongs.
-We also point to earlier literature from the statistics and economics literature dating back to the 1960s that provides important historical context for the PB inference problem in Section 4.
+We also review modern and classical literature from statistics and economics dating back to the 1960s that provides important context for the PB inference problem in Section 4.
 In Section 5, we evaluate our findings with extensive simulation studies.
-Our results show that the CC estimator offers a balance between robustness to ML model specification and statistical efficiency, making it the preferred choice for use in practice.
+In Section 6, we study the association between android fat mass and clinical and demographic covariates in the South Asian population in the UK Biobank to demonstrate the practical performance of our method.
 We close with additional discussion in Section 7.
+All in all, our theoretical and numerical results show that the CC estimator offers a balance between robustness to ML model specification and statistical efficiency, making it the preferred choice for use in PB inference problems.
 2 Problem setting & background
@@
 We focus on the case in which $\mathscr{L}$ and $\mathscr{U}$ are independent and identically distributed samples from the population of interest.
-More complex scenarios have been the focus of recent work (Angelopoulos et al., 2023a; Zrnic and Candès, 2024a).
+More complex scenarios have been the focus of recent work [Angelopoulos et al., 2023a, Zrnic and Candès, 2024a].
 Additionally, $\hat{f}$ is assumed to be pre-trained with a dataset that is independent of $\mathscr{L}$ and $\mathscr{U}$ and is therefore treated as “fixed” (i.e., the variation in estimating $\hat{f}$ is not of interest in downstream inference).
 In fact, $\hat{f}$ may be a black box function in that its training data and specification are not be available to the users.
-Approaches that address uncertainty in the estimation of $\hat{f}$ have recently been proposed in Zrnic and Candès (2024b).
+Approaches that address uncertainty in the estimation of $\hat{f}$ have recently been proposed by Zrnic and Candès [2024b].
 2.2 Parameter of interest
-Following Motwani and Witten (2023), we consider least squares estimation and aim to make inference on the population parameter
+Following Motwani and Witten [2023], we consider least squares estimation and aim to make inference on the population parameter
 $\bm{\beta}^{*}\coloneq\operatorname*{arg\,min}_{\bm{\beta}}\mathbb{E}[(Y-{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\beta})^{2}]=[\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})]^{-1}\mathbb{E}({\bf{X}}Y),$
@@
 Incorporating information from $\widehat{Y}$ should not change the parameter being estimated and the credibility of inferences should not depend on the quality of predictions from $\hat{f}$.
-The need for PB inference is highlighted by limitations of “naive” inference, in which inference on $\widehat{Y}$ is performed in lieu of inference on $Y$ (Wang et al., 2020).
+The need for PB inference is highlighted by limitations of “naive” inference, in which inference on $\widehat{Y}$ is performed in lieu of inference on $Y$ [Wang et al., 2020].
 The parameter targeted by naive inference is
@@
 The first is the proposal of Wang et al.
-(2020), which is a two-step procedure to incorporate $\widehat{Y}$ into estimation.
+[2020], which is a two-step procedure to incorporate $\widehat{Y}$ into estimation.
 First, $\mathscr{L}$ is used to model the relationship between $Y$ and $\widehat{Y}$ to address potential inaccuracies in the predictions.
-Then, this model is used together with $\mathscr{U}$ to make inference on $\bm{\beta}^{*}$, with analytical or bootstrap approaches described.
+Then, this model is used together with $\mathscr{U}$ to make inference on $\bm{\beta}^{*}$ with either analytical or bootstrap approaches.
 In their illuminating work, Motwani and Witten show that the estimator of Wang et al is consistent for $\bm{\beta}^{*}$ under the strong and unrealistic assumption that the ML model captures the true regression function.
@@
 This approach dates back to at least Robins et al.
-(1994), and has been applied in multiple settings to improve estimation efficiency outside of the PB inference problem (e.g., Chen and Chen (2000); Tsiatis (2006); Tian et al.
-(2012); Wang and Wang (2015)).
+[1994], and has been applied in multiple settings to improve estimation efficiency outside of the PB inference problem (e.g., Chen and Chen [2000], Tsiatis [2006], Tian et al.
+[2012], Wang and Wang [2015]).
 We return to this point in Section 4.
@@
 $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}=\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}\right),$
-where $\hat{\bm{\gamma}}^{\text{\tiny{lab}}}$ is the solution to $\sum_{i=1}^{n_{\text{\tiny{lab}}}}\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma})=\boldsymbol{0}$ [^1] Angelopoulos et al.
-(2023a) express $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ as $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}-\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\beta}}^{\text{\tiny{lab}}}\right)$ where $\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\beta}}^{\text{\tiny{lab}}}\right)$ is referred to as a “rectifier.” The rectifier serves the purpose of correcting potential bias in $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ due to inaccuracies in $\widehat{Y}$ for $Y$.
+where $\hat{\bm{\gamma}}^{\text{\tiny{lab}}}$ is the solution to $\sum_{i=1}^{n_{\text{\tiny{lab}}}}\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma})=\bm{0}$ [^2] Angelopoulos et al.
+[2023a] express $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ as $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}-\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\beta}}^{\text{\tiny{lab}}}\right)$ where $\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\beta}}^{\text{\tiny{lab}}}\right)$ is referred to as a “rectifier.” The rectifier serves the purpose of correcting potential bias in $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ due to inaccuracies in $\widehat{Y}$ for $Y$.
 We reframe the PPI method to study its statistical efficiency..
@@
 $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}=\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right),$
-where $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$ is the solution to $\sum_{i=1}^{n}\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma})=\boldsymbol{0}$.
+where $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$ is the solution to $\sum_{i=1}^{n}\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma})=\bm{0}$.
 The only difference between $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is that all of the data, rather than the unlabeled data, is used to estimate $\bm{\gamma}^{*}$ in the augmentation term (i.e., $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ is replaced by $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$).
@@
 All of the asymptotic results for $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ are detailed in Appendix A.
-Theorem 1 (Influence function expansion of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$)
+Theorem 1 (Influence function expansion of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$).
 Let $R_{i}$ be the indicator of whether the $i$th observation is labeled. Then,
@@
 Corollary 1.1 shows that the benefit of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ scales directly with the predictiveness of $\widehat{Y}$ for $Y$ and inversely with the proportion of labeled data, $\pi$.
-Corollary 1.1 (Comparison of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ with $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$)
-Let $\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}=\mathbb{E}\left(\epsilon_{\widehat{Y}}^{2}{\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\right)$, $\boldsymbol{\Sigma}_{Y,\widehat{Y}}=\mathbb{E}\left(\epsilon_{Y}\epsilon_{\widehat{Y}}{\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\right)$, $\epsilon_{Y}=Y-{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\beta}^{*}$ and $\epsilon_{\widehat{Y}}=\widehat{Y}-{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma}^{*}$.
+Corollary 1.1 (Comparison of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ with $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$).
+Let $\epsilon_{Y}=Y-{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\beta}^{*}$, $\epsilon_{\widehat{Y}}=\widehat{Y}-{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma}^{*}$, $\bm{\Sigma}_{\widehat{Y},\widehat{Y}}=\mathbb{E}\left(\epsilon_{\widehat{Y}}^{2}{\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\right)$, and $\bm{\Sigma}_{Y,\widehat{Y}}=\mathbb{E}\left(\epsilon_{Y}\epsilon_{\widehat{Y}}{\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\right)$.
 The difference in the asymptotic variance-covariance (aVar) of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is
-$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})$ $=n\left[\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{lab}}})-\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})\right]=\left(\frac{1}{\pi}-1\right){\bf{A}}\left(2\boldsymbol{\Sigma}_{Y,\widehat{Y}}-\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}\right){\bf{A}}.$
-$\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is therefore more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if $2\boldsymbol{\Sigma}_{Y,\widehat{Y}}\succ\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}$.
+$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})$ $=n\left[\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{lab}}})-\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})\right]=\left(\frac{1}{\pi}-1\right){\bf{A}}\left(2\bm{\Sigma}_{Y,\widehat{Y}}-\bm{\Sigma}_{\widehat{Y},\widehat{Y}}\right){\bf{A}}.$
+$\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is therefore more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if $2\bm{\Sigma}_{Y,\widehat{Y}}\succ\bm{\Sigma}_{\widehat{Y},\widehat{Y}}$.
 A sufficient condition for $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ to have lower variance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ is that
@@
 This condition is similar to the result for mean estimation provided in the Supplementary Materials of Angelopoulos et al.
-(2023a) and can be used to describe several scenarios when $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is guaranteed to have lower asymptotic variance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when the ML model captures the true regression function (see Appendix A.4).
+[2023a] and can be used to describe several scenarios when $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is guaranteed to have lower asymptotic variance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when the ML model captures the true regression function (see Appendix A.4).
 Here we consider a simple example to demonstrate when $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ gains efficiency over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ under homoskedasticity.
-Example 1 (Efficiency gain of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ under homoskedasticity)
+Example 1 (Efficiency gain of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ under homoskedasticity).
 In the homoskedastic setting where $\mathbb{E}(\epsilon_{Y}^{2}\mid{\bf{X}})=\sigma_{Y}^{2}$, $\mathbb{E}(\epsilon_{\widehat{Y}}^{2}\mid{\bf{X}})=\sigma_{\widehat{Y}}^{2}$, and $\mathbb{E}(\epsilon_{Y}\epsilon_{\widehat{Y}}\mid{\bf{X}})=\sigma_{Y,\widehat{Y}}$,
@@
 where $\rho_{Y,\widehat{Y}}=\sigma_{Y,\widehat{Y}}/\sigma_{Y}\sigma_{\widehat{Y}}$ is the correlation of the residuals for $Y$ and $\widehat{Y}$ given ${\bf{X}}$.
-$\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ therefore gains efficiency if and only if the correlation of the residuals is at least half of the ratio of residual standard deviations of $\widehat{Y}$ to that of $Y$.
 Put another way, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ improves over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ with respect to efficiency when $\widehat{Y}$ is a strong predictor for $Y$ after adjusting for ${\bf{X}}$, though $\widehat{Y}$ does not necessarily need to be derived from a ML model that captures the true regression function.
@@
 3.2 Asymptotic analysis of the CC estimator
-To further analyze the efficiency of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ we draw on a landmark paper that introduced a class of augmented inverse probability weighted (AIPW) estimators which includes all regular asymptotic linear estimators (Robins et al., 1994).
+To further analyze the efficiency of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ we draw on the landmark paper that introduced a class of augmented inverse probability weighted (AIPW) estimators which includes all regular asymptotic linear estimators [Robins et al., 1994].
 In the current setting, where the goal is to estimate linear regression coefficients $\bm{\beta}^{*}$ with $Y$ partially observed, but $\widehat{Y}$ fully observed, these estimators have influence functions of the form
@@
 By Theorem 1, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is an AIPW estimator and can be seen as an approximation to the optimal estimator with $\bm{\omega}(\widehat{Y},{\bf{X}})=\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})$.
-The PPI estimator is therefore not the most efficient estimator in its class and, from Corollary 1.1, can even be less efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when $\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}\succ 2\boldsymbol{\Sigma}_{Y,\widehat{Y}}$.
+The PPI estimator is therefore not the most efficient estimator in its class and, from Corollary 1.1, can even be less efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when $\bm{\Sigma}_{\widehat{Y},\widehat{Y}}\succ 2\bm{\Sigma}_{Y,\widehat{Y}}$.
 A PB inference strategy should enhance statistical efficiency relative to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ whenever possible and, at the very least, should not be negatively impacted by incorporating unlabeled data into estimation.
@@
 Moreover, when the model for $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$ is misspecified, an efficiency gain is not guaranteed.
-To overcome these difficulties, Chen and Chen (2000) proposed a particular choice of $\bm{\omega}(\widehat{Y},{\bf{X}})$ that is guaranteed to be at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and that does not require modeling of $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$.
-Their approach is motivated by the double-sampling design which, apart from differences in terminology and notation (i.e., they regard $\widehat{Y}$ as a mismeasured version of $Y$), is analogous to the partially labeled data setting considered here.
+To overcome these difficulties, Chen and Chen [2000] proposed a particular choice of $\bm{\omega}(\widehat{Y},{\bf{X}})$ that is guaranteed to be at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and that does not require modeling of $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$.
+Their approach is motivated by the double-sampling design which, apart from differences in terminology (i.e., they regard $\widehat{Y}$ as a mismeasured version of $Y$), is analogous to the partially labeled data setting considered here.
 The CC estimator approximates $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$ with
@@
 obtained with $\mathscr{L}$.
-Therefore, a simple way to ensure that an augmented estimator like the PPI estimator has asymptotic variance at least as low as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ is to add a weight function in front of the augmentation term.
+Therefore, a simple way to ensure that an estimator like the PPI estimator has asymptotic variance at least as low as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ is to add a weight function in front of the augmentation term.
 In fact, the framework of Robins et al.
-(1994) implies that all regular asymptotically linear estimators for $\bm{\beta}^{*}$ augment $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ with an augmentation term that incorporates information from $\widehat{Y}$.
+[1994] implies that all regular asymptotically linear estimators for $\bm{\beta}^{*}$ augment $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ with an augmentation term that incorporates information from $\widehat{Y}$.
 We present the influence function expansion for $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ in Theorem 2 to further study its asymptotic variance and compare it with $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$.
 All of the asymptotic results provided for $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ are detailed in Appendix C.
-Theorem 2 (Influence function expansion of $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$)
-Following Chen and Chen (2000),
+Theorem 2 (Influence function expansion of $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$).
+Following Chen and Chen [2000],
 $n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{CC}}}-\bm{\beta}^{*}\right)=n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{IF}}^{\text{\tiny{CC}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})+o_{p}(1),$
@@
 Together these results illustrate that the CC method is both robust to ML model specification and optimizes statistical efficiency, making it preferable to the PPI method in practice.
-Corollary 2.1 ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$)
+Corollary 2.1 ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$).
 The difference in the asymptotic variance-covariance of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is
-$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{CC}}})=\left(\frac{1}{\pi}-1\right){\bf{A}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}\boldsymbol{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}{\bf{A}}\succeq{\bf{0}}.$
+$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{CC}}})=\left(\frac{1}{\pi}-1\right){\bf{A}}\bm{\Sigma}_{Y,\widehat{Y}}\bm{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\bm{\Sigma}_{Y,\widehat{Y}}{\bf{A}}\succeq{\bf{0}}.$
 Therefore, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is guaranteed to be at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
-Corollary 2.2 ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is the most efficient estimator in its class)
+Corollary 2.2 ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is the most efficient estimator in its class).
 Let the mean squared error (MSE) of an estimator $\hat{\bm{\beta}}$ for $\bm{\beta}^{*}$ be defined as $\mathbb{E}[(\hat{\bm{\beta}}-\bm{\beta}^{*})^{{\sf\scriptscriptstyle{T}}}(\hat{\bm{\beta}}-\bm{\beta}^{*})]$.
-Analogously to the well-known result for the best linear predictor (e.g., Bickel and Doksum (2015), page 40), ${\bf{W}}^{\text{\tiny{CC}}}$ is the weight, ${\bf{W}}$, that minimizes the asymptotic MSE of estimators of the form $\hat{\bm{\beta}}^{\text{\tiny{lab}}}-{{\bf{W}}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right)$ for estimating $\bm{\beta}^{*}$.
-From Corollary 2.1, we note that in contrast to $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})$, the covariance term, $\boldsymbol{\Sigma}_{Y,\widehat{Y}}$, in $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{CC}}})$ is normalized by the residual variance of $\widehat{Y}$, $\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}$.
+Analogously to the well-known result for the best linear predictor (e.g., Bickel and Doksum [2015], page 40), ${\bf{W}}^{\text{\tiny{CC}}}$ is the weight, ${\bf{W}}$, that minimizes the asymptotic MSE of estimators of the form $\hat{\bm{\beta}}^{\text{\tiny{lab}}}-{{\bf{W}}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right)$ for estimating $\bm{\beta}^{*}$.
+From Corollary 2.1, we note that in contrast to $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})$, the covariance term, $\bm{\Sigma}_{Y,\widehat{Y}}$, in $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{CC}}})$ is normalized by the residual variance of $\widehat{Y}$, $\bm{\Sigma}_{\widehat{Y},\widehat{Y}}$.
 This normalization effectively scales the covariance according to the precision of $\widehat{Y}$ and enables $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ to always be at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
@@
 4 Related literature
-Existing PB inference literature has established connections with semi-supervised inference, semi-parametric inference, missing data, and measurement error (Angelopoulos et al., 2023a, b; Miao et al., 2023; Zrnic and Candès, 2024a, b).
-Here we draw connections between PB inference methods and methods from the economics and statistics literature dating back to the 1960s.
+Existing PB inference literature has established connections with semi-supervised inference, semi-parametric inference, missing data, and measurement error [Angelopoulos et al., 2023a, b, Miao et al., 2023, Zrnic and Candès, 2024a, b, Angelopoulos et al., 2024].
+Here we draw connections between PB inference methods and modern and classical methods from the economics and statistics literature dating back to the 1960s.
 These connections reveal that aspects of the PB inference problem have been explored in various areas, albeit with differences in motivation, terminology, and notation.
@@
 4.1 Seemingly unrelated regression & its use in statistical genetics
-Seemingly unrelated regression (SUR) was introduced by economist Arnold Zellner in 1962 for simultaneous estimation of multiple related regression models (Zellner, 1962; Swamy and Mehta, 1975; Schmidt, 1977).
+Seemingly unrelated regression (SUR) was introduced by economist Arnold Zellner in 1962 for simultaneous estimation of multiple related regression models [Zellner, 1962, Swamy and Mehta, 1975, Schmidt, 1977].
 The regression models have different outcome variables, but are assumed to have correlated errors.
@@
 $Y={\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\beta}^{*}+\epsilon\quad\mbox{and}\quad\widehat{Y}={\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma}^{*}+\epsilon^{\prime}\quad$
-where $\mathbb{E}(\epsilon\mid{\bf{X}})=0$ and $\mathbb{E}(\epsilon^{\prime}\mid{\bf{X}})=0.$ An estimator for $\bm{\beta}^{*}$ that is based on augmenting $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ was proposed by Conniffe (1985) and is given by:
+where $\mathbb{E}(\epsilon\mid{\bf{X}})=0$ and $\mathbb{E}(\epsilon^{\prime}\mid{\bf{X}})=0.$ An estimator for $\bm{\beta}^{*}$ that is based on augmenting $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ was proposed by Conniffe [1985] and is given by:
 $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}=\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\widehat{W}^{\text{\tiny{SUR}}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right)\quad\mbox{where}\quad\widehat{W}^{\text{\tiny{SUR}}}=\frac{\sum_{i=1}^{n_{\text{\tiny{lab}}}}(Y_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\hat{\bm{\beta}}^{\text{\tiny{lab}}})(\widehat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\hat{\bm{\gamma}}^{\text{\tiny{lab}}})}{\sum_{i=1}^{n_{\text{\tiny{lab}}}}(\widehat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\hat{\bm{\gamma}}^{\text{\tiny{lab}}})^{2}}.$
@@
 In the extreme case where $\widehat{Y}$ is a perfect predictor for $Y$ such that $\widehat{Y}=Y$, $\widehat{W}^{\text{\tiny{SUR}}}=1$ and $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ is simply $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$.
-In the homoskedastic case when $\text{Var}(\epsilon\mid{\bf{X}})$ and $\text{Cov}(\epsilon,\epsilon^{\prime}\mid{\bf{X}})$ are constant and do not depend on ${\bf{X}}$, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is asymptotically equivalent to $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ (Chen and Chen, 2000).
-Moreover, when the conditional distribution of $(\epsilon,\epsilon^{\prime})$ given ${\bf{X}}$ is bivariate normal with mean $\boldsymbol{0}$ and has constant variance-covariance, $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ (and thus $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ as well) is the maximum likelihood estimator of $\bm{\beta}^{*}$ and fully efficient under the assumed parametric model (Conniffe, 1985).
+In the homoskedastic case when $\text{Var}(\epsilon\mid{\bf{X}})$ and $\text{Cov}(\epsilon,\epsilon^{\prime}\mid{\bf{X}})$ are constant and do not depend on ${\bf{X}}$, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is asymptotically equivalent to $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ [Chen and Chen, 2000].
+Moreover, when the conditional distribution of $(\epsilon,\epsilon^{\prime})$ given ${\bf{X}}$ is bivariate normal with mean $\bm{0}$ and has constant variance-covariance, $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ (and thus $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ as well) is the maximum likelihood estimator of $\bm{\beta}^{*}$ and fully efficient under the assumed parametric model [Conniffe, 1985].
 In other scenarios, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ achieves a lower asymptotic variance than $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$.
 The SUR estimator was recently applied for a PB inference problem inspired by genome-wide association studies conducted with population biobanks.
-These studies often rely on pre-trained models to impute outcomes of interest (i.e., phenotypes or traits) as they are partially missing due to the time and expense of ascertainment (McCaw et al., 2024).
+These studies often rely on pre-trained models to impute outcomes of interest (i.e., phenotypes or traits) as they are partially missing due to the time and expense of ascertainment [McCaw et al., 2024].
 The proposed method, “SynSurr”, assumes the true and predicted outcomes follow a bivariate normal distribution.
@@
 Building on SynSurr, Miao et al.
-(2024) proposed the “POP-GWAS” estimator for a variety of regression models utilizing the augmentation technique.
+[2024] proposed the “POP-GWAS” estimator for a variety of regression models utilizing the augmentation technique.
 The POP-GWAS estimator is
@@
 Outside of statistical genetics, the augmented estimation approach used by the PPI and CC methods has been widely used for treatment effect estimation in the statistics and biostatistics literature (e.g., Leon et al.
-(2003); Lu and Tsiatis (2008); Tsiatis et al.
-(2008); Gilbert et al.
-(2009); Zhang and Gilbert (2010); Tian et al.
-(2012); Yang and Ding (2020)).
+[2003], Lu and Tsiatis [2008], Tsiatis et al.
+[2008], Gilbert et al.
+[2009], Zhang and Gilbert [2010], Tian et al.
+[2012], Yang and Ding [2020]).
 Many of these approaches build on methods from the survey sampling literature (e.g., Cassel et al.
-(1976); Särndal et al.
-(2003)).
+[1976], Särndal et al.
+[2003]).
 For example, Tsiatis et al.
-(2008) consider covariate adjusted analyses for comparing two treatments in randomized clinical trials.
+[2008] consider covariate adjusted analyses for comparing two treatments in randomized clinical trials.
 Letting $X$ be an indicator for randomization to the treated or control arms, the focus of inference is on $\beta$ in the model
@@
 To utilize information in baseline covariates, ${\bf{Z}}$, Tsiatis et al.
-(2008) propose an augmented estimator of the treatment effect given by
+[2008] propose an augmented estimator of the treatment effect given by
 $\widehat{\Delta}-\sum_{i=1}^{n}(X_{i}-\bar{X})[n_{0}^{-1}h^{(0)}({\bf{Z}}_{i})+n_{1}^{-1}h^{(1)}({\bf{Z}}_{i})],$ (2)
@@
 The augmentation technique has also been utilized to improve the efficiency of Monte Carlo estimation.
-Specifically, Rothery (1982) considers using a control variate, a statistic whose properties are known, to improve precision when estimating power via simulation.
+Specifically, Rothery [1982] considers using a control variate, a statistic whose properties are known, to improve precision when estimating power via simulation.
 For exposition, suppose interest lies in estimating the power of a test $T$ of $H_{0}:\theta=0$ for some scalar parameter $\theta$.
@@
 $\frac{1}{n}\sum_{i=1}^{n}\left[Y_{i}-\lambda\{Z_{i}-\mu_{Z}\}\right].$ (3)
-The variance of the augmented estimator is minimized at $\lambda^{*}=\frac{\text{Cov}(Y,Z)}{\text{Var}(Z)}$ (Rothery, 1982; Davidson and MacKinnon, 1992).
-The augmentation term reduces the variance of the unadjusted power estimate $\overline{Y}$ in proportion to the correlation between $Y$ and $Z$.
+The variance of the augmented estimator is minimized at $\lambda^{*}=\frac{\text{Cov}(Y,Z)}{\text{Var}(Z)}$ [Rothery, 1982, Davidson and MacKinnon, 1992].
+The augmentation term in 3 reduces the variance of the unadjusted power estimate $\overline{Y}$ in proportion to the correlation between $Y$ and $Z$.
 Building on work from the survey sampling literature, a similar idea was later used for mean estimation in the semi-supervised setting by Zhang et al.
-(2019).
+[2019].
 5 Numerical studies
@@
      $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$0.014 -0.006 -0.003 0.011 0.010
-Table 1: Percent biases of the various estimators. Negative values indicate underestimation, while positive values indicate overestimation of the true parameter.
-Figure 2: Confidence interval length ratio of the PB inference methods compared to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
-A ratio less than 1 indicates that the PB inference method has a confidence interval that is shorter than the confidence interval for $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
+Table 1: Percent biases of the PPI estimators ($\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$), the CC estimator ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$), the baseline estimator ($\hat{\bm{\beta}}^{\text{\tiny{lab}}}$), and the naive estimator ($\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$).
+Negative values indicate underestimation, while positive values indicate overestimation of the true parameter.
+Figure 2: Confidence interval length ratio of the PB inference methods compared to baseline estimation with the labeled data ($\hat{\bm{\beta}}^{\text{\tiny{lab}}}$).
+A ratio less than 1 indicates that the PB inference method has a confidence interval that is shorter than the confidence interval using $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
+6 UK Biobank Analysis
+Central obesity, defined as excessive accumulation of body fat around abdominal region, is strongly associated with cardiometabolic diseases such as type 2 diabetes and cardiovascular disease.
+A major challenge in studying central obesity is the limited availability of direct body composition measurements.
+Android fat mass (AFM), measured via a dual-energy X-ray absorptiometry (DEXA) scan, is a key marker for assessing central obesity in clinical practice [Zhu et al., 2022].
+However, due to its high cost, DEXA-derived AFM is usually only measured in a small subset of participants in large-scale studies.
+For example, in the UK Biobank, a prospective cohort study measuring genetic, lifestyle and health information on approximately 500,000 UK participants, DEXA-derived AFM is incompletely ascertained, particularly among populations populations historically underrepresented in genetic research [Sudlow et al., 2015].
+For instance, out of $5,397$ South Asian participants in the UK Biobank, only $288$ (5.3%) participants have undergone DEXA scans.
+[…]
+Table 2 presents point estimates, standard errors, and CI ratios (or equivalently standard errors length ratios) across the two prediction model settings for four estimators: the PPI estimators ($\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$), the CC estimator ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$), and the baseline estimator ($\hat{\bm{\beta}}^{\text{\tiny{lab}}}$).
+Echoing our theoretical and simulation study results, all methods yield similar point estimates as they are all consistent estimators for the same parameter.
+For the analysis with the strong prediction model, $\widehat{Y}$ provides additional predictive power beyond the regression of $Y$ on ${\bf{X}}$, satisfying the sufficient condition in Corollary 2.1.
+Consequently, both $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ outperform $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
+In fact, all three PB inference methods ($\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, and $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$) perform similarly, yielding substantially smaller standard errors compared to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ with average CI ratios below 0.85 for all methods.
+The analysis with the weak prediction model highlights the difference in performance between the PPI and CC methods and reinforces the findings in Corollary 2.1 and Corollary 2.2.
+Here, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ yield larger standard errors than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ for all parameters of interest, with CIs 38.5% longer on average.
+In contrast, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is no worse than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ for all parameters of interest.
+Strong Prediction Model
+Intercept Age Sex BMI
+$\hat{\bm{\beta}}^{\text{\tiny{lab}}}$
+EST -4.857 -0.003 0.813 0.173
+SE 0.308 0.004 0.070 0.009
+CI Ratio 1.00 1.00 1.00 1.00
+$\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$
+EST -4.513 -0.007 0.904 0.164
+SE 0.261 0.004 0.063 0.007
+CI Ratio 0.85 0.86 0.89 0.80
+$\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$
+EST -4.527 -0.007 0.900 0.165
+SE 0.260 0.004 0.063 0.007
+CI Ratio 0.85 0.86 0.89 0.80
+$\hat{\bm{\beta}}^{\text{\tiny{CC}}}$
+EST -4.525 -0.006 0.905 0.163
+SE 0.255 0.004 0.062 0.007
+CI Ratio 0.83 0.85 0.89 0.77
+Weak Prediction Model
+Intercept Age Sex BMI
+$\hat{\bm{\beta}}^{\text{\tiny{lab}}}$
+EST -4.857 -0.003 0.813 0.173
+SE 0.308 0.004 0.070 0.009
+CI Ratio 1.00 1.00 1.00 1.00
+$\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$
+EST -4.530 -0.006 0.835 0.165
+SE 0.416 0.006 0.098 0.012
+CI Ratio 1.35 1.43 1.40 1.36
+$\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$
+EST -4.546 -0.006 0.834 0.165
+SE 0.415 0.006 0.098 0.012
+CI Ratio 1.35 1.43 1.40 1.36
+$\hat{\bm{\beta}}^{\text{\tiny{CC}}}$
+EST -4.779 -0.004 0.819 0.171
+SE 0.294 0.004 0.072 0.008
+CI Ratio 0.95 0.96 1.03 0.93
+Table 2: UK Biobank analysis of android fat mass based on the PPI estimators ($\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$), the CC estimator ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$), the baseline estimator ($\hat{\bm{\beta}}^{\text{\tiny{lab}}}$).
+The strong prediction model has an $R^{2}$ of 0.8, while the weak prediction model has an $R^{2}$ of 0.5.
+Estimates (EST), standard errors (SE), and confidence interval (CI) ratios are reported to evaluate performance.
+A CI ratio less than 1 indicates that a method yields a confidence interval shorter than the CI using $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
 7 Discussion
 In this paper, we showed that adding a simple weight to the PPI estimator provides valid PB inference that is guaranteed to be at least as efficient as estimation using only labeled data.
-The weight is drawn from the work of Chen and Chen (2000) in the context of double sampling.
-In contrast, the original PPI estimator can be less efficient than estimation using only the labeled data, a property that can limit its practical utility.
+The weight is drawn from the work of Chen and Chen [2000] in the context of double sampling.
+In contrast, the original PPI estimator can be less efficient than estimation using only the labeled data, a property that threatens its practical utility.
 By incorporating the proposed weight, the augmented estimation approach principally balances the contribution of labeled and unlabeled data and extends the applicability of the PPI method to scenarios where leveraging all available data is essential for making valid and efficient inference.
-In addition to theoretical results that support these findings, our simulation studies demonstrate that the PPI method can result in substantial efficiency loss when the ML model is inaccurate.
+In addition to theoretical results that support these findings, our simulation studies and application to UK Biobank data demonstrate that the PPI method can result in substantial efficiency loss when the ML model is inaccurate.
 Additionally, we provide connections to several existing approaches from the economics and statistics literature that provide interesting context for the PB inference problem.
-While we focused on least squares estimation, our results extend to other settings.
-The work of Chen and Chen (2000) and Angelopoulos et al.
-(2023a) apply to a wide variety of settings beyond linear regression, such as generalized linear models.
+While we focused on least squares estimation for exposition, our results extend to other settings.
+The work of Chen and Chen [2000] and Angelopoulos et al.
+[2023a] apply to a wide variety of settings beyond linear regression, including Z-estimation problems.
 The work of Angelopoulos et al.
-(2023a) was pivotal in advancing the growing body of literature in PB inference.
-Recent work has considered settings where the ML model is not black box and where active learning is utilized to obtain the labeled data (Zrnic and Candès, 2024b, b).
+[2023a] was pivotal in advancing the growing body of literature in PB inference.
+Recent work has considered settings where the ML model is not black box and where active learning is utilized to obtain the labeled data [Zrnic and Candès, 2024b, b].
 Additionally, two concurrent pre-prints have noted potential losses in efficiency of PB inference approaches.
-The authors of the PPI paper introduced a new method, PPI++, which aims to address this limitation of the PPI estimator (Angelopoulos et al., 2023b).
+The authors of the PPI paper introduced a new method, PPI++, which aims to address this limitation of the PPI estimator [Angelopoulos et al., 2023b].
 The PPI++ method is designed to adaptively minimize the trace of the asymptotic variance-covariance of the proposed estimator to ensure that it yields an improvement over an approach based on labeled data.
-In contrast, the approach of Chen and Chen (2000) minimizes the mean square error among a class of estimators with a linear augmentation term.
+In contrast, the approach of Chen and Chen [2000] minimizes the mean square error among a class of estimators with a linear augmentation term.
 Simultaneously, Miao et al.
-(2023) introduced the PoSt-Prediction Adaptive inference (PSPA) estimator that is both robust to the accuracy of the ML model and guaranteed to be at least as efficient as estimation using only labeled data.
+[2023] introduced the PoSt-Prediction Adaptive inference (PSPA) estimator that is both robust to the accuracy of the ML model and guaranteed to be at least as efficient as estimation using only labeled data.
 Unlike the PPI++ estimator, the PSPA estimator is designed to achieve element-wise variance reduction and was shown to outperform the PPI++ approach.
 The PSPA approach is based on augmenting estimating equations with information from $\widehat{Y}$, while the CC method provides a simple and natural extension to the PPI method by adding a weight to the linear augmentation term.
-Code availability
+SUPPLEMENTARY MATERIAL
+Title:
+The supplementary materials include proofs for theorems and additional examples for Corollary 1.1 (appendix.pdf).
+Code availability:
 Our analyses can be replicated with the code at: https://github.com/SelinaS37/PBInference.
-Acknowledgments
-The authors gratefully acknowledge funding from an NSERC Discovery Grant (RGPIN-2021-03734), the Connaught Fund, the University of Toronto Data Science Institute, and the McLaughlin Center for Genetic Research.
+UK Biobank Dataset:
+This work used data from the UK Biobank study (https://www.ukbiobank.ac.uk) and our access was approved under application 64875.
 Appendix A Asymptotic properties of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$
@@
 Notation.
-We denote the outer product of a vector $\boldsymbol{v}$ as $\boldsymbol{v}^{\otimes 2}=\boldsymbol{v}\boldsymbol{v}^{{\sf\scriptscriptstyle{T}}}$.
+We denote the outer product of a vector $\bm{v}$ as $\bm{v}^{\otimes 2}=\bm{v}\bm{v}^{{\sf\scriptscriptstyle{T}}}$.
 We denote that a matrix, ${\bf{M}}$, is positive semi-definite as ${\bf{M}}\succeq{\bf{0}},$ and positive definite as ${\bf{M}}\succ{\bf{0}}.$ For any estimator $\hat{\bm{\beta}}$, we let $\text{aVar}(\hat{\bm{\beta}})$ represent its asymptotic variance-covariance matrix.
@@
 A.2 Proof of Theorem 1
-Theorem 1 (Influence function expansion of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$)
+Theorem 1 (Influence function expansion of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$).
 Let $R_{i}$ be the indicator of whether the $i$th observation is labeled. Then,
@@
 $\mathbb{E}[{\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})\hskip 2.84526pt{\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})^{{\sf\scriptscriptstyle{T}}}].$
-Proof  First, note that $n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}-\bm{\beta}^{*}\right)$ may be decomposed as
+Proof.
+First, note that $n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}-\bm{\beta}^{*}\right)$ may be decomposed as
 $n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}-\bm{\beta}^{*}\right)=n^{\frac{1}{2}}\left[\left(\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\bm{\beta}^{*}\right)-\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\bm{\gamma}^{*}\right)+\left(\hat{\bm{\gamma}}^{\text{\tiny{all}}}-\bm{\gamma}^{*}\right)\right].$
@@
 A.3 Proof of Corollary 1.1
-Corollary 1.1 (Comparison of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ with $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$)
+Corollary 1.1 (Comparison of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ with $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$).
 The difference in the asymptotic variance-covariance of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is
-$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})=\left(\frac{1}{\pi}-1\right){\bf{A}}\left(2\boldsymbol{\Sigma}_{Y,\widehat{Y}}-\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}\right){\bf{A}}.$
-$\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is therefore more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if $2\boldsymbol{\Sigma}_{Y,\widehat{Y}}\succ\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}$.
+$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})=\left(\frac{1}{\pi}-1\right){\bf{A}}\left(2\bm{\Sigma}_{Y,\widehat{Y}}-\bm{\Sigma}_{\widehat{Y},\widehat{Y}}\right){\bf{A}}.$
+$\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is therefore more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if $2\bm{\Sigma}_{Y,\widehat{Y}}\succ\bm{\Sigma}_{\widehat{Y},\widehat{Y}}$.
 A sufficient condition for $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ to have lower variance-covariance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ is that
 $\mathbb{E}\left[\left(\epsilon_{\widehat{Y}}-\epsilon_{Y}\right)^{2}\mid{\bf{X}}\right]<\mathbb{E}\left(\epsilon_{Y}^{2}\mid{\bf{X}}\right).$
-Proof  Equation (5) in the proof of Theorem 1 gives the asymptotic expansion of $n^{\frac{1}{2}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\bm{\beta})$.
+Proof.
+Equation (5) in the proof of Theorem 1 gives the asymptotic expansion of $n^{\frac{1}{2}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\bm{\beta})$.
 The standard CLT implies $n^{\frac{1}{2}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\bm{\beta})$ converges weakly to a zero-mean multivariate Gaussian distribution with variance-covariance given by
-$\mathbb{E}\left\{\left[{\bf{A}}\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\right]^{\otimes 2}\frac{R^{2}}{\pi^{2}}\right\}=\frac{1}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{Y,Y}{\bf{A}}.$
+$\mathbb{E}\left\{\left[{\bf{A}}\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\right]^{\otimes 2}\frac{R^{2}}{\pi^{2}}\right\}=\frac{1}{\pi}{\bf{A}}\bm{\Sigma}_{Y,Y}{\bf{A}}.$
 Similarly, by Theorem 1
-$\begin{split}n[\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})]=&\mathbb{E}\left\{\left[{\bf{A}}\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\right]^{\otimes 2}\frac{R^{2}}{\pi^{2}}\right\}+\mathbb{E}\left\{\left[{\bf{A}}\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})\right]^{\otimes 2}\frac{(\pi-R)^{2}}{\pi^{2}}\right\}\\ +&2\mathbb{E}\left\{\left[{\bf{A}}\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\right]\left[{\bf{A}}\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})\right]^{{\sf\scriptscriptstyle{T}}}\frac{R(\pi-R)}{\pi^{2}}\right\}\\ =&\frac{1}{\pi}{\bf{A}}\left[\boldsymbol{\Sigma}_{Y,Y}+(1-\pi)\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}-2(1-\pi)\boldsymbol{\Sigma}_{Y,\widehat{Y}}\right]{\bf{A}}.\end{split}$
+$\begin{split}n[\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})]=&\mathbb{E}\left\{\left[{\bf{A}}\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\right]^{\otimes 2}\frac{R^{2}}{\pi^{2}}\right\}+\mathbb{E}\left\{\left[{\bf{A}}\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})\right]^{\otimes 2}\frac{(\pi-R)^{2}}{\pi^{2}}\right\}\\ +&2\mathbb{E}\left\{\left[{\bf{A}}\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\right]\left[{\bf{A}}\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})\right]^{{\sf\scriptscriptstyle{T}}}\frac{R(\pi-R)}{\pi^{2}}\right\}\\ =&\frac{1}{\pi}{\bf{A}}\left[\bm{\Sigma}_{Y,Y}+(1-\pi)\bm{\Sigma}_{\widehat{Y},\widehat{Y}}-2(1-\pi)\bm{\Sigma}_{Y,\widehat{Y}}\right]{\bf{A}}.\end{split}$
 Thus,
-$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})=\left(\frac{1}{\pi}-1\right){\bf{A}}\left(2\boldsymbol{\Sigma}_{Y,\widehat{Y}}-\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}\right){\bf{A}}.$
-$\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is therefore more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if $2\boldsymbol{\Sigma}_{Y,\widehat{Y}}\succ\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}$.
+$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})=\left(\frac{1}{\pi}-1\right){\bf{A}}\left(2\bm{\Sigma}_{Y,\widehat{Y}}-\bm{\Sigma}_{\widehat{Y},\widehat{Y}}\right){\bf{A}}.$
+$\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is therefore more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if $2\bm{\Sigma}_{Y,\widehat{Y}}\succ\bm{\Sigma}_{\widehat{Y},\widehat{Y}}$.
 Additionally, if $2\mathbb{E}\left(\epsilon_{Y}\epsilon_{\widehat{Y}}\mid{\bf{X}}\right)>\mathbb{E}\left(\epsilon_{\widehat{Y}}^{2}\mid{\bf{X}}\right)$ or equivalently if $\mathbb{E}\left[\left(\epsilon_{\widehat{Y}}-\epsilon_{Y}\right)^{2}\mid{\bf{X}}\right]<\mathbb{E}\left(\epsilon_{Y}^{2}\mid{\bf{X}}\right),$ the diagonal entries of $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})$ are positive and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ will have a lower variance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
+∎
 A.4 Examples for Corollary 1.1
@@
     $\hat{f}({\bf{W}})=\mathbb{E}(Y\mid{\bf{W}})$ and ${\bf{X}}\perp Y\mid{\bf{W}}$
-Proof  For scenario 1, assume that $\hat{f}({\bf{X}})=\mathbb{E}(Y\mid{\bf{X}})$. It follows that
+Proof.
+For scenario 1, assume that $\hat{f}({\bf{X}})=\mathbb{E}(Y\mid{\bf{X}})$. It follows that
 $\bm{\gamma}^{*}=[\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})]^{-1}\mathbb{E}[{\bf{X}}\mathbb{E}(Y\mid{\bf{X}})]=\bm{\beta}^{*}$
@@
 B.1 Asymptotic expansion of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$
-Theorem S1 (Influence function expansion of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$)
+Theorem S1 (Influence function expansion of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$).
 Let $R_{i}$ be the indicator of whether the $i$th observation is labeled. Then,
@@
 $\mathbb{E}\left[{\bf{IF}}^{\text{\tiny{PPI}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})\hskip 2.84526pt{\bf{IF}}^{\text{\tiny{PPI}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})^{{\sf\scriptscriptstyle{T}}}\right].$
-Proof  The only difference in studying the asymptotic properties of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ compared to $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is that we rely on the asymptotic linear expansion of $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ rather than $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$.
+Proof.
+The only difference in studying the asymptotic properties of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ compared to $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is that we rely on the asymptotic linear expansion of $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ rather than $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$.
 Using the same argument as in the proof of Theorem 1,
@@
 B.2 Comparison of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$
-Corollary S1.1 (Comparison of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ with $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$)
+Corollary S1.1 (Comparison of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ with $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$).
 The difference in the asymptotic variance-covariance of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is
-$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}})$ $=\frac{1}{\pi}{\bf{A}}\left(2\boldsymbol{\Sigma}_{Y,\widehat{Y}}-\frac{1}{1-\pi}\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}\right){\bf{A}}.$
-$\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is therefore more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if $2\boldsymbol{\Sigma}_{Y,\widehat{Y}}\succ\frac{1}{1-\pi}\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}$.
+$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}})$ $=\frac{1}{\pi}{\bf{A}}\left(2\bm{\Sigma}_{Y,\widehat{Y}}-\frac{1}{1-\pi}\bm{\Sigma}_{\widehat{Y},\widehat{Y}}\right){\bf{A}}.$
+$\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is therefore more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if $2\bm{\Sigma}_{Y,\widehat{Y}}\succ\frac{1}{1-\pi}\bm{\Sigma}_{\widehat{Y},\widehat{Y}}$.
 A sufficient condition for $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ to have lower variance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ is that
 $\mathbb{E}\left[\left(\epsilon_{\widehat{Y}}-\epsilon_{Y}\right)^{2}\mid{\bf{X}}\right]<\mathbb{E}\left(\epsilon_{Y}^{2}\mid{\bf{X}}\right)+\left(1-\frac{1}{1-\pi}\right)\mathbb{E}\left(\epsilon_{\widehat{Y}}^{2}\mid{\bf{X}}\right).$
-Proof  From the proof for Corollary 1.1, $n[\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{lab}}})]=\frac{1}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{Y,Y}{\bf{A}}.$ By Theorem S1,
+Proof.
+From the proof for Corollary 1.1, $n[\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{lab}}})]=\frac{1}{\pi}{\bf{A}}\bm{\Sigma}_{Y,Y}{\bf{A}}.$ By Theorem S1,
 $n[\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{PPI}}})]=$ $\mathbb{E}\left\{\left[{\bf{A}}\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\right]^{\otimes 2}\frac{R^{2}}{\pi^{2}}\right\}+\mathbb{E}\left\{\left[{\bf{A}}\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})\right]^{\otimes 2}\frac{(\pi-R)^{2}}{\pi^{2}(1-\pi)^{2}}\right\}$
 $+2\mathbb{E}\left\{\left[{\bf{A}}\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\right]\left[{\bf{A}}\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})\right]^{{\sf\scriptscriptstyle{T}}}\frac{R(\pi-R)}{\pi^{2}(1-\pi)}\right\}$
-$=\frac{1}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{Y,Y}{\bf{A}}+\frac{1}{\pi(1-\pi)}{\bf{A}}\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}{\bf{A}}-\frac{2}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}{\bf{A}}$
-$=\frac{1}{\pi}{\bf{A}}\left[\boldsymbol{\Sigma}_{Y,Y}+\frac{1}{1-\pi}\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}-2\boldsymbol{\Sigma}_{Y,\widehat{Y}}\right]{\bf{A}}.$
+$=\frac{1}{\pi}{\bf{A}}\bm{\Sigma}_{Y,Y}{\bf{A}}+\frac{1}{\pi(1-\pi)}{\bf{A}}\bm{\Sigma}_{\widehat{Y},\widehat{Y}}{\bf{A}}-\frac{2}{\pi}{\bf{A}}\bm{\Sigma}_{Y,\widehat{Y}}{\bf{A}}$
+$=\frac{1}{\pi}{\bf{A}}\left[\bm{\Sigma}_{Y,Y}+\frac{1}{1-\pi}\bm{\Sigma}_{\widehat{Y},\widehat{Y}}-2\bm{\Sigma}_{Y,\widehat{Y}}\right]{\bf{A}}.$
 Thus,
-$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}})=\frac{1}{\pi}{\bf{A}}\left[2\boldsymbol{\Sigma}_{Y,\widehat{Y}}-\frac{1}{1-\pi}\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}\right]{\bf{A}}.$
-$\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is therefore more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if $2\boldsymbol{\Sigma}_{Y,\widehat{Y}}\succ\frac{1}{1-\pi}\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}$.
+$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}})=\frac{1}{\pi}{\bf{A}}\left[2\bm{\Sigma}_{Y,\widehat{Y}}-\frac{1}{1-\pi}\bm{\Sigma}_{\widehat{Y},\widehat{Y}}\right]{\bf{A}}.$
+$\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is therefore more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if $2\bm{\Sigma}_{Y,\widehat{Y}}\succ\frac{1}{1-\pi}\bm{\Sigma}_{\widehat{Y},\widehat{Y}}$.
 Note that if $2\mathbb{E}\left(\epsilon_{Y}\epsilon_{\widehat{Y}}\mid{\bf{X}}\right)>\frac{1}{1-\pi}\mathbb{E}\left(\epsilon_{\widehat{Y}}^{2}\mid{\bf{X}}\right)$ or equivalently if $\mathbb{E}\left[\left(\epsilon_{\widehat{Y}}-\epsilon_{Y}\right)^{2}\mid{\bf{X}}\right]<\mathbb{E}\left(\epsilon_{Y}^{2}\mid{\bf{X}}\right)+\left(1-\frac{1}{1-\pi}\right)\mathbb{E}\left(\epsilon_{\widehat{Y}}^{2}\mid{\bf{X}}\right),$ the diagonal entries of $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}})$ are positive and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ will have a lower variance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
+∎
 B.3 Comparison of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$
@@
 From Corollary 1.1 and Corollary S1.1, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if
-$2\boldsymbol{\Sigma}_{Y,\widehat{Y}}\succ\frac{1}{1-\pi}\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}$
+$2\bm{\Sigma}_{Y,\widehat{Y}}\succ\frac{1}{1-\pi}\bm{\Sigma}_{\widehat{Y},\widehat{Y}}$
 whereas $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if
-$2\boldsymbol{\Sigma}_{Y,\widehat{Y}}\succ\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}.$
+$2\bm{\Sigma}_{Y,\widehat{Y}}\succ\bm{\Sigma}_{\widehat{Y},\widehat{Y}}.$
 Therefore, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
@@
 $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{PPI}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})=\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})-\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}})$
-$=\left(\frac{1}{\pi}-1\right){\bf{A}}\left(2\boldsymbol{\Sigma}_{Y,\widehat{Y}}-\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}\right){\bf{A}}-\frac{1}{\pi}{\bf{A}}\left(2\boldsymbol{\Sigma}_{Y,\widehat{Y}}-\frac{1}{1-\pi}\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}\right){\bf{A}}$
-$={\bf{A}}\left(\frac{2-\pi}{1-\pi}\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}-2\boldsymbol{\Sigma}_{Y,\widehat{Y}}\right){\bf{A}}.$
-Note that if $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is less efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ then $\frac{1}{1-\pi}\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}\succ 2\boldsymbol{\Sigma}_{Y,\widehat{Y}}$, implying
-$\frac{2-\pi}{1-\pi}\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}\succ 2\boldsymbol{\Sigma}_{Y,\widehat{Y}}.$
+$=\left(\frac{1}{\pi}-1\right){\bf{A}}\left(2\bm{\Sigma}_{Y,\widehat{Y}}-\bm{\Sigma}_{\widehat{Y},\widehat{Y}}\right){\bf{A}}-\frac{1}{\pi}{\bf{A}}\left(2\bm{\Sigma}_{Y,\widehat{Y}}-\frac{1}{1-\pi}\bm{\Sigma}_{\widehat{Y},\widehat{Y}}\right){\bf{A}}$
+$={\bf{A}}\left(\frac{2-\pi}{1-\pi}\bm{\Sigma}_{\widehat{Y},\widehat{Y}}-2\bm{\Sigma}_{Y,\widehat{Y}}\right){\bf{A}}.$
+Note that if $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is less efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ then $\frac{1}{1-\pi}\bm{\Sigma}_{\widehat{Y},\widehat{Y}}\succ 2\bm{\Sigma}_{Y,\widehat{Y}}$, implying
+$\frac{2-\pi}{1-\pi}\bm{\Sigma}_{\widehat{Y},\widehat{Y}}\succ 2\bm{\Sigma}_{Y,\widehat{Y}}.$
 This confirms that when $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is less efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, it will also be less efficient than $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$.
@@
 C.1 Proof of Theorem 2
-Theorem 2 (Influence function expansion of $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$)
-Following Chen and Chen (2000),
+Theorem 2 (Influence function expansion of $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$).
+Following Chen and Chen [2000],
 $n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{CC}}}-\bm{\beta}^{*}\right)=n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{IF}}^{\text{\tiny{CC}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})+o_{p}(1)$
@@
 $\hskip 2.84526pt\mathbb{E}\left[{\bf{IF}}^{\text{\tiny{CC}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})\hskip 2.84526pt{\bf{IF}}^{\text{\tiny{CC}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})^{{\sf\scriptscriptstyle{T}}}\right]\hskip 2.84526pt.$
-Proof  The proof follows from Chen and Chen (2000), but is detailed for completeness. Recall that
-$\begin{split}\widehat{{\bf{W}}}^{\text{\tiny{CC}}}&=\widehat{{\bf{A}}}\left\{\widehat{\text{Cov}}\left[\phi(Y,{\bf{X}};\bm{\beta}^{*}),\phi(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})\right]\right\}\{\widehat{\text{Var}}[\phi(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})]\}^{-1}\widehat{{\bf{A}}}^{-1}\\ &=\widehat{{\bf{A}}}\widehat{\boldsymbol{\Sigma}}_{Y,\widehat{Y}}\widehat{\boldsymbol{\Sigma}}^{-1}_{\widehat{Y},\widehat{Y}}\widehat{{\bf{A}}}^{-1}\end{split}$
+Proof.
+The proof follows from Chen and Chen [2000], but is detailed for completeness. Recall that
+$\begin{split}\widehat{{\bf{W}}}^{\text{\tiny{CC}}}&=\widehat{{\bf{A}}}\left\{\widehat{\text{Cov}}\left[\phi(Y,{\bf{X}};\bm{\beta}^{*}),\phi(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})\right]\right\}\{\widehat{\text{Var}}[\phi(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})]\}^{-1}\widehat{{\bf{A}}}^{-1}\\ &=\widehat{{\bf{A}}}\widehat{\bm{\Sigma}}_{Y,\widehat{Y}}\widehat{\bm{\Sigma}}^{-1}_{\widehat{Y},\widehat{Y}}\widehat{{\bf{A}}}^{-1}\end{split}$
 where
-$\begin{split}&\widehat{{\bf{A}}}=n_{\text{\tiny{lab}}}\left(\sum_{i=1}^{n_{\text{\tiny{lab}}}}{\bf{X}}_{i}{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\right)^{-1},\\ &\widehat{\boldsymbol{\Sigma}}_{Y,\widehat{Y}}=n_{\text{\tiny{lab}}}^{-1}\sum_{i=1}^{n_{\text{\tiny{lab}}}}\left(Y_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\hat{\bm{\beta}}^{\text{\tiny{lab}}}\right)\left(\hat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\hat{\bm{\gamma}}^{\text{\tiny{lab}}}\right){\bf{X}}_{i}{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}},\\ &\mbox{and }\widehat{\boldsymbol{\Sigma}}_{\widehat{Y},\widehat{Y}}=n_{\text{\tiny{lab}}}^{-1}\sum_{i=1}^{n_{\text{\tiny{lab}}}}\left(\hat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\hat{\bm{\gamma}}^{\text{\tiny{lab}}}\right)^{2}{\bf{X}}_{i}{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}.\end{split}$
+$\begin{split}&\widehat{{\bf{A}}}=n_{\text{\tiny{lab}}}\left(\sum_{i=1}^{n_{\text{\tiny{lab}}}}{\bf{X}}_{i}{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\right)^{-1},\\ &\widehat{\bm{\Sigma}}_{Y,\widehat{Y}}=n_{\text{\tiny{lab}}}^{-1}\sum_{i=1}^{n_{\text{\tiny{lab}}}}\left(Y_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\hat{\bm{\beta}}^{\text{\tiny{lab}}}\right)\left(\hat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\hat{\bm{\gamma}}^{\text{\tiny{lab}}}\right){\bf{X}}_{i}{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}},\\ &\mbox{and }\widehat{\bm{\Sigma}}_{\widehat{Y},\widehat{Y}}=n_{\text{\tiny{lab}}}^{-1}\sum_{i=1}^{n_{\text{\tiny{lab}}}}\left(\hat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\hat{\bm{\gamma}}^{\text{\tiny{lab}}}\right)^{2}{\bf{X}}_{i}{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}.\end{split}$
 It follows that
-$n^{\frac{1}{2}}\left[\widehat{{\bf{W}}}^{\text{\tiny{CC}}}(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\bm{\gamma}^{*})\right]={\bf{A}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}\boldsymbol{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\left\{n^{-\frac{1}{2}}\sum_{i=1}^{n}\left[{\bf{X}}_{i}\left(\widehat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma}^{*}\right)\right]\frac{R_{i}}{\pi}\right\}+o_{p}(1)$
+$n^{\frac{1}{2}}\left[\widehat{{\bf{W}}}^{\text{\tiny{CC}}}(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\bm{\gamma}^{*})\right]={\bf{A}}\bm{\Sigma}_{Y,\widehat{Y}}\bm{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\left\{n^{-\frac{1}{2}}\sum_{i=1}^{n}\left[{\bf{X}}_{i}\left(\widehat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma}^{*}\right)\right]\frac{R_{i}}{\pi}\right\}+o_{p}(1)$
 and
-$n^{\frac{1}{2}}\left[\widehat{{\bf{W}}}^{\text{\tiny{CC}}}(\hat{\bm{\gamma}}^{\text{\tiny{all}}}-\bm{\gamma}^{*})\right]={\bf{A}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}\boldsymbol{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\left\{n^{-\frac{1}{2}}\sum_{i=1}^{n}\left[{\bf{X}}_{i}\left(\widehat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma}^{*}\right)\right]\right\}+o_{p}(1).$
+$n^{\frac{1}{2}}\left[\widehat{{\bf{W}}}^{\text{\tiny{CC}}}(\hat{\bm{\gamma}}^{\text{\tiny{all}}}-\bm{\gamma}^{*})\right]={\bf{A}}\bm{\Sigma}_{Y,\widehat{Y}}\bm{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\left\{n^{-\frac{1}{2}}\sum_{i=1}^{n}\left[{\bf{X}}_{i}\left(\widehat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma}^{*}\right)\right]\right\}+o_{p}(1).$
 Consequently,
@@
 C.2 Proof of Corollary 2.1
-Corollary 2.1 ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$)
+Corollary 2.1 ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$).
 The difference in the asymptotic covariance of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is
-$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{CC}}})=\left(\frac{1}{\pi}-1\right){\bf{A}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}\boldsymbol{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}{\bf{A}}\succeq{\bf{0}}.$
+$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{CC}}})=\left(\frac{1}{\pi}-1\right){\bf{A}}\bm{\Sigma}_{Y,\widehat{Y}}\bm{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\bm{\Sigma}_{Y,\widehat{Y}}{\bf{A}}\succeq{\bf{0}}.$
 Therefore, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is guaranteed to be at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
-Proof  From the proof for Corollary 1.1, $n[\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{lab}}})]=\frac{1}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{Y,Y}{\bf{A}}.$ By Theorem 2.1, $n[\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{CC}}})]$ is given by
-$\begin{split}&\mathbb{E}\left\{\left[{\bf{A}}\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\right]^{\otimes 2}\frac{R^{2}}{\pi^{2}}\right\}+\mathbb{E}\left\{\left[{\bf{A}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}\boldsymbol{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})\right]^{\otimes 2}\frac{(\pi-R)^{2}}{\pi^{2}}\right\}\\ +&2\mathbb{E}\left\{\left[{\bf{A}}\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\right]\left[{\bf{A}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}\boldsymbol{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})\right]^{{\sf\scriptscriptstyle{T}}}\frac{R(\pi-R)}{\pi^{2}}\right\}\\ =&\frac{1}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{Y,Y}{\bf{A}}+\frac{1-\pi}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}\boldsymbol{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}{\bf{A}}-2\frac{1-\pi}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}\boldsymbol{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}{\bf{A}}\\ =&\frac{1}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{Y,Y}{\bf{A}}-\frac{1-\pi}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}\boldsymbol{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}{\bf{A}}.\end{split}$
-Hence, $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{CC}}})=\frac{1-\pi}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}\boldsymbol{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}{\bf{A}}\succeq 0$.
+Proof.
+From the proof for Corollary 1.1, $n[\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{lab}}})]=\frac{1}{\pi}{\bf{A}}\bm{\Sigma}_{Y,Y}{\bf{A}}.$ By Theorem 2.1, $n[\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{CC}}})]$ is given by
+$\begin{split}&\mathbb{E}\left\{\left[{\bf{A}}\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\right]^{\otimes 2}\frac{R^{2}}{\pi^{2}}\right\}+\mathbb{E}\left\{\left[{\bf{A}}\bm{\Sigma}_{Y,\widehat{Y}}\bm{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})\right]^{\otimes 2}\frac{(\pi-R)^{2}}{\pi^{2}}\right\}\\ +&2\mathbb{E}\left\{\left[{\bf{A}}\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\right]\left[{\bf{A}}\bm{\Sigma}_{Y,\widehat{Y}}\bm{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})\right]^{{\sf\scriptscriptstyle{T}}}\frac{R(\pi-R)}{\pi^{2}}\right\}\\ =&\frac{1}{\pi}{\bf{A}}\bm{\Sigma}_{Y,Y}{\bf{A}}+\frac{1-\pi}{\pi}{\bf{A}}\bm{\Sigma}_{Y,\widehat{Y}}\bm{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\bm{\Sigma}_{Y,\widehat{Y}}{\bf{A}}-2\frac{1-\pi}{\pi}{\bf{A}}\bm{\Sigma}_{Y,\widehat{Y}}\bm{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\bm{\Sigma}_{Y,\widehat{Y}}{\bf{A}}\\ =&\frac{1}{\pi}{\bf{A}}\bm{\Sigma}_{Y,Y}{\bf{A}}-\frac{1-\pi}{\pi}{\bf{A}}\bm{\Sigma}_{Y,\widehat{Y}}\bm{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\bm{\Sigma}_{Y,\widehat{Y}}{\bf{A}}.\end{split}$
+Hence, $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{CC}}})=\frac{1-\pi}{\pi}{\bf{A}}\bm{\Sigma}_{Y,\widehat{Y}}\bm{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\bm{\Sigma}_{Y,\widehat{Y}}{\bf{A}}\succeq 0$.
+∎
 C.3 Proof of Corollary 2.2
-Corollary 2.2 ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is the most efficient estimator in its class)
+Corollary 2.2 ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is the most efficient estimator in its class).
 Let the mean squared error (MSE) of $\hat{\bm{\beta}}$ for estimating $\bm{\beta}^{*}$ be defined as $\mathbb{E}[(\hat{\bm{\beta}}-\bm{\beta}^{*})^{{\sf\scriptscriptstyle{T}}}(\hat{\bm{\beta}}-\bm{\beta}^{*})]$.
-Analogously to the well-known result for the best linear predictor (e.g., Bickel and Doksum (2015), page 40), ${\bf{W}}^{\text{\tiny{CC}}}$ is the weight, ${\bf{W}}$, that minimizes the asymptotic MSE of estimators of the form $\hat{\bm{\beta}}^{\text{\tiny{lab}}}-{{\bf{W}}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right)$ for estimating $\bm{\beta}^{*}$.
-Proof  Letting $\hat{\bm{\beta}}_{{\bf{W}}}=\hat{\bm{\beta}}^{\text{\tiny{lab}}}-{\bf{W}}(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}})$, the MSE of estimating $\hat{\bm{\beta}}_{{\bf{W}}}$ is
+Analogously to the well-known result for the best linear predictor (e.g., Bickel and Doksum [2015], page 40), ${\bf{W}}^{\text{\tiny{CC}}}$ is the weight, ${\bf{W}}$, that minimizes the asymptotic MSE of estimators of the form $\hat{\bm{\beta}}^{\text{\tiny{lab}}}-{{\bf{W}}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right)$ for estimating $\bm{\beta}^{*}$.
+Proof.
+Letting $\hat{\bm{\beta}}_{{\bf{W}}}=\hat{\bm{\beta}}^{\text{\tiny{lab}}}-{\bf{W}}(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}})$, the MSE of estimating $\hat{\bm{\beta}}_{{\bf{W}}}$ is
 $\mathbb{Q}({\bf{W}})=\mathbb{E}\left\{\left[\hat{\bm{\beta}}^{\text{\tiny{lab}}}-{\bf{W}}(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}})-\bm{\beta}^{*}\right]^{\otimes 2}\right\}.$
@@
 $n^{\frac{1}{2}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right)={\bf{A}}\left[n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{X}}_{i}(\widehat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma}^{*})\frac{R_{i}-\pi}{\pi}\right]+o_{p}(1),$
-we have $\text{Cov}(\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\bm{\beta}^{*},\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}})=\frac{1-\pi}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}{\bf{A}}$ and $\text{Var}(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}})=\frac{1-\pi}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}{\bf{A}}$.
+we have $\text{Cov}(\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\bm{\beta}^{*},\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}})=\frac{1-\pi}{\pi}{\bf{A}}\bm{\Sigma}_{Y,\widehat{Y}}{\bf{A}}$ and $\text{Var}(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}})=\frac{1-\pi}{\pi}{\bf{A}}\bm{\Sigma}_{\widehat{Y},\widehat{Y}}{\bf{A}}$.
 Thus,
-${\bf{W}}^{*}=\frac{1-\pi}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}{\bf{A}}\left[\frac{1-\pi}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}{\bf{A}}\right]^{-1}={\bf{A}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}\boldsymbol{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}{\bf{A}}^{-1}={\bf{W}}^{CC}.$
+${\bf{W}}^{*}=\frac{1-\pi}{\pi}{\bf{A}}\bm{\Sigma}_{Y,\widehat{Y}}{\bf{A}}\left[\frac{1-\pi}{\pi}{\bf{A}}\bm{\Sigma}_{\widehat{Y},\widehat{Y}}{\bf{A}}\right]^{-1}={\bf{A}}\bm{\Sigma}_{Y,\widehat{Y}}\bm{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}{\bf{A}}^{-1}={\bf{W}}^{CC}.$

### v4

@@
 Abstract
-From structural biology to epidemiology, predictions from machine learning (ML) models increasingly complement costly gold-standard data to enable faster, more affordable, and scalable scientific inquiry.
+From structural biology to epidemiology, predictions from machine learning (ML) models are increasingly used to complement costly gold-standard data to enable faster, more affordable, and scalable scientific inquiry.
 In response, prediction-based (PB) inference has emerged to accommodate statistical analysis using a large volume of predictions together with a small amount of gold-standard data.
 The goals of PB inference are two-fold: (i) to mitigate bias from errors in predictions and (ii) to improve efficiency relative to traditional inference using only the gold-standard data.
-Motwani and Witten [2023] recently revisited two key PB inference approaches and found that only one method, Prediction-powered Inference (PPI) proposed by Angelopoulos et al.
-[2023a], achieves (i).
-In this paper, we find that PPI does not achieve (ii).
-We revisit the double sampling literature and show that, with a simple modification, PPI can be adjusted to provide theoretically justified improvements in efficiency.
-We also contextualize PB inference with economics and statistics literature dating back to the 1960s to highlight the utility of classical methods in this contemporary problem.
-Our extensive theoretical analyses, along with an analysis of UK Biobank data, indicate that our proposal effectively mitigates bias and improves efficiency, making it preferable for use in practice.
+While early PB inference methods focused on bias, their ability to enhance efficiency remains unclear.
+We revisit a popular PB inference method and show that a simple modification can be applied to guarantee improvements in efficiency beyond yielding valid inferences when the ML predictions are imperfect.
+The utility of this approach in leveraging prediction-based outcomes to enhance efficiency is demonstrated through extensive simulation studies and an application to the UK Biobank data.
+We further contextualize the problem of PB inference through historical literature from economics and statistics to highlight perspectives from classical methods in this contemporary problem.
 Keywords: Machine learning, measurement error, missing data, prediction-based inference, semi-parametric inference
 1 Introduction
-In the rapidly evolving landscape of data science, accurately predicting an outcome is often just an initial step in scientific inquiry.
-Researchers are increasingly turning to machine learning (ML) models to obtain predictions for outcomes that are prohibitively time-consuming or expensive to directly measure, and then drawing statistical inferences from the ML model’s predictions [Hoffman et al., 2024, Fan et al., 2024, Gao et al., 2024, Miao and Lu, 2024].
-This prediction-based (PB) inference scenario was introduced by Wang et al in the context of global health and genomics applications, and is comprised of two steps outlined in Figure 1 [Wang et al., 2020].
+Researchers are increasingly turning to machine learning (ML) models to obtain predictions for outcomes that are prohibitively time-consuming or expensive to measure directly and conducting statistical analyses treating the predictions as observed data.
+1, 2, 3, 4 This prediction-based (PB) inference approach, introduced by Wang et al,5 is comprised of two steps outlined in Figure 1.
 First, the outcome of interest is imputed using a prediction from a pre-trained model.
-Then, the predictions are used to draw statistical inference on the parameter of interest, such as the association between an outcome and a set of covariates, together with a small amount of gold-standard data  [^1] PB inference has focused on outcomes that are difficult to measure.
-Recent work has considered predicted covariates (e.g., Kluger et al.
-[2025])..
+Then, the predictions are used to draw inferences about the parameters of interest, such as the association parameters between the outcome and a set of covariates, together with a small amount of gold-standard data.***PB inference has primarily focused on outcomes that are difficult to measure.
+Recent work has considered predicted covariates (e.g., Kluger et al6).
 In contrast to traditional inference that relies solely on gold-standard data, PB inference aims to leverage the predictions to improve statistical efficiency without introducing bias.
+In the last several years, applications of PB inference have arisen in diverse fields, spanning genetics, biology, and medicine.
+For example, population biobanks, such as the UK Biobank ($n\approx$ 500K individuals), collect high-throughput genomic data on large populations and offer unprecedented opportunities for genetic discovery research.
+However, many traits and disease outcomes are only partially observed due to the time and cost of measurement.
+ML models are commonly used to impute missing outcomes based on fully observed baseline data to enable downstream association studies.
+7, 8, 9, 10, 11, 12, 13, 14 Similarly, in structural biology, AlphaFold provides structural predictions without the need for expensive and labor-intensive experiments for proteomic studies.
+15, 16 In the analysis of electronic health records data, large language models can rapidly ascertain disease status at an accuracy competitive with manual chart review for epidemiological and clinical studies.
+17, 18, 19
 Figure 1: Prediction-based (PB) inference.
@@
 The focus of scientific inquiry is on the association between $Y$ and ${\bf{X}}$.
-Traditional statistical inference relies solely on the data on $Y$ and ${\bf{X}}$ while PB inference incorporates $\widehat{Y}$ to maximize the use of all available data.
+Traditional statistical inference relies solely on the data on $Y$ and ${\bf{X}}$ while PB inference incorporates $\widehat{Y}$ to make use of all of the available data.
 Dark red represents labeled observations while light red represents unlabeled observations.
-In the last several years, applications of PB inference have arisen in diverse fields, spanning genetics, biology, and medicine.
-For example, population biobanks, such as the UK Biobank ($N\approx$ 500K individuals), provide an unprecedented data source for high-throughput genetic research.
-However, many traits and disease outcomes are only partially observed due to the time and cost of measurement.
-ML models are commonly used to predict missing outcomes based on fully observed baseline data and the predictions are utilized for genetic discovery [Tong et al., 2020, Alipanahi et al., 2021, Cosentino et al., 2023, Dahl et al., 2023, An et al., 2023, McCaw et al., 2024, Liu et al., 2024, Miao et al., 2024].
-Similarly, in the field of structural biology, AlphaFold provides structural predictions without the need for expensive and labor-intensive experiments for proteomic studies [Jumper et al., 2021, Gan et al., 2022].
-In the analysis of electronic health records data, large language models can rapidly ascertain disease status at an accuracy competitive with manual chart review by experts, enabling downstream epidemiological and clinical studies [Yang et al., 2022, Zhou et al., 2022, Alsentzer et al., 2023].
-Given the relevance of PB inference, Motwani and Witten [2023] recently revisited two innovative methods in this field.
-In the context of least squares estimation, they showed that the approach proposed by Wang et al.
-[2020] provides a consistent estimator for the association between an outcome and a set of covariates only in restrictive scenarios.
-For example, when the pre-trained model perfectly captures the regression function.
-This is a significant practical limitation as it is essential that PB inference methods yield valid inference in the likely scenario that the prediction model is imperfect.
-In contrast, the prediction powered inference (PPI) approach proposed by Angelopoulos et al.
-[2023a] possesses this necessary property and yields valid inference regardless of the quality of the model.
-This finding is critical given that the PPI paper has been cited over 100 times in its first year of publication in Science.
-In this paper, we study the statistical efficiency of the PPI estimator and show that it can be outperformed by traditional inference.
-We propose a more efficient estimator that can be obtained by simply adding a weight to the PPI estimator, an approach introduced by Chen & Chen (CC) 25 years ago in the context of double sampling designs [Chen and Chen, 2000].
-We begin by formally introducing the PB inference problem and reviewing the contributions of Motwani and Witten in Section 2.
-We then study the asymptotic properties of the PPI estimator and introduce the CC estimator in Section 3.
+Given the relevance of PB inference, Motwani and Witten 20 recently revisited two seminal methods in the field.
+In the context of least squares estimation, they showed that the post-prediction inference (PostPi) method proposed by Wang et al5 provides a consistent estimator for the association between an outcome and a set of covariates only in restrictive scenarios (e.g., when the pre-trained ML model perfectly captures the underlying regression function).
+This is a significant practical limitation as it is essential that PB inference methods yield valid inference (i.e., hypothesis tests with controlled type I error and confidence intervals with nominal coverage rates) in the likely scenario that the prediction model is imperfect.
+In contrast, the prediction powered inference (PPI) method proposed by Angelopoulos et al 21 yields consistent estimation and valid inference regardless of the quality of the ML model.
+This finding is critical given that the PPI paper, published just over a year ago in the high-impact journal Science, has already been cited nearly 150 times and widely adopted in applications.
+In this paper, we build on the work of Motwani and Witten 20 and further examine the PPI estimator by assessing its statistical efficiency.
+Our contributions are two-fold.
+First, we develop a set of results that characterize when the PPI estimator gains over traditional estimators using labeled data only.
+We also show that a more efficient estimator can be obtained by adding a simple weight to the PPI estimator.
+This approach was proposed by Chen & Chen (CC) over 25 years ago in the context of double sampling designs.22 Second, we revisit literature from statistics and economics and identify work dating back to the 1960s that bears directly on current developments in PB inference.
+Our review puts classical statistical methods in conversation with this modern problem to inform future methodological advances and applications.
+We begin by formally introducing the PB inference problem and reviewing the contributions of Motwani and Witten20 in Section 2.
+We then study the asymptotic properties of the PPI estimator, showing that it can be outperformed by traditional inference methods, and introduce the CC estimator in Section 3.
 Our key finding is that the CC estimator is optimal with respect to statistical efficiency in a class of estimators to which the PPI estimator belongs.
-We also review modern and classical literature from statistics and economics dating back to the 1960s that provides important context for the PB inference problem in Section 4.
-In Section 5, we evaluate our findings with extensive simulation studies.
-In Section 6, we study the association between android fat mass and clinical and demographic covariates in the South Asian population in the UK Biobank to demonstrate the practical performance of our method.
+In Section 4, we discuss related literature and provide connections between the CC estimator and other established methods.
+We then evaluate our findings with extensive simulation studies in Section 5.
+In Section 6, we investigate the association between android fat mass and clinical and demographic covariates in the South Asian population in the UK Biobank to demonstrate the practical performance of these approaches.
 We close with additional discussion in Section 7.
-All in all, our theoretical and numerical results show that the CC estimator offers a balance between robustness to ML model specification and statistical efficiency, making it the preferred choice for use in PB inference problems.
+Overall, our theoretical and numerical results show that the CC estimator offers a balance between robustness to ML model specification and statistical efficiency, making it a preferred choice for use in practice.
 2 Problem setting & background
@@
 We let $Y$ denote the outcome of interest and ${\bf{Z}}$ a set of predictors.
-$Y$ is only partially observed and a pre-trained ML model is available to predict it on the basis of ${\bf{Z}}$.
-The ML model, $\hat{f}:\mathcal{Z}\rightarrow\mathcal{Y}$, maps from the predictor space, $\mathcal{Z}$, to the outcome space, $\mathcal{Y}$.
-The predictions of $Y$ are denoted as $\widehat{Y}=\hat{f}({\bf{Z}})$.
-We let ${\bf{X}}=(1,X_{1},\dots,X_{p})^{{\sf\scriptscriptstyle{T}}}$ be a set of covariates that are potentially associated with $Y$ and that are the focus of scientific inquiry.
+$Y$ is only partially observed and a pre-trained ML model is available to predict it over the full data given ${\bf{Z}}$, which are fully observed.
+The ML model $\hat{f}:\mathcal{Z}\rightarrow\mathcal{Y}$ maps the predictor space $\mathcal{Z}$ to the outcome space $\mathcal{Y}$ and produces predictions $\widehat{Y}=\hat{f}({\bf{Z}})$.
+We let ${\bf{X}}=(1,X_{1},\dots,X_{p})^{{\sf\scriptscriptstyle{T}}}$ denote fully observed covariates that are potentially associated with $Y$ and that are the focus of scientific inquiry.
 ${\bf{X}}$ may also be predictive of $Y$ and these covariates may or may not be included in ${\bf{Z}}$.
-We assume that ${\bf{X}}$ and ${\bf{Z}}$ are both fully observed.
 The available data therefore consists of a labeled dataset and an independent unlabeled dataset, respectively denoted as
@@
 where $n=n_{\text{\tiny{lab}}}+n_{\text{\tiny{unlab}}}$, $n_{\text{\tiny{unlab}}}$ is the number of unlabeled observations, and $n_{\text{\tiny{lab}}}/n\to\pi$ as $n\to\infty$ for $\pi\in(0,1)$.
-We focus on the case in which $\mathscr{L}$ and $\mathscr{U}$ are independent and identically distributed samples from the population of interest.
-More complex scenarios have been the focus of recent work [Angelopoulos et al., 2023a, Zrnic and Candès, 2024a].
+We focus on the case in which $\mathscr{L}$ and $\mathscr{U}$ include independent and identically distributed samples from a common population of interest.
 Additionally, $\hat{f}$ is assumed to be pre-trained with a dataset that is independent of $\mathscr{L}$ and $\mathscr{U}$ and is therefore treated as “fixed” (i.e., the variation in estimating $\hat{f}$ is not of interest in downstream inference).
-In fact, $\hat{f}$ may be a black box function in that its training data and specification are not be available to the users.
-Approaches that address uncertainty in the estimation of $\hat{f}$ have recently been proposed by Zrnic and Candès [2024b].
+$\hat{f}$ may be a black box function in that its training data and specification are not be available to the users.
+Approaches that address uncertainty in the estimation of $\hat{f}$ have recently been considered.23
 2.2 Parameter of interest
-Following Motwani and Witten [2023], we consider least squares estimation and aim to make inference on the population parameter
+Following Motwani and Witten,20 we consider least squares estimation and aim to make inferences on the population parameter
 $\bm{\beta}^{*}\coloneq\operatorname*{arg\,min}_{\bm{\beta}}\mathbb{E}[(Y-{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\beta})^{2}]=[\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})]^{-1}\mathbb{E}({\bf{X}}Y),$
 which quantifies the association between $Y$ and ${\bf{X}}$.
-Throughout, we use superscripts to denote different estimators of $\bm{\beta}^{*}$.
-For example, the simplest unbiased estimator of $\bm{\beta}^{*}$, denoted as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, uses only $\mathscr{L}$ for estimation and is the solution to the estimating equation
+The simplest consistent estimator of $\bm{\beta}^{*}$, which we denote by $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, uses only $\mathscr{L}$ for estimation and solves the estimating equation
 $\sum_{i=1}^{n_{\text{\tiny{lab}}}}\bm{\phi}(Y_{i},{\bf{X}}_{i};\bm{\beta})\coloneq\sum_{i=1}^{n_{\text{\tiny{lab}}}}{\bf{X}}_{i}(Y_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\bm{\beta})=\bm{0}.$
 2.3 Motivation for PB inference
-The goal of PB inference is to utilize the abundant realizations of $\widehat{Y}$, which can be computed in the unlabeled data, to provide valid and efficient inference on the association between $Y$ and ${\bf{X}}$.
-Motwani and Witten consider the question of validity and we study efficiency in Section 3.
-With respect to validity, PB inference should enable users to draw correct conclusions regarding the magnitude and direction of $\bm{\beta}^{*}$ and to perform reliable uncertainty quantification.
-Incorporating information from $\widehat{Y}$ should not change the parameter being estimated and the credibility of inferences should not depend on the quality of predictions from $\hat{f}$.
-The need for PB inference is highlighted by limitations of “naive” inference, in which inference on $\widehat{Y}$ is performed in lieu of inference on $Y$ [Wang et al., 2020].
-The parameter targeted by naive inference is
-$\bm{\gamma}^{*}\coloneq\operatorname*{arg\,min}_{\bm{\gamma}}\mathbb{E}[(\widehat{Y}-{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma})^{2}]=[\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})]^{-1}\mathbb{E}({\bf{X}}\widehat{Y}).$
-A least squares estimator for $\bm{\gamma}^{*}$, denoted as $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$, is the solution to
-$\sum_{i=n_{\text{\tiny{lab}}}+1}^{n}\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma})\coloneq\sum_{i=n_{\text{\tiny{lab}}}+1}^{n}{\bf{X}}_{i}(\widehat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma})=\bm{0}.$
-Critically, $\bm{\gamma}^{*}$ bears no necessary correspondence with $\bm{\beta}^{*}$ except under highly restrictive scenarios, such as when ${\bf{Z}}$ contains ${\bf{X}}$ and the ML model captures the true regression function.
-The fundamental problem is that $(\widehat{Y},{\bf{X}}^{{\sf\scriptscriptstyle{T}}})^{{\sf\scriptscriptstyle{T}}}$ has a different distribution than $(Y,{\bf{X}}^{{\sf\scriptscriptstyle{T}}})^{{\sf\scriptscriptstyle{T}}}$.
-To overcome these limitations, PB inference methods utilize $\mathscr{L}$ together with $\mathscr{U}$ to essentially balance the relative contributions of $Y$ and $\widehat{Y}$ in making inference on $\bm{\beta}^{*}$.
+Although $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ is consistent and enables valid inference for $\bm{\beta}^{*}$, it has limited precision when gold-standard data is difficult to collect and $n_{\text{\tiny{lab}}}$ is not large.
+In contrast, the predictions $\widehat{Y}$ can be computed in the full data, leading to a much larger dataset for analysis.
+However, relying on $\widehat{Y}$ for inference may result in biased estimation depending on the quality of predictions from $\hat{f}$.
+For example, in “naive” inference, $\widehat{Y}$ is directly substituted for $Y$ in downstream analysis.5 The parameter targeted by this approach is
+$\bm{\gamma}^{*}\coloneq\operatorname*{arg\,min}_{\bm{\gamma}}\mathbb{E}[(\widehat{Y}-{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma})^{2}]=[\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})]^{-1}\mathbb{E}({\bf{X}}\widehat{Y})$
+and the corresponding least squares estimator, denoted as $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$, is the solution to
+$\sum_{i=1}^{n}\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma})\coloneq\sum_{i=1}^{n}{\bf{X}}_{i}(\widehat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma})=\bm{0}.$
+As the distribution of $(\widehat{Y},{\bf{X}}^{{\sf\scriptscriptstyle{T}}})^{{\sf\scriptscriptstyle{T}}}$ may not coincide with that of $(Y,{\bf{X}}^{{\sf\scriptscriptstyle{T}}})^{{\sf\scriptscriptstyle{T}}}$, $\bm{\gamma}^{*}$ bears no necessary correspondence with $\bm{\beta}^{*}$ except under restrictive scenarios (e.g., when ${\bf{Z}}$ contains ${\bf{X}}$ and the ML model captures the true regression function).
+To overcome these limitations, PB inference methods utilize $\mathscr{L}$ and $\mathscr{U}$ to essentially balance the relative contributions of $Y$ and $\widehat{Y}$ in making inference on $\bm{\beta}^{*}$.
 2.4 PB inference methods
-Motwani and Witten revisited two innovative strategies for PB inference.
-The first is the proposal of Wang et al.
-[2020], which is a two-step procedure to incorporate $\widehat{Y}$ into estimation.
+Motwani and Witten20 revisited two seminal strategies for PB inference.
+The first is the PostPI method of Wang et al,5 which is a two-step procedure to incorporate $\widehat{Y}$ into estimation.
 First, $\mathscr{L}$ is used to model the relationship between $Y$ and $\widehat{Y}$ to address potential inaccuracies in the predictions.
 Then, this model is used together with $\mathscr{U}$ to make inference on $\bm{\beta}^{*}$ with either analytical or bootstrap approaches.
-In their illuminating work, Motwani and Witten show that the estimator of Wang et al is consistent for $\bm{\beta}^{*}$ under the strong and unrealistic assumption that the ML model captures the true regression function.
-In other scenarios, the Wang et al estimator may not be consistent for $\bm{\beta}^{*}$ even while its standard error converges to 0 as $(n_{\text{\tiny{lab}}},n_{\text{\tiny{unlab}}})\to\infty$.
-Consequently, the familiar Wald statistic will not converge in distribution under the null, resulting in type I error that is not controlled at the nominal level and confidence intervals with inadequate coverage.
-Motwani and Witten also studied the PPI method, which has the crucial property that inference remains valid regardless of the quality of the predictions.
-Stated otherwise, the PPI method does not require that $\hat{f}$ capture the true regression function for the estimator to maintain consistency or to yield reliable hypothesis testing and interval estimation.
+This approach is shown to yield a consistent estimator for $\bm{\beta}^{*}$ under the strong and unrealistic assumption that the ML model captures the true regression function.
+When this assumption is not met, the PostPI estimator may not be consistent.
+Consequently, corresponding hypothesis tests may not have controlled type I error rates and confidence intervals may not have adequate coverage.
+Motwani and Witten20 also considered the PPI method,21 which has the crucial property that inference remains valid regardless of the quality of the predictions.
+That is, the PPI method does not require that $\hat{f}$ capture the true regression function for the estimator to maintain consistency for $\bm{\beta}^{*}$.
 The proposed estimator utilizes a common statistical technique involving the augmentation of an initial consistent estimator for the parameter of interest with an additional term that converges in probability to 0.
-This approach dates back to at least Robins et al.
-[1994], and has been applied in multiple settings to improve estimation efficiency outside of the PB inference problem (e.g., Chen and Chen [2000], Tsiatis [2006], Tian et al.
-[2012], Wang and Wang [2015]).
-We return to this point in Section 4.
+This approach dates back to at least Robins et al 24 and has been applied in multiple settings to improve estimation efficiency outside of the PB inference problem.
+22, 25, 26, 27 We return to this point in our review of related literature in Section 4.
 The PPI estimator is
-$\hat{\bm{\beta}}^{\text{\tiny{PPI}}}=\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}\right),$
-where $\hat{\bm{\gamma}}^{\text{\tiny{lab}}}$ is the solution to $\sum_{i=1}^{n_{\text{\tiny{lab}}}}\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma})=\bm{0}$ [^2] Angelopoulos et al.
-[2023a] express $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ as $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}-\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\beta}}^{\text{\tiny{lab}}}\right)$ where $\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\beta}}^{\text{\tiny{lab}}}\right)$ is referred to as a “rectifier.” The rectifier serves the purpose of correcting potential bias in $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ due to inaccuracies in $\widehat{Y}$ for $Y$.
+$\hat{\bm{\beta}}^{\text{\tiny{PPI}}}=\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}\right)$
+where $\hat{\bm{\gamma}}^{\text{\tiny{lab}}}$ solves $\sum_{i=1}^{n_{\text{\tiny{lab}}}}\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma})=\bm{0}$ and $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ solves $\sum_{i=n_{\text{\tiny{lab}}}+1}^{n}\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma})=\bm{0}$Angelopoulos et al21 express $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ as $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}-\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\beta}}^{\text{\tiny{lab}}}\right)$ where $\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\beta}}^{\text{\tiny{lab}}}\right)$ is referred to as a “rectifier.” The rectifier serves the purpose of correcting potential bias in $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ due to inaccuracies in $\widehat{Y}$ for $Y$.
 We reframe the PPI method to study its statistical efficiency..
-In contrast to the approach of Wang et al, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is consistent for $\bm{\beta}^{*}$ and $n^{\frac{1}{2}}(\hat{\bm{\beta}}^{\text{\tiny{PPI}}}-\bm{\beta}^{*})$ converges weakly to a zero-mean multivariate normal distribution regardless of the quality of the prediction.
-The proposed method therefore asymptotically controls the type I error for Wald-based hypothesis testing and provides confidence intervals with the nominal coverage level.
+In contrast to the PostPI method, Motwani and Witten20 show that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is consistent for $\bm{\beta}^{*}$ and $n^{\frac{1}{2}}(\hat{\bm{\beta}}^{\text{\tiny{PPI}}}-\bm{\beta}^{*})$ converges weakly to a zero-mean multivariate normal distribution regardless of the quality of the prediction.
+This approach therefore asymptotically controls the type I error rate in hypothesis tests and produces confidence intervals with the nominal coverage level.
 3 Efficiency analysis
-As PB inference methods suppose the availability of at least a small labeled data set, a natural baseline estimator for comparison is $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, which simply regresses $Y$ on ${\bf{X}}$ within $\mathscr{L}$.
-For a PB inference method to be preferred over this baseline, it should not only yield valid inference, but also improve statistical efficiency relative to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when the prediction is informative of $Y$, and perform no worse than estimation based on $\mathscr{L}$ when $\widehat{Y}$ is uninformative of $Y$.
-For the remainder of this paper, we study the latter property and analyze an estimator very similar to $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ given by
-$\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}=\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right),$
-where $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$ is the solution to $\sum_{i=1}^{n}\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma})=\bm{0}$.
+As PB inference methods suppose the availability of at least a small labeled data set, a natural baseline for comparison is the estimator based on $\mathscr{L}$ only, $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
+For a PB inference method to be preferred over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, it should not only yield valid inference, but also achieve gains in statistical efficiency relative to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when the prediction is informative of $Y$ and perform no worse when $\widehat{Y}$ is uninformative.
+For the remainder of this paper, we study the latter property and analyze a modified version of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ given by
+$\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}=\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right).$
 The only difference between $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is that all of the data, rather than the unlabeled data, is used to estimate $\bm{\gamma}^{*}$ in the augmentation term (i.e., $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ is replaced by $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$).
 When $\pi\approx 0$, $\hat{\bm{\gamma}}^{\text{\tiny{all}}}\approx\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ so that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}\approx\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$.
-In general, it can be shown that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ gains efficiency over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ whenever $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ gains over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ can be more efficient than $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ when $\widehat{Y}$ is informative of $Y$ (see Appendix B, specifically Section B.3, for a detailed analysis).
-Additionally, estimators of this form are commonly used in the semi-parametric inference literature, which we draw upon in the subsequent section in our efficiency analysis.
+In general, it can be shown that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ gains efficiency over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ whenever $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ gains over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ can be more efficient than $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ when $\widehat{Y}$ is informative of $Y$ (see Appendix, specifically Section B.3, for a detailed analysis).
+Additionally, estimators of this form are commonly used in the semi-parametric inference literature, which we draw upon in the subsequent section in our asymptotic analysis.
 3.1 Asymptotic analysis of the PPI estimator
-To study the asymptotic variance of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, Theorem 1 provides the influence function expansion of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$.
-All of the asymptotic results for $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ are detailed in Appendix A.
+To study the asymptotic propeties of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, we first derive its influence function expansion in Theorem 1. The detailed derivations are deferred to the Appendix.
 Theorem 1 (Influence function expansion of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$).
@@
 where ${\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})={\bf{A}}\left\{\left[\ \bm{\phi}(Y_{i},{\bf{X}}_{i};\bm{\beta}^{*})\right]\frac{R_{i}}{\pi}+\left[\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma}^{*})\right]\frac{\pi-R_{i}}{\pi}\right\}$ and ${\bf{A}}=[\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})]^{-1}$.
-Therefore, $n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}-\bm{\beta}^{*}\right)$ converges weakly to a zero-mean multivariate normal distribution with variance-covariance
-$\mathbb{E}\left[{\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})\hskip 2.84526pt{\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})^{{\sf\scriptscriptstyle{T}}}\right].$
+Therefore, $n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}-\bm{\beta}^{*}\right)$ converges weakly to a zero-mean multivariate normal distribution with variance-covariance $\mathbb{E}\left[{\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})\hskip 2.84526pt{\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})^{{\sf\scriptscriptstyle{T}}}\right].$
 We use Theorem 1 to compare the asymptotic variance-covariance of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ with that of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
-Corollary 1.1 shows that the benefit of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ scales directly with the predictiveness of $\widehat{Y}$ for $Y$ and inversely with the proportion of labeled data, $\pi$.
+Corollary 1.1 shows that the benefit of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ scales directly with the predictiveness of $\widehat{Y}$ for $Y$ and inversely with the proportion of labeled data $\pi$.
 Corollary 1.1 (Comparison of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ with $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$).
@@
 The sufficient condition in Corollary 1.1 coincides with the intuition that incorporating unlabeled data into estimation can improve efficiency when $\widehat{Y}$ has additional predictive power for $Y$ beyond what is already explained by the regression model of $Y$ on ${\bf{X}}$.
-This condition is similar to the result for mean estimation provided in the Supplementary Materials of Angelopoulos et al.
-[2023a] and can be used to describe several scenarios when $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is guaranteed to have lower asymptotic variance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when the ML model captures the true regression function (see Appendix A.4).
+This condition is similar to the result for mean estimation provided in the Supplementary Materials of Angelopoulos et al 21 and can be used to describe several scenarios when $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is guaranteed to have lower asymptotic variance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when the ML model captures the true regression function (see Appendix, Section A.4).
 Here we consider a simple example to demonstrate when $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ gains efficiency over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ under homoskedasticity.
@@
 In the homoskedastic setting where $\mathbb{E}(\epsilon_{Y}^{2}\mid{\bf{X}})=\sigma_{Y}^{2}$, $\mathbb{E}(\epsilon_{\widehat{Y}}^{2}\mid{\bf{X}})=\sigma_{\widehat{Y}}^{2}$, and $\mathbb{E}(\epsilon_{Y}\epsilon_{\widehat{Y}}\mid{\bf{X}})=\sigma_{Y,\widehat{Y}}$,
-$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})=\left(\frac{1}{\pi}-1\right)\left[2\sigma_{Y,\widehat{Y}}-\sigma_{\widehat{Y}}^{2}\right]{\bf{A}},$
-and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if
+$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})=\left(\frac{1}{\pi}-1\right)\left[2\sigma_{Y,\widehat{Y}}-\sigma_{\widehat{Y}}^{2}\right]{\bf{A}}.$
+$\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if
 $\rho_{Y,\widehat{Y}}>\frac{1}{2}\frac{\sigma_{\widehat{Y}}}{\sigma_{Y}},$
 where $\rho_{Y,\widehat{Y}}=\sigma_{Y,\widehat{Y}}/\sigma_{Y}\sigma_{\widehat{Y}}$ is the correlation of the residuals for $Y$ and $\widehat{Y}$ given ${\bf{X}}$.
-Put another way, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ improves over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ with respect to efficiency when $\widehat{Y}$ is a strong predictor for $Y$ after adjusting for ${\bf{X}}$, though $\widehat{Y}$ does not necessarily need to be derived from a ML model that captures the true regression function.
+Thus $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ gains efficiency over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when $\widehat{Y}$ is a strong predictor for $Y$ after adjusting for ${\bf{X}}$, though $\widehat{Y}$ does not necessarily need to be derived from a ML model that captures the true regression function.
 The improvement is stronger with increasing correlation of the residuals between $Y$ and $\widehat{Y}$ and decreasing variability in the residuals of $\widehat{Y}$.
 3.2 Asymptotic analysis of the CC estimator
-To further analyze the efficiency of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ we draw on the landmark paper that introduced a class of augmented inverse probability weighted (AIPW) estimators which includes all regular asymptotic linear estimators [Robins et al., 1994].
-In the current setting, where the goal is to estimate linear regression coefficients $\bm{\beta}^{*}$ with $Y$ partially observed, but $\widehat{Y}$ fully observed, these estimators have influence functions of the form
+To further analyze the efficiency of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, we draw on the landmark paper that introduced a class of augmented inverse probability weighted (AIPW) estimators which includes all regular asymptotic linear estimators.
+24 In the current setting, where the goal is to estimate $\bm{\beta}^{*}$ with $Y$ partially observed and $\widehat{Y}$ fully observed, these estimators have influence functions of the form
 $\text{IF}(Y,\widehat{Y},{\bf{X}};\bm{\omega})={\bf{A}}\left[\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\frac{R}{\pi}+\bm{\omega}(\widehat{Y},{\bf{X}})\frac{\pi-R}{\pi}\right],$
@@
 $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})=\mathbb{E}[\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\mid\widehat{Y},{\bf{X}}].$
-By Theorem 1, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is an AIPW estimator and can be seen as an approximation to the optimal estimator with $\bm{\omega}(\widehat{Y},{\bf{X}})=\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})$.
+By Theorem 1, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is thus an AIPW estimator and can be seen as an approximation to the optimal estimator with $\bm{\omega}(\widehat{Y},{\bf{X}})=\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})$.
 The PPI estimator is therefore not the most efficient estimator in its class and, from Corollary 1.1, can even be less efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when $\bm{\Sigma}_{\widehat{Y},\widehat{Y}}\succ 2\bm{\Sigma}_{Y,\widehat{Y}}$.
@@
 To improve upon $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, one approach is to assume a parametric model for $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$ that can be estimated from $\mathscr{L}$ for analysis.
-However, it is often difficult to specify an appropriate parametric model for $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$, particularly when the relationship between $Y$ and $\widehat{Y}$ is unknown.
+However, it is often difficult to specify an appropriate model for $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$, particularly when the relationship between $Y$ and $\widehat{Y}$ is unknown.
 Moreover, when the model for $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$ is misspecified, an efficiency gain is not guaranteed.
-To overcome these difficulties, Chen and Chen [2000] proposed a particular choice of $\bm{\omega}(\widehat{Y},{\bf{X}})$ that is guaranteed to be at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and that does not require modeling of $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$.
+To overcome these difficulties, Chen and Chen 22 proposed a particular choice of $\bm{\omega}(\widehat{Y},{\bf{X}})$ that is guaranteed to be at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and that does not require modeling of $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$.
 Their approach is motivated by the double-sampling design which, apart from differences in terminology (i.e., they regard $\widehat{Y}$ as a mismeasured version of $Y$), is analogous to the partially labeled data setting considered here.
@@
 $\bm{\omega}^{\text{\tiny{CC}}}(\widehat{Y},{\bf{X}})=\text{Cov}[\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*}),\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})]\{\text{Var}[\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})]\}^{-1}\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*}).$ (1)
-The estimator with this influence function can be written as
+The corresponding estimator can be written as
 $\hat{\bm{\beta}}^{\text{\tiny{CC}}}=\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\widehat{{\bf{W}}}^{\text{\tiny{CC}}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right),$
@@
 obtained with $\mathscr{L}$.
-Therefore, a simple way to ensure that an estimator like the PPI estimator has asymptotic variance at least as low as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ is to add a weight function in front of the augmentation term.
-In fact, the framework of Robins et al.
-[1994] implies that all regular asymptotically linear estimators for $\bm{\beta}^{*}$ augment $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ with an augmentation term that incorporates information from $\widehat{Y}$.
+Therefore, a simple way to ensure that an estimator with the same structure as the PPI estimator has asymptotic variance no greater than that of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ is to add a weight function in front of the augmentation term.
 We present the influence function expansion for $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ in Theorem 2 to further study its asymptotic variance and compare it with $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$.
-All of the asymptotic results provided for $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ are detailed in Appendix C.
 Theorem 2 (Influence function expansion of $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$).
-Following Chen and Chen [2000],
+Following Chen and Chen 22,
 $n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{CC}}}-\bm{\beta}^{*}\right)=n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{IF}}^{\text{\tiny{CC}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})+o_{p}(1),$
 where ${\bf{IF}}^{\text{\tiny{CC}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})={\bf{A}}\left[\bm{\phi}(Y_{i},{\bf{X}}_{i};\bm{\beta}^{*})\frac{R_{i}}{\pi}+\bm{\omega}^{\text{\tiny{CC}}}(\widehat{Y}_{i},{\bf{X}}_{i})\frac{\pi-R_{i}}{\pi}\right]$ and $\bm{\omega}^{\text{\tiny{CC}}}(\widehat{Y},{\bf{X}})$ is defined in (1).
-Therefore, $n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{CC}}}-\bm{\beta}^{*}\right)$ converges weakly to a zero-mean multivariate normal distribution with variance-covariance
-$\hskip 2.84526pt\mathbb{E}\left[{\bf{IF}}^{\text{\tiny{CC}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})\hskip 2.84526pt{\bf{IF}}^{\text{\tiny{CC}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})^{{\sf\scriptscriptstyle{T}}}\right]\hskip 2.84526pt.$
+Therefore, $n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{CC}}}-\bm{\beta}^{*}\right)$ converges weakly to a zero-mean multivariate normal distribution with variance-covariance $\hskip 2.84526pt\mathbb{E}\left[{\bf{IF}}^{\text{\tiny{CC}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})\hskip 2.84526pt{\bf{IF}}^{\text{\tiny{CC}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})^{{\sf\scriptscriptstyle{T}}}\right]\hskip 2.84526pt.$
 The subsequent corollaries apply Theorem 2 to show our two key results: (1) $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and (2) $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ has the highest asymptotic efficiency within the class of estimators with a weighted linear augmentation term.
-Together these results illustrate that the CC method is both robust to ML model specification and optimizes statistical efficiency, making it preferable to the PPI method in practice.
+These results suggest that the CC method is preferable to the PPI method for gaining efficiency.
 Corollary 2.1 ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$).
@@
 Let the mean squared error (MSE) of an estimator $\hat{\bm{\beta}}$ for $\bm{\beta}^{*}$ be defined as $\mathbb{E}[(\hat{\bm{\beta}}-\bm{\beta}^{*})^{{\sf\scriptscriptstyle{T}}}(\hat{\bm{\beta}}-\bm{\beta}^{*})]$.
-Analogously to the well-known result for the best linear predictor (e.g., Bickel and Doksum [2015], page 40), ${\bf{W}}^{\text{\tiny{CC}}}$ is the weight, ${\bf{W}}$, that minimizes the asymptotic MSE of estimators of the form $\hat{\bm{\beta}}^{\text{\tiny{lab}}}-{{\bf{W}}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right)$ for estimating $\bm{\beta}^{*}$.
+Analogously to the well-known result for the best linear predictor (e.g.,Bickel and Doksum,28 page 40), ${\bf{W}}^{\text{\tiny{CC}}}$ is the weight, ${\bf{W}}$, that minimizes the asymptotic MSE of estimators of the form $\hat{\bm{\beta}}^{\text{\tiny{lab}}}-{{\bf{W}}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right)$ for estimating $\bm{\beta}^{*}$.
 From Corollary 2.1, we note that in contrast to $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})$, the covariance term, $\bm{\Sigma}_{Y,\widehat{Y}}$, in $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{CC}}})$ is normalized by the residual variance of $\widehat{Y}$, $\bm{\Sigma}_{\widehat{Y},\widehat{Y}}$.
 This normalization effectively scales the covariance according to the precision of $\widehat{Y}$ and enables $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ to always be at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
-Coming back to the example from the previous section, we can gain more intuition as to why $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ possesses this property while $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ does not.
+Returning to the previous example, we can gain more intuition as to why $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ possesses this property while $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ does not.
 Example 1 continued (Efficiency gain of $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ under homoskedasticity) We revisit the case of homoskedasticity and recall that
@@
 4 Related literature
-Existing PB inference literature has established connections with semi-supervised inference, semi-parametric inference, missing data, and measurement error [Angelopoulos et al., 2023a, b, Miao et al., 2023, Zrnic and Candès, 2024a, b, Angelopoulos et al., 2024].
-Here we draw connections between PB inference methods and modern and classical methods from the economics and statistics literature dating back to the 1960s.
-These connections reveal that aspects of the PB inference problem have been explored in various areas, albeit with differences in motivation, terminology, and notation.
-Given the close relationship between the PPI and CC methods, future research in PB inference could benefit from a more comprehensive review of these and other related approaches than what is provided here.
-4.1 Seemingly unrelated regression & its use in statistical genetics
-Seemingly unrelated regression (SUR) was introduced by economist Arnold Zellner in 1962 for simultaneous estimation of multiple related regression models [Zellner, 1962, Swamy and Mehta, 1975, Schmidt, 1977].
-The regression models have different outcome variables, but are assumed to have correlated errors.
-SUR leverages the correlation of the error terms to produce more efficient estimators compared to estimating each regression independently.
-In the current setting, the SUR problem can be stated formally as
+Before evaluating the empirical properties of the CC estimator, we provide a brief overview of several closely related areas.
+Existing PB inference literature has established connections with semi-supervised inference, semi-parametric inference, missing data, and measurement error.
+21, 29, 30, 31, 23, 32, 33 We identify additional connections between PB inference methods and modern and classical methods from the economics and statistics literature dating back to the 1960s.
+These connections underscore that aspects of the PB inference problem have been previously examined, albeit with differences in motivation, terminology, and notation.
+Given the close relationship between the PPI and CC methods, ongoing research in PB inference would benefit from a more comprehensive review of these and other related approaches.
+4.1 Seemingly unrelated regression
+Seemingly unrelated regression (SUR) was introduced by economist Arnold Zellner in 1962 for simultaneous estimation of multiple related regression models.
+34, 35, 36 The regression models have different outcome variables, but are assumed to have correlated errors.
+SUR leverages the correlation between the error terms to produce more efficient estimators compared to estimating each regression separately.
+In the current setting, the SUR problem considers the regression models
 $Y={\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\beta}^{*}+\epsilon\quad\mbox{and}\quad\widehat{Y}={\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma}^{*}+\epsilon^{\prime}\quad$
-where $\mathbb{E}(\epsilon\mid{\bf{X}})=0$ and $\mathbb{E}(\epsilon^{\prime}\mid{\bf{X}})=0.$ An estimator for $\bm{\beta}^{*}$ that is based on augmenting $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ was proposed by Conniffe [1985] and is given by:
+where $\mathbb{E}(\epsilon\mid{\bf{X}})=0$ and $\mathbb{E}(\epsilon^{\prime}\mid{\bf{X}})=0.$ An estimator for $\bm{\beta}^{*}$ that is based on augmenting $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ was proposed by Conniffe 37 and is given by:
 $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}=\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\widehat{W}^{\text{\tiny{SUR}}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right)\quad\mbox{where}\quad\widehat{W}^{\text{\tiny{SUR}}}=\frac{\sum_{i=1}^{n_{\text{\tiny{lab}}}}(Y_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\hat{\bm{\beta}}^{\text{\tiny{lab}}})(\widehat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\hat{\bm{\gamma}}^{\text{\tiny{lab}}})}{\sum_{i=1}^{n_{\text{\tiny{lab}}}}(\widehat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\hat{\bm{\gamma}}^{\text{\tiny{lab}}})^{2}}.$
-Note that $\widehat{W}^{\text{\tiny{SUR}}}$ is an estimator of the scalar $W^{\text{\tiny{SUR}}}=\frac{\mathbb{E}[\text{Cov}(\epsilon,\epsilon^{\prime}\mid{\bf{X}})]}{\mathbb{E}[\text{Var}(\epsilon^{\prime}\mid{\bf{X}})]}$.
+Note that $\widehat{W}^{\text{\tiny{SUR}}}$ is an estimator of $W^{\text{\tiny{SUR}}}=\frac{\mathbb{E}[\text{Cov}(\epsilon,\epsilon^{\prime}\mid{\bf{X}})]}{\mathbb{E}[\text{Var}(\epsilon^{\prime}\mid{\bf{X}})]}$.
 Similar to the PPI and CC methods, the contribution of the augmentation term to estimation increases with the predictiveness of $\widehat{Y}$ for $Y$.
-In the extreme case where $\widehat{Y}$ is a perfect predictor for $Y$ such that $\widehat{Y}=Y$, $\widehat{W}^{\text{\tiny{SUR}}}=1$ and $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ is simply $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$.
-In the homoskedastic case when $\text{Var}(\epsilon\mid{\bf{X}})$ and $\text{Cov}(\epsilon,\epsilon^{\prime}\mid{\bf{X}})$ are constant and do not depend on ${\bf{X}}$, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is asymptotically equivalent to $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ [Chen and Chen, 2000].
-Moreover, when the conditional distribution of $(\epsilon,\epsilon^{\prime})$ given ${\bf{X}}$ is bivariate normal with mean $\bm{0}$ and has constant variance-covariance, $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ (and thus $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ as well) is the maximum likelihood estimator of $\bm{\beta}^{*}$ and fully efficient under the assumed parametric model [Conniffe, 1985].
-In other scenarios, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ achieves a lower asymptotic variance than $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$.
-The SUR estimator was recently applied for a PB inference problem inspired by genome-wide association studies conducted with population biobanks.
-These studies often rely on pre-trained models to impute outcomes of interest (i.e., phenotypes or traits) as they are partially missing due to the time and expense of ascertainment [McCaw et al., 2024].
-The proposed method, “SynSurr”, assumes the true and predicted outcomes follow a bivariate normal distribution.
+In the extreme case where $\widehat{Y}$ is a perfect predictor for $Y$ such that $\widehat{Y}=Y$, $\widehat{W}^{\text{\tiny{SUR}}}=1$, and $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ is simply $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$.
+In the homoskedastic case when $\text{Var}(\epsilon\mid{\bf{X}})$ and $\text{Cov}(\epsilon,\epsilon^{\prime}\mid{\bf{X}})$ are constant and do not depend on ${\bf{X}}$, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is asymptotically equivalent to $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$.
+22 Moreover, when the conditional distribution of $(\epsilon,\epsilon^{\prime})$ given ${\bf{X}}$ is bivariate normal with mean $\bm{0}$ and has constant variance-covariance, $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ (and thus $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ as well) is the maximum likelihood estimator of $\bm{\beta}^{*}$ and fully efficient under the assumed parametric model.
+37 In other scenarios, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ achieves a lower asymptotic variance than $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$.
+The SUR estimator was recently applied in a PB inference problem inspired by genome-wide association studies conducted with population biobanks.
+These studies often rely on pre-trained models to impute outcomes of interest (i.e., phenotypes or traits) as they are partially missing due to the time and expense of ascertainment.
+12 The proposed method, “SynSurr”, assumes the true and predicted outcomes follow a bivariate normal distribution.
 The SynSurr estimator is equivalent to $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ and asymptotically equivalent to $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ when $\text{Var}(\epsilon\mid{\bf{X}})$ and $\text{Cov}(\epsilon,\epsilon^{\prime}\mid{\bf{X}})$ are constant.
-Building on SynSurr, Miao et al.
-[2024] proposed the “POP-GWAS” estimator for a variety of regression models utilizing the augmentation technique.
+Building on SynSurr, Miao et al 14 proposed the “POP-GWAS” estimator for a variety of regression models utilizing the augmentation technique.
 The POP-GWAS estimator is
 $\hat{\bm{\beta}}^{\text{\tiny{POP}}}=\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\widehat{W}^{\text{\tiny{POP}}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}\right)\quad\mbox{where}\quad\widehat{W}^{\text{\tiny{POP}}}=\frac{n_{\text{\tiny{unlab}}}}{n}\widehat{\text{Corr}}(Y,\widehat{Y}\mid{\bf{X}}).$
-In contrast to the SUR estimator, $\hat{\bm{\beta}}^{\text{\tiny{POP}}}$ uses $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ instead of $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$ for the augmentation, and the weight, $\widehat{W}^{\text{\tiny{POP}}}$ is not generally a consistent estimator of $W^{\text{SUR}}$.
+In contrast to the SUR estimator, $\hat{\bm{\beta}}^{\text{\tiny{POP}}}$ uses $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ instead of $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$ for the augmentation, and the weight, $\widehat{W}^{\text{\tiny{POP}}}$, is not generally a consistent estimator of $W^{\text{SUR}}$.
 $\hat{\bm{\beta}}^{\text{\tiny{POP}}}$ is therefore less efficient than $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ when the residuals are bivariate normal with constant variance-covariance and less efficient than $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ in general.
 4.2 Augmented estimation for quantifying a treatment effect
-Outside of statistical genetics, the augmented estimation approach used by the PPI and CC methods has been widely used for treatment effect estimation in the statistics and biostatistics literature (e.g., Leon et al.
-[2003], Lu and Tsiatis [2008], Tsiatis et al.
-[2008], Gilbert et al.
-[2009], Zhang and Gilbert [2010], Tian et al.
-[2012], Yang and Ding [2020]).
-Many of these approaches build on methods from the survey sampling literature (e.g., Cassel et al.
-[1976], Särndal et al.
-[2003]).
-For example, Tsiatis et al.
-[2008] consider covariate adjusted analyses for comparing two treatments in randomized clinical trials.
+The augmented estimation approach used by the PPI and CC methods has also been widely used for treatment effect estimation.38, 39, 40, 41, 42, 26, 43 Many of these approaches build on methods from survey sampling.44, 45 For example, Tsiatis et al 40 consider covariate adjusted analyses for comparing two treatments in randomized clinical trials.
 Letting $X$ be an indicator for randomization to the treated or control arms, the focus of inference is on $\beta$ in the model
 $\mathbb{E}(Y\mid X)=\alpha+\beta X$
-in order to estimate the treatment effect $\Delta=\mathbb{E}(Y\mid X=1)-\mathbb{E}(Y\mid X=0)$.
+to estimate the treatment effect $\Delta=\mathbb{E}(Y\mid X=1)-\mathbb{E}(Y\mid X=0)$.
 A standard estimator of $\Delta$ is $\widehat{\Delta}=\overline{Y}^{1}-\overline{Y}^{0}$, where $\overline{Y}^{1}=n_{1}^{-1}\sum_{i=1}^{n}Y_{i}X_{i}$ and $\overline{Y}^{0}=n_{0}^{-1}\sum_{i=1}^{n}Y_{i}(1-X_{i})$ are the empirical means of $Y$ in the treated and control groups of sizes $n_{1}$ and $n_{0}$, respectively.
-To utilize information in baseline covariates, ${\bf{Z}}$, Tsiatis et al.
-[2008] propose an augmented estimator of the treatment effect given by
+To utilize information from baseline covariates ${\bf{Z}}$ for improving efficiency, Tsiatis et al 40 propose an augmented estimator of the treatment effect given by
 $\widehat{\Delta}-\sum_{i=1}^{n}(X_{i}-\bar{X})[n_{0}^{-1}h^{(0)}({\bf{Z}}_{i})+n_{1}^{-1}h^{(1)}({\bf{Z}}_{i})],$ (2)
@@
 As $X$ and ${\bf{Z}}$ are independent due to randomization, the augmentation term in 2 converges in probability to 0, analogously to the augmentation terms in the PPI and CC estimators.
-In the context of PB inference, pre-trained models could be used for predicting $Y$ on the basis of ${\bf{Z}}$ in the augmentation term.
-4.3 Improving precision of Monte Carlo simulations with control variates
+The gain in efficiency similarly arises from the potential correlation between $Y$ and ${\bf{Z}}$.
+This technique has been applied in numerous settings, most recently for the analysis of time-to-event outcomes46 and recurrent events.47
+4.3 Control variates in Monte Carlo simulations
 The augmentation technique has also been utilized to improve the efficiency of Monte Carlo estimation.
-Specifically, Rothery [1982] considers using a control variate, a statistic whose properties are known, to improve precision when estimating power via simulation.
-For exposition, suppose interest lies in estimating the power of a test $T$ of $H_{0}:\theta=0$ for some scalar parameter $\theta$.
+Specifically, Rothery 48 considers using a control variate, a statistic whose properties are known, to improve precision when estimating power via simulation.
+Suppose interest lies in estimating the power of a test $T$ of $H_{0}:\theta=0$ for some scalar parameter $\theta$.
 Let $S$ denote an existing test of the same hypothesis whose power $\mu_{Z}=\mathbb{E}_{\theta}(Z)$ is known.
@@
 $\frac{1}{n}\sum_{i=1}^{n}\left[Y_{i}-\lambda\{Z_{i}-\mu_{Z}\}\right].$ (3)
-The variance of the augmented estimator is minimized at $\lambda^{*}=\frac{\text{Cov}(Y,Z)}{\text{Var}(Z)}$ [Rothery, 1982, Davidson and MacKinnon, 1992].
-The augmentation term in 3 reduces the variance of the unadjusted power estimate $\overline{Y}$ in proportion to the correlation between $Y$ and $Z$.
-Building on work from the survey sampling literature, a similar idea was later used for mean estimation in the semi-supervised setting by Zhang et al.
-[2019].
+The variance of the augmented estimator is minimized at $\lambda^{*}=\frac{\text{Cov}(Y,Z)}{\text{Var}(Z)}$.
+48, 49 The augmentation term in 3 reduces the variance of the unadjusted power estimate $\overline{Y}$ in proportion to the correlation between $Y$ and $Z$.
+Building on work from the survey sampling literature, a similar idea was later applied for semi-supervised estimation of means.50
 5 Numerical studies
 5.1 Data generation
-We perform several simulation studies to compare the performance of the PPI and CC methods.
-In all settings, we generate two datasets: (i) a training dataset with realizations of $(Y,{\bf{Z}}^{{\sf\scriptscriptstyle{T}}},X_{1})^{{\sf\scriptscriptstyle{T}}}$ to train the ML model and (ii) an independent analysis dataset for evaluating the PB inference methods containing a labeled dataset with realizations of $(Y,{\bf{Z}}^{{\sf\scriptscriptstyle{T}}},X_{1})^{{\sf\scriptscriptstyle{T}}}$ and an unlabeled dataset with realizations of $({\bf{Z}}^{{\sf\scriptscriptstyle{T}}},X_{1})^{{\sf\scriptscriptstyle{T}}}$.
-We use 10,000 replications of the analysis dataset for each scenario and a single training set as $\hat{f}$ is assumed to be fixed.
-The size of the training set is set to 10,000 and the size of the analysis dataset is also $n=10,000$.
-We let $n_{\text{\tiny{lab}}}=1000,2000,3000,4000,5000$ and $n_{\text{\tiny{unlab}}}=n-n_{\text{\tiny{lab}}}$.
-For both the training and analysis datasets, $({\bf{Z}}^{{\sf\scriptscriptstyle{T}}},X_{1})^{{\sf\scriptscriptstyle{T}}}$ are generated from a multivariate normal distribution.
-Each element of $({\bf{Z}}^{{\sf\scriptscriptstyle{T}}},X_{1})^{{\sf\scriptscriptstyle{T}}}$ has a mean of 0, variance of 1, and covariance of 0.4 with the remaining variables.
-The goal is to conduct inference on the marginal association of $Y$ and $X_{1}$ in a linear model.
-The parameter of interest is therefore $\beta^{*}_{1}$ defined as
-$(\beta^{*}_{0},\beta^{*}_{1})=\operatorname*{arg\,min}_{\bm{\beta}}\mathbb{E}[(Y-(\beta_{0}+\beta_{1}X_{1}))^{2}].$
-We consider two primary scenarios for evaluating the PB inference methods. In both scenarios, the outcome is generated as
-$Y=\theta_{X_{1}}X_{1}+\theta_{\text{\tiny{INT}}}X_{1}W_{1}+{\bf{W}}^{{\sf\scriptscriptstyle{T}}}\bm{\theta}+\epsilon$
-where ${\bf{W}}^{{\sf\scriptscriptstyle{T}}}=(W_{1},\dots,W_{q})$ and the error term $\epsilon$ is simulated independently from a standard normal distribution.
-We varied the coefficients of $\theta_{X_{1}}$, $\theta_{\text{\tiny{INT}}}$, and $\bm{\theta}$ as well as $q$ and the composition of ${\bf{Z}}$.
-Scenario 1
+We next consider two primary scenarios to evaluate the performance of the PB inference methods and to demonstrate the implications of our theoretical results.
+In both settings, the generative model for the outcome was
+$Y=\theta_{X_{1}}X_{1}+\theta_{\text{\tiny{INT}}}X_{1}W_{1}+{\bf{W}}^{\top}\bm{\theta}+\epsilon$
+where ${\bf{W}}=(W_{1},\dots,W_{q})^{{\sf\scriptscriptstyle{T}}}$ and the error term $\epsilon$ was simulated independently from a standard normal distribution.
+We generated $({\bf{W}}^{\top},X_{1})^{{\sf\scriptscriptstyle{T}}}$ from a mean-zero multivariate normal distribution with a compound symmetric covariance structure with unit variance and pairwise correlation of 0.4.
+Our objective was to perform inference on the marginal effect of $X_{1}$ on $Y$ in a linear model.
+Specifically, we target the parameter $\beta_{1}^{*}$, defined as
+$(\beta_{0}^{*},\beta_{1}^{*})=\operatorname*{arg\,min}_{\bm{\beta}}\mathbb{E}\left[(Y-(\beta_{0}+\beta_{1}X_{1}))^{2}\right].$
+The predictors to train the ML model, ${\bf{Z}}$, were constructed based on ${\bf{W}}$ and $X_{1}$.
+The scenarios described below were obtained by varying the values of $\theta_{X_{1}}$, $\theta_{\text{\tiny{INT}}}$, and $\bm{\theta}$, as well as the dimension $q$ of ${\bf{W}}$ and the composition of the predictors ${\bf{Z}}$.
+Scenario 1 (Strong and Moderate Prediction Models)
 We consider two sub-scenarios to illustrate Corollary 1.1, that is, that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when the sufficient condition holds.
@@
 The $R^{2}$ of $\widehat{Y}$ for $Y$ is 0.9 and 0.6 in Scenarios 1a and 1b, respectively.
-Scenario 2
+Scenario 2 (Weak Prediction Models)
 We consider two sub-scenarios to illustrate the results in Corollary 1.1 and Corollary 2.1, which show that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ can be less efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ with a weak prediction model while $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is at least as efficient $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
@@
 The $R^{2}$ of $\widehat{Y}$ for $Y$ is 0.5 and 0.4 in Scenarios 2a and 2b, respectively.
-We evaluate the performance of the PPI and CC methods: $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, and $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$.
-We also include two comparison methods: $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, serving as the baseline, and $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$, highlighting the limitations of naive inference.
-We present the percent bias for each estimator.
-For the PPI and CC methods, we present the confidence interval (CI) length ratio, computed as the mean of the ratio of the length of the PB inference estimator’s normal-based CI relative to the length of the normal-based CI for $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ (which is simply the ratio of the asymptotic standard errors).
-The CI length ratio has a baseline value of 1, which indicates that the CI of the PB inference method is the same length as the CI for $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
-A CI length ratio less than 1 indicates that the PB inference estimator has lower standard error than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
+In all settings, we generated two datasets from the same underlying distribution: (i) a training dataset to train the ML model and (ii) an independent analysis dataset for evaluating the PB inference methods containing a labeled dataset and an unlabeled dataset.
+We used 10,000 replications of the analysis dataset for each scenario and a single training set as $\hat{f}$ is assumed to be fixed.
+The size of the training set was 100,000 and the size of the analysis dataset was $n=10,000$.
+We varied the proportion of labeled data in the analysis dataset by letting $n_{\text{\tiny{lab}}}=300,600,1000$ or $2000$ in each setting.
+We evaluated the performance of the PPI and CC methods: $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, and $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$.
+We also included three comparison methods: $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ to serve as the baseline, $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$ to highlight the limitations of naive inference, and a PB inference method from a concurrent pre-print, the PoSt-Prediction Adaptive (PSPA) method.30 The PSPA method is a generalization of the PPI++ method, an extension of the original PPI estimator that adaptively minimizes the trace of the asymptotic variance-covariance of the PPI estimator.
+The PSPA estimator, denoted as $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$, is designed to achieve element-wise variance reduction relative to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ by incorporating the information from $\widehat{Y}$ directly into the estimating equation for $\bm{\beta}^{*}$.
+We present the percent bias for each estimator as well as the coverage probability (CP) of the 95% confidence intervals (CIs).
+For the four PB inference methods ($\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$, and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$), we also present the relative efficiency (RE) defined as the ratio of the mean squared error (MSE) of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ relative to the MSE of the PB inference estimator.
 5.2 Results
 Table 1 presents the percent bias of the various estimators across the different scenarios and labeled dataset sizes.
-With the exception of $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$, the methods exhibit minimal bias as they are all consistent estimators for $\beta^{*}_{1}$.
+With the exception of $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$, the methods exhibit minimal bias as they are all consistent estimators for the target parameter.
 The limitations of naive inference are particularly apparent in Scenarios 2a and 2b, which do not include $X_{1}$ in ML model training.
-$\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ exhibits an average percent bias of -47% in Scenario 2a and -64% in Scenario 2b.
-Figure 2 displays the CI length ratio across the different settings.
-In Scenario 1, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ and $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ perform similarly, with CI length ratios below 1, indicating that the PB inference methods have reduced standard error relative to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
+$\hat{\bm{\gamma}}^{\text{\tiny{all}}}$ exhibits an average percent bias of -47% in Scenario 2a and -64% in Scenario 2b.
+Table 2 summarizes the CPs of the 95% CIs for all methods across the various scenarios and labeled dataset sizes.
+As expected, $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and the four PB inference estimators ($\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$, and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$) maintain coverage near the nominal level in all settings.
+In contrast, the 95% CIs based on $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$ achieve adequate coverage only in Scenario 1a, where a strong prediction model yields negligible bias in estimating $\beta_{1}^{*}$.
+Coverage drops to 92% in Scenario 1b and falls to 0% in Scenarios 2a and 2b, where bias is substantial.
+These results underscore the risks of relying on naive inference in practice and highlight the necessity of using reliable PB inference methods.
+Figure 2 displays the relative efficiency across the different settings.
+In Scenario 1, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ perform similarly, with RE significantly above 1, indicating that the PB inference methods have lower variance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
 Comparing Scenario 1a and 1b, we note a significant decrease in the gains in statistical efficiency of the PB inference methods due to the weaker prediction model used in Scenario 1b.
 However, all methods are at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, as the sufficient condition in Corollary 1.1 is satisfied.
-Additionally, as the labeled set size increases and hence $\pi=n_{\text{\tiny{lab}}}/n$ increases, the efficiency gain decreases as expected.
-$\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is always more efficient than $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, particularly when the labeled dataset size is larger, due to the difference in the augmentation terms of these two estimators.
-In this scenario, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is asymptotically equivalent to $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$.
-Scenario 2 highlights the differences in performance between $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ and $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$, which underscore our findings in Corollary 1.1 and Corollary 2.1.
-In both sub-scenarios, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ consistently outperforms both $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ and is at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
-In Scenario 2a, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ achieves an average CI length ratio of 0.9 across the various sample sizes, indicating a gain in efficiency relative to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
-In Scenario 2b, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is comparable to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ for the larger labeled dataset sizes as there is not much to gain from the unlabeled data due to its relatively small size and the fact that $\widehat{Y}$ is a weak prediction of $Y$.
-In contrast, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ are outperformed by $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ in nearly all scenarios.
-In Scenario 2a, these estimators have CIs up to 30% longer than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ for the larger labeled dataset sizes.
-This behavior is exacerbated in Scenario 2b, where the CI of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is up to 50% longer than the CI for $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and up to 72% longer for $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$.
-Labeled set size
-Method 1000 2000 3000 4000 5000
-Scenario 1a
-     $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$0.181 0.124 0.650 0.430 0.165
-     $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ 0.035 -0.042 0.015 0.005 -0.005
-     $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ 0.004 0.010 -0.004 -0.018 -0.001
-     $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ 0.007 0.000 0.002 -0.009 -0.003
-     $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ 0.007 0.000 0.002 -0.009 -0.003
-Scenario 1b
-     $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$1.070 0.181 0.426 -0.423 -1.205
-     $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ 0.008 0.002 -0.016 -0.001 0.004
-     $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ 0.007 -0.001 -0.003 -0.008 -0.010
-     $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ 0.007 $0$0.007 -0.006 -0.003
-     $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ 0.006 $0$0.007 -0.006 -0.003
-Scenario 2a
-     $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$46.980 -47.523 -45.946 -46.725 -47.695
-     $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$0.033 -0.023 0.027 0.004 0.005
-     $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$0.023 -0.009 -0.003 -0.032 -0.003
-     $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$0.024 -0.012 0.006 -0.018 0.001
-     $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$0.030 -0.016 0.016 -0.007 0.003
-Scenario 2b
-     $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$64.272 -64.115 -63.307 -63.759 -64.627
-     $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$0.012 -0.003 0.002 0.012 0.013
-     $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ 0.000 0.001 -0.006 0.012 0.009
-     $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$0.001 $0$0.004 0.012 0.010
-     $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$0.014 -0.006 -0.003 0.011 0.010
-Table 1: Percent biases of the PPI estimators ($\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$), the CC estimator ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$), the baseline estimator ($\hat{\bm{\beta}}^{\text{\tiny{lab}}}$), and the naive estimator ($\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$).
-Negative values indicate underestimation, while positive values indicate overestimation of the true parameter.
-Figure 2: Confidence interval length ratio of the PB inference methods compared to baseline estimation with the labeled data ($\hat{\bm{\beta}}^{\text{\tiny{lab}}}$).
-A ratio less than 1 indicates that the PB inference method has a confidence interval that is shorter than the confidence interval using $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
+Additionally, as the labeled set size increases and hence $n_{\text{\tiny{lab}}}/n$ increases, the efficiency gain decreases as expected.
+$\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is always at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ in Scenarios 1a and 1b, and more efficient in Scenario 1a when the labeled dataset size is larger due to the difference in the augmentation terms of these two estimators.
+Scenario 2 highlights the performance differences between $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ and $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$, which underscore our findings in Corollary 1.1 and Corollary 2.1.
+In both sub-scenarios, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ consistently outperforms both $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ and is substantially more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
+In Scenario 2a, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ achieves an average RE of 1.27 across the various sample sizes, indicating a significant gain over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
+In Scenario 2b, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ again demonstrates superior RE compared to all other methods and remains more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
+$\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ offers a modest efficiency gain in Scenario 2a and comparable performance to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ in Scenario 2b.
+Both $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ are outperformed by $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ in nearly all scenarios.
+In Scenario 2a, while $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ remains comparable to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, the RE for $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is as low as 0.88 with $n_{\text{\tiny{lab}}}=2000$.
+Both $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ demonstrate a substantial loses in efficiency in Scenario 2b, with $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ exhibiting a RE of 0.63 with $n_{\text{\tiny{lab}}}=2000$.
+Figure 2: Relative efficiency of the PB inference methods compared to baseline estimation with the labeled data ($\hat{\bm{\beta}}^{\text{\tiny{lab}}}$).
+A ratio greater than 1 indicates that the PB inference method has a smaller mean squared error (MSE) than that of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
 6 UK Biobank Analysis
-Central obesity, defined as excessive accumulation of body fat around abdominal region, is strongly associated with cardiometabolic diseases such as type 2 diabetes and cardiovascular disease.
-A major challenge in studying central obesity is the limited availability of direct body composition measurements.
-Android fat mass (AFM), measured via a dual-energy X-ray absorptiometry (DEXA) scan, is a key marker for assessing central obesity in clinical practice [Zhu et al., 2022].
-However, due to its high cost, DEXA-derived AFM is usually only measured in a small subset of participants in large-scale studies.
-For example, in the UK Biobank, a prospective cohort study measuring genetic, lifestyle and health information on approximately 500,000 UK participants, DEXA-derived AFM is incompletely ascertained, particularly among populations populations historically underrepresented in genetic research [Sudlow et al., 2015].
-For instance, out of $5,397$ South Asian participants in the UK Biobank, only $288$ (5.3%) participants have undergone DEXA scans.
-[…]
-Table 2 presents point estimates, standard errors, and CI ratios (or equivalently standard errors length ratios) across the two prediction model settings for four estimators: the PPI estimators ($\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$), the CC estimator ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$), and the baseline estimator ($\hat{\bm{\beta}}^{\text{\tiny{lab}}}$).
-Echoing our theoretical and simulation study results, all methods yield similar point estimates as they are all consistent estimators for the same parameter.
-For the analysis with the strong prediction model, $\widehat{Y}$ provides additional predictive power beyond the regression of $Y$ on ${\bf{X}}$, satisfying the sufficient condition in Corollary 2.1.
+We also evaluated the performance of the PB inference methods in an analysis focused on central obesity.
+Central obesity is excessive accumulation of body fat around the abdominal region and is strongly associated with cardiometabolic diseases such as type 2 diabetes and cardiovascular disease.
+A major challenge in its study is the limited availability of direct body composition measurements.
+Android fat mass (AFM), measured via a dual-energy X-ray absorptiometry (DEXA) scan, is a key marker for assessing central obesity in clinical practice.
+51 However, due to its high cost, DEXA-derived AFM is usually only measured in a small subset of participants in large-scale studies.
+For example, in the UK Biobank,52 a prospective cohort study measuring genetic, lifestyle, and health information on approximately 500,000 participants in the UK, DEXA-derived AFM is incompletely ascertained, particularly among populations historically underrepresented in genetic research.
+52 For instance, out of $5,397$ South Asian participants in the UK Biobank, only $288$ (5.3%) participants have undergone DEXA scans.
+Our goal is to examine associations between AFM and age, sex, and body mass index (BMI) among the South Asian population.
+Previous research has indicated that this population is at higher risk of central obesity and that it is associated with cardiometabolic diseases at lower BMI thresholds than European populations.
+53 Since DEXA-derived AFM is missing for most South Asian participants, it provides a natural opportunity to leverage ML-predicted AFM in downstream analyses.
+To illustrate the impact of different ML models on the efficiency of the PB inference methods, we trained two random forest models in the White British population ($32,460$ participants) and used them to predict missing AFM values in the South Asian population.
+We built a “strong” prediction model that incorporated both the covariates of interest (i.e., age, sex, and BMI) as well as additional predictors, including smoking status, alcohol consumption, two bioelectrical impedance measurements of the legs, physical activity scores from the International Physical Activity Questionnaire, and sedentary behavior quantified as total hours spent driving, using a computer, and watching television.
+We built another “weak” prediction model that included only the additional predictors.
+The strong prediction model achieved an $R^{2}$ of 0.8 while the weak prediction model had a substantially lower $R^{2}$ of 0.5 in the South Asian subpopulation.
+Table 3 presents point estimates, standard errors, and RE, defined as the ratio of the variance of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ to the variance of the given estimator, across the two prediction model settings for five estimators: the baseline estimator ($\hat{\bm{\beta}}^{\text{\tiny{lab}}}$), the PPI estimators ($\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$), the PSPA estimator ($\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$), and the CC estimator ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$).
+We do not include naive inference in our analyses due to the limitations demonstrated in our simulation studies.
+Echoing our theoretical and simulation study results, all methods yield similar point estimates as they are all consistent estimators of the same parameter.
+In the analysis with the strong prediction model, the predictions provide information on central obesity beyond what is captured by the covariates of interest, satisfying the sufficient condition in Corollary 2.1.
 Consequently, both $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ outperform $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
-In fact, all three PB inference methods ($\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, and $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$) perform similarly, yielding substantially smaller standard errors compared to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ with average CI ratios below 0.85 for all methods.
-The analysis with the weak prediction model highlights the difference in performance between the PPI and CC methods and reinforces the findings in Corollary 2.1 and Corollary 2.2.
-Here, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ yield larger standard errors than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ for all parameters of interest, with CIs 38.5% longer on average.
-In contrast, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is no worse than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ for all parameters of interest.
-Strong Prediction Model
-Intercept Age Sex BMI
-$\hat{\bm{\beta}}^{\text{\tiny{lab}}}$
-EST -4.857 -0.003 0.813 0.173
-SE 0.308 0.004 0.070 0.009
-CI Ratio 1.00 1.00 1.00 1.00
-$\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$
-EST -4.513 -0.007 0.904 0.164
-SE 0.261 0.004 0.063 0.007
-CI Ratio 0.85 0.86 0.89 0.80
-$\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$
-EST -4.527 -0.007 0.900 0.165
-SE 0.260 0.004 0.063 0.007
-CI Ratio 0.85 0.86 0.89 0.80
-$\hat{\bm{\beta}}^{\text{\tiny{CC}}}$
-EST -4.525 -0.006 0.905 0.163
-SE 0.255 0.004 0.062 0.007
-CI Ratio 0.83 0.85 0.89 0.77
-Weak Prediction Model
-Intercept Age Sex BMI
-$\hat{\bm{\beta}}^{\text{\tiny{lab}}}$
-EST -4.857 -0.003 0.813 0.173
-SE 0.308 0.004 0.070 0.009
-CI Ratio 1.00 1.00 1.00 1.00
-$\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$
-EST -4.530 -0.006 0.835 0.165
-SE 0.416 0.006 0.098 0.012
-CI Ratio 1.35 1.43 1.40 1.36
-$\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$
-EST -4.546 -0.006 0.834 0.165
-SE 0.415 0.006 0.098 0.012
-CI Ratio 1.35 1.43 1.40 1.36
-$\hat{\bm{\beta}}^{\text{\tiny{CC}}}$
-EST -4.779 -0.004 0.819 0.171
-SE 0.294 0.004 0.072 0.008
-CI Ratio 0.95 0.96 1.03 0.93
-Table 2: UK Biobank analysis of android fat mass based on the PPI estimators ($\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$), the CC estimator ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$), the baseline estimator ($\hat{\bm{\beta}}^{\text{\tiny{lab}}}$).
-The strong prediction model has an $R^{2}$ of 0.8, while the weak prediction model has an $R^{2}$ of 0.5.
-Estimates (EST), standard errors (SE), and confidence interval (CI) ratios are reported to evaluate performance.
-A CI ratio less than 1 indicates that a method yields a confidence interval shorter than the CI using $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
+In fact, all four PB inference methods ($\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$) yield substantially smaller standard errors compared to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
+Among the four PB inference methods, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ show comparable performance, with average REs of 1.39, 1.39, and 1.47, respectively.
+$\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ exhibits higher efficiency gains with an average RE of 1.68.
+The analysis with the weak prediction model highlights the differences in performance between the PPI and CC methods and reinforces the findings in Corollary 2.1 and Corollary 2.2.
+Here, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ yield larger standard errors than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ for all parameters of interest, with an average RE of 0.52.
+In contrast, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$, as well as $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$, outperform $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, with average RE values of 1.28 and 1.16, respectively.
+Overall, the CC estimator is the best performing method in terms of efficiency, underscoring its utility in practice.
 7 Discussion
-In this paper, we showed that adding a simple weight to the PPI estimator provides valid PB inference that is guaranteed to be at least as efficient as estimation using only labeled data.
-The weight is drawn from the work of Chen and Chen [2000] in the context of double sampling.
-In contrast, the original PPI estimator can be less efficient than estimation using only the labeled data, a property that threatens its practical utility.
-By incorporating the proposed weight, the augmented estimation approach principally balances the contribution of labeled and unlabeled data and extends the applicability of the PPI method to scenarios where leveraging all available data is essential for making valid and efficient inference.
-In addition to theoretical results that support these findings, our simulation studies and application to UK Biobank data demonstrate that the PPI method can result in substantial efficiency loss when the ML model is inaccurate.
-Additionally, we provide connections to several existing approaches from the economics and statistics literature that provide interesting context for the PB inference problem.
-While we focused on least squares estimation for exposition, our results extend to other settings.
-The work of Chen and Chen [2000] and Angelopoulos et al.
-[2023a] apply to a wide variety of settings beyond linear regression, including Z-estimation problems.
-The work of Angelopoulos et al.
-[2023a] was pivotal in advancing the growing body of literature in PB inference.
-Recent work has considered settings where the ML model is not black box and where active learning is utilized to obtain the labeled data [Zrnic and Candès, 2024b, b].
-Additionally, two concurrent pre-prints have noted potential losses in efficiency of PB inference approaches.
-The authors of the PPI paper introduced a new method, PPI++, which aims to address this limitation of the PPI estimator [Angelopoulos et al., 2023b].
-The PPI++ method is designed to adaptively minimize the trace of the asymptotic variance-covariance of the proposed estimator to ensure that it yields an improvement over an approach based on labeled data.
-In contrast, the approach of Chen and Chen [2000] minimizes the mean square error among a class of estimators with a linear augmentation term.
-Simultaneously, Miao et al.
-[2023] introduced the PoSt-Prediction Adaptive inference (PSPA) estimator that is both robust to the accuracy of the ML model and guaranteed to be at least as efficient as estimation using only labeled data.
-Unlike the PPI++ estimator, the PSPA estimator is designed to achieve element-wise variance reduction and was shown to outperform the PPI++ approach.
-The PSPA approach is based on augmenting estimating equations with information from $\widehat{Y}$, while the CC method provides a simple and natural extension to the PPI method by adding a weight to the linear augmentation term.
+In this paper, we bridge classical statistical methods with a modern problem to equip practitioners and methodologists with a rigorous approach for PB inference.
+We find that incorporating a simple weight in the PPI estimator previously proposed for studies with double-sampling designs22 yields valid inference that is guaranteed to be at least as efficient as estimation using only labeled data.
+Our work addresses efficiency shortcomings of the original PPI estimator which can undermine its practical value.
+Alongside theoretical results supporting these findings, our simulation studies and application to UK Biobank data verify that the PPI method can suffer substantial efficiency loss when the ML model is inaccurate.
+Additionally, we draw important connections to several established approaches from the economics and statistics literature to provide context for ongoing developments in PB inference.
+Our work is not without limitations.
+First, we focused our analyses on least squares estimation following Motwani and Witten.20 Further work is needed to analyze other statistical procedures (e.g., Z-estimation problems in general).
+Second, we assumed that the labeled and unlabeled data are drawn from the same underlying distribution or equivalently, that $Y$ is missing completely at random.
+This assumption may not hold in practice, particularly when samples cannot be randomly selected for labeling due to study design constraints or the use of alternative sampling strategies to enhance efficiency (e.g., stratified sampling).
+However, the CC estimator can be extended to settings where $Y$ is missing at random by incorporating inverse probability weighting and regression adjustment.
+Lastly, we focused on the traditional missing data scenario where the proportion of labeled data, $n_{\text{\tiny{lab}}}/n$, is bounded away from 0.
+Analyses of semi-supervised settings where $\lim_{n_{\text{\tiny{lab}}}\to\infty}n_{\text{\tiny{lab}}}/n=0$ would be of interest to formally model scenarios where the size of labeled and unlabeled are of different orders of magnitude.
 SUPPLEMENTARY MATERIAL
-Title:
-The supplementary materials include proofs for theorems and additional examples for Corollary 1.1 (appendix.pdf).
+The supplementary materials include proofs for theorems and additional examples for Corollary 1.1.
 Code availability:
@@
 This work used data from the UK Biobank study (https://www.ukbiobank.ac.uk) and our access was approved under application 64875.
-Appendix A Asymptotic properties of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$
-A.1 Preliminaries
-Notation.
-We denote the outer product of a vector $\bm{v}$ as $\bm{v}^{\otimes 2}=\bm{v}\bm{v}^{{\sf\scriptscriptstyle{T}}}$.
-We denote that a matrix, ${\bf{M}}$, is positive semi-definite as ${\bf{M}}\succeq{\bf{0}},$ and positive definite as ${\bf{M}}\succ{\bf{0}}.$ For any estimator $\hat{\bm{\beta}}$, we let $\text{aVar}(\hat{\bm{\beta}})$ represent its asymptotic variance-covariance matrix.
-We define the difference in asymptotic variance-covariance matrices between two estimators, $\hat{\bm{\beta}}_{1}$ and $\hat{\bm{\beta}}_{2}$, as $\Delta_{\text{aVar}}(\hat{\bm{\beta}}_{1},\hat{\bm{\beta}}_{2})=n[\text{aVar}(\hat{\bm{\beta}}_{1})-\text{aVar}(\hat{\bm{\beta}}_{2})].$
-Basic assumptions.
-We make the standard assumption that $\text{Var}({\bf{X}})\succ{\bf{0}}$.
-We also assume that ${\bf{X}}$ has compact support and that $(Y_{i},\widehat{Y}_{i},{\bf{X}}^{{\sf\scriptscriptstyle{T}}})^{{\sf\scriptscriptstyle{T}}}$ have finite first and second moments.
-A.2 Proof of Theorem 1
-Theorem 1 (Influence function expansion of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$).
-Let $R_{i}$ be the indicator of whether the $i$th observation is labeled. Then,
-$n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}-\bm{\beta}^{*}\right)=n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})+o_{p}(1)$
-where ${\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})={\bf{A}}\left\{\left[\ \bm{\phi}(Y_{i},{\bf{X}}_{i};\bm{\beta}^{*})\right]\frac{R_{i}}{\pi}+\left[\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma}^{*})\right]\frac{\pi-R_{i}}{\pi}\right\}$ and ${\bf{A}}=[\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})]^{-1}$.
-Therefore, $n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}-\bm{\beta}^{*}\right)$ converges weakly to a zero-mean multivariate normal distribution with variance-covariance
-$\mathbb{E}[{\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})\hskip 2.84526pt{\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})^{{\sf\scriptscriptstyle{T}}}].$
-Proof.
-First, note that $n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}-\bm{\beta}^{*}\right)$ may be decomposed as
-$n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}-\bm{\beta}^{*}\right)=n^{\frac{1}{2}}\left[\left(\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\bm{\beta}^{*}\right)-\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\bm{\gamma}^{*}\right)+\left(\hat{\bm{\gamma}}^{\text{\tiny{all}}}-\bm{\gamma}^{*}\right)\right].$
-We consider the asymptotic linear expansions of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, $\hat{\bm{\gamma}}^{\text{\tiny{lab}}}$, and $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$.
-Since $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ is the familiar least squares estimator,
-$n_{\text{\tiny{lab}}}^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\bm{\beta}^{*}\right)=\left(n_{\text{\tiny{lab}}}^{-1}\sum_{i=1}^{n_{\text{\tiny{lab}}}}{\bf{X}}_{i}{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\right)^{-1}n_{\text{\tiny{lab}}}^{-\frac{1}{2}}\sum_{i=1}^{n_{\text{\tiny{lab}}}}{\bf{X}}_{i}(Y_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\bm{\beta}^{*}).$ (4)
-By Slutsky’s theorem and the assumption that $\lim_{n\to\infty}n_{\text{\tiny{lab}}}/n=\pi$, equation (4) can be rewritten as
-$\begin{split}&n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\bm{\beta}^{*}\right)=n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{A}}\left\{\left[{\bf{X}}_{i}\left(Y_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\bm{\beta}^{*}\right)\right]\frac{R_{i}}{\pi}\right\}+o_{p}(1).\end{split}$ (5)
-Similar arguments can also be used to show that
-$n^{\frac{1}{2}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\bm{\gamma}^{*}\right)=n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{A}}\left\{\left[{\bf{X}}_{i}\left(\widehat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma}^{*}\right)\right]\frac{R_{i}}{\pi}\right\}+o_{p}(1)$
-and
-$n^{\frac{1}{2}}\left(\hat{\bm{\gamma}}^{\text{\tiny{all}}}-\bm{\gamma}^{*}\right)=n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{A}}\left\{\left[{\bf{X}}_{i}\left(\widehat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma}^{*}\right)\right]\right\}+o_{p}(1).$
-Consequently,
-$n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}-\bm{\beta}^{*}\right)$ $=n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{A}}\left\{\left[\bm{\phi}(Y_{i},{\bf{X}}_{i};\bm{\beta}^{*})\frac{R_{i}}{\pi}+\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i},\bm{\gamma}^{*})\frac{\pi-R_{i}}{\pi}\right]\right\}+o_{p}(1)$
-$=n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})+o_{p}(1).$
-It follows from the standard central limit theorem (CLT) that $n^{\frac{1}{2}}(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}-\bm{\beta}^{*})$ converges weakly to a zero mean multivariate Gaussian distribution with variance-covariance given by $\mathbb{E}\left[{\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})\hskip 2.84526pt{\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})^{{\sf\scriptscriptstyle{T}}}\right].$
-A.3 Proof of Corollary 1.1
-Corollary 1.1 (Comparison of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ with $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$).
-The difference in the asymptotic variance-covariance of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is
-$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})=\left(\frac{1}{\pi}-1\right){\bf{A}}\left(2\bm{\Sigma}_{Y,\widehat{Y}}-\bm{\Sigma}_{\widehat{Y},\widehat{Y}}\right){\bf{A}}.$
-$\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is therefore more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if $2\bm{\Sigma}_{Y,\widehat{Y}}\succ\bm{\Sigma}_{\widehat{Y},\widehat{Y}}$.
-A sufficient condition for $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ to have lower variance-covariance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ is that
-$\mathbb{E}\left[\left(\epsilon_{\widehat{Y}}-\epsilon_{Y}\right)^{2}\mid{\bf{X}}\right]<\mathbb{E}\left(\epsilon_{Y}^{2}\mid{\bf{X}}\right).$
-Proof.
-Equation (5) in the proof of Theorem 1 gives the asymptotic expansion of $n^{\frac{1}{2}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\bm{\beta})$.
-The standard CLT implies $n^{\frac{1}{2}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\bm{\beta})$ converges weakly to a zero-mean multivariate Gaussian distribution with variance-covariance given by
-$\mathbb{E}\left\{\left[{\bf{A}}\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\right]^{\otimes 2}\frac{R^{2}}{\pi^{2}}\right\}=\frac{1}{\pi}{\bf{A}}\bm{\Sigma}_{Y,Y}{\bf{A}}.$
-Similarly, by Theorem 1
-$\begin{split}n[\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})]=&\mathbb{E}\left\{\left[{\bf{A}}\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\right]^{\otimes 2}\frac{R^{2}}{\pi^{2}}\right\}+\mathbb{E}\left\{\left[{\bf{A}}\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})\right]^{\otimes 2}\frac{(\pi-R)^{2}}{\pi^{2}}\right\}\\ +&2\mathbb{E}\left\{\left[{\bf{A}}\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\right]\left[{\bf{A}}\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})\right]^{{\sf\scriptscriptstyle{T}}}\frac{R(\pi-R)}{\pi^{2}}\right\}\\ =&\frac{1}{\pi}{\bf{A}}\left[\bm{\Sigma}_{Y,Y}+(1-\pi)\bm{\Sigma}_{\widehat{Y},\widehat{Y}}-2(1-\pi)\bm{\Sigma}_{Y,\widehat{Y}}\right]{\bf{A}}.\end{split}$
-Thus,
-$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})=\left(\frac{1}{\pi}-1\right){\bf{A}}\left(2\bm{\Sigma}_{Y,\widehat{Y}}-\bm{\Sigma}_{\widehat{Y},\widehat{Y}}\right){\bf{A}}.$
-$\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is therefore more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if $2\bm{\Sigma}_{Y,\widehat{Y}}\succ\bm{\Sigma}_{\widehat{Y},\widehat{Y}}$.
-Additionally, if $2\mathbb{E}\left(\epsilon_{Y}\epsilon_{\widehat{Y}}\mid{\bf{X}}\right)>\mathbb{E}\left(\epsilon_{\widehat{Y}}^{2}\mid{\bf{X}}\right)$ or equivalently if $\mathbb{E}\left[\left(\epsilon_{\widehat{Y}}-\epsilon_{Y}\right)^{2}\mid{\bf{X}}\right]<\mathbb{E}\left(\epsilon_{Y}^{2}\mid{\bf{X}}\right),$ the diagonal entries of $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})$ are positive and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ will have a lower variance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
-∎
-A.4 Examples for Corollary 1.1
-We provide several scenarios where the sufficient condition from Corollary 1.1 is satisfied and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
-These scenarios characterize settings where the predictions contain more information about $Y$ beyond what is explained by the linear regression model based on ${\bf{X}}$ (i.e., the focus of statistical inference.)
-Examples of Corollary 1.1.
-Let ${\bf{Z}}=({\bf{X}}^{{\sf\scriptscriptstyle{T}}},{\bf{W}}^{{\sf\scriptscriptstyle{T}}})^{{\sf\scriptscriptstyle{T}}}$ where ${\bf{W}}$ are predictors of $Y$ beyond ${\bf{X}}$.
-$\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ in the following scenarios where $\hat{f}$ perfectly estimates the true regression function:
-  1. 1.
-    $\hat{f}({\bf{X}})=\mathbb{E}(Y\mid{\bf{X}})$ and $\mathbb{E}(Y\mid{\bf{X}})\neq{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\beta}^{*}$
-  2. 2.
-    $\hat{f}({\bf{Z}})=\mathbb{E}(Y\mid{\bf{Z}})$, ${\bf{W}}\perp Y\mid{\bf{X}}$, and $\mathbb{E}(Y\mid{\bf{X}})\neq{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\beta}^{*}$
-  3. 3.
-    $\hat{f}({\bf{Z}})=\mathbb{E}(Y\mid{\bf{Z}})$ and ${\bf{W}}\not\perp Y\mid{\bf{X}}$
-  4. 4.
-    $\hat{f}({\bf{W}})=\mathbb{E}(Y\mid{\bf{W}})$ and ${\bf{X}}\perp Y\mid{\bf{W}}$
-Proof.
-For scenario 1, assume that $\hat{f}({\bf{X}})=\mathbb{E}(Y\mid{\bf{X}})$. It follows that
-$\bm{\gamma}^{*}=[\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})]^{-1}\mathbb{E}[{\bf{X}}\mathbb{E}(Y\mid{\bf{X}})]=\bm{\beta}^{*}$
-so that
-$\mathbb{E}\left[\left(\epsilon_{\widehat{Y}}-\epsilon_{Y}\right)^{2}\mid{\bf{X}}\right]$ $=\mathbb{E}\left\{[\mathbb{E}(Y\mid{\bf{X}})-Y]^{2}\mid{\bf{X}}\right\}$ $<\mathbb{E}\left(\epsilon_{Y}^{2}\mid{\bf{X}}\right)=\mathbb{E}\left[(Y-{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\beta}^{*})^{2}\mid{\bf{X}}\right].$
-This inequality is strict provided that $\mathbb{E}(Y\mid{\bf{X}})\neq{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\beta}^{*}$.
-If the regression model is also correctly specified (i.e., $\mathbb{E}(Y\mid{\bf{X}})={\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\beta}^{*}$), then $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ are asymptotically equivalent.
-For scenario 2, assume that $\hat{f}({\bf{Z}})=\mathbb{E}(Y\mid{\bf{Z}})$ where ${\bf{Z}}=({\bf{X}}^{{\sf\scriptscriptstyle{T}}},{\bf{W}}^{{\sf\scriptscriptstyle{T}}})^{{\sf\scriptscriptstyle{T}}}$ and ${\bf{W}}\perp Y\mid{\bf{X}}$.
-Similar to scenario 1,
-$\bm{\gamma}^{*}=[\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})]^{-1}\mathbb{E}[{\bf{X}}\mathbb{E}(Y\mid{\bf{X}},{\bf{W}})]=[\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})]^{-1}\mathbb{E}[{\bf{X}}\mathbb{E}(Y\mid{\bf{X}})]=\bm{\beta}^{*}.$
-Therefore,
-$\mathbb{E}\left[\left(\epsilon_{\widehat{Y}}-\epsilon_{Y}\right)^{2}\mid{\bf{X}}\right]$ $=\mathbb{E}\left\{[\mathbb{E}(Y\mid{\bf{Z}})-Y]^{2}\mid{\bf{X}}\right\}$
-$=\mathbb{E}\left\{[\mathbb{E}(Y\mid{\bf{X}})-Y]^{2}\mid{\bf{X}}\right\}=\text{Var}(Y\mid{\bf{X}})\leq\mathbb{E}\left(\epsilon_{Y}^{2}\mid{\bf{X}}\right).$
-The inequality is strict when $\mathbb{E}(Y\mid{\bf{X}})\neq{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\beta}^{*}$ and is an equality if the regression model is correctly specified.
-For scenario 3, again assume that $\hat{f}({\bf{Z}})=\mathbb{E}(Y\mid{\bf{Z}})$ so that $\bm{\gamma}^{*}=\bm{\beta}^{*}$. It then follows that
-$\mathbb{E}\left[\left(\epsilon_{\widehat{Y}}-\epsilon_{Y}\right)^{2}\mid{\bf{X}}\right]$ $=\mathbb{E}\left\{[\mathbb{E}(Y\mid{\bf{Z}})-Y]^{2}\mid{\bf{X}}\right\}$
-$=\text{Var}(Y\mid{\bf{X}})-\text{Var}[\mathbb{E}(Y\mid{\bf{X}},{\bf{W}})\mid{\bf{X}}]$
-$<\text{Var}(Y\mid{\bf{X}})\leq\mathbb{E}\left(\epsilon_{Y}^{2}\mid{\bf{X}}\right).$
-The equality is strict as we assume ${\bf{W}}\not\perp Y\mid{\bf{X}}$.
-Lastly, for scenario 4, assume that $\hat{f}({\bf{W}})=E(Y\mid{\bf{W}})$ and ${\bf{X}}\perp Y\mid{\bf{W}}$. Then,
-$\bm{\gamma}^{*}=[\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})]^{-1}\mathbb{E}[{\bf{X}}\mathbb{E}(Y\mid{\bf{W}})]=[\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})]^{-1}\mathbb{E}[{\bf{X}}\mathbb{E}(Y\mid{\bf{W}},{\bf{X}})]=\bm{\beta}^{*}.$
-Analogously to the argument for scenario 3,
-$\mathbb{E}\left[\left(\epsilon_{\widehat{Y}}-\epsilon_{Y}\right)^{2}\mid{\bf{X}}\right]$ $=\mathbb{E}\left\{[\mathbb{E}(Y\mid{\bf{W}})-Y]^{2}\mid{\bf{X}}\right\}$
-$=\text{Var}(Y\mid{\bf{X}})-\text{Var}[\mathbb{E}(Y\mid{\bf{W}})\mid{\bf{X}}]$
-$<\text{Var}(Y\mid{\bf{X}})\leq\mathbb{E}\left(\epsilon_{Y}^{2}\mid{\bf{X}}\right).$
-Appendix B Asymptotic properties of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$
-B.1 Asymptotic expansion of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$
-Theorem S1 (Influence function expansion of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$).
-Let $R_{i}$ be the indicator of whether the $i$th observation is labeled. Then,
-$n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{PPI}}}-\bm{\beta}^{*}\right)=n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{IF}}^{\text{\tiny{PPI}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})+o_{p}(1)$
-where ${\bf{IF}}^{\text{\tiny{PPI}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})={\bf{A}}\left\{\left[\bm{\phi}(Y_{i},{\bf{X}}_{i};\bm{\beta}^{*})\frac{R_{i}}{\pi}+\bm{\omega}^{\text{\tiny{PPI}}}(\widehat{Y}_{i},{\bf{X}}_{i},\bm{\gamma}^{*})\frac{\pi-R_{i}}{\pi}\right]\right\}$, and
-$\bm{\omega}^{\text{\tiny{PPI}}}(\widehat{Y}_{i},{\bf{X}}_{i},\bm{\gamma}^{*})=\frac{\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma}^{*})}{1-\pi}$.
-Therefore, $n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{PPI}}}-\bm{\beta}^{*}\right)$ converges weakly to a zero-mean multivariate normal distribution with variance-covariance
-$\mathbb{E}\left[{\bf{IF}}^{\text{\tiny{PPI}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})\hskip 2.84526pt{\bf{IF}}^{\text{\tiny{PPI}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})^{{\sf\scriptscriptstyle{T}}}\right].$
-Proof.
-The only difference in studying the asymptotic properties of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ compared to $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is that we rely on the asymptotic linear expansion of $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ rather than $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$.
-Using the same argument as in the proof of Theorem 1,
-$\begin{split}n^{\frac{1}{2}}\left(\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}-\bm{\gamma}^{*}\right)&=n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{A}}\left\{\left[{\bf{X}}_{i}\left(\widehat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma}^{*}\right)\right]\frac{1-R_{i}}{1-\pi}\right\}+o_{p}(1).\end{split}$
-Consequently,
-$n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{PPI}}}-\bm{\beta}^{*}\right)$ $=n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{A}}\left\{\left[\bm{\phi}(Y_{i},{\bf{X}}_{i};\bm{\beta}^{*})\frac{R_{i}}{\pi}+\frac{\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma}^{*})}{1-\pi}\frac{\pi-R_{i}}{\pi}\right]\right\}+o_{p}(1)$
-$=n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{IF}}^{\text{\tiny{PPI}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})+o_{p}(1).$
-The standard CLT implies $n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{PPI}}}-\bm{\beta}^{*}\right)$ converges weakly to a zero-mean multivariate normal with variance-covariance $\mathbb{E}\left[{\bf{IF}}^{\text{\tiny{PPI}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})\hskip 2.84526pt{\bf{IF}}^{\text{\tiny{PPI}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})^{{\sf\scriptscriptstyle{T}}}\right].$
-B.2 Comparison of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$
-Corollary S1.1 (Comparison of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ with $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$).
-The difference in the asymptotic variance-covariance of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is
-$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}})$ $=\frac{1}{\pi}{\bf{A}}\left(2\bm{\Sigma}_{Y,\widehat{Y}}-\frac{1}{1-\pi}\bm{\Sigma}_{\widehat{Y},\widehat{Y}}\right){\bf{A}}.$
-$\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is therefore more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if $2\bm{\Sigma}_{Y,\widehat{Y}}\succ\frac{1}{1-\pi}\bm{\Sigma}_{\widehat{Y},\widehat{Y}}$.
-A sufficient condition for $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ to have lower variance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ is that
-$\mathbb{E}\left[\left(\epsilon_{\widehat{Y}}-\epsilon_{Y}\right)^{2}\mid{\bf{X}}\right]<\mathbb{E}\left(\epsilon_{Y}^{2}\mid{\bf{X}}\right)+\left(1-\frac{1}{1-\pi}\right)\mathbb{E}\left(\epsilon_{\widehat{Y}}^{2}\mid{\bf{X}}\right).$
-Proof.
-From the proof for Corollary 1.1, $n[\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{lab}}})]=\frac{1}{\pi}{\bf{A}}\bm{\Sigma}_{Y,Y}{\bf{A}}.$ By Theorem S1,
-$n[\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{PPI}}})]=$ $\mathbb{E}\left\{\left[{\bf{A}}\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\right]^{\otimes 2}\frac{R^{2}}{\pi^{2}}\right\}+\mathbb{E}\left\{\left[{\bf{A}}\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})\right]^{\otimes 2}\frac{(\pi-R)^{2}}{\pi^{2}(1-\pi)^{2}}\right\}$
-$+2\mathbb{E}\left\{\left[{\bf{A}}\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\right]\left[{\bf{A}}\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})\right]^{{\sf\scriptscriptstyle{T}}}\frac{R(\pi-R)}{\pi^{2}(1-\pi)}\right\}$
-$=\frac{1}{\pi}{\bf{A}}\bm{\Sigma}_{Y,Y}{\bf{A}}+\frac{1}{\pi(1-\pi)}{\bf{A}}\bm{\Sigma}_{\widehat{Y},\widehat{Y}}{\bf{A}}-\frac{2}{\pi}{\bf{A}}\bm{\Sigma}_{Y,\widehat{Y}}{\bf{A}}$
-$=\frac{1}{\pi}{\bf{A}}\left[\bm{\Sigma}_{Y,Y}+\frac{1}{1-\pi}\bm{\Sigma}_{\widehat{Y},\widehat{Y}}-2\bm{\Sigma}_{Y,\widehat{Y}}\right]{\bf{A}}.$
-Thus,
-$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}})=\frac{1}{\pi}{\bf{A}}\left[2\bm{\Sigma}_{Y,\widehat{Y}}-\frac{1}{1-\pi}\bm{\Sigma}_{\widehat{Y},\widehat{Y}}\right]{\bf{A}}.$
-$\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is therefore more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if $2\bm{\Sigma}_{Y,\widehat{Y}}\succ\frac{1}{1-\pi}\bm{\Sigma}_{\widehat{Y},\widehat{Y}}$.
-Note that if $2\mathbb{E}\left(\epsilon_{Y}\epsilon_{\widehat{Y}}\mid{\bf{X}}\right)>\frac{1}{1-\pi}\mathbb{E}\left(\epsilon_{\widehat{Y}}^{2}\mid{\bf{X}}\right)$ or equivalently if $\mathbb{E}\left[\left(\epsilon_{\widehat{Y}}-\epsilon_{Y}\right)^{2}\mid{\bf{X}}\right]<\mathbb{E}\left(\epsilon_{Y}^{2}\mid{\bf{X}}\right)+\left(1-\frac{1}{1-\pi}\right)\mathbb{E}\left(\epsilon_{\widehat{Y}}^{2}\mid{\bf{X}}\right),$ the diagonal entries of $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}})$ are positive and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ will have a lower variance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
-∎
-B.3 Comparison of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$
-Recall from Theorem 1 and S1 that
-$n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}-\bm{\beta}^{*}\right)=n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{A}}\left\{\left[\bm{\phi}(Y_{i},{\bf{X}}_{i};\bm{\beta}^{*})\frac{R_{i}}{\pi}+\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i},\bm{\gamma}^{*})\frac{\pi-R_{i}}{\pi}\right]\right\}+o_{p}(1)\quad\mbox{and}$
-$n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{PPI}}}-\bm{\beta}^{*}\right)=n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{A}}\left\{\left[\bm{\phi}(Y_{i},{\bf{X}}_{i};\bm{\beta}^{*})\frac{R_{i}}{\pi}+\bm{\omega}^{\text{\tiny{PPI}}}(\widehat{Y}_{i},{\bf{X}}_{i},\bm{\gamma}^{*})\frac{\pi-R_{i}}{\pi}\right]\right\}+o_{p}(1).$
-From Corollary 1.1 and Corollary S1.1, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if
-$2\bm{\Sigma}_{Y,\widehat{Y}}\succ\frac{1}{1-\pi}\bm{\Sigma}_{\widehat{Y},\widehat{Y}}$
-whereas $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if
-$2\bm{\Sigma}_{Y,\widehat{Y}}\succ\bm{\Sigma}_{\widehat{Y},\widehat{Y}}.$
-Therefore, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
-Moreover, the difference between asymptotic variance-covariance can be derived from Corollary 1.1 and S1.1 as
-$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{PPI}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})=\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})-\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}})$
-$=\left(\frac{1}{\pi}-1\right){\bf{A}}\left(2\bm{\Sigma}_{Y,\widehat{Y}}-\bm{\Sigma}_{\widehat{Y},\widehat{Y}}\right){\bf{A}}-\frac{1}{\pi}{\bf{A}}\left(2\bm{\Sigma}_{Y,\widehat{Y}}-\frac{1}{1-\pi}\bm{\Sigma}_{\widehat{Y},\widehat{Y}}\right){\bf{A}}$
-$={\bf{A}}\left(\frac{2-\pi}{1-\pi}\bm{\Sigma}_{\widehat{Y},\widehat{Y}}-2\bm{\Sigma}_{Y,\widehat{Y}}\right){\bf{A}}.$
-Note that if $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is less efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ then $\frac{1}{1-\pi}\bm{\Sigma}_{\widehat{Y},\widehat{Y}}\succ 2\bm{\Sigma}_{Y,\widehat{Y}}$, implying
-$\frac{2-\pi}{1-\pi}\bm{\Sigma}_{\widehat{Y},\widehat{Y}}\succ 2\bm{\Sigma}_{Y,\widehat{Y}}.$
-This confirms that when $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is less efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, it will also be less efficient than $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$.
-A sufficient condition for $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ to be more efficient than $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is
-$\mathbb{E}[\epsilon_{Y}^{2}\mid{\bf{X}}]<\mathbb{E}[(\epsilon_{\widehat{Y}}-\epsilon_{Y})^{2}\mid{\bf{X}}]+\frac{1}{1-\pi}\mathbb{E}[\epsilon_{\widehat{Y}}^{2}\mid{\bf{X}}].$
-Appendix C Asymptotic properties of $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$
-C.1 Proof of Theorem 2
-Theorem 2 (Influence function expansion of $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$).
-Following Chen and Chen [2000],
-$n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{CC}}}-\bm{\beta}^{*}\right)=n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{IF}}^{\text{\tiny{CC}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})+o_{p}(1)$
-where ${\bf{IF}}^{\text{\tiny{CC}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})={\bf{A}}\left[\bm{\phi}(Y_{i},{\bf{X}}_{i};\bm{\beta}^{*})\frac{R_{i}}{\pi}+\bm{\omega}^{\text{\tiny{CC}}}(\widehat{Y}_{i},{\bf{X}}_{i})\frac{\pi-R_{i}}{\pi}\right]$ and $\bm{\omega}^{\text{\tiny{CC}}}(\widehat{Y},{\bf{X}})$ is defined in (1).
-Therefore, $n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{CC}}}-\bm{\beta}^{*}\right)$ converges weakly to a zero-mean multivariate normal distribution with variance-covariance
-$\hskip 2.84526pt\mathbb{E}\left[{\bf{IF}}^{\text{\tiny{CC}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})\hskip 2.84526pt{\bf{IF}}^{\text{\tiny{CC}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})^{{\sf\scriptscriptstyle{T}}}\right]\hskip 2.84526pt.$
-Proof.
-The proof follows from Chen and Chen [2000], but is detailed for completeness. Recall that
-$\begin{split}\widehat{{\bf{W}}}^{\text{\tiny{CC}}}&=\widehat{{\bf{A}}}\left\{\widehat{\text{Cov}}\left[\phi(Y,{\bf{X}};\bm{\beta}^{*}),\phi(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})\right]\right\}\{\widehat{\text{Var}}[\phi(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})]\}^{-1}\widehat{{\bf{A}}}^{-1}\\ &=\widehat{{\bf{A}}}\widehat{\bm{\Sigma}}_{Y,\widehat{Y}}\widehat{\bm{\Sigma}}^{-1}_{\widehat{Y},\widehat{Y}}\widehat{{\bf{A}}}^{-1}\end{split}$
-where
-$\begin{split}&\widehat{{\bf{A}}}=n_{\text{\tiny{lab}}}\left(\sum_{i=1}^{n_{\text{\tiny{lab}}}}{\bf{X}}_{i}{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\right)^{-1},\\ &\widehat{\bm{\Sigma}}_{Y,\widehat{Y}}=n_{\text{\tiny{lab}}}^{-1}\sum_{i=1}^{n_{\text{\tiny{lab}}}}\left(Y_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\hat{\bm{\beta}}^{\text{\tiny{lab}}}\right)\left(\hat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\hat{\bm{\gamma}}^{\text{\tiny{lab}}}\right){\bf{X}}_{i}{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}},\\ &\mbox{and }\widehat{\bm{\Sigma}}_{\widehat{Y},\widehat{Y}}=n_{\text{\tiny{lab}}}^{-1}\sum_{i=1}^{n_{\text{\tiny{lab}}}}\left(\hat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\hat{\bm{\gamma}}^{\text{\tiny{lab}}}\right)^{2}{\bf{X}}_{i}{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}.\end{split}$
-It follows that
-$n^{\frac{1}{2}}\left[\widehat{{\bf{W}}}^{\text{\tiny{CC}}}(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\bm{\gamma}^{*})\right]={\bf{A}}\bm{\Sigma}_{Y,\widehat{Y}}\bm{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\left\{n^{-\frac{1}{2}}\sum_{i=1}^{n}\left[{\bf{X}}_{i}\left(\widehat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma}^{*}\right)\right]\frac{R_{i}}{\pi}\right\}+o_{p}(1)$
-and
-$n^{\frac{1}{2}}\left[\widehat{{\bf{W}}}^{\text{\tiny{CC}}}(\hat{\bm{\gamma}}^{\text{\tiny{all}}}-\bm{\gamma}^{*})\right]={\bf{A}}\bm{\Sigma}_{Y,\widehat{Y}}\bm{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\left\{n^{-\frac{1}{2}}\sum_{i=1}^{n}\left[{\bf{X}}_{i}\left(\widehat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma}^{*}\right)\right]\right\}+o_{p}(1).$
-Consequently,
-$n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{CC}}}-\bm{\beta}^{*}\right)=$ $n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{A}}\left[\bm{\phi}(Y_{i},{\bf{X}}_{i};\bm{\beta}^{*})\frac{R_{i}}{\pi}+\bm{\omega}^{\text{\tiny{CC}}}(\widehat{Y}_{i},{\bf{X}}_{i})\frac{\pi-R_{i}}{\pi}\right]+o_{p}(1).$
-$=$ $n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{IF}}^{\text{\tiny{CC}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})+o_{p}(1).$
-The standard CLT implies $n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{CC}}}-\bm{\beta}^{*}\right)$ converges weakly to a zero-mean multivariate normal distribution with variance-covariance
-$\mathbb{E}\left[{\bf{IF}}^{\text{\tiny{CC}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})\hskip 2.84526pt{\bf{IF}}^{\text{\tiny{CC}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})^{{\sf\scriptscriptstyle{T}}}\right].$
-C.2 Proof of Corollary 2.1
-Corollary 2.1 ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$).
-The difference in the asymptotic covariance of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is
-$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{CC}}})=\left(\frac{1}{\pi}-1\right){\bf{A}}\bm{\Sigma}_{Y,\widehat{Y}}\bm{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\bm{\Sigma}_{Y,\widehat{Y}}{\bf{A}}\succeq{\bf{0}}.$
-Therefore, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is guaranteed to be at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
-Proof.
-From the proof for Corollary 1.1, $n[\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{lab}}})]=\frac{1}{\pi}{\bf{A}}\bm{\Sigma}_{Y,Y}{\bf{A}}.$ By Theorem 2.1, $n[\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{CC}}})]$ is given by
-$\begin{split}&\mathbb{E}\left\{\left[{\bf{A}}\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\right]^{\otimes 2}\frac{R^{2}}{\pi^{2}}\right\}+\mathbb{E}\left\{\left[{\bf{A}}\bm{\Sigma}_{Y,\widehat{Y}}\bm{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})\right]^{\otimes 2}\frac{(\pi-R)^{2}}{\pi^{2}}\right\}\\ +&2\mathbb{E}\left\{\left[{\bf{A}}\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\right]\left[{\bf{A}}\bm{\Sigma}_{Y,\widehat{Y}}\bm{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})\right]^{{\sf\scriptscriptstyle{T}}}\frac{R(\pi-R)}{\pi^{2}}\right\}\\ =&\frac{1}{\pi}{\bf{A}}\bm{\Sigma}_{Y,Y}{\bf{A}}+\frac{1-\pi}{\pi}{\bf{A}}\bm{\Sigma}_{Y,\widehat{Y}}\bm{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\bm{\Sigma}_{Y,\widehat{Y}}{\bf{A}}-2\frac{1-\pi}{\pi}{\bf{A}}\bm{\Sigma}_{Y,\widehat{Y}}\bm{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\bm{\Sigma}_{Y,\widehat{Y}}{\bf{A}}\\ =&\frac{1}{\pi}{\bf{A}}\bm{\Sigma}_{Y,Y}{\bf{A}}-\frac{1-\pi}{\pi}{\bf{A}}\bm{\Sigma}_{Y,\widehat{Y}}\bm{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\bm{\Sigma}_{Y,\widehat{Y}}{\bf{A}}.\end{split}$
-Hence, $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{CC}}})=\frac{1-\pi}{\pi}{\bf{A}}\bm{\Sigma}_{Y,\widehat{Y}}\bm{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\bm{\Sigma}_{Y,\widehat{Y}}{\bf{A}}\succeq 0$.
-∎
-C.3 Proof of Corollary 2.2
-Corollary 2.2 ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is the most efficient estimator in its class).
-Let the mean squared error (MSE) of $\hat{\bm{\beta}}$ for estimating $\bm{\beta}^{*}$ be defined as $\mathbb{E}[(\hat{\bm{\beta}}-\bm{\beta}^{*})^{{\sf\scriptscriptstyle{T}}}(\hat{\bm{\beta}}-\bm{\beta}^{*})]$.
-Analogously to the well-known result for the best linear predictor (e.g., Bickel and Doksum [2015], page 40), ${\bf{W}}^{\text{\tiny{CC}}}$ is the weight, ${\bf{W}}$, that minimizes the asymptotic MSE of estimators of the form $\hat{\bm{\beta}}^{\text{\tiny{lab}}}-{{\bf{W}}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right)$ for estimating $\bm{\beta}^{*}$.
-Proof.
-Letting $\hat{\bm{\beta}}_{{\bf{W}}}=\hat{\bm{\beta}}^{\text{\tiny{lab}}}-{\bf{W}}(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}})$, the MSE of estimating $\hat{\bm{\beta}}_{{\bf{W}}}$ is
-$\mathbb{Q}({\bf{W}})=\mathbb{E}\left\{\left[\hat{\bm{\beta}}^{\text{\tiny{lab}}}-{\bf{W}}(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}})-\bm{\beta}^{*}\right]^{\otimes 2}\right\}.$
-Thus, the optimal choice of ${\bf{W}}$ that minimizes the MSE is
-${\bf{W}}^{*}=\mathbb{E}[(\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\bm{\beta}^{*})(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}})]\left\{\mathbb{E}\left[(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}})^{\otimes 2}\right]\right\}^{-1}.$
-Note that the first term is the covariance between $\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\bm{\beta}^{*}$ and $\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}$, and the second term is the inverse variance-covariance of $\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}$.
-Since
-$n^{\frac{1}{2}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right)={\bf{A}}\left[n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{X}}_{i}(\widehat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma}^{*})\frac{R_{i}-\pi}{\pi}\right]+o_{p}(1),$
-we have $\text{Cov}(\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\bm{\beta}^{*},\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}})=\frac{1-\pi}{\pi}{\bf{A}}\bm{\Sigma}_{Y,\widehat{Y}}{\bf{A}}$ and $\text{Var}(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}})=\frac{1-\pi}{\pi}{\bf{A}}\bm{\Sigma}_{\widehat{Y},\widehat{Y}}{\bf{A}}$.
-Thus,
-${\bf{W}}^{*}=\frac{1-\pi}{\pi}{\bf{A}}\bm{\Sigma}_{Y,\widehat{Y}}{\bf{A}}\left[\frac{1-\pi}{\pi}{\bf{A}}\bm{\Sigma}_{\widehat{Y},\widehat{Y}}{\bf{A}}\right]^{-1}={\bf{A}}\bm{\Sigma}_{Y,\widehat{Y}}\bm{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}{\bf{A}}^{-1}={\bf{W}}^{CC}.$
+ACKNOWLEDGMENTS
+J.
+Gronsbell gratefully acknowledges funding from an NSERC Discovery Grant (RGPIN-2021-03734), the Connaught Fund, the University of Toronto Data Science Institute, and the McLaughlin Center for Genetic Research.

### v5

@@
 Abstract
-From structural biology to epidemiology, predictions from machine learning (ML) models are increasingly used to complement costly gold-standard data to enable faster, more affordable, and scalable scientific inquiry.
+From structural biology to epidemiology, predictions from machine learning (ML) models increasingly complement costly gold-standard data to enable faster, more affordable, and scalable scientific inquiry.
 In response, prediction-based (PB) inference has emerged to accommodate statistical analysis using a large volume of predictions together with a small amount of gold-standard data.
-The goals of PB inference are two-fold: (i) to mitigate bias from errors in predictions and (ii) to improve efficiency relative to traditional inference using only the gold-standard data.
-While early PB inference methods focused on bias, their ability to enhance efficiency remains unclear.
-We revisit a popular PB inference method and show that a simple modification can be applied to guarantee improvements in efficiency beyond yielding valid inferences when the ML predictions are imperfect.
-The utility of this approach in leveraging prediction-based outcomes to enhance efficiency is demonstrated through extensive simulation studies and an application to the UK Biobank data.
-We further contextualize the problem of PB inference through historical literature from economics and statistics to highlight perspectives from classical methods in this contemporary problem.
+The goals of PB inference are two-fold: (i) to mitigate bias from errors in predictions and (ii) to improve efficiency relative to classical inference using only the gold-standard data.
+While early PB inference methods focused on bias, their ability to enhance efficiency remains a focus of ongoing research.
+We revisit a foundational PB inference method and show that a simple modification can be applied to guarantee provable improvements in efficiency.
+In doing so, we establish new connections between augmented inverse probability weighted estimators (AIPW) and several recently proposed PB inference methods with a similar focus.
+The utility of our proposal, which leverages prediction-based outcomes to enhance efficiency, is demonstrated through extensive simulation studies and an application to real data from the UK Biobank.
+Further, we contextualize PB inference by drawing connections to historical literature from economics and statistics, highlighting how classic methods directly inform this contemporary problem.
+Correspondence to:
+Jessica Gronsbell
+Postal address: 700 University Ave, Toronto, ON, Canada, M5G 1Z5
+Email: j.gronsbell@utoronto.ca.
+Telephone number: 416-978-3452
 Keywords: Machine learning, measurement error, missing data, prediction-based inference, semi-parametric inference
 1 Introduction
-Researchers are increasingly turning to machine learning (ML) models to obtain predictions for outcomes that are prohibitively time-consuming or expensive to measure directly and conducting statistical analyses treating the predictions as observed data.
+Researchers are increasingly turning to machine learning (ML) models to obtain predictions for outcomes that are prohibitively time-consuming or expensive to measure directly and then conducting statistical analyses which treat the predictions as observed data.
 1, 2, 3, 4 This prediction-based (PB) inference approach, introduced by Wang et al,5 is comprised of two steps outlined in Figure 1.
@@
 Recent work has considered predicted covariates (e.g., Kluger et al6).
-In contrast to traditional inference that relies solely on gold-standard data, PB inference aims to leverage the predictions to improve statistical efficiency without introducing bias.
+In contrast to classical inference that relies solely on gold-standard data, PB inference aims to leverage the predictions to improve statistical efficiency without introducing bias.
 In the last several years, applications of PB inference have arisen in diverse fields, spanning genetics, biology, and medicine.
-For example, population biobanks, such as the UK Biobank ($n\approx$ 500K individuals), collect high-throughput genomic data on large populations and offer unprecedented opportunities for genetic discovery research.
-However, many traits and disease outcomes are only partially observed due to the time and cost of measurement.
-ML models are commonly used to impute missing outcomes based on fully observed baseline data to enable downstream association studies.
-7, 8, 9, 10, 11, 12, 13, 14 Similarly, in structural biology, AlphaFold provides structural predictions without the need for expensive and labor-intensive experiments for proteomic studies.
-15, 16 In the analysis of electronic health records data, large language models can rapidly ascertain disease status at an accuracy competitive with manual chart review for epidemiological and clinical studies.
+For example, population biobanks, such as the UK Biobank ($n\approx$ 500K individuals), collect high-throughput genomic data on large populations and offer unprecedented opportunities for genetic discovery.
+However, many phenotypes are only partially observed due to the time and expense of measurement.
+ML models are frequently used to impute missing outcomes based on fully observed baseline data to enable downstream association studies with phenotypes of interest.
+7, 8, 9, 10, 11, 12, 13, 14 Similarly, in structural biology, AlphaFold provides protein conformation predictions without the need for expensive and labor-intensive experiments.
+15, 16 In the analysis of electronic health records data, large language models can rapidly ascertain disease status at an accuracy competitive with manual chart review by trained annotators, facilitating clinical and epidemiological studies.
 17, 18, 19
@@
 The focus of scientific inquiry is on the association between $Y$ and $\bf{X}$}.
-Traditional statistical inference relies solely on the data on $Y$ and ${\bf{X}}$ while PB inference incorporates $\widehat{Y}$ to make use of all of the available data.
+Classical statistical inference relies solely on the data on $Y$ and $\bf{X}$} while PB inference incorporates $\widehat{Y}$ to make use of all of the available data.
 Dark red represents labeled observations while light red represents unlabeled observations.
@@
 In the context of least squares estimation, they showed that the post-prediction inference (PostPi) method proposed by Wang et al5 provides a consistent estimator for the association between an outcome and a set of covariates only in restrictive scenarios (e.g., when the pre-trained ML model perfectly captures the underlying regression function).
-This is a significant practical limitation as it is essential that PB inference methods yield valid inference (i.e., hypothesis tests with controlled type I error and confidence intervals with nominal coverage rates) in the likely scenario that the prediction model is imperfect.
-In contrast, the prediction powered inference (PPI) method proposed by Angelopoulos et al 21 yields consistent estimation and valid inference regardless of the quality of the ML model.
+This is a significant practical limitation.
+PB inference methods should guarantee valid inference (i.e., hypothesis tests with controlled type I error and confidence intervals with nominal coverage rates) in the common scenario that the prediction model is imperfect.
+In contrast to PostPi, the prediction powered inference (PPI) method proposed by Angelopoulos et al 21 yields consistent estimation and valid inference regardless of the quality of the ML model.
 This finding is critical given that the PPI paper, published just over a year ago in the high-impact journal Science, has already been cited nearly 150 times and widely adopted in applications.
-In this paper, we build on the work of Motwani and Witten 20 and further examine the PPI estimator by assessing its statistical efficiency.
-Our contributions are two-fold.
-First, we develop a set of results that characterize when the PPI estimator gains over traditional estimators using labeled data only.
-We also show that a more efficient estimator can be obtained by adding a simple weight to the PPI estimator.
-This approach was proposed by Chen & Chen (CC) over 25 years ago in the context of double sampling designs.22 Second, we revisit literature from statistics and economics and identify work dating back to the 1960s that bears directly on current developments in PB inference.
-Our review puts classical statistical methods in conversation with this modern problem to inform future methodological advances and applications.
-We begin by formally introducing the PB inference problem and reviewing the contributions of Motwani and Witten20 in Section 2.
-We then study the asymptotic properties of the PPI estimator, showing that it can be outperformed by traditional inference methods, and introduce the CC estimator in Section 3.
-Our key finding is that the CC estimator is optimal with respect to statistical efficiency in a class of estimators to which the PPI estimator belongs.
-In Section 4, we discuss related literature and provide connections between the CC estimator and other established methods.
+In this paper, we further study the PPI estimator by assessing its statistical efficiency.
+Our contributions are three-fold.
+First, we make new contributions to the literature identifying inefficiencies of the PPI method.
+22, 23, 24 We derive a set of results that characterize when the PPI estimator gains over classical estimation using labeled data only in the context of Z-estimation.
+Second, we show that a more efficient estimator can be obtained by adding a simple weight to the PPI estimator.
+This approach was proposed by Chen & Chen (CC) over 25 years ago in the context of double sampling designs 25 and its application within the PB inference setting is new.
+Third, we draw important connections between augmented inverse probability weighted (AIPW) estimators, our proposal, and several recently proposed methods.
+In doing so, we identify settings in which the CC estimator is theoretically guaranteed to improve over comparable approaches.
+Building on these findings, we also revisit literature from statistics and economics to identify work dating back to the 1960s that bears directly on current developments in PB inference.
+Overall, our contributions put classical statistical methods in conversation with this modern problem to inform future methodological developments and applications.
+We begin by formally introducing the PB inference problem, reviewing the contributions of Motwani and Witten20 and summarizing existing work in Section 2.
+We then study the asymptotic properties of the PPI estimator, showing that it can be outperformed by classical inference methods and introduce the CC estimator in Section 3.
+Our key finding is that the CC estimator is optimal with respect to statistical efficiency in the class of estimators to which the PPI estimator belongs.
+In Section 4, we discuss related literature and provide connections between the CC estimator, other PB inference methods, and AIPW estimators.
+We also develop a set of results that compare the efficiency of the CC estimator with several PB inference methods and identify scenarios where the CC estimator is theoretically optimal.
 We then evaluate our findings with extensive simulation studies in Section 5.
@@
 2.1 Notation and Assumptions
-We let $Y$ denote the outcome of interest and ${\bf{Z}}$ a set of predictors.
-$Y$ is only partially observed and a pre-trained ML model is available to predict it over the full data given ${\bf{Z}}$, which are fully observed.
+We let $Y$ denote the outcome of interest and $\bf{Z}$} its predictors.
+$Y$ is only partially observed and a pre-trained ML model is available to predict it over the full data given $\bf{Z}$}, which is fully observed.
 The ML model $\hat{f}:\mathcal{Z}\rightarrow\mathcal{Y}$ maps the predictor space $\mathcal{Z}$ to the outcome space $\mathcal{Y}$ and produces predictions $\widehat{Y}=\hat{f}({\bf{Z}})$.
@@
 The available data therefore consists of a labeled dataset and an independent unlabeled dataset, respectively denoted as
-=$\mathscr{L}=\{(Y_{i},{\bf{Z}}_{i}^{{\sf\scriptscriptstyle{T}}},{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}})^{{\sf\scriptscriptstyle{T}}}\mid i=1,\dots,n_{\text{\tiny{lab}}}\}\quad\mbox{and}\quad\mathscr{U}=\{({\bf{Z}}_{i}^{{\sf\scriptscriptstyle{T}}},{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}})^{{\sf\scriptscriptstyle{T}}}\mid i=n_{\text{\tiny{lab}}}+1,\dots,n\},$
-where $n=n_{\text{\tiny{lab}}}+n_{\text{\tiny{unlab}}}$, $n_{\text{\tiny{unlab}}}$ is the number of unlabeled observations, and $n_{\text{\tiny{lab}}}/n\to\pi$ as $n\to\infty$ for $\pi\in(0,1)$.
+$\mathscr{L}=\{(Y_{i},{\bf{Z}}_{i}^{{\sf\scriptscriptstyle{T}}},{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}})^{{\sf\scriptscriptstyle{T}}}\mid i=1,\dots,n_{\text{\tiny{lab}}}\}\quad\mbox{and}\quad\mathscr{U}=\{({\bf{Z}}_{i}^{{\sf\scriptscriptstyle{T}}},{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}})^{{\sf\scriptscriptstyle{T}}}\mid i=n_{\text{\tiny{lab}}}+1,\dots,n\},$
+where $n=n_{\text{\tiny{lab}}}+n_{\text{\tiny{unlab}}}$, $n_{\text{\tiny{unlab}}}$ is the number of unlabeled observations and $n_{\text{\tiny{lab}}}/n\to\pi$ as $n\to\infty$ for $\pi\in(0,1)$.
 We focus on the case in which $\mathscr{L}$ and $\mathscr{U}$ include independent and identically distributed samples from a common population of interest.
@@
 $\hat{f}$ may be a black box function in that its training data and specification are not be available to the users.
-Approaches that address uncertainty in the estimation of $\hat{f}$ have recently been considered.23
+Approaches that address uncertainty in the estimation of $\hat{f}$ have recently been considered.26
 2.2 Parameter of interest
-Following Motwani and Witten,20 we consider least squares estimation and aim to make inferences on the population parameter
-$\bm{\beta}^{*}\coloneq\operatorname*{arg\,min}_{\bm{\beta}}\mathbb{E}[(Y-{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\beta})^{2}]=[\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})]^{-1}\mathbb{E}({\bf{X}}Y),$
-which quantifies the association between $Y$ and ${\bf{X}}$.
-The simplest consistent estimator of $\bm{\beta}^{*}$, which we denote by $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, uses only $\mathscr{L}$ for estimation and solves the estimating equation
-$\sum_{i=1}^{n_{\text{\tiny{lab}}}}\bm{\phi}(Y_{i},{\bf{X}}_{i};\bm{\beta})\coloneq\sum_{i=1}^{n_{\text{\tiny{lab}}}}{\bf{X}}_{i}(Y_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\bm{\beta})=\bm{0}.$
+We focus on Z-estimation and aim to make inferences on the population parameter $\bm{\beta}^{*}$ that is the solution to the estimating equation
+$\mathbb{E}[\bm{\phi}(Y,{\bf{X}};\bm{\beta})]=\bm{0}.$
+$\bm{\phi}(\cdot)$ is a deterministic function that satisfies conditions C1-C4 in the Appendix.
+The simplest consistent estimator of $\bm{\beta}^{*}$, which we denote by $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, uses only $\mathscr{L}$ for estimation and is the solution to the estimating equation
+$\sum_{i=1}^{n_{\text{\tiny{lab}}}}\bm{\phi}(Y_{i},{\bf{X}}_{i};\bm{\beta})=\bm{0}.$
 2.3 Motivation for PB inference
@@
 However, relying on $\widehat{Y}$ for inference may result in biased estimation depending on the quality of predictions from $\hat{f}$.
-For example, in “naive” inference, $\widehat{Y}$ is directly substituted for $Y$ in downstream analysis.5 The parameter targeted by this approach is
-$\bm{\gamma}^{*}\coloneq\operatorname*{arg\,min}_{\bm{\gamma}}\mathbb{E}[(\widehat{Y}-{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma})^{2}]=[\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})]^{-1}\mathbb{E}({\bf{X}}\widehat{Y})$
-and the corresponding least squares estimator, denoted as $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$, is the solution to
-$\sum_{i=1}^{n}\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma})\coloneq\sum_{i=1}^{n}{\bf{X}}_{i}(\widehat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma})=\bm{0}.$
-As the distribution of $(\widehat{Y},{\bf{X}}^{{\sf\scriptscriptstyle{T}}})^{{\sf\scriptscriptstyle{T}}}$ may not coincide with that of $(Y,{\bf{X}}^{{\sf\scriptscriptstyle{T}}})^{{\sf\scriptscriptstyle{T}}}$, $\bm{\gamma}^{*}$ bears no necessary correspondence with $\bm{\beta}^{*}$ except under restrictive scenarios (e.g., when ${\bf{Z}}$ contains ${\bf{X}}$ and the ML model captures the true regression function).
+For example, in “naive” inference, $\widehat{Y}$ is directly substituted for $Y$ in downstream analysis.5 The parameter $\bm{\gamma}^{*}$ targeted by this approach solves
+$\mathbb{E}[\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma})]=\bm{0}.$
+The corresponding estimator based on $\mathscr{L}$ and $\mathscr{U}$, denoted as $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$, is the solution to
+$\sum_{i=1}^{n}\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma})=\bm{0}.$
+As the distribution of $(\widehat{Y},{\bf{X}}^{{\sf\scriptscriptstyle{T}}})^{{\sf\scriptscriptstyle{T}}}$ may not coincide with that of $(Y,{\bf{X}}^{{\sf\scriptscriptstyle{T}}})^{{\sf\scriptscriptstyle{T}}}$, $\bm{\gamma}^{*}$ bears no necessary correspondence with $\bm{\beta}^{*}$ except under restrictive scenarios.
+For example, $\bm{\gamma}^{*}$ coincides with $\bm{\beta}^{*}$ when $\bf{Z}$} contains $\bf{X}$}, the ML model captures the true regression function $\mathbb{E}(Y\mid{\bf{Z}})$, and the focus of inference is on a parameter from a linear regression model.
 To overcome these limitations, PB inference methods utilize $\mathscr{L}$ and $\mathscr{U}$ to essentially balance the relative contributions of $Y$ and $\widehat{Y}$ in making inference on $\bm{\beta}^{*}$.
 2.4 PB inference methods
-Motwani and Witten20 revisited two seminal strategies for PB inference.
-The first is the PostPI method of Wang et al,5 which is a two-step procedure to incorporate $\widehat{Y}$ into estimation.
+In the context of least squares estimation, Motwani and Witten20 revisited two seminal strategies for PB inference.
+One of these is the PostPI method of Wang et al,5 which was the first modern method proposed for PB inference.
+PostPI consists of two steps to incorporate $\widehat{Y}$ into inference for $\bm{\beta}^{*}$.
 First, $\mathscr{L}$ is used to model the relationship between $Y$ and $\widehat{Y}$ to address potential inaccuracies in the predictions.
-Then, this model is used together with $\mathscr{U}$ to make inference on $\bm{\beta}^{*}$ with either analytical or bootstrap approaches.
-This approach is shown to yield a consistent estimator for $\bm{\beta}^{*}$ under the strong and unrealistic assumption that the ML model captures the true regression function.
+Then, this model is used together with $\mathscr{U}$ to make inference on $\bm{\beta}^{*}$ with either analytical or bootstrap methods.
+This approach was shown to yield a consistent estimator for $\bm{\beta}^{*}$ under the strong and arguably unrealistic assumption that the ML model captures the true regression function.
 When this assumption is not met, the PostPI estimator may not be consistent.
 Consequently, corresponding hypothesis tests may not have controlled type I error rates and confidence intervals may not have adequate coverage.
-Motwani and Witten20 also considered the PPI method,21 which has the crucial property that inference remains valid regardless of the quality of the predictions.
-That is, the PPI method does not require that $\hat{f}$ capture the true regression function for the estimator to maintain consistency for $\bm{\beta}^{*}$.
-The proposed estimator utilizes a common statistical technique involving the augmentation of an initial consistent estimator for the parameter of interest with an additional term that converges in probability to 0.
-This approach dates back to at least Robins et al 24 and has been applied in multiple settings to improve estimation efficiency outside of the PB inference problem.
-22, 25, 26, 27 We return to this point in our review of related literature in Section 4.
-The PPI estimator is
+Motwani and Witten20 also studied the PPI method 21 which was proposed a few years after PostPI.
+Unlike its predecessor, PPI has the crucial property that inference remains valid regardless of the quality of the predictions.
+That is, $\hat{f}$ need not capture the true regression function for the estimator to maintain consistency for $\bm{\beta}^{*}$.
+The PPI estimator utilizes a common statistical technique that augments an initial consistent estimator for the target parameter based solely on $\mathscr{L}$ with an additional term that is estimable from both $\mathscr{L}$ and $\mathscr{U}$.
+The augmentation term converges in probability to zero in order to improve efficiency without introducing bias.
+This approach dates back to at least Robins et al 27 and has been applied in multiple settings to improve estimation efficiency outside of the PB inference problem.
+25, 28, 29, 30, 31 We return to this point in our review of related literature in Section 4.
+Formally, the PPI estimator is
 $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}=\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}\right)$
@@
 In contrast to the PostPI method, Motwani and Witten20 show that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is consistent for $\bm{\beta}^{*}$ and $n^{\frac{1}{2}}(\hat{\bm{\beta}}^{\text{\tiny{PPI}}}-\bm{\beta}^{*})$ converges weakly to a zero-mean multivariate normal distribution regardless of the quality of the prediction.
+Analogous results for Z-estimation problems are established in the original PPI paper (e.g., see Theorem S1 in the Supplementary Materials).
 This approach therefore asymptotically controls the type I error rate in hypothesis tests and produces confidence intervals with the nominal coverage level.
+However, for a PB inference method to be preferred in practice over classical inference based on $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, it should not only yield valid inference, but also be more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when $\widehat{Y}$ is informative of $Y$ and equally efficient when $\widehat{Y}$ is uninformative of $Y$.
+Two recent works have empirically identified that the PPI method can be less efficient than classical inference and have proposed extensions of the original method that are theoretically guaranteed to outperform classical inference.
+The first such extension, called PPI++, abandons the augmentation-based approach of PPI and instead directly augments the estimating equation for $\bm{\beta}^{*}$ with information from $\widehat{Y}$.
+A tuning parameter is used that controls the amount of information from $\widehat{Y}$ that is used for estimation and is cleverly designed to minimize the trace of the resulting estimator’s asymptotic covariance matrix.
+Subsequently, Miao et al.
+23 proposed Post-Prediction Adaptive inference (PSPA), which adopts a similar strategy, but instead produces an estimator with minimal element-wise variance.
+While both PPI++ and PSPA offer provable improvements over classical inference using only $\mathscr{L}$, their relationships to the original PPI framework remain underexplored.
+Furthermore, the conditions under which these methods are optimal have yet to be fully established.
+We revisit these proposals and clarify some of these connections in Section 4.
 3 Efficiency analysis
-As PB inference methods suppose the availability of at least a small labeled data set, a natural baseline for comparison is the estimator based on $\mathscr{L}$ only, $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
-For a PB inference method to be preferred over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, it should not only yield valid inference, but also achieve gains in statistical efficiency relative to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when the prediction is informative of $Y$ and perform no worse when $\widehat{Y}$ is uninformative.
-For the remainder of this paper, we study the latter property and analyze a modified version of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ given by
+For the remainder of this paper, we study a modified version of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ given by
 $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}=\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right).$
@@
 When $\pi\approx 0$, $\hat{\bm{\gamma}}^{\text{\tiny{all}}}\approx\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ so that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}\approx\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$.
-In general, it can be shown that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ gains efficiency over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ whenever $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ gains over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ can be more efficient than $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ when $\widehat{Y}$ is informative of $Y$ (see Appendix, specifically Section B.3, for a detailed analysis).
+In general, it can be shown that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ gains efficiency over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ whenever $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ gains over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ can be more efficient than $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ when $\widehat{Y}$ is informative of $Y$ (see Appendix, specifically Section B.3, for a detailed analysis).
 Additionally, estimators of this form are commonly used in the semi-parametric inference literature, which we draw upon in the subsequent section in our asymptotic analysis.
+We provide general results for Z-estimation and use linear regression as a running example throughout our discussion to further cultivate intuition.
 3.1 Asymptotic analysis of the PPI estimator
-To study the asymptotic propeties of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, we first derive its influence function expansion in Theorem 1. The detailed derivations are deferred to the Appendix.
+To study the asymptotic properties of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, we present its influence function expansion in Theorem 1.
+The detailed derivations are deferred to the Appendix and follow those provided for the original PPI estimator.
 Theorem 1 (Influence function expansion of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$).
@@
 $n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}-\bm{\beta}^{*}\right)=n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})+o_{p}(1),$
-where ${\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})={\bf{A}}\left\{\left[\ \bm{\phi}(Y_{i},{\bf{X}}_{i};\bm{\beta}^{*})\right]\frac{R_{i}}{\pi}+\left[\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma}^{*})\right]\frac{\pi-R_{i}}{\pi}\right\}$ and ${\bf{A}}=[\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})]^{-1}$.
+where ${\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})={\bf{A}}\left\{\left[\ \bm{\phi}(Y_{i},{\bf{X}}_{i};\bm{\beta}^{*})\right]\frac{R_{i}}{\pi}+\left[\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma}^{*})\right]\frac{\pi-R_{i}}{\pi}\right\}$ and ${\bf{A}}=\mathbb{E}\left[\frac{\partial\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})}{\partial\bm{\beta}}\right]^{-1}$.
 Therefore, $n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}-\bm{\beta}^{*}\right)$ converges weakly to a zero-mean multivariate normal distribution with variance-covariance $\mathbb{E}\left[{\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})\hskip 2.84526pt{\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})^{{\sf\scriptscriptstyle{T}}}\right].$
-We use Theorem 1 to compare the asymptotic variance-covariance of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ with that of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
-Corollary 1.1 shows that the benefit of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ scales directly with the predictiveness of $\widehat{Y}$ for $Y$ and inversely with the proportion of labeled data $\pi$.
+We use Theorem 1 to compare the asymptotic variance of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ with that of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
+The result in Corollary 1.1 shows that the improvement of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ increases with the proportion of unlabeled data provided that $\widehat{Y}$ is of sufficient quality.
 Corollary 1.1 (Comparison of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ with $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$).
+Let ${\bf{C}}_{11}=\mathbb{E}\left[\bm{\phi}(Y,X;\bm{\beta}^{*})\bm{\phi}(Y,X;\bm{\beta}^{*})^{{\sf\scriptscriptstyle{T}}}\right]$, ${\bf{C}}_{12}=\mathbb{E}\left[\bm{\phi}(Y,X;\bm{\beta}^{*})\bm{\phi}(\widehat{Y},X;\bm{\beta}^{*})^{{\sf\scriptscriptstyle{T}}}\right]$, ${\bf{C}}_{22}=\mathbb{E}\left[\bm{\phi}(\widehat{Y},X;\bm{\gamma}^{*})\bm{\phi}(\widehat{Y},X;\bm{\gamma}^{*})^{{\sf\scriptscriptstyle{T}}}\right]$, and ${\bf{B}}=\mathbb{E}\left[\frac{\partial\bm{\phi}(\widehat{Y},X;\bm{\gamma}^{*})}{\partial\bm{\gamma}}\right]^{-1}$.
+The difference in the asymptotic variance-covariance (aVar) of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is
+$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})$ $=n\left[\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{lab}}})-\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})\right]=\left(\frac{1}{\pi}-1\right)\left(2{\bf{A}}{\bf{C}}_{12}-{\bf{B}}{\bf{C}}_{22}\right){\bf{B}}.$
+$\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is therefore more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if $2{\bf{A}}{\bf{C}}_{12}\succ{\bf{B}}{\bf{C}}_{22}$.
+We further break down this result in the case of a linear regression model.
+Example 1 (Linear Regression).
 Let $\epsilon_{Y}=Y-{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\beta}^{*}$, $\epsilon_{\widehat{Y}}=\widehat{Y}-{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma}^{*}$, $\bm{\Sigma}_{\widehat{Y},\widehat{Y}}=\mathbb{E}\left(\epsilon_{\widehat{Y}}^{2}{\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\right)$, and $\bm{\Sigma}_{Y,\widehat{Y}}=\mathbb{E}\left(\epsilon_{Y}\epsilon_{\widehat{Y}}{\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\right)$.
-The difference in the asymptotic variance-covariance (aVar) of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is
+Following Corollary 1.1, the difference in the asymptotic variance-covariance (aVar) of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is
 $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})$ $=n\left[\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{lab}}})-\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})\right]=\left(\frac{1}{\pi}-1\right){\bf{A}}\left(2\bm{\Sigma}_{Y,\widehat{Y}}-\bm{\Sigma}_{\widehat{Y},\widehat{Y}}\right){\bf{A}}.$
 $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is therefore more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if $2\bm{\Sigma}_{Y,\widehat{Y}}\succ\bm{\Sigma}_{\widehat{Y},\widehat{Y}}$.
-A sufficient condition for $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ to have lower variance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ is that
+Moreover, a sufficient condition for $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ to have lower variance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ is that
 $\mathbb{E}\left[\left(\epsilon_{\widehat{Y}}-\epsilon_{Y}\right)^{2}\mid{\bf{X}}\right]<\mathbb{E}\left(\epsilon_{Y}^{2}\mid{\bf{X}}\right).$
-The sufficient condition in Corollary 1.1 coincides with the intuition that incorporating unlabeled data into estimation can improve efficiency when $\widehat{Y}$ has additional predictive power for $Y$ beyond what is already explained by the regression model of $Y$ on ${\bf{X}}$.
+This condition coincides with the intuition that incorporating unlabeled data into estimation can improve efficiency when $\widehat{Y}$ has additional predictive power for $Y$ beyond what is already explained by the regression model of $Y$ on $\bf{X}$}.
 This condition is similar to the result for mean estimation provided in the Supplementary Materials of Angelopoulos et al 21 and can be used to describe several scenarios when $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is guaranteed to have lower asymptotic variance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when the ML model captures the true regression function (see Appendix, Section A.4).
-Here we consider a simple example to demonstrate when $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ gains efficiency over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ under homoskedasticity.
-Example 1 (Efficiency gain of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ under homoskedasticity).
-In the homoskedastic setting where $\mathbb{E}(\epsilon_{Y}^{2}\mid{\bf{X}})=\sigma_{Y}^{2}$, $\mathbb{E}(\epsilon_{\widehat{Y}}^{2}\mid{\bf{X}})=\sigma_{\widehat{Y}}^{2}$, and $\mathbb{E}(\epsilon_{Y}\epsilon_{\widehat{Y}}\mid{\bf{X}})=\sigma_{Y,\widehat{Y}}$,
+This insight can be better understood in the case of homoskedasticity.
+Specifically, suppose that $\mathbb{E}(\epsilon_{Y}^{2}\mid{\bf{X}})=\sigma_{Y}^{2}$, $\mathbb{E}(\epsilon_{\widehat{Y}}^{2}\mid{\bf{X}})=\sigma_{\widehat{Y}}^{2}$, and $\mathbb{E}(\epsilon_{Y}\epsilon_{\widehat{Y}}\mid{\bf{X}})=\sigma_{Y,\widehat{Y}}$,
 $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})=\left(\frac{1}{\pi}-1\right)\left[2\sigma_{Y,\widehat{Y}}-\sigma_{\widehat{Y}}^{2}\right]{\bf{A}}.$
@@
 $\rho_{Y,\widehat{Y}}>\frac{1}{2}\frac{\sigma_{\widehat{Y}}}{\sigma_{Y}},$
-where $\rho_{Y,\widehat{Y}}=\sigma_{Y,\widehat{Y}}/\sigma_{Y}\sigma_{\widehat{Y}}$ is the correlation of the residuals for $Y$ and $\widehat{Y}$ given ${\bf{X}}$.
-Thus $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ gains efficiency over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when $\widehat{Y}$ is a strong predictor for $Y$ after adjusting for ${\bf{X}}$, though $\widehat{Y}$ does not necessarily need to be derived from a ML model that captures the true regression function.
+where $\rho_{Y,\widehat{Y}}=\sigma_{Y,\widehat{Y}}/\sigma_{Y}\sigma_{\widehat{Y}}$ is the partial correlation of the residuals for $Y$ and $\widehat{Y}$ given $\bf{X}$}.
+Thus, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ gains efficiency over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when $\widehat{Y}$ is a strong predictor for $Y$ after adjusting for $\bf{X}$}, though $\widehat{Y}$ does not necessarily need to be derived from a ML model that captures the true regression function.
 The improvement is stronger with increasing correlation of the residuals between $Y$ and $\widehat{Y}$ and decreasing variability in the residuals of $\widehat{Y}$.
 3.2 Asymptotic analysis of the CC estimator
-To further analyze the efficiency of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, we draw on the landmark paper that introduced a class of augmented inverse probability weighted (AIPW) estimators which includes all regular asymptotic linear estimators.
-24 In the current setting, where the goal is to estimate $\bm{\beta}^{*}$ with $Y$ partially observed and $\widehat{Y}$ fully observed, these estimators have influence functions of the form
-$\text{IF}(Y,\widehat{Y},{\bf{X}};\bm{\omega})={\bf{A}}\left[\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\frac{R}{\pi}+\bm{\omega}(\widehat{Y},{\bf{X}})\frac{\pi-R}{\pi}\right],$
-with $\bm{\omega}(\widehat{Y},{\bf{X}})$ being a function of $\widehat{Y}$ and ${\bf{X}}$. The optimal choice of $\bm{\omega}(\widehat{Y},{\bf{X}})$ with respect to asymptotic variance is
+To further analyze the efficiency of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, we draw on the landmark paper that introduced a class of augmented inverse probability weighted (AIPW) estimators which includes all regular asymptotic linear estimators.27 While connections between the PPI estimator and AIPW have been made in the context of mean estimation21, we explictly leverage this framework to introduce a simple modification of the original PPI estimator with theoretical guarantees of improved efficiency.
+In the current setting, where the goal is to estimate $\bm{\beta}^{*}$ with $Y$ partially observed and $\widehat{Y}$ fully observed, AIPW estimators have influence functions of the form
+${\bf{IF}}(Y,\widehat{Y},{\bf{X}};\bm{\omega})={\bf{A}}\left[\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\frac{R}{\pi}+\bm{\omega}(\widehat{Y},{\bf{X}})\frac{\pi-R}{\pi}\right],$
+with $\bm{\omega}(\widehat{Y},{\bf{X}})$ being a function of $\widehat{Y}$ and $\bf{X}$}. The optimal choice of $\bm{\omega}(\widehat{Y},{\bf{X}})$ with respect to asymptotic variance27 is
 $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})=\mathbb{E}[\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\mid\widehat{Y},{\bf{X}}].$
-By Theorem 1, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is thus an AIPW estimator and can be seen as an approximation to the optimal estimator with $\bm{\omega}(\widehat{Y},{\bf{X}})=\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})$.
-The PPI estimator is therefore not the most efficient estimator in its class and, from Corollary 1.1, can even be less efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when $\bm{\Sigma}_{\widehat{Y},\widehat{Y}}\succ 2\bm{\Sigma}_{Y,\widehat{Y}}$.
-A PB inference strategy should enhance statistical efficiency relative to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ whenever possible and, at the very least, should not be negatively impacted by incorporating unlabeled data into estimation.
-This raises concerns about the practical utility of the PPI estimator.
+By Theorem 1, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is an AIPW estimator and can be seen as an approximation to the optimal estimator with $\bm{\omega}(\widehat{Y},{\bf{X}})=\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})$.
+This is a good approximation, for example, when $\bf{Z}$} contains $\bf{X}$}, the ML model captures the true regression function, and the focus is in fitting a generalized linear model with the canonical link function, $g(\cdot)$.
+In this case, $\bm{\phi}(Y,X;\bm{\beta})={\bf{X}}[Y-g(\bm{\beta}^{{\sf\scriptscriptstyle{T}}}{\bf{X}})]$ and $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})=\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})$.
+However, this is also a scenario where naive inference will yield consistent estimation that is more efficient than inference based on the PPI estimator.
+More broadly, the PPI estimator is not the most efficient estimator in its class and, from Corollary 1.1, can even be less efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when the quality of the prediction is insufficient.
+A PB inference strategy should enhance statistical efficiency relative to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ whenever possible and, at the very least, should not lose efficiency from incorporating unlabeled data into estimation.
 To improve upon $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, one approach is to assume a parametric model for $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$ that can be estimated from $\mathscr{L}$ for analysis.
@@
 Moreover, when the model for $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$ is misspecified, an efficiency gain is not guaranteed.
-To overcome these difficulties, Chen and Chen 22 proposed a particular choice of $\bm{\omega}(\widehat{Y},{\bf{X}})$ that is guaranteed to be at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and that does not require modeling of $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$.
+To overcome these difficulties, Chen and Chen 25 proposed a particular choice of $\bm{\omega}(\widehat{Y},{\bf{X}})$ that is guaranteed to be at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and that does not require modeling of $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$.
 Their approach is motivated by the double-sampling design which, apart from differences in terminology (i.e., they regard $\widehat{Y}$ as a mismeasured version of $Y$), is analogous to the partially labeled data setting considered here.
@@
 $\bm{\omega}^{\text{\tiny{CC}}}(\widehat{Y},{\bf{X}})=\text{Cov}[\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*}),\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})]\{\text{Var}[\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})]\}^{-1}\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*}).$ (1)
-The corresponding estimator can be written as
+Relating this estimator back to the PPI estimator, the high-level idea is analogous, but $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$ is approximated by the linear projection of $\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})$ on $\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})$ rather than $\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})$ itself, which effectively protects against potential inadequacies in $\widehat{Y}$ that can lead to efficiency loss.
+The corresponding estimator, which we call the CC estimator, can be written as
 $\hat{\bm{\beta}}^{\text{\tiny{CC}}}=\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\widehat{{\bf{W}}}^{\text{\tiny{CC}}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right),$
 where $\widehat{{\bf{W}}}^{\text{\tiny{CC}}}$ is a consistent estimator of
-${{\bf{W}}}^{\text{\tiny{CC}}}={\bf{A}}{\text{Cov}}[\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*}),\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})]\{{\text{Var}}[\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})]\}^{-1}{\bf{A}}^{-1}$
-obtained with $\mathscr{L}$.
+${{\bf{W}}}^{\text{\tiny{CC}}}={\bf{A}}{\text{Cov}}[\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*}),\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})]\{{\text{Var}}[\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})]\}^{-1}{\bf{B}}^{-1}$
+obtained with $\mathscr{L}$ and $\mathscr{U}$.
 Therefore, a simple way to ensure that an estimator with the same structure as the PPI estimator has asymptotic variance no greater than that of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ is to add a weight function in front of the augmentation term.
@@
 Theorem 2 (Influence function expansion of $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$).
-Following Chen and Chen 22,
+Following Chen and Chen 25,
 $n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{CC}}}-\bm{\beta}^{*}\right)=n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{IF}}^{\text{\tiny{CC}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})+o_{p}(1),$
@@
 Therefore, $n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{CC}}}-\bm{\beta}^{*}\right)$ converges weakly to a zero-mean multivariate normal distribution with variance-covariance $\hskip 2.84526pt\mathbb{E}\left[{\bf{IF}}^{\text{\tiny{CC}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})\hskip 2.84526pt{\bf{IF}}^{\text{\tiny{CC}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})^{{\sf\scriptscriptstyle{T}}}\right]\hskip 2.84526pt.$
-The subsequent corollaries apply Theorem 2 to show our two key results: (1) $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and (2) $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ has the highest asymptotic efficiency within the class of estimators with a weighted linear augmentation term.
-These results suggest that the CC method is preferable to the PPI method for gaining efficiency.
+The subsequent corollaries apply Theorem 2 to show our two key results:(1) $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and (2) $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is optimal with respect to efficiency in the class of estimators with a weighted linear augmentation term.
+These results are formally stated in Corollaries 2.1 and 2.2 and show that the CC method is preferable to the PPI method.
 Corollary 2.1 ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$).
 The difference in the asymptotic variance-covariance of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is
-$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{CC}}})=\left(\frac{1}{\pi}-1\right){\bf{A}}\bm{\Sigma}_{Y,\widehat{Y}}\bm{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\bm{\Sigma}_{Y,\widehat{Y}}{\bf{A}}\succeq{\bf{0}}.$
+$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{CC}}})=\left(\frac{1}{\pi}-1\right){\bf{A}}{\bf{C}}_{12}{\bf{C}}_{22}^{-1}{\bf{C}}_{12}{\bf{B}}\succeq{\bf{0}}.$
 Therefore, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is guaranteed to be at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
@@
 Let the mean squared error (MSE) of an estimator $\hat{\bm{\beta}}$ for $\bm{\beta}^{*}$ be defined as $\mathbb{E}[(\hat{\bm{\beta}}-\bm{\beta}^{*})^{{\sf\scriptscriptstyle{T}}}(\hat{\bm{\beta}}-\bm{\beta}^{*})]$.
-Analogously to the well-known result for the best linear predictor (e.g.,Bickel and Doksum,28 page 40), ${\bf{W}}^{\text{\tiny{CC}}}$ is the weight, ${\bf{W}}$, that minimizes the asymptotic MSE of estimators of the form $\hat{\bm{\beta}}^{\text{\tiny{lab}}}-{{\bf{W}}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right)$ for estimating $\bm{\beta}^{*}$.
-From Corollary 2.1, we note that in contrast to $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})$, the covariance term, $\bm{\Sigma}_{Y,\widehat{Y}}$, in $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{CC}}})$ is normalized by the residual variance of $\widehat{Y}$, $\bm{\Sigma}_{\widehat{Y},\widehat{Y}}$.
-This normalization effectively scales the covariance according to the precision of $\widehat{Y}$ and enables $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ to always be at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
-Returning to the previous example, we can gain more intuition as to why $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ possesses this property while $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ does not.
-Example 1 continued (Efficiency gain of $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ under homoskedasticity) We revisit the case of homoskedasticity and recall that
+${\bf{W}}^{\text{\tiny{CC}}}$ is the weight, $\bf{W}$}, that minimizes the MSE of estimators of the form $\hat{\bm{\beta}}^{\text{\tiny{lab}}}-{{\bf{W}}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right)$ for estimating $\bm{\beta}^{*}$.
+We return to the example of linear regression to further break down Corollary 2.1.
+Example 1 Continued (Linear Regression).
+For linear regression, the difference in the asymptotic variance-covariance of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is
+$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{CC}}})=\left(\frac{1}{\pi}-1\right){\bf{A}}\bm{\Sigma}_{Y,\widehat{Y}}\bm{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\bm{\Sigma}_{Y,\widehat{Y}}{\bf{A}}\succeq{\bf{0}}$
+while
+$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})=\left(\frac{1}{\pi}-1\right){\bf{A}}\left(2\bm{\Sigma}_{Y,\widehat{Y}}-\bm{\Sigma}_{\widehat{Y},\widehat{Y}}\right){\bf{A}}.$
+In contrast to $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})$, the covariance term, $\bm{\Sigma}_{Y,\widehat{Y}}$, in $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{CC}}})$ is normalized by the residual variance of $\widehat{Y}$, $\bm{\Sigma}_{\widehat{Y},\widehat{Y}}$.
+This normalization scales the covariance according to the precision of $\widehat{Y}$ and enables $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ to always be at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
+Returning to the example of homoskedasticity, we can gain more intuition as to why $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ always improves over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ while $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ does not.
+In this setting,
 $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})$ $=\left(\frac{1}{\pi}-1\right)\left(2\sigma_{Y,\widehat{Y}}-\sigma_{\widehat{Y}}^{2}\right){\bf{A}}$
-$=\left(\frac{1}{\pi}-1\right)\left[-\left(\sigma_{Y}\rho_{Y,\widehat{Y}}-\sigma_{\widehat{Y}}\right)^{2}+\sigma_{Y}^{2}\rho_{Y,\widehat{Y}}^{2}\right]{\bf{A}}.$
-Theorem 2 can be used to show that
+$=\left(\frac{1}{\pi}-1\right)\left[-\left(\sigma_{Y}\rho_{Y,\widehat{Y}}-\sigma_{\widehat{Y}}\right)^{2}+\sigma_{Y}^{2}\rho_{Y,\widehat{Y}}^{2}\right]{\bf{A}}$
+and
 $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{CC}}})$ $=\left(\frac{1}{\pi}-1\right)\left(\frac{\sigma_{Y,\widehat{Y}}^{2}}{\sigma_{\widehat{Y},\widehat{Y}}^{2}}\right){\bf{A}}=\left(\frac{1}{\pi}-1\right)\sigma_{Y}^{2}\rho_{Y,\widehat{Y}}^{2}{\bf{A}}.$
-$\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ provided that $\rho_{Y,\widehat{Y}}>0$ and the gain in efficiency scales directly with $\rho_{Y,\widehat{Y}}$ and $\sigma_{Y}^{2}$.
+$\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is therefore more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ provided that $\rho_{Y,\widehat{Y}}>0$ and the gain in efficiency scales directly with $\rho_{Y,\widehat{Y}}$ and $\sigma_{Y}^{2}$.
 In contrast, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ will only be more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ provided that $\rho_{Y,\widehat{Y}}>\frac{1}{2}\frac{\sigma_{\widehat{Y}}}{\sigma_{Y}}$ and the gain in efficiency will only scale in this manner with the additional condition that $\rho_{Y,\widehat{Y}}=\frac{\sigma_{\widehat{Y}}}{\sigma_{Y}}$.
@@
 Before evaluating the empirical properties of the CC estimator, we provide a brief overview of several closely related areas.
-Existing PB inference literature has established connections with semi-supervised inference, semi-parametric inference, missing data, and measurement error.
-21, 29, 30, 31, 23, 32, 33 We identify additional connections between PB inference methods and modern and classical methods from the economics and statistics literature dating back to the 1960s.
-These connections underscore that aspects of the PB inference problem have been previously examined, albeit with differences in motivation, terminology, and notation.
+Existing PB inference literature has noted connections with semi-supervised inference, semi-parametric inference, missing data, and measurement error.
+21, 22, 23, 32, 26, 33, 34 Here we establish additional connections between augmentation-based estimators and recently proposed PB inference approaches.
+We also review some modern and classical methods from the economics and statistics literature dating back to the 1960s that directly relate to PB inference.
+These connections underscore that aspects of this modern problem have been previously examined, albeit with differences in motivation, terminology, and notation.
 Given the close relationship between the PPI and CC methods, ongoing research in PB inference would benefit from a more comprehensive review of these and other related approaches.
+4.1 Extensions of PPI
+We first return to the PPI++ and PSPA methods introduced in Section 2.4 that similarly aim to improve the efficiency of the PPI estimator.
+In contrast to the PPI and CC estimators that augment $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ with information from $\widehat{Y}$, PPI++ and PSPA directly incorporate $\widehat{Y}$ into the estimation equation for $\bm{\beta}^{*}$.
+Specifically, they aim to find the solution to the augmented estimating equation
+$\sum_{i=1}^{n_{\text{\tiny{lab}}}}\bm{\phi}(Y_{i},{\bf{X}}_{i};\bm{\beta})-{\bf{V}}\left\{\sum_{i=1}^{n_{\text{\tiny{lab}}}}\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta})-\sum_{i=1}^{n}\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta})\right\}=0$ (2)
+where $\bf{V}$} is a matrix that controls the contribution of $\widehat{Y}$ into estimation of $\bm{\beta}^{*}$Similar to the PPI estimator, the authors propose to use only the unlabeled data in the augmentation term.
+We provide further discussion of this point in the Appendix..
+For PPI++, ${\bf{V}}^{\text{\tiny{PPI++}}}=\lambda{\bf{I}}$, where $\lambda$ is selected to minimize the trace of the variance-covariance matrix of the corresponding estimator.
+For PSPA, ${\bf{V}}^{\text{\tiny{PSPA}}}=\text{diag}({\bf{v}})$ where $\bf{v}$} is instead selected to minimize element-wise variance.
+While both of these approaches fall within the AIPW framework by construction, we again return to our example of linear regression to connect these approaches to augmentation-based estimators and, in turn, show that the CC method is preferred over both PSPA and PPI++.
+We leave further comparisons of the three methods, particularly for general Z-estimation problems, to the Appendix and highlight some of our findings in our numerical studies in Section 5.
+Example 1 Continued (Linear Regression). The PSPA and PPI++ estimators are asymptotically equivalent to augmentation-based estimators. That is,
+$\hat{\bm{\beta}}^{\text{\tiny{PPI++}}}\approx\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\widehat{{\bf{W}}}^{\text{\tiny{PPI++}}}(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}})\quad\mbox{and}\quad\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}\approx\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\widehat{{\bf{W}}}^{\text{\tiny{PSPA}}}(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}})$
+where $\widehat{{\bf{W}}}^{\text{\tiny{PPI++}}}={\bf{A}}\widehat{{\bf{V}}}^{\text{\tiny{PPI++}}}{\bf{A}}^{-1}$, $\widehat{{\bf{W}}}^{\text{\tiny{PSPA}}}={\bf{A}}\widehat{{\bf{V}}}^{\text{\tiny{PSPA}}}{\bf{A}}^{-1}$, and $\widehat{{\bf{V}}}^{\text{\tiny{PPI++}}}$ and $\widehat{{\bf{V}}}^{\text{\tiny{PSPA}}}$ are defined in the Appendix.
+From Corollary 2.2, the CC estimator is asymptotically at least as efficient as both PPI++ and PSPA.
+In fact, both can be viewed as special cases of the CC estimator: PSPA restricts $\bf{W}$} to be diagonal, while PPI++ further constrains the diagonal elements to be equal.
 4.2 Seemingly unrelated regression
-Seemingly unrelated regression (SUR) was introduced by economist Arnold Zellner in 1962 for simultaneous estimation of multiple related regression models.
-34, 35, 36 The regression models have different outcome variables, but are assumed to have correlated errors.
+The augmentation-based representation not only unifies the PPI++, PSPA, and CC estimators under a common framework, but also connects to several other estimators developed for linear regression, both historically and in recent work.
+In particular, the framework underlies seemingly unrelated regression (SUR), a method introduced by economist Arnold Zellner in 1962 for the joint estimation of multiple related regression models.
+35, 36, 37.
+Analogously to the current setting, the regression models have different outcome variables, but may have correlated errors.
 SUR leverages the correlation between the error terms to produce more efficient estimators compared to estimating each regression separately.
-In the current setting, the SUR problem considers the regression models
-$Y={\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\beta}^{*}+\epsilon\quad\mbox{and}\quad\widehat{Y}={\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma}^{*}+\epsilon^{\prime}\quad$
-where $\mathbb{E}(\epsilon\mid{\bf{X}})=0$ and $\mathbb{E}(\epsilon^{\prime}\mid{\bf{X}})=0.$ An estimator for $\bm{\beta}^{*}$ that is based on augmenting $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ was proposed by Conniffe 37 and is given by:
+More formally, SUR considers the regression models
+$Y={\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\beta}_{0}+\epsilon\quad\mbox{and}\quad\widehat{Y}={\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma}_{0}+\epsilon^{\prime}\quad$
+where $\mathbb{E}(\epsilon\mid{\bf{X}})=0$ and $\mathbb{E}(\epsilon^{\prime}\mid{\bf{X}})=0.$ An estimator for $\bm{\beta}_{0}$ that is based on augmenting $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ was proposed by Conniffe 38 and is given by:
 $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}=\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\widehat{W}^{\text{\tiny{SUR}}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right)\quad\mbox{where}\quad\widehat{W}^{\text{\tiny{SUR}}}=\frac{\sum_{i=1}^{n_{\text{\tiny{lab}}}}(Y_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\hat{\bm{\beta}}^{\text{\tiny{lab}}})(\widehat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\hat{\bm{\gamma}}^{\text{\tiny{lab}}})}{\sum_{i=1}^{n_{\text{\tiny{lab}}}}(\widehat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\hat{\bm{\gamma}}^{\text{\tiny{lab}}})^{2}}.$
@@
 In the homoskedastic case when $\text{Var}(\epsilon\mid{\bf{X}})$ and $\text{Cov}(\epsilon,\epsilon^{\prime}\mid{\bf{X}})$ are constant and do not depend on $\bf{X}$}, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is asymptotically equivalent to $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$.
-22 Moreover, when the conditional distribution of $(\epsilon,\epsilon^{\prime})$ given ${\bf{X}}$ is bivariate normal with mean $\bm{0}$ and has constant variance-covariance, $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ (and thus $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ as well) is the maximum likelihood estimator of $\bm{\beta}^{*}$ and fully efficient under the assumed parametric model.
-37 In other scenarios, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ achieves a lower asymptotic variance than $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$.
+25 Moreover, when the conditional distribution of $(\epsilon,\epsilon^{\prime})$ given $\bf{X}$} is bivariate normal with mean $\bm{0}$ and has constant variance-covariance, $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ (and thus $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ as well) is the maximum likelihood estimator of $\bm{\beta}_{0}$ and fully efficient under the assumed parametric model.
+38 In other scenarios, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ achieves a lower asymptotic variance than $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ due to Corollary 2.2.
 The SUR estimator was recently applied in a PB inference problem inspired by genome-wide association studies conducted with population biobanks.
 These studies often rely on pre-trained models to impute outcomes of interest (i.e., phenotypes or traits) as they are partially missing due to the time and expense of ascertainment.
-12 The proposed method, “SynSurr”, assumes the true and predicted outcomes follow a bivariate normal distribution.
+12 The proposed method, SynSurr, assumes the true and predicted outcomes follow a bivariate normal distribution.
 The SynSurr estimator is equivalent to $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ and asymptotically equivalent to $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ when $\text{Var}(\epsilon\mid{\bf{X}})$ and $\text{Cov}(\epsilon,\epsilon^{\prime}\mid{\bf{X}})$ are constant.
-Building on SynSurr, Miao et al 14 proposed the “POP-GWAS” estimator for a variety of regression models utilizing the augmentation technique.
+Building on SynSurr, Miao et al 14 proposed the POP-GWAS estimator for a variety of regression models utilizing the augmentation technique.
 The POP-GWAS estimator is
@@
 4.3 Augmented estimation for quantifying a treatment effect
-The augmented estimation approach used by the PPI and CC methods has also been widely used for treatment effect estimation.38, 39, 40, 41, 42, 26, 43 Many of these approaches build on methods from survey sampling.44, 45 For example, Tsiatis et al 40 consider covariate adjusted analyses for comparing two treatments in randomized clinical trials.
-Letting $X$ be an indicator for randomization to the treated or control arms, the focus of inference is on $\beta$ in the model
-$\mathbb{E}(Y\mid X)=\alpha+\beta X$
-to estimate the treatment effect $\Delta=\mathbb{E}(Y\mid X=1)-\mathbb{E}(Y\mid X=0)$.
+The augmented estimation approach has also been widely used for treatment effect estimation.39, 40, 41, 42, 43, 29, 44 Many of these approaches build on methods from survey sampling.45, 46 For example, Tsiatis et al 41 consider covariate adjusted analyses for comparing two treatments in randomized clinical trials.
+The goal is to estimate the treatment effect $\Delta=\mathbb{E}(Y\mid X=1)-\mathbb{E}(Y\mid X=0)$.
 A standard estimator of $\Delta$ is $\widehat{\Delta}=\overline{Y}^{1}-\overline{Y}^{0}$, where $\overline{Y}^{1}=n_{1}^{-1}\sum_{i=1}^{n}Y_{i}X_{i}$ and $\overline{Y}^{0}=n_{0}^{-1}\sum_{i=1}^{n}Y_{i}(1-X_{i})$ are the empirical means of $Y$ in the treated and control groups of sizes $n_{1}$ and $n_{0}$, respectively.
-To utilize information from baseline covariates ${\bf{Z}}$ for improving efficiency, Tsiatis et al 40 propose an augmented estimator of the treatment effect given by
+To utilize information from baseline covariates $\bf{Z}$} for improving efficiency, Tsiatis et al 41 propose an augmented estimator of the treatment effect given by
 $\widehat{\Delta}-\sum_{i=1}^{n}(X_{i}-\bar{X})[n_{0}^{-1}h^{(0)}({\bf{Z}}_{i})+n_{1}^{-1}h^{(1)}({\bf{Z}}_{i})],$ (3)
 where $\bar{X}=n^{-1}\sum_{i=1}^{n}X_{i}$ and $h^{(k)}({\bf{Z}})$ are arbitrary scalar functions of $\bf{Z}$} for $k=0,1$.
-As $X$ and ${\bf{Z}}$ are independent due to randomization, the augmentation term in 2 converges in probability to 0, analogously to the augmentation terms in the PPI and CC estimators.
+As $X$ and $\bf{Z}$} are independent due to randomization, the augmentation term in 3 converges in probability to 0, analogously to the augmentation terms in the PPI and CC estimators.
 The gain in efficiency similarly arises from the potential correlation between $Y$ and $\bf{Z}$}.
-This technique has been applied in numerous settings, most recently for the analysis of time-to-event outcomes46 and recurrent events.47
+This technique has been applied in numerous settings, most recently for the analysis of time-to-event outcomes47 and recurrent events.31
 4.4 Control variates in Monte Carlo simulations
-The augmentation technique has also been utilized to improve the efficiency of Monte Carlo estimation.
-Specifically, Rothery 48 considers using a control variate, a statistic whose properties are known, to improve precision when estimating power via simulation.
+Lastly, the augmentation technique has also been utilized to improve the efficiency of Monte Carlo estimation.
+In 1982, Rothery 48 considered using a control variate, a statistic whose properties are known, to improve precision when estimating power via simulation.
 Suppose interest lies in estimating the power of a test $T$ of $H_{0}:\theta=0$ for some scalar parameter $\theta$.
@@
 The variance of the augmented estimator is minimized at $\lambda^{*}=\frac{\text{Cov}(Y,Z)}{\text{Var}(Z)}$.
-48, 49 The augmentation term in 3 reduces the variance of the unadjusted power estimate $\overline{Y}$ in proportion to the correlation between $Y$ and $Z$.
-Building on work from the survey sampling literature, a similar idea was later applied for semi-supervised estimation of means.50
+48, 49 The augmentation term in 4 reduces the variance of the unadjusted power estimate $\overline{Y}$ in proportion to the correlation between $Y$ and $Z$.
+Building on work from the survey sampling literature, a similar idea was later applied for semi-supervised estimation of means 50 and for causal effect effect estimation with multiple observational data sources44.
 5 Numerical studies
 5.1 Data generation
-We next consider two primary scenarios to evaluate the performance of the PB inference methods and to demonstrate the implications of our theoretical results.
-In both settings, the generative model for the outcome was
-$Y=\theta_{X_{1}}X_{1}+\theta_{\text{\tiny{INT}}}X_{1}W_{1}+{\bf{W}}^{\top}\bm{\theta}+\epsilon$
-where ${\bf{W}}=(W_{1},\dots,W_{q})^{{\sf\scriptscriptstyle{T}}}$ and the error term $\epsilon$ was simulated independently from a standard normal distribution.
-We generated $({\bf{W}}^{\top},X_{1})^{{\sf\scriptscriptstyle{T}}}$ from a mean-zero multivariate normal distribution with a compound symmetric covariance structure with unit variance and pairwise correlation of 0.4.
-Our objective was to perform inference on the marginal effect of $X_{1}$ on $Y$ in a linear model.
-Specifically, we target the parameter $\beta_{1}^{*}$, defined as
-$(\beta_{0}^{*},\beta_{1}^{*})=\operatorname*{arg\,min}_{\bm{\beta}}\mathbb{E}\left[(Y-(\beta_{0}+\beta_{1}X_{1}))^{2}\right].$
-The predictors to train the ML model, ${\bf{Z}}$, were constructed based on ${\bf{W}}$ and $X_{1}$.
-The scenarios described below were obtained by varying the values of $\theta_{X_{1}}$, $\theta_{\text{\tiny{INT}}}$, and $\bm{\theta}$, as well as the dimension $q$ of ${\bf{W}}$ and the composition of the predictors ${\bf{Z}}$.
-Scenario 1 (Strong and Moderate Prediction Models)
-We consider two sub-scenarios to illustrate Corollary 1.1, that is, that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when the sufficient condition holds.
-In both sub-scenarios, we let $q=10$, $\theta_{X_{1}}=1$, $\theta_{\text{\tiny{INT}}}=0$, and $\bm{\theta}=(1,1,0.5,0.5,0.25,{\bf{0}}_{1\times 5})^{{\sf\scriptscriptstyle{T}}}$.
-In Scenario 1a, we obtain a strong prediction model by regressing $Y$ on ${\bf{Z}}=(X_{1},W_{1},\dots,W_{10})^{{\sf\scriptscriptstyle{T}}}$ while in Scenario 1b we obtain a weaker prediction model by regressing $Y$ on ${\bf{Z}}=(X_{1},W_{5},X_{1}W_{5})^{{\sf\scriptscriptstyle{T}}}$.
-The $R^{2}$ of $\widehat{Y}$ for $Y$ is 0.9 and 0.6 in Scenarios 1a and 1b, respectively.
-Scenario 2 (Weak Prediction Models)
-We consider two sub-scenarios to illustrate the results in Corollary 1.1 and Corollary 2.1, which show that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ can be less efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ with a weak prediction model while $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is at least as efficient $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
-In both scenarios, we mimic a common scenario where the prediction model does not contain $X_{1}$.
-In Scenario 2a, we let $q=9$, $\theta_{X_{1}}=1$, $\theta_{\text{\tiny{INT}}}=0$, and $\bm{\theta}=(0.2,0.1\times{\bf{1}}_{1\times 8})^{{\sf\scriptscriptstyle{T}}}$.
-We obtain the prediction model by regressing $Y$ on ${\bf{Z}}=(W_{1},\dots,W_{9})^{{\sf\scriptscriptstyle{T}}}$.
-In Scenario 2b, we let $q=3$, $\theta_{X_{1}}=2$, $\theta_{\text{\tiny{INT}}}=0.25$, and $\bm{\theta}=(0.25,0.25,0.25)^{{\sf\scriptscriptstyle{T}}}$.
-We obtain the prediction model by regressing $Y$ on ${\bf{Z}}=(W_{1},W_{2},W_{3},W_{2}W_{1},W_{3}W_{1},W_{3}W_{2})^{{\sf\scriptscriptstyle{T}}}$.
-The $R^{2}$ of $\widehat{Y}$ for $Y$ is 0.5 and 0.4 in Scenarios 2a and 2b, respectively.
-In all settings, we generated two datasets from the same underlying distribution: (i) a training dataset to train the ML model and (ii) an independent analysis dataset for evaluating the PB inference methods containing a labeled dataset and an unlabeled dataset.
-We used 10,000 replications of the analysis dataset for each scenario and a single training set as $\hat{f}$ is assumed to be fixed.
-The size of the training set was 100,000 and the size of the analysis dataset was $n=10,000$.
-We varied the proportion of labeled data in the analysis dataset by letting $n_{\text{\tiny{lab}}}=300,600,1000$ or $2000$ in each setting.
-We evaluated the performance of the PPI and CC methods: $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, and $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$.
-We also included three comparison methods: $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ to serve as the baseline, $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$ to highlight the limitations of naive inference, and a PB inference method from a concurrent pre-print, the PoSt-Prediction Adaptive (PSPA) method.30 The PSPA method is a generalization of the PPI++ method, an extension of the original PPI estimator that adaptively minimizes the trace of the asymptotic variance-covariance of the PPI estimator.
-The PSPA estimator, denoted as $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$, is designed to achieve element-wise variance reduction relative to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ by incorporating the information from $\widehat{Y}$ directly into the estimating equation for $\bm{\beta}^{*}$.
-We present the percent bias for each estimator as well as the coverage probability (CP) of the 95% confidence intervals (CIs).
-For the four PB inference methods ($\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$, and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$), we also present the relative efficiency (RE) defined as the ratio of the mean squared error (MSE) of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ relative to the MSE of the PB inference estimator.
+We next evaluate the performance of the CC method and its comparators through extensive simulation studies for both linear and logistic regression.
+Although the data is generated in a manner similar to prior work in PB inference 22, 51, we consider several additional mechanisms for generating predictions.
+In all scenarios, we generated four covariates, ${\bf{W}}=(W_{1},W_{2},W_{3},W_{4})$, from a zero-mean multivariate normal distribution with a compound symmetric covariance structure with unit variance and pairwise correlation of 0.4.
+For linear regression, the true outcome was generated as
+$Y=h({\bf{W}})+\epsilon\quad\mbox{where}\quad\epsilon\sim N(0,0.75^{2})$
+and $h({\bf{W}})=1+0.1\exp(W_{1})+\sin(W_{2})+W_{3}+W_{3}^{2}+I(W_{4}>0)$. To assess performance of the PB inference methods using different predictions, we generated
+$\widehat{Y}=h({\bf{W}})+\tau\quad\mbox{where}\quad\tau\sim N\left(a+b|W_{2}|,\sigma^{2}_{\tau}\right).$
+We varied the amount of bias in the prediction by adjusting $a$ and $b$ and the amount of noise by adjusting $\sigma_{\tau}$.
+The predictions were generated with three types of errors: (i) random error where $a=b=0$ and $\sigma_{\tau}$ is a positive constant, (ii) non-random error where a=2a=-2, $b=0$, and $\sigma_{\tau}$ is a positive constant, and (iii) covariate-dependent error where $a=0$, b=3.5b=-3.5, and $\sigma_{\tau}=c|W_{2}|$ where $c$ is a positive constant.
+In all three scenarios, $\sigma_{\tau}$ was set to achieve $R^{2}=0.8$ or 0.6 to represent a high and low quality prediction.
+For logistic regression, we generated the true outcome $Y\sim\text{Bern}\left(\frac{1}{1+\exp[-h({\bf{W}})]}\right).$ The predictions were similarly generated with (i) random error by randomly flipping the true labels with probability $p_{error}$, (ii) non-random error by randomly flipping the labels with $Y=1$ with probability $p_{error}$ and those with $Y=0$ with probability $\frac{p_{error}}{20}$, and (iii) covariate-dependent error by flipping the observations with $Y=1$ with probability $I(W_{4}>0)\times\frac{p_{error}}{4}+I(W_{4}<0)\times 3p_{error}$ and those with $Y=0$ with probability $I(W_{4}>0)\times\frac{p_{error}}{4}+I(W_{4}<0)\times 2p_{error}$.
+In all three scenarios, $p_{error}$ was set to achieve an overall accuracy of either 0.9 or 0.7 to represent a high and low quality prediction.
+For both linear and logistic regression, we focus on making inference for the parameters obtained from regressing $Y$ on ${\bf{X}}=[1,W_{1},W_{3},I(W_{4}>0)]$.
+The size of the full dataset was $n=10,000$ and the proportion of labeled data was varied by letting $n_{\text{\tiny{lab}}}=300,600,1000$ or $2000$ in each setting.
+We evaluated the performance of the five aforementioned PB inference methods: $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI++}}}$, and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$.
+With the exception of the CC method, all estimators were obtained from the ipd R package.52 In the main text, we present the coverage probability (CP) of the 95% confidence intervals (CIs) to evaluate the quality of inferences and the relative efficiency (RE) to assess the improvement over classical inference for each method.
+The RE is defined as the ratio of the mean squared error (MSE) of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ relative to the MSE of the given estimator.
+In the Appendix, we include results for percent bias and comparisons of the asymptotic standard error (ASE) and empirical standard error (ESE).
+All of our results are summarized over 2,000 replications.
 5.2 Results
-Table 1 presents the percent bias of the various estimators across the different scenarios and labeled dataset sizes.
-With the exception of $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$, the methods exhibit minimal bias as they are all consistent estimators for the target parameter.
-The limitations of naive inference are particularly apparent in Scenarios 2a and 2b, which do not include $X_{1}$ in ML model training.
-$\hat{\bm{\gamma}}^{\text{\tiny{all}}}$ exhibits an average percent bias of -47% in Scenario 2a and -64% in Scenario 2b.
-Table 2 summarizes the CPs of the 95% CIs for all methods across the various scenarios and labeled dataset sizes.
-As expected, $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and the four PB inference estimators ($\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$, and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$) maintain coverage near the nominal level in all settings.
-In contrast, the 95% CIs based on $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$ achieve adequate coverage only in Scenario 1a, where a strong prediction model yields negligible bias in estimating $\beta_{1}^{*}$.
-Coverage drops to 92% in Scenario 1b and falls to 0% in Scenarios 2a and 2b, where bias is substantial.
-These results underscore the risks of relying on naive inference in practice and highlight the necessity of using reliable PB inference methods.
-Figure 2 displays the relative efficiency across the different settings.
-In Scenario 1, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ perform similarly, with RE significantly above 1, indicating that the PB inference methods have lower variance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
-Comparing Scenario 1a and 1b, we note a significant decrease in the gains in statistical efficiency of the PB inference methods due to the weaker prediction model used in Scenario 1b.
-However, all methods are at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, as the sufficient condition in Corollary 1.1 is satisfied.
-Additionally, as the labeled set size increases and hence $n_{\text{\tiny{lab}}}/n$ increases, the efficiency gain decreases as expected.
-$\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is always at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ in Scenarios 1a and 1b, and more efficient in Scenario 1a when the labeled dataset size is larger due to the difference in the augmentation terms of these two estimators.
-Scenario 2 highlights the performance differences between $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ and $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$, which underscore our findings in Corollary 1.1 and Corollary 2.1.
-In both sub-scenarios, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ consistently outperforms both $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ and is substantially more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
-In Scenario 2a, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ achieves an average RE of 1.27 across the various sample sizes, indicating a significant gain over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
-In Scenario 2b, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ again demonstrates superior RE compared to all other methods and remains more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
-$\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ offers a modest efficiency gain in Scenario 2a and comparable performance to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ in Scenario 2b.
-Both $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ are outperformed by $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ in nearly all scenarios.
-In Scenario 2a, while $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ remains comparable to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, the RE for $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is as low as 0.88 with $n_{\text{\tiny{lab}}}=2000$.
-Both $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ demonstrate a substantial loses in efficiency in Scenario 2b, with $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ exhibiting a RE of 0.63 with $n_{\text{\tiny{lab}}}=2000$.
-Figure 2: Relative efficiency of the PB inference methods compared to baseline estimation with the labeled data ($\hat{\bm{\beta}}^{\text{\tiny{lab}}}$).
-A ratio greater than 1 indicates that the PB inference method has a smaller mean squared error (MSE) than that of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
+5.2.1 Linear Regression
+For brevity, we present results for the lower quality prediction with $R^{2}=0.6$ and the higher quality prediction with $R^{2}=0.8$ in the Appendix.
+The general patterns for the PB inference methods are consistent across the settings, but with higher gains in efficiency and more pronounced differences across the methods.
+Figures 2 - 4 display the results for the three types of errors in the predictions.
+All methods provide coverage close to the nominal level across all scenarios.
+With respect to RE, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ generally have significantly higher MSE than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, with $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ offering some improvement over $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ when the proportion of labeled data is larger.
+In contrast, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI++}}}$, and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ consistently outperform $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ for all parameters and in all scenarios.
+For the prediction with random error, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ perform similarly and have REs exceeding 1.5 for all parameters, while $\hat{\bm{\beta}}^{\text{\tiny{PPI++}}}$ is substantially less efficient for the coefficient for $X_{2}$.
+For the predictions with non-random and covariate-dependent error, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ overall outperforms $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI++}}}$.
+$\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ provides substantial improvement over both methods for the coefficient for $X_{2}$ and over $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ for the coefficients for $X_{1}$ and $X_{3}$.
+In fact, $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI++}}}$ are both outperformed by $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ for the smaller labeled sample sizes for the coefficient for $X_{2}$.
+Figure 2: Coverage probability of the 95% confidence intervals for all methods and relative efficiency for the PB inference methods (PPI, PPIa, PPI++, PSPA, and $C$) for linear regression with a prediction with random error.
+Figure 3: Coverage probability of the 95% confidence intervals for all methods and relative efficiency for the PB inference methods (PPI, PPIa, PPI++, PSPA, and $C$) for linear regression with a prediction with non-random error.
+Figure 4: Coverage probability of the 95% confidence intervals for all methods and relative efficiency for the PB inference methods (PPI, PPIa, PPI++, PSPA, and $C$) for linear regression with a prediction with covariate-dependent error.
+5.2.2 Logistic Regression
+We again present results for the lower quality prediction with overall accuracy of 0.7 and the higher quality prediction with 0.9 accuracy in the Appendix.
+The general patterns are again consistent, but with higher gains in efficiency and more pronounced differences across the methods for the higher quality prediction.
+Figures 5 - 7 display the results for the three types of errors in the predictions.
+All methods provide coverage close to the nominal level across all scenarios.
+With respect to RE, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ underperform relative to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ while $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI++}}}$, and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ generally have MSE much lower than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ in all scenarios.
+For the predictions with random error, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI++}}}$, and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ perform similarly and have REs near 1.2 for all parameters.
+For the predictions with non-random error, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ outperforms $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI++}}}$ for all parameters, offering an additinal 10-20% improvement in efficiency.
+For the predictions with covariate-dependent error, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ performs similarly to $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ for all parameters and is substantially more efficient than $\hat{\bm{\beta}}^{\text{\tiny{PPI++}}}$ for the coefficient for $X_{3}$.
+Figure 5: Coverage probability of the 95% confidence intervals for all methods and relative efficiency for the PB inference methods (PPI, PPIa, PPI++, PSPA, and $C$) for logistic regression with a prediction with random error.
+Figure 6: Coverage probability of the 95% confidence intervals for all methods and relative efficiency for the PB inference methods (PPI, PPIa, PPI++, PSPA, and $C$) for logistic regression with a prediction with non-random error.
+Figure 7: Coverage probability of the 95% confidence intervals for all methods and relative efficiency for the PB inference methods (PPI, PPIa, PPI++, PSPA, and $C$) for logistic regression with a prediction with covariate-dependent error.
 6 UK Biobank Analysis
@@
 53 However, due to its high cost, DEXA-derived AFM is usually only measured in a small subset of participants in large-scale studies.
-For example, in the UK Biobank,52 a prospective cohort study measuring genetic, lifestyle, and health information on approximately 500,000 participants in the UK, DEXA-derived AFM is incompletely ascertained, particularly among populations historically underrepresented in genetic research.
+For example, in the UK Biobank,54 a prospective cohort study measuring genetic, lifestyle, and health information on approximately 500,000 participants in the UK, DEXA-derived AFM is incompletely ascertained, particularly among populations historically underrepresented in genetic research.
 54 For instance, out of $5,397$ South Asian participants in the UK Biobank, only $288$ (5.3%) participants have undergone DEXA scans.
@@
 The strong prediction model achieved an $R^{2}$ of 0.8 while the weak prediction model had a substantially lower $R^{2}$ of 0.5 in the South Asian subpopulation.
-Table 3 presents point estimates, standard errors, and RE, defined as the ratio of the variance of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ to the variance of the given estimator, across the two prediction model settings for five estimators: the baseline estimator ($\hat{\bm{\beta}}^{\text{\tiny{lab}}}$), the PPI estimators ($\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$), the PSPA estimator ($\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$), and the CC estimator ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$).
-We do not include naive inference in our analyses due to the limitations demonstrated in our simulation studies.
+Table 1 presents point estimates, standard errors, and RE, defined as the ratio of the variance of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ to the variance of the given estimator, across the two prediction model settings for five estimators: the baseline estimator ($\hat{\bm{\beta}}^{\text{\tiny{lab}}}$), the PPI estimators ($\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI++}}}$), the PSPA estimator ($\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$), and the CC estimator ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$).
 Echoing our theoretical and simulation study results, all methods yield similar point estimates as they are all consistent estimators of the same parameter.
@@
 Consequently, both $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ outperform $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
-In fact, all four PB inference methods ($\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$) yield substantially smaller standard errors compared to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
-Among the four PB inference methods, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ show comparable performance, with average REs of 1.39, 1.39, and 1.47, respectively.
+In fact, all of the PB inference methods ($\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI++}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$, and $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$) yield substantially smaller standard errors compared to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
+Among the PB inference methods, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI++}}}$, and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ show comparable performance, with average REs of 1.39, 1.39, 1.44, and 1.47, respectively.
 $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ exhibits higher efficiency gains with an average RE of 1.68.
@@
 Here, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ yield larger standard errors than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ for all parameters of interest, with an average RE of 0.52.
-In contrast, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$, as well as $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$, outperform $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, with average RE values of 1.28 and 1.16, respectively.
+In contrast, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$, as well as $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI++}}}$, outperform $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, with average RE values of 1.28, 1.29, and 1.16, respectively.
 Overall, the CC estimator is the best performing method in terms of efficiency, underscoring its utility in practice.
 7 Discussion
-In this paper, we bridge classical statistical methods with a modern problem to equip practitioners and methodologists with a rigorous approach for PB inference.
-We find that incorporating a simple weight in the PPI estimator previously proposed for studies with double-sampling designs22 yields valid inference that is guaranteed to be at least as efficient as estimation using only labeled data.
-Our work addresses efficiency shortcomings of the original PPI estimator which can undermine its practical value.
-Alongside theoretical results supporting these findings, our simulation studies and application to UK Biobank data verify that the PPI method can suffer substantial efficiency loss when the ML model is inaccurate.
-Additionally, we draw important connections to several established approaches from the economics and statistics literature to provide context for ongoing developments in PB inference.
-Our work is not without limitations.
-First, we focused our analyses on least squares estimation following Motwani and Witten.20 Further work is needed to analyze other statistical procedures (e.g., Z-estimation problems in general).
+In this paper, we expand on existing research into the efficiency of PB inference methods by making the following contributions:
+  1. 1.
+    We employed the AIPW framework to introduce the CC estimator from the double-sampling literature as a simple modification of the PPI estimator, offering theoretically guaranteed efficiency gains over classical inference.
+Additionally, we demonstrated that the CC estimator is the most efficient estimator within the class of weighted linear augmentation estimators.
+  2. 2.
+    We made numerous connections between PB inference and existing literature with direct implications for future methodology and applications.
+Of note, we provided a set of results relating newly proposed PB inference methods to the CC method.
+For linear regression, we showed that the PPI++ and PSPA estimators are asymptotically equivalent to weighted linear augmentation estimators and therefore the CC estimator is always at least as efficient as these two methods.
+This theoretical finding was corroborated in our simulated and real data analyses.
+  3. 3.
+    We conducted the most extensive simulation study to date comparing existing PB inference approaches.
+We demonstrated empirically that the CC estimator outperforms the PPI++ and PSPA estimators in numerous simulation scenarios as well as in an analysis of UK Biobank data.
+However, our work is not without limitations.
+First, we focused our analyses on Z-estimation and further work is needed to analyze other statistical procedures.
 Second, we assumed that the labeled and unlabeled data are drawn from the same underlying distribution or equivalently, that $Y$ is missing completely at random.
@@
 However, the CC estimator can be extended to settings where $Y$ is missing at random by incorporating inverse probability weighting and regression adjustment.
+Third, our current analysis assumes that only one prediction is available for analysis.
+In many practical scenarios multiple predictions with unknown accuracy are available to the analyst.
+The CC estimator can be naturally extended to accommodate multiple predictions and we provide a simulated example in the Appendix.
 Lastly, we focused on the traditional missing data scenario where the proportion of labeled data, $n_{\text{\tiny{lab}}}/n$, is bounded away from 0.
@@
 Code availability:
-Our analyses can be replicated with the code at: https://github.com/SelinaS37/PBInference.
+Our analyses can be replicated with the code at: https://github.com/jlgrons/PBInf-Simulations.
 UK Biobank Dataset:

### v6

@@
 Abstract
-From structural biology to epidemiology, predictions from machine learning (ML) models increasingly complement costly gold-standard data to enable faster, more affordable, and scalable scientific inquiry.
+From structural biology to epidemiology, predictions from machine learning (ML) models are increasingly used to complement costly gold-standard data to enable faster, more affordable, and scalable scientific inquiry.
 In response, prediction-based (PB) inference has emerged to accommodate statistical analysis using a large volume of predictions together with a small amount of gold-standard data.
-The goals of PB inference are two-fold: (i) to mitigate bias from errors in predictions and (ii) to improve efficiency relative to classical inference using only the gold-standard data.
-While early PB inference methods focused on bias, their ability to enhance efficiency remains a focus of ongoing research.
-We revisit a foundational PB inference method and show that a simple modification can be applied to guarantee provable improvements in efficiency.
-In doing so, we establish new connections between augmented inverse probability weighted estimators (AIPW) and several recently proposed PB inference methods with a similar focus.
-The utility of our proposal, which leverages prediction-based outcomes to enhance efficiency, is demonstrated through extensive simulation studies and an application to real data from the UK Biobank.
-Further, we contextualize PB inference by drawing connections to historical literature from economics and statistics, highlighting how classic methods directly inform this contemporary problem.
+The goals of PB inference are two-fold: (i) to mitigate bias from errors in predictions and (ii) to improve efficiency relative to traditional inference using only the gold-standard data.
+While early PB inference methods focused on bias, their ability to enhance efficiency remains unclear.
+We revisit a popular PB inference method and show that a simple modification can be applied to guarantee improvements in efficiency beyond yielding valid inferences when the ML predictions are imperfect.
+The utility of this approach in leveraging prediction-based outcomes to enhance efficiency is demonstrated through extensive simulation studies and an application to the UK Biobank data.
+We further contextualize the problem of PB inference through historical literature from economics and statistics to highlight perspectives from classical methods in this contemporary problem.
 Correspondence to:
@@
 1 Introduction
-Researchers are increasingly turning to machine learning (ML) models to obtain predictions for outcomes that are prohibitively time-consuming or expensive to measure directly and then conducting statistical analyses which treat the predictions as observed data.
-1, 2, 3, 4 This prediction-based (PB) inference approach, introduced by Wang et al,5 is comprised of two steps outlined in Figure 1.
+Researchers are increasingly turning to machine learning (ML) models to obtain predictions for outcomes that are prohibitively time-consuming or expensive to measure directly and conducting statistical analyses treating the predictions as observed data.
+hoffman2024we, fan2024narratives, gao2024semi, miao2024task This prediction-based (PB) inference approach, introduced by Wang et al,wang2020methods is comprised of two steps outlined in Figure 1.
 First, the outcome of interest is imputed using a prediction from a pre-trained model.
 Then, the predictions are used to draw inferences about the parameters of interest, such as the association parameters between the outcome and a set of covariates, together with a small amount of gold-standard data.***PB inference has primarily focused on outcomes that are difficult to measure.
-Recent work has considered predicted covariates (e.g., Kluger et al6).
-In contrast to classical inference that relies solely on gold-standard data, PB inference aims to leverage the predictions to improve statistical efficiency without introducing bias.
+Recent work has considered predicted covariates (e.g., Kluger et alkluger2025prediction).
+In contrast to traditional inference that relies solely on gold-standard data, PB inference aims to leverage the predictions to improve statistical efficiency without introducing bias.
 In the last several years, applications of PB inference have arisen in diverse fields, spanning genetics, biology, and medicine.
-For example, population biobanks, such as the UK Biobank ($n\approx$ 500K individuals), collect high-throughput genomic data on large populations and offer unprecedented opportunities for genetic discovery.
-However, many phenotypes are only partially observed due to the time and expense of measurement.
-ML models are frequently used to impute missing outcomes based on fully observed baseline data to enable downstream association studies with phenotypes of interest.
-7, 8, 9, 10, 11, 12, 13, 14 Similarly, in structural biology, AlphaFold provides protein conformation predictions without the need for expensive and labor-intensive experiments.
-15, 16 In the analysis of electronic health records data, large language models can rapidly ascertain disease status at an accuracy competitive with manual chart review by trained annotators, facilitating clinical and epidemiological studies.
-17, 18, 19
+For example, population biobanks, such as the UK Biobank ($n\approx$ 500K individuals), collect high-throughput genomic data on large populations and offer unprecedented opportunities for genetic discovery research.
+However, many traits and disease outcomes are only partially observed due to the time and cost of measurement.
+ML models are commonly used to impute missing outcomes based on fully observed baseline data to enable downstream association studies.
+tong2020augmented, alipanahi2021glaucoma, cosentino2023copd, dahl2023phenotype, an2023autocomplete, mccaw2024synthetic, liu2024semiparametric, miao2024popgwas Similarly, in structural biology, AlphaFold provides structural predictions without the need for expensive and labor-intensive experiments for proteomic studies.
+jumper2021highly, gan2022activation In the analysis of electronic health records data, large language models can rapidly ascertain disease status at an accuracy competitive with manual chart review for epidemiological and clinical studies.
+yang2022large, zhou2022cancerbert, alsentzer2023zero
 Figure 1: Prediction-based (PB) inference.
@@
 The focus of scientific inquiry is on the association between $Y$ and $\bf{X}$}.
-Classical statistical inference relies solely on the data on $Y$ and $\bf{X}$} while PB inference incorporates $\widehat{Y}$ to make use of all of the available data.
+Traditional statistical inference relies solely on the data on $Y$ and $\bf{X}$} while PB inference incorporates $\widehat{Y}$ to make use of all of the available data.
 Dark red represents labeled observations while light red represents unlabeled observations.
-Given the relevance of PB inference, Motwani and Witten 20 recently revisited two seminal methods in the field.
-In the context of least squares estimation, they showed that the post-prediction inference (PostPi) method proposed by Wang et al5 provides a consistent estimator for the association between an outcome and a set of covariates only in restrictive scenarios (e.g., when the pre-trained ML model perfectly captures the underlying regression function).
-This is a significant practical limitation.
-PB inference methods should guarantee valid inference (i.e., hypothesis tests with controlled type I error and confidence intervals with nominal coverage rates) in the common scenario that the prediction model is imperfect.
-In contrast to PostPi, the prediction powered inference (PPI) method proposed by Angelopoulos et al 21 yields consistent estimation and valid inference regardless of the quality of the ML model.
+Given the relevance of PB inference, Motwani and Witten motwani2023revisiting recently revisited two seminal methods in the field.
+In the context of least squares estimation, they showed that the post-prediction inference (PostPi) method proposed by Wang et alwang2020methods provides a consistent estimator for the association between an outcome and a set of covariates only in restrictive scenarios (e.g., when the pre-trained ML model perfectly captures the underlying regression function).
+This is a significant practical limitation as it is essential that PB inference methods yield valid inference (i.e., hypothesis tests with controlled type I error and confidence intervals with nominal coverage rates) in the likely scenario that the prediction model is imperfect.
+In contrast, the prediction powered inference (PPI) method proposed by Angelopoulos et al angelopoulos2023prediction yields consistent estimation and valid inference regardless of the quality of the ML model.
 This finding is critical given that the PPI paper, published just over a year ago in the high-impact journal Science, has already been cited nearly 150 times and widely adopted in applications.
-In this paper, we further study the PPI estimator by assessing its statistical efficiency.
-Our contributions are three-fold.
-First, we make new contributions to the literature identifying inefficiencies of the PPI method.
-22, 23, 24 We derive a set of results that characterize when the PPI estimator gains over classical estimation using labeled data only in the context of Z-estimation.
-Second, we show that a more efficient estimator can be obtained by adding a simple weight to the PPI estimator.
-This approach was proposed by Chen & Chen (CC) over 25 years ago in the context of double sampling designs 25 and its application within the PB inference setting is new.
-Third, we draw important connections between augmented inverse probability weighted (AIPW) estimators, our proposal, and several recently proposed methods.
-In doing so, we identify settings in which the CC estimator is theoretically guaranteed to improve over comparable approaches.
-Building on these findings, we also revisit literature from statistics and economics to identify work dating back to the 1960s that bears directly on current developments in PB inference.
-Overall, our contributions put classical statistical methods in conversation with this modern problem to inform future methodological developments and applications.
-We begin by formally introducing the PB inference problem, reviewing the contributions of Motwani and Witten20 and summarizing existing work in Section 2.
-We then study the asymptotic properties of the PPI estimator, showing that it can be outperformed by classical inference methods and introduce the CC estimator in Section 3.
-Our key finding is that the CC estimator is optimal with respect to statistical efficiency in the class of estimators to which the PPI estimator belongs.
-In Section 4, we discuss related literature and provide connections between the CC estimator, other PB inference methods, and AIPW estimators.
-We also develop a set of results that compare the efficiency of the CC estimator with several PB inference methods and identify scenarios where the CC estimator is theoretically optimal.
+In this paper, we build on the work of Motwani and Witten motwani2023revisiting and further examine the PPI estimator by assessing its statistical efficiency.
+Our contributions are two-fold.
+First, we develop a set of results that characterize when the PPI estimator gains over traditional estimators using labeled data only.
+We also show that a more efficient estimator can be obtained by adding a simple weight to the PPI estimator.
+This approach was proposed by Chen & Chen (CC) over 25 years ago in the context of double sampling designs.chen2000unified Second, we revisit literature from statistics and economics and identify work dating back to the 1960s that bears directly on current developments in PB inference.
+Our review puts classical statistical methods in conversation with this modern problem to inform future methodological advances and applications.
+We begin by formally introducing the PB inference problem and reviewing the contributions of Motwani and Wittenmotwani2023revisiting in Section 2.
+We then study the asymptotic properties of the PPI estimator, showing that it can be outperformed by traditional inference methods, and introduce the CC estimator in Section 3.
+Our key finding is that the CC estimator is optimal with respect to statistical efficiency in a class of estimators to which the PPI estimator belongs.
+In Section 4, we discuss related literature and provide connections between the CC estimator and other established methods.
 We then evaluate our findings with extensive simulation studies in Section 5.
@@
 2.1 Notation and Assumptions
-We let $Y$ denote the outcome of interest and $\bf{Z}$} its predictors.
-$Y$ is only partially observed and a pre-trained ML model is available to predict it over the full data given $\bf{Z}$}, which is fully observed.
+We let $Y$ denote the outcome of interest and $\bf{Z}$} a set of predictors.
+$Y$ is only partially observed and a pre-trained ML model is available to predict it over the full data given $\bf{Z}$}, which are fully observed.
 The ML model $\hat{f}:\mathcal{Z}\rightarrow\mathcal{Y}$ maps the predictor space $\mathcal{Z}$ to the outcome space $\mathcal{Y}$ and produces predictions $\widehat{Y}=\hat{f}({\bf{Z}})$.
@@
 $\mathscr{L}=\{(Y_{i},{\bf{Z}}_{i}^{{\sf\scriptscriptstyle{T}}},{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}})^{{\sf\scriptscriptstyle{T}}}\mid i=1,\dots,n_{\text{\tiny{lab}}}\}\quad\mbox{and}\quad\mathscr{U}=\{({\bf{Z}}_{i}^{{\sf\scriptscriptstyle{T}}},{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}})^{{\sf\scriptscriptstyle{T}}}\mid i=n_{\text{\tiny{lab}}}+1,\dots,n\},$
-where $n=n_{\text{\tiny{lab}}}+n_{\text{\tiny{unlab}}}$, $n_{\text{\tiny{unlab}}}$ is the number of unlabeled observations and $n_{\text{\tiny{lab}}}/n\to\pi$ as $n\to\infty$ for $\pi\in(0,1)$.
+where $n=n_{\text{\tiny{lab}}}+n_{\text{\tiny{unlab}}}$, $n_{\text{\tiny{unlab}}}$ is the number of unlabeled observations, and $n_{\text{\tiny{lab}}}/n\to\pi$ as $n\to\infty$ for $\pi\in(0,1)$.
 We focus on the case in which $\mathscr{L}$ and $\mathscr{U}$ include independent and identically distributed samples from a common population of interest.
@@
 $\hat{f}$ may be a black box function in that its training data and specification are not be available to the users.
-Approaches that address uncertainty in the estimation of $\hat{f}$ have recently been considered.26
+Approaches that address uncertainty in the estimation of $\hat{f}$ have recently been considered.zrnic2024cross
 2.2 Parameter of interest
-We focus on Z-estimation and aim to make inferences on the population parameter $\bm{\beta}^{*}$ that is the solution to the estimating equation
-$\mathbb{E}[\bm{\phi}(Y,{\bf{X}};\bm{\beta})]=\bm{0}.$
-$\bm{\phi}(\cdot)$ is a deterministic function that satisfies conditions C1-C4 in the Appendix.
-The simplest consistent estimator of $\bm{\beta}^{*}$, which we denote by $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, uses only $\mathscr{L}$ for estimation and is the solution to the estimating equation
-$\sum_{i=1}^{n_{\text{\tiny{lab}}}}\bm{\phi}(Y_{i},{\bf{X}}_{i};\bm{\beta})=\bm{0}.$
+Following Motwani and Witten,motwani2023revisiting we consider least squares estimation and aim to make inferences on the population parameter
+$\bm{\beta}^{*}\coloneq\operatorname*{arg\,min}_{\bm{\beta}}\mathbb{E}[(Y-{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\beta})^{2}]=[\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})]^{-1}\mathbb{E}({\bf{X}}Y),$
+which quantifies the association between $Y$ and $\bf{X}$}.
+The simplest consistent estimator of $\bm{\beta}^{*}$, which we denote by $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, uses only $\mathscr{L}$ for estimation and solves the estimating equation
+$\sum_{i=1}^{n_{\text{\tiny{lab}}}}\bm{\phi}(Y_{i},{\bf{X}}_{i};\bm{\beta})\coloneq\sum_{i=1}^{n_{\text{\tiny{lab}}}}{\bf{X}}_{i}(Y_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\bm{\beta})=\bm{0}.$
 2.3 Motivation for PB inference
@@
 However, relying on $\widehat{Y}$ for inference may result in biased estimation depending on the quality of predictions from $\hat{f}$.
-For example, in “naive” inference, $\widehat{Y}$ is directly substituted for $Y$ in downstream analysis.5 The parameter $\bm{\gamma}^{*}$ targeted by this approach solves
-$\mathbb{E}[\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma})]=\bm{0}.$
-The corresponding estimator based on $\mathscr{L}$ and $\mathscr{U}$, denoted as $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$, is the solution to
-$\sum_{i=1}^{n}\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma})=\bm{0}.$
-As the distribution of $(\widehat{Y},{\bf{X}}^{{\sf\scriptscriptstyle{T}}})^{{\sf\scriptscriptstyle{T}}}$ may not coincide with that of $(Y,{\bf{X}}^{{\sf\scriptscriptstyle{T}}})^{{\sf\scriptscriptstyle{T}}}$, $\bm{\gamma}^{*}$ bears no necessary correspondence with $\bm{\beta}^{*}$ except under restrictive scenarios.
-For example, $\bm{\gamma}^{*}$ coincides with $\bm{\beta}^{*}$ when $\bf{Z}$} contains $\bf{X}$}, the ML model captures the true regression function $\mathbb{E}(Y\mid{\bf{Z}})$, and the focus of inference is on a parameter from a linear regression model.
+For example, in “naive” inference, $\widehat{Y}$ is directly substituted for $Y$ in downstream analysis.wang2020methods The parameter targeted by this approach is
+[…]
+and the corresponding least squares estimator, denoted as $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$, is the solution to
+$\sum_{i=1}^{n}\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma})\coloneq\sum_{i=1}^{n}{\bf{X}}_{i}(\widehat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma})=\bm{0}.$
+As the distribution of $(\widehat{Y},{\bf{X}}^{{\sf\scriptscriptstyle{T}}})^{{\sf\scriptscriptstyle{T}}}$ may not coincide with that of $(Y,{\bf{X}}^{{\sf\scriptscriptstyle{T}}})^{{\sf\scriptscriptstyle{T}}}$, $\bm{\gamma}^{*}$ bears no necessary correspondence with $\bm{\beta}^{*}$ except under restrictive scenarios (e.g., when $\bf{Z}$} contains $\bf{X}$} and the ML model captures the true regression function).
 To overcome these limitations, PB inference methods utilize $\mathscr{L}$ and $\mathscr{U}$ to essentially balance the relative contributions of $Y$ and $\widehat{Y}$ in making inference on $\bm{\beta}^{*}$.
 2.4 PB inference methods
-In the context of least squares estimation, Motwani and Witten20 revisited two seminal strategies for PB inference.
-One of these is the PostPI method of Wang et al,5 which was the first modern method proposed for PB inference.
-PostPI consists of two steps to incorporate $\widehat{Y}$ into inference for $\bm{\beta}^{*}$.
+Motwani and Wittenmotwani2023revisiting revisited two seminal strategies for PB inference.
+The first is the PostPI method of Wang et al,wang2020methods which is a two-step procedure to incorporate $\widehat{Y}$ into estimation.
 First, $\mathscr{L}$ is used to model the relationship between $Y$ and $\widehat{Y}$ to address potential inaccuracies in the predictions.
-Then, this model is used together with $\mathscr{U}$ to make inference on $\bm{\beta}^{*}$ with either analytical or bootstrap methods.
-This approach was shown to yield a consistent estimator for $\bm{\beta}^{*}$ under the strong and arguably unrealistic assumption that the ML model captures the true regression function.
+Then, this model is used together with $\mathscr{U}$ to make inference on $\bm{\beta}^{*}$ with either analytical or bootstrap approaches.
+This approach is shown to yield a consistent estimator for $\bm{\beta}^{*}$ under the strong and unrealistic assumption that the ML model captures the true regression function.
 When this assumption is not met, the PostPI estimator may not be consistent.
 Consequently, corresponding hypothesis tests may not have controlled type I error rates and confidence intervals may not have adequate coverage.
-Motwani and Witten20 also studied the PPI method 21 which was proposed a few years after PostPI.
-Unlike its predecessor, PPI has the crucial property that inference remains valid regardless of the quality of the predictions.
-That is, $\hat{f}$ need not capture the true regression function for the estimator to maintain consistency for $\bm{\beta}^{*}$.
-The PPI estimator utilizes a common statistical technique that augments an initial consistent estimator for the target parameter based solely on $\mathscr{L}$ with an additional term that is estimable from both $\mathscr{L}$ and $\mathscr{U}$.
-The augmentation term converges in probability to zero in order to improve efficiency without introducing bias.
-This approach dates back to at least Robins et al 27 and has been applied in multiple settings to improve estimation efficiency outside of the PB inference problem.
-25, 28, 29, 30, 31 We return to this point in our review of related literature in Section 4.
-Formally, the PPI estimator is
+Motwani and Wittenmotwani2023revisiting also considered the PPI method,angelopoulos2023prediction which has the crucial property that inference remains valid regardless of the quality of the predictions.
+That is, the PPI method does not require that $\hat{f}$ capture the true regression function for the estimator to maintain consistency for $\bm{\beta}^{*}$.
+The proposed estimator utilizes a common statistical technique involving the augmentation of an initial consistent estimator for the parameter of interest with an additional term that converges in probability to 0.
+This approach dates back to at least Robins et al robins1994estimation and has been applied in multiple settings to improve estimation efficiency outside of the PB inference problem.
+chen2000unified, tsiatis2006, tian2012covariate, wang2015semiparametric We return to this point in our review of related literature in Section 4.
+The PPI estimator is
 $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}=\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}\right)$
-where $\hat{\bm{\gamma}}^{\text{\tiny{lab}}}$ solves $\sum_{i=1}^{n_{\text{\tiny{lab}}}}\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma})=\bm{0}$ and $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ solves $\sum_{i=n_{\text{\tiny{lab}}}+1}^{n}\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma})=\bm{0}$Angelopoulos et al21 express $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ as $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}-\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\beta}}^{\text{\tiny{lab}}}\right)$ where $\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\beta}}^{\text{\tiny{lab}}}\right)$ is referred to as a “rectifier.” The rectifier serves the purpose of correcting potential bias in $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ due to inaccuracies in $\widehat{Y}$ for $Y$.
+where $\hat{\bm{\gamma}}^{\text{\tiny{lab}}}$ solves $\sum_{i=1}^{n_{\text{\tiny{lab}}}}\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma})=\bm{0}$ and $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ solves $\sum_{i=n_{\text{\tiny{lab}}}+1}^{n}\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma})=\bm{0}$Angelopoulos et alangelopoulos2023prediction express $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ as $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}-\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\beta}}^{\text{\tiny{lab}}}\right)$ where $\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\beta}}^{\text{\tiny{lab}}}\right)$ is referred to as a “rectifier.” The rectifier serves the purpose of correcting potential bias in $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ due to inaccuracies in $\widehat{Y}$ for $Y$.
 We reframe the PPI method to study its statistical efficiency..
-In contrast to the PostPI method, Motwani and Witten20 show that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is consistent for $\bm{\beta}^{*}$ and $n^{\frac{1}{2}}(\hat{\bm{\beta}}^{\text{\tiny{PPI}}}-\bm{\beta}^{*})$ converges weakly to a zero-mean multivariate normal distribution regardless of the quality of the prediction.
-Analogous results for Z-estimation problems are established in the original PPI paper (e.g., see Theorem S1 in the Supplementary Materials).
+In contrast to the PostPI method, Motwani and Wittenmotwani2023revisiting show that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is consistent for $\bm{\beta}^{*}$ and $n^{\frac{1}{2}}(\hat{\bm{\beta}}^{\text{\tiny{PPI}}}-\bm{\beta}^{*})$ converges weakly to a zero-mean multivariate normal distribution regardless of the quality of the prediction.
 This approach therefore asymptotically controls the type I error rate in hypothesis tests and produces confidence intervals with the nominal coverage level.
-However, for a PB inference method to be preferred in practice over classical inference based on $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, it should not only yield valid inference, but also be more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when $\widehat{Y}$ is informative of $Y$ and equally efficient when $\widehat{Y}$ is uninformative of $Y$.
-Two recent works have empirically identified that the PPI method can be less efficient than classical inference and have proposed extensions of the original method that are theoretically guaranteed to outperform classical inference.
-The first such extension, called PPI++, abandons the augmentation-based approach of PPI and instead directly augments the estimating equation for $\bm{\beta}^{*}$ with information from $\widehat{Y}$.
-A tuning parameter is used that controls the amount of information from $\widehat{Y}$ that is used for estimation and is cleverly designed to minimize the trace of the resulting estimator’s asymptotic covariance matrix.
-Subsequently, Miao et al.
-23 proposed Post-Prediction Adaptive inference (PSPA), which adopts a similar strategy, but instead produces an estimator with minimal element-wise variance.
-While both PPI++ and PSPA offer provable improvements over classical inference using only $\mathscr{L}$, their relationships to the original PPI framework remain underexplored.
-Furthermore, the conditions under which these methods are optimal have yet to be fully established.
-We revisit these proposals and clarify some of these connections in Section 4.
 3 Efficiency analysis
-For the remainder of this paper, we study a modified version of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ given by
+As PB inference methods suppose the availability of at least a small labeled data set, a natural baseline for comparison is the estimator based on $\mathscr{L}$ only, $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
+For a PB inference method to be preferred over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, it should not only yield valid inference, but also achieve gains in statistical efficiency relative to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when the prediction is informative of $Y$ and perform no worse when $\widehat{Y}$ is uninformative.
+For the remainder of this paper, we study the latter property and analyze a modified version of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ given by
 $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}=\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right).$
@@
 When $\pi\approx 0$, $\hat{\bm{\gamma}}^{\text{\tiny{all}}}\approx\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ so that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}\approx\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$.
-In general, it can be shown that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ gains efficiency over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ whenever $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ gains over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ can be more efficient than $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ when $\widehat{Y}$ is informative of $Y$ (see Appendix, specifically Section B.3, for a detailed analysis).
+In general, it can be shown that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ gains efficiency over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ whenever $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ gains over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ can be more efficient than $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ when $\widehat{Y}$ is informative of $Y$ (see Appendix, specifically Section B.3, for a detailed analysis).
 Additionally, estimators of this form are commonly used in the semi-parametric inference literature, which we draw upon in the subsequent section in our asymptotic analysis.
-We provide general results for Z-estimation and use linear regression as a running example throughout our discussion to further cultivate intuition.
 3.1 Asymptotic analysis of the PPI estimator
-To study the asymptotic properties of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, we present its influence function expansion in Theorem 1.
-The detailed derivations are deferred to the Appendix and follow those provided for the original PPI estimator.
+To study the asymptotic propeties of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, we first derive its influence function expansion in Theorem 1. The detailed derivations are deferred to the Appendix.
 Theorem 1 (Influence function expansion of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$).
@@
 $n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}-\bm{\beta}^{*}\right)=n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})+o_{p}(1),$
-where ${\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})={\bf{A}}\left\{\left[\ \bm{\phi}(Y_{i},{\bf{X}}_{i};\bm{\beta}^{*})\right]\frac{R_{i}}{\pi}+\left[\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma}^{*})\right]\frac{\pi-R_{i}}{\pi}\right\}$ and ${\bf{A}}=\mathbb{E}\left[\frac{\partial\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})}{\partial\bm{\beta}}\right]^{-1}$.
+where ${\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})={\bf{A}}\left\{\left[\ \bm{\phi}(Y_{i},{\bf{X}}_{i};\bm{\beta}^{*})\right]\frac{R_{i}}{\pi}+\left[\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma}^{*})\right]\frac{\pi-R_{i}}{\pi}\right\}$ and ${\bf{A}}=[\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})]^{-1}$.
 Therefore, $n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}-\bm{\beta}^{*}\right)$ converges weakly to a zero-mean multivariate normal distribution with variance-covariance $\mathbb{E}\left[{\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})\hskip 2.84526pt{\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})^{{\sf\scriptscriptstyle{T}}}\right].$
-We use Theorem 1 to compare the asymptotic variance of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ with that of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
-The result in Corollary 1.1 shows that the improvement of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ increases with the proportion of unlabeled data provided that $\widehat{Y}$ is of sufficient quality.
+We use Theorem 1 to compare the asymptotic variance-covariance of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ with that of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
+Corollary 1.1 shows that the benefit of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ scales directly with the predictiveness of $\widehat{Y}$ for $Y$ and inversely with the proportion of labeled data $\pi$.
 Corollary 1.1 (Comparison of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ with $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$).
-Let ${\bf{C}}_{11}=\mathbb{E}\left[\bm{\phi}(Y,X;\bm{\beta}^{*})\bm{\phi}(Y,X;\bm{\beta}^{*})^{{\sf\scriptscriptstyle{T}}}\right]$, ${\bf{C}}_{12}=\mathbb{E}\left[\bm{\phi}(Y,X;\bm{\beta}^{*})\bm{\phi}(\widehat{Y},X;\bm{\beta}^{*})^{{\sf\scriptscriptstyle{T}}}\right]$, ${\bf{C}}_{22}=\mathbb{E}\left[\bm{\phi}(\widehat{Y},X;\bm{\gamma}^{*})\bm{\phi}(\widehat{Y},X;\bm{\gamma}^{*})^{{\sf\scriptscriptstyle{T}}}\right]$, and ${\bf{B}}=\mathbb{E}\left[\frac{\partial\bm{\phi}(\widehat{Y},X;\bm{\gamma}^{*})}{\partial\bm{\gamma}}\right]^{-1}$.
+Let $\epsilon_{Y}=Y-{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\beta}^{*}$, $\epsilon_{\widehat{Y}}=\widehat{Y}-{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma}^{*}$, $\bm{\Sigma}_{\widehat{Y},\widehat{Y}}=\mathbb{E}\left(\epsilon_{\widehat{Y}}^{2}{\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\right)$, and $\bm{\Sigma}_{Y,\widehat{Y}}=\mathbb{E}\left(\epsilon_{Y}\epsilon_{\widehat{Y}}{\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\right)$.
 The difference in the asymptotic variance-covariance (aVar) of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is
-$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})$ $=n\left[\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{lab}}})-\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})\right]=\left(\frac{1}{\pi}-1\right)\left(2{\bf{A}}{\bf{C}}_{12}-{\bf{B}}{\bf{C}}_{22}\right){\bf{B}}.$
-$\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is therefore more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if $2{\bf{A}}{\bf{C}}_{12}\succ{\bf{B}}{\bf{C}}_{22}$.
-We further break down this result in the case of a linear regression model.
-Example 1 (Linear Regression).
-Let $\epsilon_{Y}=Y-{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\beta}^{*}$, $\epsilon_{\widehat{Y}}=\widehat{Y}-{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma}^{*}$, $\bm{\Sigma}_{\widehat{Y},\widehat{Y}}=\mathbb{E}\left(\epsilon_{\widehat{Y}}^{2}{\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\right)$, and $\bm{\Sigma}_{Y,\widehat{Y}}=\mathbb{E}\left(\epsilon_{Y}\epsilon_{\widehat{Y}}{\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\right)$.
-Following Corollary 1.1, the difference in the asymptotic variance-covariance (aVar) of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is
 $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})$ $=n\left[\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{lab}}})-\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})\right]=\left(\frac{1}{\pi}-1\right){\bf{A}}\left(2\bm{\Sigma}_{Y,\widehat{Y}}-\bm{\Sigma}_{\widehat{Y},\widehat{Y}}\right){\bf{A}}.$
 $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is therefore more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if $2\bm{\Sigma}_{Y,\widehat{Y}}\succ\bm{\Sigma}_{\widehat{Y},\widehat{Y}}$.
-Moreover, a sufficient condition for $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ to have lower variance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ is that
+A sufficient condition for $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ to have lower variance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ is that
 $\mathbb{E}\left[\left(\epsilon_{\widehat{Y}}-\epsilon_{Y}\right)^{2}\mid{\bf{X}}\right]<\mathbb{E}\left(\epsilon_{Y}^{2}\mid{\bf{X}}\right).$
-This condition coincides with the intuition that incorporating unlabeled data into estimation can improve efficiency when $\widehat{Y}$ has additional predictive power for $Y$ beyond what is already explained by the regression model of $Y$ on $\bf{X}$}.
-This condition is similar to the result for mean estimation provided in the Supplementary Materials of Angelopoulos et al 21 and can be used to describe several scenarios when $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is guaranteed to have lower asymptotic variance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when the ML model captures the true regression function (see Appendix, Section A.4).
-This insight can be better understood in the case of homoskedasticity.
-Specifically, suppose that $\mathbb{E}(\epsilon_{Y}^{2}\mid{\bf{X}})=\sigma_{Y}^{2}$, $\mathbb{E}(\epsilon_{\widehat{Y}}^{2}\mid{\bf{X}})=\sigma_{\widehat{Y}}^{2}$, and $\mathbb{E}(\epsilon_{Y}\epsilon_{\widehat{Y}}\mid{\bf{X}})=\sigma_{Y,\widehat{Y}}$,
+The sufficient condition in Corollary 1.1 coincides with the intuition that incorporating unlabeled data into estimation can improve efficiency when $\widehat{Y}$ has additional predictive power for $Y$ beyond what is already explained by the regression model of $Y$ on $\bf{X}$}.
+This condition is similar to the result for mean estimation provided in the Supplementary Materials of Angelopoulos et al angelopoulos2023prediction and can be used to describe several scenarios when $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is guaranteed to have lower asymptotic variance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when the ML model captures the true regression function (see Appendix, Section A.4).
+Here we consider a simple example to demonstrate when $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ gains efficiency over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ under homoskedasticity.
+Example 1 (Efficiency gain of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ under homoskedasticity).
+In the homoskedastic setting where $\mathbb{E}(\epsilon_{Y}^{2}\mid{\bf{X}})=\sigma_{Y}^{2}$, $\mathbb{E}(\epsilon_{\widehat{Y}}^{2}\mid{\bf{X}})=\sigma_{\widehat{Y}}^{2}$, and $\mathbb{E}(\epsilon_{Y}\epsilon_{\widehat{Y}}\mid{\bf{X}})=\sigma_{Y,\widehat{Y}}$,
 $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})=\left(\frac{1}{\pi}-1\right)\left[2\sigma_{Y,\widehat{Y}}-\sigma_{\widehat{Y}}^{2}\right]{\bf{A}}.$
@@
 $\rho_{Y,\widehat{Y}}>\frac{1}{2}\frac{\sigma_{\widehat{Y}}}{\sigma_{Y}},$
-where $\rho_{Y,\widehat{Y}}=\sigma_{Y,\widehat{Y}}/\sigma_{Y}\sigma_{\widehat{Y}}$ is the partial correlation of the residuals for $Y$ and $\widehat{Y}$ given $\bf{X}$}.
-Thus, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ gains efficiency over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when $\widehat{Y}$ is a strong predictor for $Y$ after adjusting for $\bf{X}$}, though $\widehat{Y}$ does not necessarily need to be derived from a ML model that captures the true regression function.
+where $\rho_{Y,\widehat{Y}}=\sigma_{Y,\widehat{Y}}/\sigma_{Y}\sigma_{\widehat{Y}}$ is the correlation of the residuals for $Y$ and $\widehat{Y}$ given $\bf{X}$}.
+Thus $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ gains efficiency over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when $\widehat{Y}$ is a strong predictor for $Y$ after adjusting for $\bf{X}$}, though $\widehat{Y}$ does not necessarily need to be derived from a ML model that captures the true regression function.
 The improvement is stronger with increasing correlation of the residuals between $Y$ and $\widehat{Y}$ and decreasing variability in the residuals of $\widehat{Y}$.
 3.2 Asymptotic analysis of the CC estimator
-To further analyze the efficiency of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, we draw on the landmark paper that introduced a class of augmented inverse probability weighted (AIPW) estimators which includes all regular asymptotic linear estimators.27 While connections between the PPI estimator and AIPW have been made in the context of mean estimation21, we explictly leverage this framework to introduce a simple modification of the original PPI estimator with theoretical guarantees of improved efficiency.
-In the current setting, where the goal is to estimate $\bm{\beta}^{*}$ with $Y$ partially observed and $\widehat{Y}$ fully observed, AIPW estimators have influence functions of the form
-${\bf{IF}}(Y,\widehat{Y},{\bf{X}};\bm{\omega})={\bf{A}}\left[\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\frac{R}{\pi}+\bm{\omega}(\widehat{Y},{\bf{X}})\frac{\pi-R}{\pi}\right],$
-with $\bm{\omega}(\widehat{Y},{\bf{X}})$ being a function of $\widehat{Y}$ and $\bf{X}$}. The optimal choice of $\bm{\omega}(\widehat{Y},{\bf{X}})$ with respect to asymptotic variance27 is
+To further analyze the efficiency of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, we draw on the landmark paper that introduced the class of augmented inverse probability weighted (AIPW) estimators.
+robins1994estimation In the current setting, where the goal is to estimate $\bm{\beta}^{*}$ with $Y$ partially observed and $\widehat{Y}$ fully observed, these estimators have influence functions of the form
+$\text{IF}(Y,\widehat{Y},{\bf{X}};\bm{\omega})={\bf{A}}\left[\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\frac{R}{\pi}+\bm{\omega}(\widehat{Y},{\bf{X}})\frac{\pi-R}{\pi}\right],$
+with $\bm{\omega}(\widehat{Y},{\bf{X}})$ being a function of $\widehat{Y}$ and $\bf{X}$}. The optimal choice of $\bm{\omega}(\widehat{Y},{\bf{X}})$ with respect to asymptotic variance is
 $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})=\mathbb{E}[\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\mid\widehat{Y},{\bf{X}}].$
-By Theorem 1, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is an AIPW estimator and can be seen as an approximation to the optimal estimator with $\bm{\omega}(\widehat{Y},{\bf{X}})=\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})$.
-This is a good approximation, for example, when $\bf{Z}$} contains $\bf{X}$}, the ML model captures the true regression function, and the focus is in fitting a generalized linear model with the canonical link function, $g(\cdot)$.
-In this case, $\bm{\phi}(Y,X;\bm{\beta})={\bf{X}}[Y-g(\bm{\beta}^{{\sf\scriptscriptstyle{T}}}{\bf{X}})]$ and $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})=\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})$.
-However, this is also a scenario where naive inference will yield consistent estimation that is more efficient than inference based on the PPI estimator.
-More broadly, the PPI estimator is not the most efficient estimator in its class and, from Corollary 1.1, can even be less efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when the quality of the prediction is insufficient.
-A PB inference strategy should enhance statistical efficiency relative to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ whenever possible and, at the very least, should not lose efficiency from incorporating unlabeled data into estimation.
+By Theorem 1, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is thus an AIPW estimator and can be seen as an approximation to the optimal estimator with $\bm{\omega}(\widehat{Y},{\bf{X}})=\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})$.
+The PPI estimator is therefore not the most efficient estimator in its class and, from Corollary 1.1, can even be less efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when $\bm{\Sigma}_{\widehat{Y},\widehat{Y}}\succ 2\bm{\Sigma}_{Y,\widehat{Y}}$.
+A PB inference strategy should enhance statistical efficiency relative to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ whenever possible and, at the very least, should not be negatively impacted by incorporating unlabeled data into estimation.
+This raises concerns about the practical utility of the PPI estimator.
 To improve upon $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, one approach is to assume a parametric model for $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$ that can be estimated from $\mathscr{L}$ for analysis.
@@
 Moreover, when the model for $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$ is misspecified, an efficiency gain is not guaranteed.
-To overcome these difficulties, Chen and Chen 25 proposed a particular choice of $\bm{\omega}(\widehat{Y},{\bf{X}})$ that is guaranteed to be at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and that does not require modeling of $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$.
+To overcome these difficulties, Chen and Chen chen2000unified proposed a particular choice of $\bm{\omega}(\widehat{Y},{\bf{X}})$ that is guaranteed to be at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and that does not require modeling of $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$.
 Their approach is motivated by the double-sampling design which, apart from differences in terminology (i.e., they regard $\widehat{Y}$ as a mismeasured version of $Y$), is analogous to the partially labeled data setting considered here.
@@
 $\bm{\omega}^{\text{\tiny{CC}}}(\widehat{Y},{\bf{X}})=\text{Cov}[\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*}),\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})]\{\text{Var}[\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})]\}^{-1}\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*}).$ (1)
-Relating this estimator back to the PPI estimator, the high-level idea is analogous, but $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$ is approximated by the linear projection of $\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})$ on $\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})$ rather than $\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})$ itself, which effectively protects against potential inadequacies in $\widehat{Y}$ that can lead to efficiency loss.
-The corresponding estimator, which we call the CC estimator, can be written as
+The corresponding estimator can be written as
 $\hat{\bm{\beta}}^{\text{\tiny{CC}}}=\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\widehat{{\bf{W}}}^{\text{\tiny{CC}}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right),$
 where $\widehat{{\bf{W}}}^{\text{\tiny{CC}}}$ is a consistent estimator of
-${{\bf{W}}}^{\text{\tiny{CC}}}={\bf{A}}{\text{Cov}}[\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*}),\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})]\{{\text{Var}}[\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})]\}^{-1}{\bf{B}}^{-1}$
-obtained with $\mathscr{L}$ and $\mathscr{U}$.
+${{\bf{W}}}^{\text{\tiny{CC}}}={\bf{A}}{\text{Cov}}[\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*}),\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})]\{{\text{Var}}[\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})]\}^{-1}{\bf{A}}^{-1}$
+obtained with $\mathscr{L}$.
 Therefore, a simple way to ensure that an estimator with the same structure as the PPI estimator has asymptotic variance no greater than that of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ is to add a weight function in front of the augmentation term.
@@
 Theorem 2 (Influence function expansion of $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$).
-Following Chen and Chen 25,
+Following Chen and Chen chen2000unified,
 $n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{CC}}}-\bm{\beta}^{*}\right)=n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{IF}}^{\text{\tiny{CC}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})+o_{p}(1),$
@@
 Therefore, $n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{CC}}}-\bm{\beta}^{*}\right)$ converges weakly to a zero-mean multivariate normal distribution with variance-covariance $\hskip 2.84526pt\mathbb{E}\left[{\bf{IF}}^{\text{\tiny{CC}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})\hskip 2.84526pt{\bf{IF}}^{\text{\tiny{CC}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})^{{\sf\scriptscriptstyle{T}}}\right]\hskip 2.84526pt.$
-The subsequent corollaries apply Theorem 2 to show our two key results:(1) $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and (2) $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is optimal with respect to efficiency in the class of estimators with a weighted linear augmentation term.
-These results are formally stated in Corollaries 2.1 and 2.2 and show that the CC method is preferable to the PPI method.
+The subsequent corollaries apply Theorem 2 to show our two key results: (1) $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and (2) $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ has the highest asymptotic efficiency within the class of estimators with a weighted linear augmentation term.
+These results suggest that the CC method is preferable to the PPI method for gaining efficiency.
 Corollary 2.1 ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$).
 The difference in the asymptotic variance-covariance of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is
-$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{CC}}})=\left(\frac{1}{\pi}-1\right){\bf{A}}{\bf{C}}_{12}{\bf{C}}_{22}^{-1}{\bf{C}}_{12}{\bf{B}}\succeq{\bf{0}}.$
+$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{CC}}})=\left(\frac{1}{\pi}-1\right){\bf{A}}\bm{\Sigma}_{Y,\widehat{Y}}\bm{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\bm{\Sigma}_{Y,\widehat{Y}}{\bf{A}}\succeq{\bf{0}}.$
 Therefore, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is guaranteed to be at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
@@
 Let the mean squared error (MSE) of an estimator $\hat{\bm{\beta}}$ for $\bm{\beta}^{*}$ be defined as $\mathbb{E}[(\hat{\bm{\beta}}-\bm{\beta}^{*})^{{\sf\scriptscriptstyle{T}}}(\hat{\bm{\beta}}-\bm{\beta}^{*})]$.
-${\bf{W}}^{\text{\tiny{CC}}}$ is the weight, $\bf{W}$}, that minimizes the MSE of estimators of the form $\hat{\bm{\beta}}^{\text{\tiny{lab}}}-{{\bf{W}}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right)$ for estimating $\bm{\beta}^{*}$.
-We return to the example of linear regression to further break down Corollary 2.1.
-Example 1 Continued (Linear Regression).
-For linear regression, the difference in the asymptotic variance-covariance of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is
-$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{CC}}})=\left(\frac{1}{\pi}-1\right){\bf{A}}\bm{\Sigma}_{Y,\widehat{Y}}\bm{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\bm{\Sigma}_{Y,\widehat{Y}}{\bf{A}}\succeq{\bf{0}}$
-while
-$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})=\left(\frac{1}{\pi}-1\right){\bf{A}}\left(2\bm{\Sigma}_{Y,\widehat{Y}}-\bm{\Sigma}_{\widehat{Y},\widehat{Y}}\right){\bf{A}}.$
-In contrast to $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})$, the covariance term, $\bm{\Sigma}_{Y,\widehat{Y}}$, in $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{CC}}})$ is normalized by the residual variance of $\widehat{Y}$, $\bm{\Sigma}_{\widehat{Y},\widehat{Y}}$.
-This normalization scales the covariance according to the precision of $\widehat{Y}$ and enables $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ to always be at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
-Returning to the example of homoskedasticity, we can gain more intuition as to why $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ always improves over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ while $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ does not.
-In this setting,
+Analogously to the well-known result for the best linear predictor (e.g.,Bickel and Doksum,bickel2015mathematical page 40), ${\bf{W}}^{\text{\tiny{CC}}}$ is the weight, $\bf{W}$}, that minimizes the asymptotic MSE of estimators of the form $\hat{\bm{\beta}}^{\text{\tiny{lab}}}-{{\bf{W}}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right)$ for estimating $\bm{\beta}^{*}$.
+From Corollary 2.1, we note that in contrast to $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})$, the covariance term, $\bm{\Sigma}_{Y,\widehat{Y}}$, in $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{CC}}})$ is normalized by the residual variance of $\widehat{Y}$, $\bm{\Sigma}_{\widehat{Y},\widehat{Y}}$.
+This normalization effectively scales the covariance according to the precision of $\widehat{Y}$ and enables $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ to always be at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
+Returning to the previous example, we can gain more intuition as to why $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ possesses this property while $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ does not.
+Example 1 continued (Efficiency gain of $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ under homoskedasticity) We revisit the case of homoskedasticity and recall that
 $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})$ $=\left(\frac{1}{\pi}-1\right)\left(2\sigma_{Y,\widehat{Y}}-\sigma_{\widehat{Y}}^{2}\right){\bf{A}}$
-$=\left(\frac{1}{\pi}-1\right)\left[-\left(\sigma_{Y}\rho_{Y,\widehat{Y}}-\sigma_{\widehat{Y}}\right)^{2}+\sigma_{Y}^{2}\rho_{Y,\widehat{Y}}^{2}\right]{\bf{A}}$
-and
+$=\left(\frac{1}{\pi}-1\right)\left[-\left(\sigma_{Y}\rho_{Y,\widehat{Y}}-\sigma_{\widehat{Y}}\right)^{2}+\sigma_{Y}^{2}\rho_{Y,\widehat{Y}}^{2}\right]{\bf{A}}.$
+Theorem 2 can be used to show that
 $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{CC}}})$ $=\left(\frac{1}{\pi}-1\right)\left(\frac{\sigma_{Y,\widehat{Y}}^{2}}{\sigma_{\widehat{Y},\widehat{Y}}^{2}}\right){\bf{A}}=\left(\frac{1}{\pi}-1\right)\sigma_{Y}^{2}\rho_{Y,\widehat{Y}}^{2}{\bf{A}}.$
-$\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is therefore more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ provided that $\rho_{Y,\widehat{Y}}>0$ and the gain in efficiency scales directly with $\rho_{Y,\widehat{Y}}$ and $\sigma_{Y}^{2}$.
+$\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ provided that $\rho_{Y,\widehat{Y}}>0$ and the gain in efficiency scales directly with $\rho_{Y,\widehat{Y}}$ and $\sigma_{Y}^{2}$.
 In contrast, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ will only be more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ provided that $\rho_{Y,\widehat{Y}}>\frac{1}{2}\frac{\sigma_{\widehat{Y}}}{\sigma_{Y}}$ and the gain in efficiency will only scale in this manner with the additional condition that $\rho_{Y,\widehat{Y}}=\frac{\sigma_{\widehat{Y}}}{\sigma_{Y}}$.
@@
 Before evaluating the empirical properties of the CC estimator, we provide a brief overview of several closely related areas.
-Existing PB inference literature has noted connections with semi-supervised inference, semi-parametric inference, missing data, and measurement error.
-21, 22, 23, 32, 26, 33, 34 Here we establish additional connections between augmentation-based estimators and recently proposed PB inference approaches.
-We also review some modern and classical methods from the economics and statistics literature dating back to the 1960s that directly relate to PB inference.
-These connections underscore that aspects of this modern problem have been previously examined, albeit with differences in motivation, terminology, and notation.
+Existing PB inference literature has established connections with semi-supervised inference, semi-parametric inference, missing data, and measurement error.
+angelopoulos2023prediction, angelopoulos2023ppi++, miao2023assumption, zrnic2024active, zrnic2024cross, angelopoulos_note_2024, ji2025predictions We identify additional connections between PB inference methods and modern and classical methods from the economics and statistics literature dating back to the 1960s.
+These connections underscore that aspects of the PB inference problem have been previously examined, albeit with differences in motivation, terminology, and notation.
 Given the close relationship between the PPI and CC methods, ongoing research in PB inference would benefit from a more comprehensive review of these and other related approaches.
-4.1 Extensions of PPI
-We first return to the PPI++ and PSPA methods introduced in Section 2.4 that similarly aim to improve the efficiency of the PPI estimator.
-In contrast to the PPI and CC estimators that augment $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ with information from $\widehat{Y}$, PPI++ and PSPA directly incorporate $\widehat{Y}$ into the estimation equation for $\bm{\beta}^{*}$.
-Specifically, they aim to find the solution to the augmented estimating equation
-$\sum_{i=1}^{n_{\text{\tiny{lab}}}}\bm{\phi}(Y_{i},{\bf{X}}_{i};\bm{\beta})-{\bf{V}}\left\{\sum_{i=1}^{n_{\text{\tiny{lab}}}}\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta})-\sum_{i=1}^{n}\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta})\right\}=0$ (2)
-where $\bf{V}$} is a matrix that controls the contribution of $\widehat{Y}$ into estimation of $\bm{\beta}^{*}$Similar to the PPI estimator, the authors propose to use only the unlabeled data in the augmentation term.
-We provide further discussion of this point in the Appendix..
-For PPI++, ${\bf{V}}^{\text{\tiny{PPI++}}}=\lambda{\bf{I}}$, where $\lambda$ is selected to minimize the trace of the variance-covariance matrix of the corresponding estimator.
-For PSPA, ${\bf{V}}^{\text{\tiny{PSPA}}}=\text{diag}({\bf{v}})$ where $\bf{v}$} is instead selected to minimize element-wise variance.
-While both of these approaches fall within the AIPW framework by construction, we again return to our example of linear regression to connect these approaches to augmentation-based estimators and, in turn, show that the CC method is preferred over both PSPA and PPI++.
-We leave further comparisons of the three methods, particularly for general Z-estimation problems, to the Appendix and highlight some of our findings in our numerical studies in Section 5.
-Example 1 Continued (Linear Regression). The PSPA and PPI++ estimators are asymptotically equivalent to augmentation-based estimators. That is,
-$\hat{\bm{\beta}}^{\text{\tiny{PPI++}}}\approx\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\widehat{{\bf{W}}}^{\text{\tiny{PPI++}}}(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}})\quad\mbox{and}\quad\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}\approx\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\widehat{{\bf{W}}}^{\text{\tiny{PSPA}}}(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}})$
-where $\widehat{{\bf{W}}}^{\text{\tiny{PPI++}}}={\bf{A}}\widehat{{\bf{V}}}^{\text{\tiny{PPI++}}}{\bf{A}}^{-1}$, $\widehat{{\bf{W}}}^{\text{\tiny{PSPA}}}={\bf{A}}\widehat{{\bf{V}}}^{\text{\tiny{PSPA}}}{\bf{A}}^{-1}$, and $\widehat{{\bf{V}}}^{\text{\tiny{PPI++}}}$ and $\widehat{{\bf{V}}}^{\text{\tiny{PSPA}}}$ are defined in the Appendix.
-From Corollary 2.2, the CC estimator is asymptotically at least as efficient as both PPI++ and PSPA.
-In fact, both can be viewed as special cases of the CC estimator: PSPA restricts $\bf{W}$} to be diagonal, while PPI++ further constrains the diagonal elements to be equal.
 4.1 Seemingly unrelated regression
-The augmentation-based representation not only unifies the PPI++, PSPA, and CC estimators under a common framework, but also connects to several other estimators developed for linear regression, both historically and in recent work.
-In particular, the framework underlies seemingly unrelated regression (SUR), a method introduced by economist Arnold Zellner in 1962 for the joint estimation of multiple related regression models.
-35, 36, 37.
-Analogously to the current setting, the regression models have different outcome variables, but may have correlated errors.
+Seemingly unrelated regression (SUR) was introduced by economist Arnold Zellner in 1962 for simultaneous estimation of multiple related regression models.
+zellner1962efficient, swamy1975bayesian, schmidt1977estimation The regression models have different outcome variables, but are assumed to have correlated errors.
 SUR leverages the correlation between the error terms to produce more efficient estimators compared to estimating each regression separately.
-More formally, SUR considers the regression models
-$Y={\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\beta}_{0}+\epsilon\quad\mbox{and}\quad\widehat{Y}={\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma}_{0}+\epsilon^{\prime}\quad$
-where $\mathbb{E}(\epsilon\mid{\bf{X}})=0$ and $\mathbb{E}(\epsilon^{\prime}\mid{\bf{X}})=0.$ An estimator for $\bm{\beta}_{0}$ that is based on augmenting $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ was proposed by Conniffe 38 and is given by:
+In the current setting, the SUR problem considers the regression models
+$Y={\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\beta}^{*}+\epsilon\quad\mbox{and}\quad\widehat{Y}={\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma}^{*}+\epsilon^{\prime}\quad$
+where $\mathbb{E}(\epsilon\mid{\bf{X}})=0$ and $\mathbb{E}(\epsilon^{\prime}\mid{\bf{X}})=0.$ An estimator for $\bm{\beta}^{*}$ that is based on augmenting $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ was proposed by Conniffe conniffe1985estimating and is given by:
 $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}=\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\widehat{W}^{\text{\tiny{SUR}}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right)\quad\mbox{where}\quad\widehat{W}^{\text{\tiny{SUR}}}=\frac{\sum_{i=1}^{n_{\text{\tiny{lab}}}}(Y_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\hat{\bm{\beta}}^{\text{\tiny{lab}}})(\widehat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\hat{\bm{\gamma}}^{\text{\tiny{lab}}})}{\sum_{i=1}^{n_{\text{\tiny{lab}}}}(\widehat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\hat{\bm{\gamma}}^{\text{\tiny{lab}}})^{2}}.$
@@
 In the homoskedastic case when $\text{Var}(\epsilon\mid{\bf{X}})$ and $\text{Cov}(\epsilon,\epsilon^{\prime}\mid{\bf{X}})$ are constant and do not depend on $\bf{X}$}, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is asymptotically equivalent to $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$.
-25 Moreover, when the conditional distribution of $(\epsilon,\epsilon^{\prime})$ given $\bf{X}$} is bivariate normal with mean $\bm{0}$ and has constant variance-covariance, $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ (and thus $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ as well) is the maximum likelihood estimator of $\bm{\beta}_{0}$ and fully efficient under the assumed parametric model.
-38 In other scenarios, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ achieves a lower asymptotic variance than $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ due to Corollary 2.2.
+chen2000unified Moreover, when the conditional distribution of $(\epsilon,\epsilon^{\prime})$ given $\bf{X}$} is bivariate normal with mean $\bm{0}$ and has constant variance-covariance, $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ (and thus $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ as well) is the maximum likelihood estimator of $\bm{\beta}^{*}$ and fully efficient under the assumed parametric model.
+conniffe1985estimating In other scenarios, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ achieves a lower asymptotic variance than $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$.
 The SUR estimator was recently applied in a PB inference problem inspired by genome-wide association studies conducted with population biobanks.
 These studies often rely on pre-trained models to impute outcomes of interest (i.e., phenotypes or traits) as they are partially missing due to the time and expense of ascertainment.
-12 The proposed method, SynSurr, assumes the true and predicted outcomes follow a bivariate normal distribution.
+mccaw2024synthetic The proposed method, “SynSurr”, assumes the true and predicted outcomes follow a bivariate normal distribution.
 The SynSurr estimator is equivalent to $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ and asymptotically equivalent to $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ when $\text{Var}(\epsilon\mid{\bf{X}})$ and $\text{Cov}(\epsilon,\epsilon^{\prime}\mid{\bf{X}})$ are constant.
-Building on SynSurr, Miao et al 14 proposed the POP-GWAS estimator for a variety of regression models utilizing the augmentation technique.
+Building on SynSurr, Miao et al miao2024popgwas proposed the “POP-GWAS” estimator for a variety of regression models utilizing the augmentation technique.
 The POP-GWAS estimator is
@@
 4.2 Augmented estimation for quantifying a treatment effect
-The augmented estimation approach has also been widely used for treatment effect estimation.39, 40, 41, 42, 43, 29, 44 Many of these approaches build on methods from survey sampling.45, 46 For example, Tsiatis et al 41 consider covariate adjusted analyses for comparing two treatments in randomized clinical trials.
-The goal is to estimate the treatment effect $\Delta=\mathbb{E}(Y\mid X=1)-\mathbb{E}(Y\mid X=0)$.
+The augmented estimation approach used by the PPI and CC methods has also been widely used for treatment effect estimation.leon2003semiparametric, lu2008improving, tsiatis2008covariate, gilbert2009efficient, zhang2010increasing, tian2012covariate, yang2020combining Many of these approaches build on methods from survey sampling.cassel1976some, sarndal2003model For example, Tsiatis et al tsiatis2008covariate consider covariate adjusted analyses for comparing two treatments in randomized clinical trials.
+Letting $X$ be an indicator for randomization to the treated or control arms, the focus of inference is on $\beta$ in the model
+$\mathbb{E}(Y\mid X)=\alpha+\beta X$
+to estimate the treatment effect $\Delta=\mathbb{E}(Y\mid X=1)-\mathbb{E}(Y\mid X=0)$.
 A standard estimator of $\Delta$ is $\widehat{\Delta}=\overline{Y}^{1}-\overline{Y}^{0}$, where $\overline{Y}^{1}=n_{1}^{-1}\sum_{i=1}^{n}Y_{i}X_{i}$ and $\overline{Y}^{0}=n_{0}^{-1}\sum_{i=1}^{n}Y_{i}(1-X_{i})$ are the empirical means of $Y$ in the treated and control groups of sizes $n_{1}$ and $n_{0}$, respectively.
-To utilize information from baseline covariates $\bf{Z}$} for improving efficiency, Tsiatis et al 41 propose an augmented estimator of the treatment effect given by
+To utilize information from baseline covariates $\bf{Z}$} for improving efficiency, Tsiatis et al tsiatis2008covariate propose an augmented estimator of the treatment effect given by
 $\widehat{\Delta}-\sum_{i=1}^{n}(X_{i}-\bar{X})[n_{0}^{-1}h^{(0)}({\bf{Z}}_{i})+n_{1}^{-1}h^{(1)}({\bf{Z}}_{i})],$ (2)
 where $\bar{X}=n^{-1}\sum_{i=1}^{n}X_{i}$ and $h^{(k)}({\bf{Z}})$ are arbitrary scalar functions of $\bf{Z}$} for $k=0,1$.
-As $X$ and $\bf{Z}$} are independent due to randomization, the augmentation term in 3 converges in probability to 0, analogously to the augmentation terms in the PPI and CC estimators.
+As $X$ and $\bf{Z}$} are independent due to randomization, the augmentation term in 2 converges in probability to 0, analogously to the augmentation terms in the PPI and CC estimators.
 The gain in efficiency similarly arises from the potential correlation between $Y$ and $\bf{Z}$}.
-This technique has been applied in numerous settings, most recently for the analysis of time-to-event outcomes47 and recurrent events.31
+This technique has been applied in numerous settings, most recently for the analysis of time-to-event outcomeszhang2025unified and recurrent events.gronsbell2024nonparametric
 4.3 Control variates in Monte Carlo simulations
-Lastly, the augmentation technique has also been utilized to improve the efficiency of Monte Carlo estimation.
-In 1982, Rothery 48 considered using a control variate, a statistic whose properties are known, to improve precision when estimating power via simulation.
+The augmentation technique has also been utilized to improve the efficiency of Monte Carlo estimation.
+Specifically, Rothery rothery1982 considers using a control variate, a statistic whose properties are known, to improve precision when estimating power via simulation.
 Suppose interest lies in estimating the power of a test $T$ of $H_{0}:\theta=0$ for some scalar parameter $\theta$.
@@
 The variance of the augmented estimator is minimized at $\lambda^{*}=\frac{\text{Cov}(Y,Z)}{\text{Var}(Z)}$.
-48, 49 The augmentation term in 4 reduces the variance of the unadjusted power estimate $\overline{Y}$ in proportion to the correlation between $Y$ and $Z$.
-Building on work from the survey sampling literature, a similar idea was later applied for semi-supervised estimation of means 50 and for causal effect effect estimation with multiple observational data sources44.
+rothery1982, davidson1992 The augmentation term in 3 reduces the variance of the unadjusted power estimate $\overline{Y}$ in proportion to the correlation between $Y$ and $Z$.
+Building on work from the survey sampling literature, a similar idea was later applied for semi-supervised estimation of means.zhang2019semi
 5 Numerical studies
 5.1 Data generation
-We next evaluate the performance of the CC method and its comparators through extensive simulation studies for both linear and logistic regression.
-Although the data is generated in a manner similar to prior work in PB inference 22, 51, we consider several additional mechanisms for generating predictions.
-In all scenarios, we generated four covariates, ${\bf{W}}=(W_{1},W_{2},W_{3},W_{4})$, from a zero-mean multivariate normal distribution with a compound symmetric covariance structure with unit variance and pairwise correlation of 0.4.
-For linear regression, the true outcome was generated as
-$Y=h({\bf{W}})+\epsilon\quad\mbox{where}\quad\epsilon\sim N(0,0.75^{2})$
-and $h({\bf{W}})=1+0.1\exp(W_{1})+\sin(W_{2})+W_{3}+W_{3}^{2}+I(W_{4}>0)$. To assess performance of the PB inference methods using different predictions, we generated
-$\widehat{Y}=h({\bf{W}})+\tau\quad\mbox{where}\quad\tau\sim N\left(a+b|W_{2}|,\sigma^{2}_{\tau}\right).$
-We varied the amount of bias in the prediction by adjusting $a$ and $b$ and the amount of noise by adjusting $\sigma_{\tau}$.
-The predictions were generated with three types of errors: (i) random error where $a=b=0$ and $\sigma_{\tau}$ is a positive constant, (ii) non-random error where a=2a=-2, $b=0$, and $\sigma_{\tau}$ is a positive constant, and (iii) covariate-dependent error where $a=0$, b=3.5b=-3.5, and $\sigma_{\tau}=c|W_{2}|$ where $c$ is a positive constant.
-In all three scenarios, $\sigma_{\tau}$ was set to achieve $R^{2}=0.8$ or 0.6 to represent a high and low quality prediction.
-For logistic regression, we generated the true outcome $Y\sim\text{Bern}\left(\frac{1}{1+\exp[-h({\bf{W}})]}\right).$ The predictions were similarly generated with (i) random error by randomly flipping the true labels with probability $p_{error}$, (ii) non-random error by randomly flipping the labels with $Y=1$ with probability $p_{error}$ and those with $Y=0$ with probability $\frac{p_{error}}{20}$, and (iii) covariate-dependent error by flipping the observations with $Y=1$ with probability $I(W_{4}>0)\times\frac{p_{error}}{4}+I(W_{4}<0)\times 3p_{error}$ and those with $Y=0$ with probability $I(W_{4}>0)\times\frac{p_{error}}{4}+I(W_{4}<0)\times 2p_{error}$.
-In all three scenarios, $p_{error}$ was set to achieve an overall accuracy of either 0.9 or 0.7 to represent a high and low quality prediction.
-For both linear and logistic regression, we focus on making inference for the parameters obtained from regressing $Y$ on ${\bf{X}}=[1,W_{1},W_{3},I(W_{4}>0)]$.
-The size of the full dataset was $n=10,000$ and the proportion of labeled data was varied by letting $n_{\text{\tiny{lab}}}=300,600,1000$ or $2000$ in each setting.
-We evaluated the performance of the five aforementioned PB inference methods: $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI++}}}$, and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$.
-With the exception of the CC method, all estimators were obtained from the ipd R package.52 In the main text, we present the coverage probability (CP) of the 95% confidence intervals (CIs) to evaluate the quality of inferences and the relative efficiency (RE) to assess the improvement over classical inference for each method.
-The RE is defined as the ratio of the mean squared error (MSE) of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ relative to the MSE of the given estimator.
-In the Appendix, we include results for percent bias and comparisons of the asymptotic standard error (ASE) and empirical standard error (ESE).
-All of our results are summarized over 2,000 replications.
+We next consider two primary scenarios to evaluate the performance of the PB inference methods and to demonstrate the implications of our theoretical results.
+In both settings, the generative model for the outcome was
+$Y=\theta_{X_{1}}X_{1}+\theta_{\text{\tiny{INT}}}X_{1}W_{1}+{\bf{W}}^{\top}\bm{\theta}+\epsilon$
+where ${\bf{W}}=(W_{1},\dots,W_{q})^{{\sf\scriptscriptstyle{T}}}$ and the error term $\epsilon$ was simulated independently from a standard normal distribution.
+We generated $({\bf{W}}^{\top},X_{1})^{{\sf\scriptscriptstyle{T}}}$ from a mean-zero multivariate normal distribution with a compound symmetric covariance structure with unit variance and pairwise correlation of 0.4.
+Our objective was to perform inference on the marginal effect of $X_{1}$ on $Y$ in a linear model.
+Specifically, we target the parameter $\beta_{1}^{*}$, defined as
+$(\beta_{0}^{*},\beta_{1}^{*})=\operatorname*{arg\,min}_{\bm{\beta}}\mathbb{E}\left[(Y-(\beta_{0}+\beta_{1}X_{1}))^{2}\right].$
+The predictors to train the ML model, $\bf{Z}$}, were constructed based on $\bf{W}$} and $X_{1}$.
+The scenarios described below were obtained by varying the values of $\theta_{X_{1}}$, $\theta_{\text{\tiny{INT}}}$, and $\bm{\theta}$, as well as the dimension $q$ of $\bf{W}$} and the composition of the predictors $\bf{Z}$}.
+Scenario 1 (Strong and Moderate Prediction Models)
+We consider two sub-scenarios to illustrate Corollary 1.1, that is, that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when the sufficient condition holds.
+In both sub-scenarios, we let $q=10$, $\theta_{X_{1}}=1$, $\theta_{\text{\tiny{INT}}}=0$, and $\bm{\theta}=(1,1,0.5,0.5,0.25,{\bf{0}}_{1\times 5})^{{\sf\scriptscriptstyle{T}}}$.
+In Scenario 1a, we obtain a strong prediction model by regressing $Y$ on ${\bf{Z}}=(X_{1},W_{1},\dots,W_{10})^{{\sf\scriptscriptstyle{T}}}$ while in Scenario 1b we obtain a weaker prediction model by regressing $Y$ on ${\bf{Z}}=(X_{1},W_{5},X_{1}W_{5})^{{\sf\scriptscriptstyle{T}}}$.
+The $R^{2}$ of $\widehat{Y}$ for $Y$ is 0.9 and 0.6 in Scenarios 1a and 1b, respectively.
+Scenario 2 (Weak Prediction Models)
+We consider two sub-scenarios to illustrate the results in Corollary 1.1 and Corollary 2.1, which show that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ can be less efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ with a weak prediction model while $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is at least as efficient $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
+In both scenarios, we mimic a common scenario where the prediction model does not contain $X_{1}$.
+In Scenario 2a, we let $q=9$, $\theta_{X_{1}}=1$, $\theta_{\text{\tiny{INT}}}=0$, and $\bm{\theta}=(0.2,0.1\times{\bf{1}}_{1\times 8})^{{\sf\scriptscriptstyle{T}}}$.
+We obtain the prediction model by regressing $Y$ on ${\bf{Z}}=(W_{1},\dots,W_{9})^{{\sf\scriptscriptstyle{T}}}$.
+In Scenario 2b, we let $q=3$, $\theta_{X_{1}}=2$, $\theta_{\text{\tiny{INT}}}=0.25$, and $\bm{\theta}=(0.25,0.25,0.25)^{{\sf\scriptscriptstyle{T}}}$.
+We obtain the prediction model by regressing $Y$ on ${\bf{Z}}=(W_{1},W_{2},W_{3},W_{2}W_{1},W_{3}W_{1},W_{3}W_{2})^{{\sf\scriptscriptstyle{T}}}$.
+The $R^{2}$ of $\widehat{Y}$ for $Y$ is 0.5 and 0.4 in Scenarios 2a and 2b, respectively.
+In all settings, we generated two datasets from the same underlying distribution: (i) a training dataset to train the ML model and (ii) an independent analysis dataset for evaluating the PB inference methods containing a labeled dataset and an unlabeled dataset.
+We used 10,000 replications of the analysis dataset for each scenario and a single training set as $\hat{f}$ is assumed to be fixed.
+The size of the training set was 100,000 and the size of the analysis dataset was $n=10,000$.
+We varied the proportion of labeled data in the analysis dataset by letting $n_{\text{\tiny{lab}}}=300,600,1000$ or $2000$ in each setting.
+We evaluated the performance of the PPI and CC methods: $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, and $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$.
+We also included three comparison methods: $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ to serve as the baseline, $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$ to highlight the limitations of naive inference, and a PB inference method from a concurrent pre-print, the PoSt-Prediction Adaptive (PSPA) method.miao2023assumption The PSPA method is a generalization of the PPI++ method, an extension of the original PPI estimator that adaptively minimizes the trace of the asymptotic variance-covariance of the PPI estimator.
+The PSPA estimator, denoted as $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$, is designed to achieve element-wise variance reduction relative to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ by incorporating the information from $\widehat{Y}$ directly into the estimating equation for $\bm{\beta}^{*}$.
+We present the percent bias for each estimator as well as the coverage probability (CP) of the 95% confidence intervals (CIs).
+For the four PB inference methods ($\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$, and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$), we also present the relative efficiency (RE) defined as the ratio of the mean squared error (MSE) of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ relative to the MSE of the PB inference estimator.
 5.2 Results
-5.2.1 Linear Regression
-For brevity, we present results for the lower quality prediction with $R^{2}=0.6$ and the higher quality prediction with $R^{2}=0.8$ in the Appendix.
-The general patterns for the PB inference methods are consistent across the settings, but with higher gains in efficiency and more pronounced differences across the methods.
-Figures 2 - 4 display the results for the three types of errors in the predictions.
-All methods provide coverage close to the nominal level across all scenarios.
-With respect to RE, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ generally have significantly higher MSE than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, with $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ offering some improvement over $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ when the proportion of labeled data is larger.
-In contrast, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI++}}}$, and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ consistently outperform $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ for all parameters and in all scenarios.
-For the prediction with random error, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ perform similarly and have REs exceeding 1.5 for all parameters, while $\hat{\bm{\beta}}^{\text{\tiny{PPI++}}}$ is substantially less efficient for the coefficient for $X_{2}$.
-For the predictions with non-random and covariate-dependent error, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ overall outperforms $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI++}}}$.
-$\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ provides substantial improvement over both methods for the coefficient for $X_{2}$ and over $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ for the coefficients for $X_{1}$ and $X_{3}$.
-In fact, $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI++}}}$ are both outperformed by $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ for the smaller labeled sample sizes for the coefficient for $X_{2}$.
-Figure 2: Coverage probability of the 95% confidence intervals for all methods and relative efficiency for the PB inference methods (PPI, PPIa, PPI++, PSPA, and $C$) for linear regression with a prediction with random error.
-Figure 3: Coverage probability of the 95% confidence intervals for all methods and relative efficiency for the PB inference methods (PPI, PPIa, PPI++, PSPA, and $C$) for linear regression with a prediction with non-random error.
-Figure 4: Coverage probability of the 95% confidence intervals for all methods and relative efficiency for the PB inference methods (PPI, PPIa, PPI++, PSPA, and $C$) for linear regression with a prediction with covariate-dependent error.
-5.2.2 Logistic Regression
-We again present results for the lower quality prediction with overall accuracy of 0.7 and the higher quality prediction with 0.9 accuracy in the Appendix.
-The general patterns are again consistent, but with higher gains in efficiency and more pronounced differences across the methods for the higher quality prediction.
-Figures 5 - 7 display the results for the three types of errors in the predictions.
-All methods provide coverage close to the nominal level across all scenarios.
-With respect to RE, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ underperform relative to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ while $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI++}}}$, and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ generally have MSE much lower than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ in all scenarios.
-For the predictions with random error, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI++}}}$, and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ perform similarly and have REs near 1.2 for all parameters.
-For the predictions with non-random error, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ outperforms $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI++}}}$ for all parameters, offering an additinal 10-20% improvement in efficiency.
-For the predictions with covariate-dependent error, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ performs similarly to $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ for all parameters and is substantially more efficient than $\hat{\bm{\beta}}^{\text{\tiny{PPI++}}}$ for the coefficient for $X_{3}$.
-Figure 5: Coverage probability of the 95% confidence intervals for all methods and relative efficiency for the PB inference methods (PPI, PPIa, PPI++, PSPA, and $C$) for logistic regression with a prediction with random error.
-Figure 6: Coverage probability of the 95% confidence intervals for all methods and relative efficiency for the PB inference methods (PPI, PPIa, PPI++, PSPA, and $C$) for logistic regression with a prediction with non-random error.
-Figure 7: Coverage probability of the 95% confidence intervals for all methods and relative efficiency for the PB inference methods (PPI, PPIa, PPI++, PSPA, and $C$) for logistic regression with a prediction with covariate-dependent error.
+Table 1 presents the percent bias of the various estimators across the different scenarios and labeled dataset sizes.
+With the exception of $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$, the methods exhibit minimal bias as they are all consistent estimators for the target parameter.
+The limitations of naive inference are particularly apparent in Scenarios 2a and 2b, which do not include $X_{1}$ in ML model training.
+$\hat{\bm{\gamma}}^{\text{\tiny{all}}}$ exhibits an average percent bias of -47% in Scenario 2a and -64% in Scenario 2b.
+Table 2 summarizes the CPs of the 95% CIs for all methods across the various scenarios and labeled dataset sizes.
+As expected, $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and the four PB inference estimators ($\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$, and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$) maintain coverage near the nominal level in all settings.
+In contrast, the 95% CIs based on $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$ achieve adequate coverage only in Scenario 1a, where a strong prediction model yields negligible bias in estimating $\beta_{1}^{*}$.
+Coverage drops to 92% in Scenario 1b and falls to 0% in Scenarios 2a and 2b, where bias is substantial.
+These results underscore the risks of relying on naive inference in practice and highlight the necessity of using reliable PB inference methods.
+Figure 2 displays the relative efficiency across the different settings.
+In Scenario 1, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ perform similarly, with RE significantly above 1, indicating that the PB inference methods have lower variance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
+Comparing Scenario 1a and 1b, we note a significant decrease in the gains in statistical efficiency of the PB inference methods due to the weaker prediction model used in Scenario 1b.
+However, all methods are at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, as the sufficient condition in Corollary 1.1 is satisfied.
+Additionally, as the labeled set size increases and hence $n_{\text{\tiny{lab}}}/n$ increases, the efficiency gain decreases as expected.
+$\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is always at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ in Scenarios 1a and 1b, and more efficient in Scenario 1a when the labeled dataset size is larger due to the difference in the augmentation terms of these two estimators.
+Scenario 2 highlights the performance differences between $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ and $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$, which underscore our findings in Corollary 1.1 and Corollary 2.1.
+In both sub-scenarios, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ consistently outperforms both $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ and is substantially more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
+In Scenario 2a, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ achieves an average RE of 1.27 across the various sample sizes, indicating a significant gain over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
+In Scenario 2b, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ again demonstrates superior RE compared to all other methods and remains more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
+$\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ offers a modest efficiency gain in Scenario 2a and comparable performance to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ in Scenario 2b.
+Both $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ are outperformed by $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ in nearly all scenarios.
+In Scenario 2a, while $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ remains comparable to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, the RE for $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is as low as 0.88 with $n_{\text{\tiny{lab}}}=2000$.
+Both $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ demonstrate a substantial loses in efficiency in Scenario 2b, with $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ exhibiting a RE of 0.63 with $n_{\text{\tiny{lab}}}=2000$.
+Figure 2: Relative efficiency of the PB inference methods compared to baseline estimation with the labeled data ($\hat{\bm{\beta}}^{\text{\tiny{lab}}}$).
+A ratio greater than 1 indicates that the PB inference method has a smaller mean squared error (MSE) than that of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
 6 UK Biobank Analysis
@@
 Android fat mass (AFM), measured via a dual-energy X-ray absorptiometry (DEXA) scan, is a key marker for assessing central obesity in clinical practice.
-53 However, due to its high cost, DEXA-derived AFM is usually only measured in a small subset of participants in large-scale studies.
-For example, in the UK Biobank,54 a prospective cohort study measuring genetic, lifestyle, and health information on approximately 500,000 participants in the UK, DEXA-derived AFM is incompletely ascertained, particularly among populations historically underrepresented in genetic research.
-54 For instance, out of $5,397$ South Asian participants in the UK Biobank, only $288$ (5.3%) participants have undergone DEXA scans.
+zhu_dxa-derived_2022 However, due to its high cost, DEXA-derived AFM is usually only measured in a small subset of participants in large-scale studies.
+For example, in the UK Biobank,sudlow_uk_2015 a prospective cohort study measuring genetic, lifestyle, and health information on approximately 500,000 participants in the UK, DEXA-derived AFM is incompletely ascertained, particularly among populations historically underrepresented in genetic research.
+sudlow_uk_2015 For instance, out of $5,397$ South Asian participants in the UK Biobank, only $288$ (5.3%) participants have undergone DEXA scans.
 Our goal is to examine associations between AFM and age, sex, and body mass index (BMI) among the South Asian population.
 Previous research has indicated that this population is at higher risk of central obesity and that it is associated with cardiometabolic diseases at lower BMI thresholds than European populations.
-55 Since DEXA-derived AFM is missing for most South Asian participants, it provides a natural opportunity to leverage ML-predicted AFM in downstream analyses.
+sniderman_why_2007 Since DEXA-derived AFM is missing for most South Asian participants, it provides a natural opportunity to leverage ML-predicted AFM in downstream analyses.
 To illustrate the impact of different ML models on the efficiency of the PB inference methods, we trained two random forest models in the White British population ($32,460$ participants) and used them to predict missing AFM values in the South Asian population.
@@
 The strong prediction model achieved an $R^{2}$ of 0.8 while the weak prediction model had a substantially lower $R^{2}$ of 0.5 in the South Asian subpopulation.
-Table 1 presents point estimates, standard errors, and RE, defined as the ratio of the variance of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ to the variance of the given estimator, across the two prediction model settings for five estimators: the baseline estimator ($\hat{\bm{\beta}}^{\text{\tiny{lab}}}$), the PPI estimators ($\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI++}}}$), the PSPA estimator ($\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$), and the CC estimator ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$).
+Table 3 presents point estimates, standard errors, and RE, defined as the ratio of the variance of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ to the variance of the given estimator, across the two prediction model settings for five estimators: the baseline estimator ($\hat{\bm{\beta}}^{\text{\tiny{lab}}}$), the PPI estimators ($\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$), the PSPA estimator ($\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$), and the CC estimator ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$).
+We do not include naive inference in our analyses due to the limitations demonstrated in our simulation studies.
 Echoing our theoretical and simulation study results, all methods yield similar point estimates as they are all consistent estimators of the same parameter.
@@
 Consequently, both $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ outperform $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
-In fact, all of the PB inference methods ($\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI++}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$, and $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$) yield substantially smaller standard errors compared to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
-Among the PB inference methods, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI++}}}$, and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ show comparable performance, with average REs of 1.39, 1.39, 1.44, and 1.47, respectively.
+In fact, all four PB inference methods ($\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$) yield substantially smaller standard errors compared to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
+Among the four PB inference methods, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ show comparable performance, with average REs of 1.39, 1.39, and 1.47, respectively.
 $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ exhibits higher efficiency gains with an average RE of 1.68.
@@
 Here, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ yield larger standard errors than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ for all parameters of interest, with an average RE of 0.52.
-In contrast, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$, as well as $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI++}}}$, outperform $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, with average RE values of 1.28, 1.29, and 1.16, respectively.
+In contrast, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$, as well as $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$, outperform $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, with average RE values of 1.28 and 1.16, respectively.
 Overall, the CC estimator is the best performing method in terms of efficiency, underscoring its utility in practice.
 7 Discussion
-In this paper, we expand on existing research into the efficiency of PB inference methods by making the following contributions:
-  1. 1.
-    We employed the AIPW framework to introduce the CC estimator from the double-sampling literature as a simple modification of the PPI estimator, offering theoretically guaranteed efficiency gains over classical inference.
-Additionally, we demonstrated that the CC estimator is the most efficient estimator within the class of weighted linear augmentation estimators.
-  2. 2.
-    We made numerous connections between PB inference and existing literature with direct implications for future methodology and applications.
-Of note, we provided a set of results relating newly proposed PB inference methods to the CC method.
-For linear regression, we showed that the PPI++ and PSPA estimators are asymptotically equivalent to weighted linear augmentation estimators and therefore the CC estimator is always at least as efficient as these two methods.
-This theoretical finding was corroborated in our simulated and real data analyses.
-  3. 3.
-    We conducted the most extensive simulation study to date comparing existing PB inference approaches.
-We demonstrated empirically that the CC estimator outperforms the PPI++ and PSPA estimators in numerous simulation scenarios as well as in an analysis of UK Biobank data.
-However, our work is not without limitations.
-First, we focused our analyses on Z-estimation and further work is needed to analyze other statistical procedures.
+In this paper, we bridge classical statistical methods with a modern problem to equip practitioners and methodologists with a rigorous approach for PB inference.
+We find that incorporating a simple weight in the PPI estimator previously proposed for studies with double-sampling designschen2000unified yields valid inference that is guaranteed to be at least as efficient as estimation using only labeled data.
+Our work addresses efficiency shortcomings of the original PPI estimator which can undermine its practical value.
+Alongside theoretical results supporting these findings, our simulation studies and application to UK Biobank data verify that the PPI method can suffer substantial efficiency loss when the ML model is inaccurate.
+Additionally, we draw important connections to several established approaches from the economics and statistics literature to provide context for ongoing developments in PB inference.
+Our work is not without limitations.
+First, we focused our analyses on least squares estimation following Motwani and Witten.motwani2023revisiting Further work is needed to analyze other statistical procedures (e.g., Z-estimation problems in general).
 Second, we assumed that the labeled and unlabeled data are drawn from the same underlying distribution or equivalently, that $Y$ is missing completely at random.
@@
 However, the CC estimator can be extended to settings where $Y$ is missing at random by incorporating inverse probability weighting and regression adjustment.
-Third, our current analysis assumes that only one prediction is available for analysis.
-In many practical scenarios multiple predictions with unknown accuracy are available to the analyst.
-The CC estimator can be naturally extended to accommodate multiple predictions and we provide a simulated example in the Appendix.
 Lastly, we focused on the traditional missing data scenario where the proportion of labeled data, $n_{\text{\tiny{lab}}}/n$, is bounded away from 0.
@@
 Code availability:
-Our analyses can be replicated with the code at: https://github.com/jlgrons/PBInf-Simulations.
+Our analyses can be replicated with the code at: https://github.com/SelinaS37/PBInference.
 UK Biobank Dataset:
@@
 Gronsbell gratefully acknowledges funding from an NSERC Discovery Grant (RGPIN-2021-03734), the Connaught Fund, the University of Toronto Data Science Institute, and the McLaughlin Center for Genetic Research.
+Labeled set size
+Method 300 600 1000 2000
+Scenario 1a
+   $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$ -0.110 -0.110 -0.110 -0.110
+   $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ -0.006 0.014 0.003 -0.020
+   $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ -0.029 -0.017 -0.014 -0.001
+   $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ -0.028 -0.015 -0.012 -0.005
+   $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ -0.029 -0.015 -0.012 -0.005
+   $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ -0.025 -0.014 -0.011 -0.004
+Scenario 1b
+   $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$ -0.245 -0.245 -0.245 -0.245
+   $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ 0.034 -0.004 0.010 0.005
+   $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ 0.012 -0.038 -0.015 -0.016
+   $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ 0.013 -0.036 -0.013 -0.012
+   $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ 0.014 -0.035 -0.012 -0.012
+   $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ 0.008 -0.035 -0.012 -0.011
+Scenario 2a
+   $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$ -47.017 -47.017 -47.017 -47.017
+   $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ -0.034 -0.027 -0.011 0.001
+   $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ 0.025 0.015 0.024 0.018
+   $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ 0.023 0.013 0.021 0.015
+   $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ -0.002 -0.004 0.005 0.009
+   $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ -0.095 -0.048 -0.019 0.003
+Scenario 2b
+   $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$ -63.899 -63.899 -63.899 -63.899
+   $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ -0.039 -0.013 -0.013 -0.006
+   $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ 0.019 0.024 -0.008 0.007
+   $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ 0.017 0.022 -0.008 0.004
+   $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ -0.036 -0.010 -0.018 -0.006
+   $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ -0.081 -0.030 -0.025 -0.009
+Table 1: Percent biases of the PPI estimators ($\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$), the CC estimator ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$), the PSPA estimator ($\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$), the baseline estimator ($\hat{\bm{\beta}}^{\text{\tiny{lab}}}$), and the naive estimator ($\hat{\bm{\gamma}}^{\text{\tiny{all}}}$) for $n=10,000$ and various labeled dataset sizes.
+Labeled set size
+Method 300 600 1000 2000
+Scenario 1a
+   $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$ 0.947 0.947 0.947 0.947
+   $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ 0.951 0.953 0.949 0.951
+   $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ 0.947 0.947 0.952 0.949
+   $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ 0.947 0.947 0.951 0.947
+   $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ 0.944 0.945 0.951 0.946
+   $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ 0.949 0.946 0.951 0.948
+Scenario 1b
+   $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$ 0.916 0.916 0.916 0.916
+   $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ 0.949 0.952 0.950 0.949
+   $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ 0.948 0.949 0.947 0.947
+   $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ 0.947 0.948 0.947 0.946
+   $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ 0.946 0.947 0.947 0.945
+   $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ 0.949 0.949 0.948 0.946
+Scenario 2a
+   $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$ 0.000 0.000 0.000 1 0.000
+   $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ 0.950 0.948 0.949 0.949
+   $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ 0.946 0.946 0.946 0.948
+   $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ 0.948 0.953 0.956 0.963
+   $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ 0.945 0.948 0.946 0.949
+   $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ 0.943 0.944 0.946 0.946
+Scenario 2b
+   $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$ 0.000 0.000 0.000 0.000
+   $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ 0.928 0.933 0.929 0.935
+   $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ 0.944 0.950 0.954 0.949
+   $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ 0.949 0.957 0.965 0.973
+   $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ 0.939 0.945 0.946 0.948
+   $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ 0.936 0.942 0.943 0.947
+Table 2: Coverage probabilities of the 95% confidence intervals based on the PPI estimators ($\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$), the CC estimator ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$), the PSPA estimator ($\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$), the baseline estimator ($\hat{\bm{\beta}}^{\text{\tiny{lab}}}$), and the naive estimator ($\hat{\bm{\gamma}}^{\text{\tiny{all}}}$) for $n=10,000$ and various labeled dataset sizes.
+Strong Prediction Model
+Intercept Age Sex BMI
+$\hat{\bm{\beta}}^{\text{\tiny{lab}}}$
+EST -4.857 -0.003 0.813 0.173
+SE 0.3080 0.0041 0.0702 0.0087
+RE 1.00 1.00 1.00 1.00
+$\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$
+EST -4.528 -0.007 0.901 0.165
+SE 0.2602 0.0036 0.0628 0.0070
+RE 1.40 1.36 1.25 1.54
+$\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$
+EST -4.541 -0.007 0.897 0.165
+SE 0.2596 0.0035 0.0628 0.0069
+RE 1.41 1.36 1.26 1.56
+$\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$
+EST -4.454 -0.009 0.888 0.167
+SE 0.2747 0.0036 0.0592 0.0063
+RE 1.22 1.36 1.40 1.87
+$\hat{\bm{\beta}}^{\text{\tiny{CC}}}$
+EST -4.553 -0.006 0.903 0.163
+SE 0.2590 0.0034 0.0570 0.0057
+RE 1.41 1.53 1.52 2.27
+Weak Prediction Model
+Intercept Age Sex BMI
+$\hat{\bm{\beta}}^{\text{\tiny{lab}}}$
+EST -4.857 -0.003 0.813 0.173
+SE 0.3080 0.0041 0.0702 0.0087
+RE 1.00 1.00 1.00 1.00
+$\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$
+EST -4.557 -0.006 0.836 0.166
+SE 0.4175 0.0060 0.0982 0.0118
+RE 0.54 0.48 0.51 0.54
+$\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$
+EST -4.572 -0.006 0.835 0.166
+SE 0.4170 0.0059 0.0981 0.0117
+RE 0.55 0.49 0.51 0.54
+$\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$
+EST -4.771 -0.001 0.835 0.161
+SE 0.2978 0.0044 0.0667 0.0070
+RE 1.07 0.90 1.11 1.55
+$\hat{\bm{\beta}}^{\text{\tiny{CC}}}$
+EST -4.790 -0.004 0.822 0.172
+SE 0.2941 0.0038 0.0652 0.0068
+RE 1.10 1.21 1.16 1.64
+Table 3: UK Biobank analysis of android fat mass based on the baseline estimator ($\hat{\bm{\beta}}^{\text{\tiny{lab}}}$), the PPI estimators ($\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$), the PSPA estimator ($\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$), and the CC estimator ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$).
+The strong prediction model has an $R^{2}$ of 0.8, while the weak prediction model has an $R^{2}$ of 0.5.
+Estimates (EST), standard errors (SE), and relative efficiency (RE) are reported to evaluate performance.
+RE is defined as the ratio of the variance of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ to the variance of the given estimator.
+An RE higher than 1 indicates that a method yields a smaller variance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.